A Complete Bibliography of Publications in Ecological Modelling (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

03 July 2024 Version 1.44

Title word cross-reference

1 [WSW23]. $142.20 [Pas21]. 3 [CGVC21, PMP^{+}21, PM22, VRRI^{+}24]. $84.99 [Mea20]. _{10} [MBMV20, NSS^{+}23, SBF^{+}21]. _{2} [AAR23, CCH20, GAMN^{+}23, HKSM23, PG23, YLLL23, ZFT20, ZBH^{+}22]. _{4} [GAMN^{+}23]. _{mix} [XLS20]. a [LKP^{+}20, WCSX22]. ϵ [WSW23]. I [RAB21]. k [WSW23]. r/K [MF20]. R_{0} [WHE^{+}24a, WHE^{+}24b].

-D [PM22, WSW23].

0304 [SGTS23]. 0304-3800 [SGTS23]. 0304-3800/109690 [SGTS23].

1 [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20g, Ano20f, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21g, Ano21f, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, MPL^{+}20]. 109165 [CVTS20a]. 109198. [SE20a]. 109265 [HES^{+}23]. 109690 [SGTS23]. 12/110244 [AAAR24]. 15 [AGS^{+}22, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20s, Ano20r, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21s, Ano21r, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x].

2 [CSZY^{+}20, Cor20, MLP^{+}20]. 2.0 [KMH^{+}21]. 2020 [Ano20a, Ano20m, Ano20b, Ano20n, Ano20c, Ano20o, Ano20d, Ano20p, Ano20e, Ano20q, Ano20g, Ano20s, Ano20f, Ano20r, Ano20h, Ano20t, Ano20i, Ano20u, Ano20j, Ano20v, Ano20k, Ano20w, Ano20l, Ano20x]. 2021 [Ano21a, Ano21m, Ano21b, Ano21n, Ano21c, Ano21o, Ano21d, Ano21p, Ano21e, Ano21q, Ano21g, Ano21s, Ano21f, Ano21r, Ano21h, Ano21t, Ano21i, Ano21u, Ano21j, Ano21v, Ano21k, Ano21w, Ano21l, Ano21x]. 2022 [Ano22a, Ano22b, Ano22c, Ano22p, Ano22q, Ano22s, Ano22r, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, PKB^{+}23]. 2023 [Ano23a, Ano23b, Ano23c, Ano23p, Ano23q, Ano23s, Ano23r, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, RFHR24]. 2024 [Ano24a, Ano24b, Ano24k, Ano24l, Ano24n, Ano24m, Ano24o, Ano24p]. 21st [YCK20]. 239 [DMdR24].

3 [Mea20, NYH20, SNM^{+}23, XLS20]. 3-PG [NYH20, SNM^{+}23, XLS20]. 359 [BS20]. 3800/109690 [SGTS23]. 3Worlds [GDF22].

4 [Pas21]. 431 [CVTS20a, SE20a]. 435 [HES^{+}23]. 443 [BdBG^{+}21b]. 457 [AGS^{+}22]. 458 [SGTS23]. 471 [KHH^{+}24]. 476 [AAAR24]. 489 [WHE^{+}24a].

978 [Mea20, Pas21]. 978-0-7503-2430-4 [Pas21]. 978-3-030-24528-3 [Mea20].

ABACUS [ACZ^{+}21]. abalone [DLZ^{+}21]. Abatus [APGD^{+}21]. Abies [SKW23, WYBX23]. Ability [EHS^{+}20]. abiotic [DAB^{+}20]. above [MK21]. aboveground [HCL^{+}24, HLH^{+}21, MMdP^{+}21]. absence [CIL^{+}20]. absences [FSY22]. abstract [SvV21]. abundance [CDJ^{+}22, CPTD22, FGM^{+}24, HLZ^{+}22, KLNMG22, PSM22, SWF21, SDP23, WWP20, YCI20]. Acadian [SLLW21]. Acari [AFdSM23]. accelerated [ABH^{+}20]. acceleration [KDC^{+}21]. accelerometers [DBH^{+}22]. access [PF23]. acclimation [TMNA20]. according [MOS21]. Accounting [JS22, BARF20, BDA^{+}20, CAS^{+}24, DS24, HGBK23, JJN22, MBFS21, NND^{+}22]. accretion [BMAME^{+}23]. accumulated [SSM24]. accumulation [BMAME^{+}23, BRBT22, CJM^{+}23, JMPB24, SLHB^{+}23]. Accuracy [Pie23, ALJ^{+}22, BBIP24, GWH20, MVAD20, WHdBB21, YCI20]. accurately [GSMD24]. achieve [VRW^{+}21]. acid [BPL^{+}20, WLQ^{+}23]. acidification [RRCZ20]. acoustic [GSH^{+}22, LDR^{+}20, PDD^{+}22]. acquisition [RRP^{+}23]. across [ATP20, ASL^{+}20, CKK^{+}24c, EHS^{+}20, EVT^{+}21, GCS^{+}21, LMW^{+}23a, LPHO^{+}20, MKS21, MF20, RFJ^{+}22, RLS^{+}24, YYS^{+}22]. action [SRÁMH^{+}20]. active [dRMJdS23, VMM^{+}21]. activity [ABH^{+}20, CDB^{+}20]. acuminata [PCN^{+}23]. acute [MMS^{+}24]. adapt [CN20]. Adaptability [KHS^{+}23]. adaptation [GHK23, MPA^{+}24, RMK21]. Adaptations [Koo24b]. Adapting [CMGD^{+}22, ERP21, OTR^{+}21]. Adaptive [JST^{+}22, AVP^{+}23, ASKR24, SLC^{+}23, SGPSzC22, SC24b]. adding [MVAD20]. Addition [KKF^{+}23, SMM^{+}22, ZMF^{+}20]. additive [CIL^{+}20, KWC24, MEPJ21, UGSG24, WSL^{+}22]. Addressing [RAC^{+}24, AAR23]. adjacent [CAFT22]. adjusted [ALJ^{+}22]. adjusting [TBZ^{+}22]. ADM1 [PFM^{+}21]. Administrative [BM22a]. adsorption [MTT22]. adults [AJJ24]. advanced [COJ23, KHS^{+}23, LE23]. Advancing [KAT^{+}22]. advantage [WZLC24]. advection [FMRMAS24]. Aedes [ABZ22, HLB^{+}20, VMM^{+}21]. aerial [DP20]. aestivum [GGS^{+}22]. affect [CLPZ20, GMM^{+}24, MJS24, PNGESS21, PCL24, VSB22]. affected [LSS^{+}21, PAPŁ^{+}23]. affecting [LDXJ22, SOP22, YG21, ZYM^{+}21]. affects [AHH22, Che22, GZW^{+}20, SGCFB22, WZWF20]. affordability [ZHC23]. Afforestation [WXT^{+}20, MTBB22]. Africa [COS23, LAK^{+}23, MDB^{+}20]. African [DDG^{+}21, MTW^{+}22, MD23, WMA21]. africana [DDG^{+}21]. after [LZL22]. Agadir [RKN^{+}21]. against [MKN23]. Age [FGM^{+}24, BMM22, DIS23, MMC^{+}21, PSBE23, RRSG21, XFK21]. age-0 [PSBE23]. Agent [BdBG^{+}21a, CNNS^{+}21, GGMR20, KLPJ24, MMMS23, NMC^{+}20, PR21, AGS^{+}21, AGS^{+}22, ASKR24, BGK^{+}20, BAM23, BTC^{+}22, CXZ^{+}23, CDNN^{+}20, CSB22, CSMP21, DDG^{+}21, DBN^{+}22, EBMK20, HBC^{+}22, JBF22, KE24, KS22, LTR^{+}24, LE23, LL22, LZW20, MSLMM22, PBS21, STQ^{+}22, SJA^{+}22, SLM24, TLL^{+}23, VVH^{+}20, WLY^{+}23, WWF^{+}23, ZMG^{+}21, ZOJ^{+}23, ZL24, BdBG^{+}21b]. agent- [VVH^{+}20]. Agent-based [BdBG^{+}21a, CNNS^{+}21, GGMR20, KLPJ24, MMMS23, NMC^{+}20, PR21, AGS^{+}21, AGS^{+}22, ASKR24, BGK^{+}20, BAM23, BTC^{+}22, CXZ^{+}23, CDNN^{+}20, CSB22, CSMP21, DDG^{+}21, DBN^{+}22, EBMK20, HBC^{+}22, KE24, KS22, LTR^{+}24, LE23, LL22, LZW20, MSLMM22, PBS21, STQ^{+}22, SJA^{+}22, SLM24, TLL^{+}23, WWF^{+}23, ZMG^{+}21, ZOJ^{+}23, ZL24, BdBG^{+}21b]. agents [CDC^{+}22]. AgentSeal [CNNS^{+}21]. ages [Jia21]. agglomerations [MFZ23]. aggregate [JSD^{+}22]. Aggregating [PDD^{+}22, NPK^{+}23]. aggregation [XLM^{+}23, YDLT24]. agri [OEC24]. agri-environment [OEC24]. Agricultural [Pas21, PHL20, RGD20, BAR24, COJ23, DYMD22, JMPB24, MAD^{+}20, MEGA21, MGdM^{+}20, MTT22, NMV^{+}23, NBMD21, SJA^{+}22, TLC^{+}22, TQM^{+}24, TCS21, TSBE23, TK20b, USA21, VBH20, WGMR23, YG21, ZBW^{+}24]. agriculture [LPA^{+}21, MFPIBGES23, SD24, SLY^{+}21, WrJP^{+}22]. agro [KMH^{+}21, PBF^{+}23, TBCV22, TBZ^{+}22]. agro-ecosystem [PBF^{+}23, TBCV22]. agro-food [KMH^{+}21]. agro-pastoral [TBZ^{+}22]. Agroecological [KKP^{+}21, MAD^{+}20, OSPZ22]. agroecosystem [BMRF21, CJM^{+}23]. agroecosystems [ZMK^{+}22]. agronomic [BCMV20]. AHP [SD24]. AHP-based [SD24]. aimed [CM21, Yao21]. air [CZ21, FWK^{+}20, LBB20, SBF^{+}21]. airborne [Laz23]. akashiwo [SIOM24]. al [NBU^{+}23, RFHR24]. Alaska [LMW^{+}23b, WA20]. alba [SKW23]. albacares [DPC20]. albopictus [ABZ22, HLB^{+}20, VMM^{+}21]. album [GRAA22]. Alces [KAMS22]. Aleppo [BAZ22]. Alewife [NGA^{+}20]. algae [KNP^{+}21, KBAC20]. Algal [XZZ^{+}22, AWW24, BS21, CMKF22, CLH23, DVVH21, KDH^{+}21, KSJ22, PTG^{+}20, RFPRC23, ZGG22, ZWZ^{+}20]. algorithm [Gat21]. algorithmic [dSMPA^{+}22]. algorithms [CAK^{+}24, SDH21, ZMF^{+}20]. alien [SBS^{+}20, SBS^{+}21]. Allee [VBH20]. alleviating [MPA^{+}24]. allocation [ABRG21, nCLpW^{+}23, LBS20, NAM22]. Allometric [KKH^{+}24]. Allometry [Abe20]. Allosaurus [KHR23, PR23]. Alluvial [WWF^{+}23]. along [BMS^{+}20, HK23, Lef23, LBGD21, LM21, NSS^{+}23, OMP^{+}23, PA20, SGTS21, SGTS23]. Alosa [NGA^{+}20]. alpine [ALS^{+}21, LGK^{+}23, LGFT^{+}20]. alter [PTD^{+}23]. altered [CWY^{+}24, GRCSA^{+}20, KNP^{+}21, LM21]. alternative [JvVvHH24, PTBFP^{+}20, SD24]. alternatives [Koo20b]. altissima [dOPdMdS^{+}21]. altitudinal [ZLL^{+}23]. altricial [CC20]. always [GAC^{+}21, Pér24]. Amazon [dOPdMdS^{+}21, RFF^{+}23]. Amazonia [JMC^{+}22]. Amazonian [RRP^{+}23]. Ambient [XWZ^{+}22]. Amblyseius [TCG^{+}20]. America [PMB^{+}22]. American [ARSC22, BHC22, HBMR24]. americanus [BHC22]. Ammospiza [HBP^{+}21]. among [ABM20, CKK^{+}24a, LLZM20, Oku21, RBGB23]. Amount [MLV24, DL21]. Ampelisca [DZBB24]. Amphipoda [DZBB24]. amplification [CLHZ22]. anadromous [DSHB23]. analogues [PR21]. analyses [BGCJ21, HKG^{+}24, MBG^{+}22, SC21]. analysing [HHFB23, Jag20]. Analysis [AP22, BDP^{+}23, CP20, RUTY23, USA21, WF24, WAHA^{+}24, AG22, AHO^{+}20, AAC20, BPB^{+}20, BCW^{+}23, BBSHC22, BLSR^{+}20, CWF^{+}21, CGVC21, DPA21, DS24, ESCA22, EGAS23, ENARM^{+}23, FEPR24, HRW^{+}24, HCL^{+}24, HEL^{+}21, HLM^{+}20, IA23, JGG23, JLLS23, KT22, KKF^{+}23, KHS^{+}23, KUK^{+}21, LL22, LLZM20, LDZ21, LZL22, LHJ^{+}24, LM23, LLW^{+}21, LAC^{+}24, MFZ23, MPC^{+}22, MMRBLN20, MLV24, NGA^{+}24, NDNN22, NPS^{+}23, OFH^{+}20, PK23, PMZ^{+}21, PDW23, PFM^{+}21, PHK23, PPSS22, QS23, RGB^{+}23, SQDM21, SDSP^{+}21, TGB^{+}23, TK20a, TQM^{+}24, TK20b, WWC^{+}22, WSHZ24, WFW^{+}24, WCF^{+}22, Yao21, ZMK^{+}22, ZWL^{+}21, ZZWF21, dGZHCA24, vDC20, BTC^{+}22]. Analytical [CLH23, JZHT23, Mas23, NND^{+}22, UGSG24]. Analyzing [RBEC21, SGPSzC22]. Angelica [ZWL^{+}21]. Animal [RLS^{+}24, ABB23, Ano20-50, AARS23, BLB24b, BVR^{+}20, BTC^{+}22, CDNN^{+}20, DBH^{+}22, JGG23, JVGC21, KLA^{+}20, KHH^{+}22, KHH^{+}24, KLNMG22, MSG23, MKS21, WMA21, dKS22]. animal-borne [DBH^{+}22]. Animal-vectored [RLS^{+}24]. animals [HI22]. ANN [LNK^{+}21]. Announcing [Fat21]. annual [HH23, RLM23]. answer [CBSB23]. antagonistic [LBL22, YL22]. Antarctic [BMB21, GSM^{+}20]. anthesis [KFM^{+}23]. Anthropocene [SMP20]. anthropogenic [MCMP^{+}24, MFH^{+}22, MAB23, SGRDA^{+}24]. antipatharian [MTL^{+}23]. Antipathella [MTL^{+}23]. APEX [AJG22b, PR21, AJG22a]. aphid [BPW^{+}21]. aphids [Hop23]. Apis [JSL^{+}20, LBT^{+}22]. Apostle [SOV21].

App [dSMPA^{+}22]. appearance [PTF23]. Application [AG22, BGA^{+}23, CMGD^{+}22, CJM^{+}23, DMG22, JB22, LCP^{+}20, MDB^{+}20, RWF22, RFCEL22, RFPRC23, SRB^{+}20, YHC^{+}21, YCI20, dMT23, AM22, ABR24, AACCIS20, BSPT23b, CH22, CAH22, DBV^{+}22, DP20, GSC^{+}23, HLS^{+}23, JZJ^{+}20, KRM^{+}22, MKR20, MEGA21, MFF^{+}24, MPA^{+}24, MKB^{+}24, OKS^{+}21, PG23, PTRP23, RJY^{+}20, RRSG21, WYJ^{+}22, ZZZD22, dVDC20, DYMD22, MTP^{+}20]. Applications [MLD^{+}23, BAR20, DGSM21, GRS^{+}20, KCP^{+}23, Laz23, MBMV20, PPSS22, RAC^{+}24, YRK^{+}20]. applied [BM22b, CAS^{+}24, KABL21, NMGFOB21]. apply [CXZ^{+}23]. applying [BVR^{+}20, RAB21]. Approach [JZS^{+}21, LB21, AZBdAI23, BdCCdSAdS23, BLB^{+}20, CB20, CMS^{+}21, DIS23, DPA21, DSC^{+}21, DPC20, EON23, FCFCM24, FFM^{+}23, FMRMAS24, GHbS21, GZK^{+}20, HBP^{+}21, HKB^{+}20, HBRK24, ILLM20, KTB^{+}23, KWG^{+}22b, KGBL24, Kos22, LTR^{+}24, LG20, LL22, LL21, LLB^{+}23, Lov24, ŁJ21, LYM^{+}23, dSMPA^{+}22, MDP21, MTBB22, MVC^{+}22, MSS24, Mur24, MKB^{+}24, NA22, NAM22, RK20, RBMRR20, RLdM20, SBS^{+}21, SRÁMH^{+}20, SLYL^{+}20, SCB^{+}20, SRTS20, SWF21, ST24, SWW^{+}23, SKH^{+}20, SZL^{+}20, TPD^{+}24, TSdOCV21, TSBE23, WST^{+}23, WBD^{+}22b, ZJEG24, dGZHCA24, dAWG23]. approaches [DTR^{+}22, HPMB20, LBB20, RBEH21]. approaching [PDG^{+}20]. appropriate [Mas23]. Approximate [BSR^{+}23]. approximation [WSL^{+}22]. April [Ano20a, Ano20m, Ano21a, Ano21m, Ano22a, Ano23a, Ano24a]. APSIM [CHR^{+}23, EHS^{+}20, HRW^{+}24, OHH^{+}21]. APSIM-Potato [OHH^{+}21]. APSIM-wheat [HRW^{+}24]. Apulia [GSC^{+}23]. AquaCrop [PAdS^{+}24]. Aquaculture [VLGG^{+}20, FMH^{+}20, GLA^{+}20, HLS^{+}23, HK23, LUV^{+}23, LLC^{+}24, WT20, YRK^{+}20]. aquatic [ADR21, DK22, KE24, LYC^{+}21, NKT^{+}20, OKS^{+}21, RBEH21, SKH^{+}20, Yao21]. aquatic-terrestrial [OKS^{+}21]. aqueous [MTT22]. arable [HSP^{+}22]. arboviral [LDS^{+}23]. archipelagos [RMV^{+}24]. architectures [WrJP^{+}22]. archival [HBMR24]. Arctic [DSHB23, JOE23, TLL^{+}23]. Arctodiaptomus [TBP^{+}23]. Area [DMdR24, WXT^{+}20, XWX^{+}24, AFdSM23, BBCD23, DHW^{+}20, GJC^{+}20, HKSM23, HH23, IA23, JS22, JHL23, LLW^{+}21, MZLRSZ21, OBFH22, RPT21, SBF^{+}21, SRK^{+}20, ST24, VPV21, WWC^{+}22, ZLZ^{+}24, MTP^{+}20, RGD20]. areas [AXMJ22, ALS^{+}21, GCS^{+}21, KT22, MRR^{+}20, PWN^{+}24, PAS^{+}22, RBLS23, RFdL^{+}24, RRFLV^{+}22, SZPW23, VRW^{+}21, ZBW^{+}24]. Argentina [AZMD^{+}22]. Argentinean [BAZ22]. arguments [AAM^{+}22]. arid [FS22, GLK^{+}22, LZF^{+}24, LHJ^{+}24, MPSS24, NWO21, PAdS^{+}24, RSASP22, RS23, WXX^{+}23, WFW^{+}24]. aridity [HH23]. arise [CR24]. arising [AP22]. armed [Gat21]. aromatic [ZYZ22]. arrangement [CGB^{+}23]. arrangements [PAdS^{+}24]. arrays [PDD^{+}22]. Arrhenius [MBMV20]. Artemia [KDC^{+}21]. Artemisia [SZL^{+}21]. arthropod [VMM^{+}21]. arthropods [ZBW^{+}24]. article [Fat21]. artificial [BSPT23a, BSPT23b, BLB^{+}20, DLY^{+}22, MS20, SNM^{+}23]. ascendency [LZSQ20]. Asia [SKPL24, ZLLZ24, ZFC^{+}23]. asiaeorientalis [DZC^{+}23a, DZC^{+}23b]. Asian [NO24]. Assembling [CdCAC^{+}20]. assembly [QZSV21]. assess [CMS^{+}21, FDSM20, JSL^{+}20, LGFT^{+}20, RRFLV^{+}22]. assessed [BPL^{+}20]. Assessing [BARF20, CHR^{+}23, CLW23, GSA24, HHBF24, JPMA^{+}23, LLB^{+}23, LHJ^{+}24, OHH^{+}21, PR22, RMV^{+}24, SCB^{+}20, TMFV22, TCS21, WA20, WWCT24, CC23, DBK20, DBDR22, JSD^{+}22, LTR^{+}24, MGK^{+}23, MM21, ZOJ^{+}23]. Assessment [BZZT22, GRCSA^{+}20, ILLM20, AFG24, ACZ^{+}21, BM22b, DIS23, EOÖ21, For24, FLG^{+}23, GMM^{+}24, HCL^{+}24, KBSS23, LBI^{+}20, LGK^{+}23, LLW^{+}21, LCMB21, MPC^{+}22, MBFS21, MFH^{+}22, MSS24, OZSU22, PBF^{+}23, RSASP22, RBLS23, RS23, RFHR23, SLY^{+}21, TBCV22, TWL^{+}22, TKO^{+}22, VPAV22, WAHA^{+}24, XWX^{+}24, Yao21, ZHJQ23, ZYM^{+}21]. assessments [CB20]. assisted [MKEH20]. association [LW23, PPSS22]. associations [KCS22]. assumptions [BHC22]. asynchrony [HSL^{+}22]. Athias [TCG^{+}20]. Athias-Henriot [TCG^{+}20]. Atlantic [dSRDC^{+}23, DTD^{+}23, HK23, RKN^{+}21, VRRI^{+}24]. Atlantis [SHF24, BLSR^{+}20, CBG^{+}22, PKB^{+}23, dGZHCA24]. atmosphere [BLB^{+}20, BGC^{+}24, GXZ^{+}24, KKP^{+}21]. attention [PDW23]. attentional [vWVH20]. attenuation [SBWB24]. audience [CVTS20a, CVTS20b]. augmentation [RMA^{+}24]. augmented [LAK20]. August [Ano20b, Ano20n, Ano21b, Ano21n, Ano22b, Ano23b, Ano24b]. Aurelia [YSFX23]. aurita [YSFX23]. Australia [CH22, FEPR24, YKZ^{+}23]. australis [YXY^{+}20]. author [CYP^{+}20]. autochthonous [LDS^{+}23]. autocorrelated [DMM^{+}23]. automata [JB22, LZW20, SLC^{+}20, ZHWZ23]. Automatic [LKP^{+}20]. automaton [YXY^{+}20]. automobile [IA23]. Autoregressive [DMG22]. availability [BLLL21, CLW23, JS22, KAB^{+}23b, NWO21, VVH^{+}20, WSN^{+}21, ZHC23]. Available [RBVL^{+}20]. average [PSG^{+}22]. averaged [Che21, HMC20]. avian [WSN^{+}21]. Avicennia [LK20]. avifauna [ADR21]. avoid [Koo24a, Uch23]. avoidance [CLH23, PF23].

B [DWH^{+}21]. backed [vdMvDSL20]. Background [SC24a, RPT21, SBMG24, WSM24, XWZ^{+}22]. Bacteria [BMBB24]. bacterial [ZL24]. bacterium [LPSF24]. Baetidae [KFJS20]. baited [KCHZ21]. Baiyangdian [ZYZ22]. balance [HCP^{+}20, LGA20, RPARM20, ZFT20]. balanced [YL22]. Balearic [EWE20]. ballistics [PM22]. Baltic [BNN20, BNAR^{+}21, SRK^{+}20]. bamboo [ZCJ^{+}20]. banana [CMSD^{+}22]. bandit [Gat21]. bank [BGR20, LKO20]. bar [Tsy22]. barbarum [SWHZ22, WSHZ24]. Barcelona [MTP^{+}20]. Barents [NHGP23]. bargaining [SACE^{+}20]. barnyardgrass [MVC^{+}22]. Barrier [ABR24, IA23]. Barro [FG21]. Based [LB21, Por21, ASHK22, AADHA^{+}23, AAM^{+}22, AGS^{+}21, AGS^{+}22, AWW24, ASKR24, AHO^{+}20, APGD^{+}21, AAC20, BdSG^{+}23, BMB24, BDC^{+}22, BPN^{+}20, BGS^{+}20, BGK^{+}20, BSCH23, BdBG^{+}21a, BAM23, BM21, BTC^{+}22, CMHZ23, CXZ^{+}23, CVTS20a, CVTS20b, CDB^{+}20, CAK^{+}24, CDNN^{+}20, CNNS^{+}21, CGVC21, CPTD22, CSB22, CSMP21, CLG20, DK22, DPA21, DSC^{+}21, DDG^{+}21, DBN^{+}22, DAB^{+}20, DWG22, DWH^{+}21, ESCA22, EWE20, EBMK20, EV22, EL22, ENARM^{+}23, GSA24, GGMR20, GRCSA^{+}20, GSRB22, HBMR24, HLB^{+}20, HKG^{+}24, HBC^{+}22, HYY^{+}24, HEL^{+}21, HGBK23, HBRK24, HCPG20, HFP20, HFP22, JvVvHH24, JHL23, KFM^{+}23, KCP^{+}23, KE24, KS22, Koo24a, KWG^{+}23, KTB^{+}21, KLPJ24, KUK^{+}21, LTR^{+}24, LGFT^{+}20, LE23, LL22, LZSQ20, LBI^{+}20, LDZ21, LJP^{+}22, LQH24, LZW20, LCW^{+}21, Lov24, LLW^{+}21, LYL^{+}24, MFZ23, MSG21a, MSG23, MKS21, MDT^{+}23, MCMP^{+}24, MRR^{+}20, MLJ^{+}23, Mea20, MSLMM22]. based [MTZA23, MMMS23, MPCG21, MM21, Mur24, NWO21, NO23, NS20, NMC^{+}20, NLMM23, OKS^{+}21, PR21, PDW23, PFM^{+}21, PAS23, PBS21, PDD^{+}22, Pér24, PMP^{+}21, PM22, RPG20, RFJ^{+}22, RBMRR20, RFdL^{+}24, RRFLV^{+}22, STQ^{+}22, SJA^{+}22, SD24, SLY^{+}21, SWHZ22, SLM24, SZ20, SPV^{+}21, TWL^{+}22, TLL^{+}23, TSUY23, TPD^{+}24, TBM^{+}24, VN23, VVH^{+}20, VSS^{+}22, WHE^{+}24a, WHE^{+}24b, WBW^{+}20, WRC21, WGAH22, WZLC24, WYBX23, WBD^{+}22a, WWF^{+}23, WKL^{+}21, WrJP^{+}22, XLM^{+}23, XHS^{+}24, YDLT24, ZMG^{+}21, ZMS20, ZOJ^{+}23, ZYZ22, ZZWF21, ZPZ^{+}22, ZHJQ23, ZZQ23, ZL24, BdBG^{+}21b, RRSG21, ZYM^{+}21, Mea20]. baseline [ABB23]. BASGRA N [HCP^{+}20]. Basin [CWY^{+}24, TGRM23, nCLpW^{+}23, RSASP22, SD24, XHS^{+}24, ZGMFESD23, LUT^{+}24, MP23, XHS^{+}24]. basis [JZS^{+}21, KLA^{+}20]. bass [LBI^{+}20, WBW^{+}20, WBD^{+}22a]. bat [PM22]. bathypelagic [CGRR^{+}24]. battle [MKN23]. Bay [HLS^{+}23, LUV^{+}23, SZL^{+}20, DHW^{+}20, DEMJ24, FMH^{+}20, HML^{+}22, HSD^{+}21, PHK23, dGZHCA24]. Bayesian [BSR^{+}23, GZK^{+}20, BRSW24, CB20, GML^{+}20, HZ22, HEL^{+}21, KDH^{+}21, MKB^{+}24, RMA^{+}24, SDH21, TDI^{+}20, VSS^{+}22, WSL^{+}22, WBD^{+}22b, ZWL^{+}21, ZS20]. bayesian-network [GZK^{+}20]. bays [BRBT22]. be [FSY22]. beachrock [SGP^{+}20]. bear [CLW23, GCS^{+}21]. bee [BSR^{+}23, CSCT23, HCL^{+}22, LBT^{+}22]. beech [LhNA^{+}21, LhNA^{+}22, SKW23]. beet [AAM^{+}22]. beetle [SC23]. BEFANA [MPC^{+}22]. before [LZL22]. behavior [DAB^{+}20, EWE20, JB22, JZHT23, KCHZ21, KS22, LCP^{+}20, LJP^{+}22, PAPŁ^{+}23, WFD23]. behavioral [CMFC23, DBH^{+}22, FTH24, MBB20]. behaviors [CAH22, GSH^{+}22, YDLT24]. behaviour [ASKR24, LHDBN20, dAWG23]. behavioural [HBMR24, JPMA^{+}23, JRD^{+}22]. behaviours [JVGC21]. behind [DPC20]. Beibu [ZHJQ23]. Beijing [LZF21, LW23]. being [GMS^{+}23, HL23, YPS^{+}21]. below [MK21]. belowground [BPN^{+}20]. Belt [LMW^{+}23a]. beluga [MCB^{+}23]. Benchmark [FHP^{+}22]. Benchmarking [UGSG24]. beneficial [XHZH24]. benefits [CTL22, SWL^{+}23]. Bengal [SD24]. Benthic [WLLW20, BAAN24, SBS^{+}21, SINS22, TRMB20]. Benthic-pelagic [WLLW20, SINS22]. bentho [NT23]. bentho-pelagic [NT23]. Bering [OCT^{+}23]. Bernard [PVS^{+}21]. Bertalanffy [LK24a]. best [BM22b, CSMP21, VPV21]. better [DAQ20, HKY24, JJN22, MJdORPG23]. between [ABZ22, AS22, AM21, BLB^{+}24a, BAR20, BBIP24, BLB^{+}20, BG22, CWY^{+}24, CMKF22, CZ21, CLY^{+}23, CL24, CMA23, DYMD22, DSC^{+}21, GVG^{+}23, GSM^{+}20, KKP^{+}21, LRA^{+}20, LZL23, LK24b, LGA^{+}21, MYACC22, MFPIBGES23, MPMT23, MGWG22, PWN^{+}24, PPSS22, RUTY23, RP22, SLLW21, TLL^{+}23, Vas20, WZHP21, YZZF22, YD20, ZMK^{+}22, ZSS21, dKS22]. Beyond [MBMV20]. BGC [CAK^{+}24, SNP^{+}23]. Bhutan [CNSK21]. bias [AJJ24, RGD^{+}22]. biased [AP22, BLP20, BBIP24, JMM^{+}20, RSC^{+}24]. bibliographic [DK22]. Bifurcation [SQDM21, NDNN22]. bighorn [DGSM21]. Bilberry [ML20]. bimodal [GDC24]. bio [ACFM22, TMFV22]. bio-energetic [TMFV22]. bio-geochemical [ACFM22]. bioaccumulation [GMM20, MLP^{+}20, MPL^{+}20]. bioactive [LBI^{+}20]. BioBaM [KMH^{+}21]. BioBaM-GHG [KMH^{+}21]. biochar [SMM^{+}22]. bioclimatic [ALJ^{+}22, BRC^{+}21]. biocontrol [ABC^{+}22, CH22, JBF22, dAWG23]. biodiverse [WGMR23]. biodiversity [BRC^{+}21, BSM23, CBW22, DSC^{+}21, DWG22, DLT^{+}22, GAM20, LG20, MPMT23, MPC^{+}22, MAGR^{+}22, MGdM^{+}20, MBD^{+}22, MPA^{+}24, PBF^{+}23, PNGESS21, TCS21]. biodiversity-ecosystem [MPC^{+}22]. Bioeconomic [PHK23]. bioenergetic [LLB^{+}23]. bioenergetics [DHW^{+}20, HMVR21, HML^{+}22, MFF^{+}24]. bioenergy [ESCA22, LG20, MEGA21]. bioenergy-driven [LG20]. Biogeochemical [CL24, SMP20, BPL^{+}20, DMdR24, ESZ^{+}23, GGE^{+}23, NBML21, SGTS21, SGTS23, Tsy22]. biogeography [ERP21]. biogeomorphology [ZZG^{+}24]. Biological [BMRF21, BPW^{+}21, CC23, CXZ^{+}23, FMH^{+}20, KCP^{+}23, LRA^{+}20, LAC^{+}24, MBR^{+}23, MBMV20, SvV21, SE21, TCG^{+}20, XLS20, XGT^{+}24]. biomass [ABRG21, AG22, AKY^{+}22, ACZ^{+}21, CJM^{+}23, DBV^{+}22, DS24, ERP21, GSR^{+}23, HCL^{+}24, HLH^{+}21, KHH^{+}22, KHH^{+}24, KLPJ24, LYC^{+}21, LCLG20, LUT^{+}24, MMdP^{+}21, MBS22, NK23, RWF22, SRB^{+}20, WHL^{+}23, ZZŠ^{+}22]. biome [DDDC21, SNP^{+}23]. biome-BGC [SNP^{+}23]. biophysical [GML^{+}20, HZ22, MTW^{+}22, RMV^{+}24, TK20b]. biophysics [RFHR23]. biorefinery [CAS^{+}24]. bioremediation [GLA^{+}20]. biosphere [FMSR20]. biotic [BLA^{+}24, DAB^{+}20, GSJC22, OMP^{+}23]. biovolume [DAQ20]. biphasic [CMG20]. bird [ASL^{+}20, CPTD22, WWCT24, YDLT24]. birds [CC20, GAC^{+}21, OEC24]. birth [NTHC23]. Biscay [DHW^{+}20, dGZHCA24]. bivalve [FMH^{+}20, HMGK20, TMFV22]. bivalves [CMGD^{+}22]. bivalvia [RMV^{+}24]. black [LK20, vdMvDSL20]. black-backed [vdMvDSL20]. blackspot [DTD^{+}23]. blessing [CBSB23]. blight [HEL^{+}21]. bloom [AWW24, BS21, CR24, HMC20, KSJ22, PTG^{+}20, SINS22, XZZ^{+}22]. blooms [CMKF22, CLH23, DVVH21, ELS23, JvVvHH24, LWZL23, RFCEL22, RFPRC23, ZGG22, ZWZ^{+}20]. Blue [TGRM23, CMFC23]. boar [KT22, LP22, RRFLV^{+}22]. Board [Ano20y, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano20-48, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j]. bobwhite [STW^{+}23]. Bogota [RBEC21]. Bohai [GSJ^{+}20, YD20]. bollworm [WXH^{+}22]. boosted [DBN^{+}22, YCI20]. border [JHL23]. boreal [BWT^{+}24, HLH^{+}21, LPS^{+}21, LMW^{+}23b, ML20, MLV24, PCL24, RPB^{+}23]. borer [BMRF21, MBR^{+}23]. borer-parasitoid [BMRF21]. borne [AV22, DBH^{+}22, WGK^{+}21]. both [CBSB23, WGMR23]. botrana [AZMD^{+}22, AZAD^{+}23, CdCAC^{+}20, PSM22]. bottom [vdHNH^{+}20]. BOUKHALEF [JB22]. boundaries [BMB24, ZYK22]. boundary [BM22a, SvV21]. bounds [LKO20]. box [LGG^{+}22, PGB^{+}20, SBR20]. box-model [PGB^{+}20]. boxes [VJK^{+}23]. boxwood [LGG^{+}22]. branch [KLS20]. branchiopods [Koo24b]. Bray [RP22]. Brazil [dSBDH23, GSA24, MRR^{+}20, OBFH22, dMT23, dPGDdFS24]. Brazilian [BdCCdSAdS23, MMdP^{+}21, dRMJdS23]. breeding [CPTD22, GAC^{+}21, RM20]. Brian [Des21]. Bridging [BMS^{+}20, CSBS21, RP22, ZJEG24]. broad [ASL^{+}20, CVTS20a, CVTS20b, TLL^{+}23]. broad-scale [TLL^{+}23]. broadleaved [WLT^{+}23]. broadleaved-Korean [WLT^{+}23]. brood [CMFC23]. brood-dependent [CMFC23]. brown [LG20, NDNN22, RWF22, SYL^{+}24]. browsing [BLB^{+}24a, HBRK24]. Brucella [LGFT^{+}20]. Bt [WXH^{+}22, dAWG23]. Bti [MMS^{+}24]. budburst [PGdAM^{+}20]. Budget [FCP^{+}20, GLA^{+}20, LFN^{+}20, VLGG^{+}20, vdMvDSL20, CDB^{+}20, DLZ^{+}21, HMGK20, JMHH22, KD20, Koo20b, LAK20, MKS21, Mat23, RPG20, RJY^{+}20, RRCZ20, VVH^{+}20, YRK^{+}20, DPC20, KGBL24, RLH24]. budget-based [VVH^{+}20]. budget-individual [CDB^{+}20]. budgets [PDG^{+}20]. budworm [SCG^{+}23, STRC24, cNKF24]. bugs [KGBL24]. Building [GZW^{+}20, JLH^{+}21, BTC^{+}22]. built [AIA^{+}23]. Bulgaria [SGRDA^{+}24]. Bumble [CSCT23, BSR^{+}23]. bumblebees [CDNN^{+}20]. Bunda [MDP21]. buoyancy [ABR24]. buoyancy-driven [ABR24]. burial [WSZ^{+}24]. burned [HH23]. Burning [LMW^{+}23b]. bycatch [Ino22]. BZI [SDP21].

C3 [Mas23]. cactus [PAdS^{+}24]. cadmium [GMM20]. cage [ABC^{+}22]. Calanoida [TBP^{+}23]. Calanus [BVM^{+}21]. calculation [RBMRR20]. calculations [AAM^{+}22]. Calgary [MMMS23]. calibrate [FFM^{+}23]. Calibration [BSR^{+}23, PFM^{+}21, PAdS^{+}24, FHP^{+}22, HEL^{+}21, JS22, LKP^{+}20, LPHO^{+}20, MPL^{+}20, OFH^{+}20, SNM^{+}23, SDH21, VSS^{+}22]. Callinectes [CMFC23]. Cambodia [TKO^{+}22]. camera [HI22]. CaMP [BWT^{+}24, BST^{+}20]. Camptotheca [PCN^{+}23]. CAMPUS [FSZ^{+}20, FZS^{+}20]. CAMPUS-S [FSZ^{+}20, FZS^{+}20]. Can [BLB^{+}24a, GSM^{+}20, vOVvdW23, CLHZ22, GMM^{+}24, KN20, KHH^{+}22, KHH^{+}24, LGG^{+}22, Lef23, ŁKPK22, MVAD20, PCL24]. Canada [BWT^{+}24, DMdR24, MPPM23, MTBB22, RPB^{+}23, SCG^{+}23]. Canadian [BST^{+}20, BWT^{+}24]. Canis [HKG^{+}24]. canopies [PNGESS21, PMP^{+}21, ZOJ^{+}23]. canopy [CWW^{+}23, CGVC21, HKY24, HKSM23, KBAC20, LWL24, MK21, MLJ^{+}23]. canyons [GZW^{+}20]. capabilities [GLH24]. capacity [HP20, MPSS24, PGB^{+}20, TWL^{+}22, YD20, ZHJQ23]. Cape [MDB^{+}20, HBP^{+}21]. capita [SGA22]. capital [BARF20, BDA^{+}20, GPN^{+}23, TBCV22]. capture [DMM^{+}23, GZW^{+}20]. capture-recapture [DMM^{+}23]. capturing [YYS^{+}22]. Carapa [dOPdMdS^{+}21]. Carbon [GZW^{+}20, RMO22, ACFM22, AKY^{+}22, BKTM20, BMAME^{+}23, BST^{+}20, BWT^{+}24, BLB^{+}20, BGC^{+}24, CM21, CMHZ23, CLW^{+}24, CSB22, CLG20, DIS23, DBV^{+}22, EL22, EON23, FEPR24, FSZ^{+}20, FZS^{+}20, HLM^{+}20, JLV^{+}21, JMPB24, JMC^{+}22, KKP^{+}21, KAB^{+}23b, KKPR20, LMW^{+}23a, LSS^{+}21, MLD^{+}23, MTBB22, MLV24, MSG^{+}21b, NKT^{+}20, PZH^{+}21, PGSL^{+}20, RRP^{+}23, RFF^{+}23, SPPC21, SNP^{+}23, WF24, WZLC24, WYBX23, XLS20, XLM^{+}23, YA24, YJC^{+}24, ZFT20]. carbon-neutral [EON23]. carbon-nitrogen-water [PGSL^{+}20]. carbonyl [CWW^{+}23]. carcass [PM22]. careful [BM22a]. Caretta [CRZ^{+}23, II24, MZLRSZ21]. caribou [SDD^{+}20]. Carlo [PM22, SDH21]. carnivore [GCB^{+}22, PBS21]. Carnosaurs [PR21]. carps [LWE^{+}22]. carrying [GJC^{+}20, HP20, MPSS24, PGB^{+}20, TWL^{+}22, YD20, ZHJQ23]. cascade [DYW^{+}22, WLZ^{+}23a]. cascades [BDC^{+}22, BNN20]. Case [OBFH22, SPPC21, ZLML20, AFdSM23, AIA^{+}23, BCW^{+}23, BH21, BGK^{+}20, BG22, BARF20, CL21, ETAAH21, GRGHM21, GSJ^{+}20, HNC^{+}21, JLLS23, LZSQ20, LDZ21, LZF^{+}24, LZL23, LDX^{+}23, LZL22, LPA^{+}21, LYL^{+}24, LCMB21, MMS^{+}24, MS20, Men22, MVC^{+}22, OHH^{+}21, OCT^{+}23, PHK23, PSM22, PTNR^{+}20, RFJ^{+}22, RKN^{+}21, RGD20, RAB21, SLC^{+}23, ST24, SWP^{+}23, SZL^{+}20, TLH^{+}20, VMM^{+}21, WBD^{+}22b, XWZ^{+}22, XHS^{+}24, XMWL23, YKZ^{+}23, ZMK^{+}22, ZWL^{+}21, ZHJQ23, ZZQ23, PMP^{+}21]. cassava [SYL^{+}24]. Catalonia [LDS^{+}23]. Catastrophic [Uch23]. catch [Lov24, LCMB21]. catch-and-release [Lov24]. catch-at-length [LCMB21]. catchment [BPL^{+}20, COS23, ZMK^{+}22]. cattle [CPAT20]. causal [AIA^{+}23]. caused [HBRK24, YB22]. causes [NGA^{+}20, RM20]. causing [EBMK20]. Cautions [ZFP21]. Cautious [MOS21]. cavities [VJK^{+}23]. CCG [XLM^{+}23]. CCG-MRIO [XLM^{+}23]. Cd [MTT22]. cedar [HKSM23]. cell [OSH^{+}21]. cellular [JB22, LZW20, SLC^{+}20, YXY^{+}20, ZHWZ23]. CEmCA [MMMS23]. Central [RSC^{+}21, RSC^{+}23, BAAN24, CNNS^{+}21, LZF21, MLL20, ZLLZ24, CCH20, DSC^{+}21, Pre22, ZFT20, ZFC^{+}23]. central-place [CNNS^{+}21]. centric [XLM^{+}23]. centroid [BM22a, MYACC22]. century [YCK20, HCL^{+}24]. cepedianum [PSBE23]. CERES [PTF23]. certain [LKO20]. CH [GAMN^{+}23]. Chagas [EBMK20]. chain [LM23]. chalcographus [OFH^{+}20]. Challenges [AGS^{+}21, AGS^{+}22, HL23, LSG21]. Change [HHI^{+}22, BPF24, BOLC^{+}22, BNAR^{+}21, BAR24, COJ23, CDN^{+}23, COS23, COPM23, CNSK21, CMS^{+}21, EBS23, FS22, HLH^{+}21, JLLS23, KWC24, LBGD21, LSS^{+}21, MTBB22, MLV24, MBD^{+}22, MLP^{+}20, MPL^{+}20, MPA^{+}24, MSS24, NGA^{+}20, NYH20, PK23, PTD^{+}23, PR22, RSASP22, RSC^{+}23, RBVL^{+}20, SLC^{+}23, SMD^{+}20, SWP^{+}23, SSB22, SGRDA^{+}24, SMV^{+}20, TKO^{+}22, WMA21, WLT^{+}23, XLS20, YCK20, YLLL23, YPS^{+}21, ZLLZ24, ZXK^{+}24, ZHWZ23]. Changes [AACCIS20, BS21, MSG^{+}21b, NYH20, BdCCdSAdS23, dSBDH23, CGRR^{+}24, CL21, CPTD22, EL22, EON23, FAC22, HKSM23, LHJ^{+}24, LLW^{+}21, MMdP^{+}21, NK23, PAS23, QS23, SBS^{+}20, WWC^{+}22, WSFZ21, YZZF22, YB22, YPS^{+}21, ZFT20]. Changing [LDS^{+}23, APGD^{+}21, BHC22, CN20, HSF^{+}20, HCX^{+}21, KWG^{+}22b, MCB^{+}23, OBFH22, TGRM23, TMNA20, TCS21, WSN^{+}21]. Channel [DHW^{+}20]. channelized [LEC^{+}23]. Chaos [GZ23]. Chaotic [CR24, BMRF21]. CHAPY [OFH^{+}20]. char [DSHB23]. character [CMA23]. characteristics [AAC20, LZF21, MFZ23, MBS22, SGH^{+}23]. Characterization [HCW^{+}22]. characterizations [HMDO20]. characterize [OTR^{+}21]. Characterizing [dAdSCA20]. Chattonella [SIOM24]. cheating [LL22]. Chelonia [SMV^{+}20]. chemicals [LBI^{+}20]. Chenopodium [GRAA22]. chensiensis [WYBX23]. Chernobyl [VSK^{+}20]. Chesapeake [DEMJ24, HSD^{+}21, PHK23]. chicken [FC21]. chicks [vdMvDSL20]. Chile [BAAN24]. chimpanzees [CDC^{+}22]. China [BCW^{+}23, CWY^{+}24, COJ23, CDN^{+}23, CL21, CMHZ23, GJC^{+}20, GSJ^{+}20, HLH^{+}21, HLS^{+}23, JLLS23, KHS^{+}23, LZSQ20, LZF^{+}24, LZL23, LPHO^{+}20, LZF21, LZL22, LGL^{+}24, LYL^{+}24, LW23, MFZ23, PWN^{+}24, PZH^{+}21, QS23, WHW^{+}23, SWP^{+}23, SZL^{+}21, SRB^{+}20, SZL^{+}20, SML^{+}20, TBZ^{+}22, TLH^{+}20, TGL^{+}22, WWC^{+}22, WXH^{+}22, WLZ^{+}23a, WYBX23, WLZ^{+}23b, WLT^{+}23, XLS20, XWZ^{+}22, XWX^{+}24, XHZH24, YLLL23, YD20, ZSS21, ZZW^{+}22, ZBH^{+}22, ZBW^{+}24, ZLML20, ZXK^{+}24]. chinensis [LGL^{+}24, YRK^{+}20]. Chinese [CZ21, ZLLZ24]. Chinook [BLLL21, HP20, MBB20, WDW^{+}22]. Chironomidae [MMS^{+}24]. chlorophyll [LKP^{+}20, WCSX22, YYS^{+}22]. Choice [AHH22, CSMP21, KTB^{+}23, PMB^{+}22]. Chondrichthyes [RSC^{+}21]. Chongqing [JLLS23]. chromosome [BPB^{+}20]. Chronic [HCW^{+}22, BGK^{+}20, BRSW24, KS22, TBM^{+}24]. circles [ZHC23]. circuit [XHS^{+}24]. circulation [KiAM^{+}20]. CiteSpace [ZZWF21]. cities [LZF21, MKR20, NSS^{+}23, YD20]. citizen [EMH^{+}22, JMM^{+}20, KAMS22, KRM^{+}22, LJRM^{+}20, SLB^{+}23, VMH^{+}21]. City [SBF^{+}21, EON23, RPARM20, XLM^{+}23, ZZQ23]. city-centric [XLM^{+}23]. Cladophora [KBAC20, KAT^{+}22]. Clair [KWM^{+}24, MSL^{+}21]. clarify [ZHC23]. class [BSCH23]. classification [Lia23, PSG^{+}22, WrJP^{+}22]. classified [dVDC20]. Clay [KFJS20, MTT22]. cleaner [JMHH22, LBN^{+}22]. Climate [HHI^{+}22, PTD^{+}23, AHH22, AMP^{+}23, BPF24, BOLC^{+}22, BS20, BNAR^{+}21, CWY^{+}24, COJ23, CDN^{+}23, CNSK21, CMS^{+}21, CN20, Dav22, DMG22, ELS23, EBS23, FS22, HMC20, HLH^{+}21, JST^{+}22, KNP^{+}21, LK20, LBGD21, MMdP^{+}21, MTBB22, MLV24, MBD^{+}22, MPA^{+}24, MSS24, NHGP23, NYH20, OBFH22, PK23, QS23, RSASP22, RSC^{+}23, RBVL^{+}20, SLC^{+}23, SMD^{+}20, SSB22, SGRDA^{+}24, SMV^{+}20, TGRM23, TCS21, TKO^{+}22, WHL^{+}23, WSHZ24, WKL^{+}21, WLT^{+}23, XLS20, YCK20, YLLL23, ZZŠ^{+}22, ZRMW^{+}20, ZXK^{+}24]. climate-driven [NHGP23]. climates [BPW^{+}21, Hop23, JS22]. Climatic [DMdR24, CLW23, CAK^{+}23, CAK^{+}24, LHJ^{+}24, TGL^{+}22]. climatological [CDN^{+}23]. clone [AXMJ22]. closure [OHR^{+}23]. CLUE [COS23, KMP^{+}23]. CLUE-S [KMP^{+}23, COS23]. clues [BMBB24]. CMEMS [ACZ^{+}21]. CMIP5 [ZFT20]. CNN [ZHWZ23]. CNN-LSTM [ZHWZ23]. co [DAG^{+}21, Kos22, MJS24, GAMN^{+}23, HKSM23, KKPR20, PG23, YLLL23, ZFT20]. co-existing [Kos22]. co-harvest [DAG^{+}21]. co-invaders [MJS24]. coast [BWS^{+}20, HMVR21, HK23, RKN^{+}21]. coastal [BRBT22, CMGD^{+}22, CL24, KPLH24, LLB^{+}23, MKEH20, NSS^{+}23, OTR^{+}21, OBFH22, PZH^{+}21, SBF^{+}21, SRK^{+}20, TWL^{+}22, Vas23, WCSX22, ZZG^{+}24]. coasts [OTR^{+}21]. cod [JOE23]. code [MTT22]. coefficient [MBMV20, RFdL^{+}24]. Coefficients [PPSS22]. Coexistence [SE22, AM21, CKC20, LGG^{+}22, Pér24, YL22]. coextinction [BANG20]. coffee [CGB^{+}23, KTB^{+}21]. cognitive [GLH24]. cohort [KK20]. cohorts [GSMD24]. coinfections [SYL^{+}24]. colchicus [KLPJ24]. cold [GSR^{+}23, PA20]. Coleoptera [AJJ24]. collaborative [BJKT21]. collapses [SLC^{+}23]. Collective [AZBdAI23, SLM24, WLY^{+}23]. collision [BLB24b]. Cololabis [KiAM^{+}20]. Colombia [RBEC21]. colonization [SKH^{+}20]. colony [HCL^{+}22]. color [CBG^{+}22]. Colorado [FG21]. Columbia [MBB20]. column [Lef23, VRRI^{+}24]. combination [HPMB20]. combinatorial [HBC^{+}22]. Combined [SM22a, WKL^{+}21, Dav22, LBGD21, LZL22, LYM^{+}23, TKO^{+}22]. Combining [BBCD23, BVR^{+}20, CDNN^{+}20, HCL^{+}24, ZMS20, BOLC^{+}22, WWP20]. commensal [JBF22]. Comment [CYP^{+}20]. Comments [CCJK^{+}20]. commercial [VRRI^{+}24, WBW^{+}20]. commercialization [MLD^{+}23]. common [CDB^{+}20, MHP^{+}23, MLP^{+}20, MPL^{+}20, YXY^{+}20, GXZ^{+}24]. commons [AZBdAI23]. communities [CDJ^{+}22, CSMP21, JRD^{+}22, KCS22, PBF^{+}23, YL22]. community [AADHA^{+}23, CWF^{+}21, Che22, CSB22, DvDA24, For20, LDZ21, LSL21, MGHT20, MPCG21, MPA^{+}24, SOV21, SDSP^{+}21, VSB22, WCF^{+}22, Wu24, YXH20, YEL^{+}21, ZMS20, LL22]. community-level [DvDA24]. compactness [WST^{+}23, ZYK22]. COMPADRE [TK20a]. Comparative [EOÖ21, MFZ23, TCG^{+}20, Koo20a, Koo24b, KBSS23, TQM^{+}24, ZMK^{+}22]. compare [BAAN24, CAB^{+}21]. Comparing [ALK20, BRSW24, CKC^{+}24, SRTS20, WHL^{+}23, WSL^{+}22, WT20, GAM20, LAC^{+}24]. Comparison [CMHZ23, GLK^{+}22, PCN^{+}23, ZMF^{+}20, LZL22, PTF23, RSWS23, SDH21, TSdOCV21, WGAH22]. compartment [KWG^{+}22a]. compatibility [AVP^{+}23]. compensation [BANG20]. Competing [PBH22]. Competition [BdSG^{+}23, CMA23, AM21, BLB^{+}24a, BPN^{+}20, NMGFOB21, Pre22, RPG20, RMK21, WSN^{+}21, YB22]. Competitive [GSRB22, MYH^{+}20]. complex [AGS^{+}21, AGS^{+}22, AIA^{+}23, BP21, BNN20, BLSR^{+}20, CGVC21, EHS^{+}20, KCS22, KLH^{+}24, Kle23, NK23, RBMRR20, SGPSzC22, SLLW21, SBR20, TK20b, XHS^{+}24]. complexity [BA20, ESZ^{+}23, GPN^{+}23, SBS^{+}21, SC24a, VRS20, WLWL23, WTT^{+}21]. complexity-stability [WLWL23]. components [SGA22]. Composing [Kle23]. composition [BPF24, HES^{+}20, HES^{+}23, HLH^{+}21, SHT21, WLT^{+}23]. compost [SSJC23]. compost-derived [SSJC23]. Comprehensive [CLY^{+}23, CMHZ23, KAB23a, WAHA^{+}24]. Computation [BSR^{+}23]. Computational [Laz23, CMS^{+}21, ESZ^{+}23, MDP21, GCSlR22]. Computer [MTT22, CGVC21]. computing [DRZ^{+}24, GSJC22]. concentration [BH20, KDC^{+}21]. concentrations [EOÖ21]. concentric [ZHC23]. concept [BWCD23, BRBT22, DP20, Gon23, SBM^{+}22, ZHC23]. Conceptual [AAR23, AMK21, BA20, CSBS21, GSMD24]. conceptualization [PGP^{+}22]. Concerns [RFHR23]. concinna [GSM^{+}20]. conclusions [ML20]. concolor [CB20]. concurrently [WXH^{+}22]. condition [CGB^{+}23, LHJ^{+}24]. conditional [RFPRC23]. conditions [AZMD^{+}22, AS22, APGD^{+}21, GGMR20, KCHZ21, LDXJ22, LUT^{+}24, MLP^{+}20, MPL^{+}20, PYD^{+}21a, PYD^{+}21b, SE21, SE22, SBWB24, TGL^{+}22, WSHZ24, WBD^{+}22b]. conductance [Mas23]. Conducting [LGA20]. Confidence [SBFBDM24, BLSR^{+}20, LK24b]. configuration [LZL23, MYACC22, OHH^{+}21, dMT23]. configurations [CHR^{+}23, MSLM^{+}24, RWF22]. conflict [KT22]. confluence [CFL^{+}22, DZC^{+}23a, DZC^{+}23b]. congeners [Kos22]. conifer [CLW^{+}24]. connections [Jor22]. Connectivity [DLT^{+}22, AGP20, JHKT^{+}23, LTR^{+}24, MP23, MS23, RMV^{+}24, SZPW23, SWW^{+}23, SKH^{+}20, TA20]. Consequences [AVRC^{+}20, JvVvHH24, Pre22]. conservation [AADHA^{+}23, AAAR23, AAAR24, BMS^{+}20, BWCD23, CXZ^{+}23, CAH22, DSC^{+}21, DWG22, GHbS21, HMGK20, HPMB20, LRA^{+}20, MS23, MGdM^{+}20, RBLS23, RSC^{+}23, RBEC21, RAC^{+}24, SACE^{+}20, TCS21, VBH20, WGMR23, WWCT24, ZYK22]. Conserving [XCH^{+}21]. consider [MPL^{+}20]. considerations [DMPLHGL20, WST^{+}23]. Considering [OTR^{+}21, BANG20, DYMD22, LZD^{+}24, NKT^{+}20, dAWG23]. Consistence [YZZF22]. consistency [SBFBDM24]. consistent [Gri20, LPHO^{+}20]. consolidated [ST24]. conspecific [WZHP21]. constrained [CIL^{+}20]. constraints [CDJ^{+}22, MVAD20, YEL^{+}21]. Construction [XHS^{+}24, ZHJQ23, CCJK^{+}20, SBMF23, WLZ^{+}23a]. consumer [CKC20, CBSB23, WLWL23]. consumer-resource [CBSB23]. consumption [CZ21, EL22, HBC^{+}22, HWS^{+}23, YZZF22]. contact [TLL^{+}23]. Contagion [LCM24]. contamination [LLW^{+}21, MLP^{+}20]. contemporary [PTD^{+}23]. content [BPK^{+}20, MGK^{+}23, SPV^{+}20]. Context [FEPR24, DGSM21, HHBF24, MLP^{+}20, MPL^{+}20, PGB^{+}20, RPG20, RLH24, VVH^{+}20, VSK^{+}20, YZZF22]. context-dependent [DGSM21]. contexts [KCS22, MVC^{+}22]. contiguity [WST^{+}23]. continent [WWF^{+}23]. continental [RFCEL22]. continuous [MKB^{+}24]. continuum [MT23, MF20]. Contrasting [JJN22, AADHA^{+}23, EHS^{+}20]. contribute [VPV21]. contributing [PYD^{+}21a, PYD^{+}21b]. Contribution [DHW^{+}20, AGP20, FSZ^{+}20, FZS^{+}20, LHJ^{+}24, MYH^{+}20, SGA22, TLC^{+}22, WLZ^{+}23b]. contributions [vDC20]. control [BPW^{+}21, BPB^{+}20, BMRF21, BCMV20, CAB^{+}21, Fer23, HZ22, HLB^{+}20, JvVvHH24, KSJ22, LRA^{+}20, MBR^{+}23, MD23, RBMRR20, SYL^{+}24, SE21, SRK^{+}20, TCG^{+}20, WLQ^{+}23, XGT^{+}24]. control-based [RBMRR20]. controlling [BLB^{+}20]. Controls [MBS22, KDH^{+}21, MBB20]. conventional [AAM^{+}22, KBSS23, MAD^{+}20, TSdOCV21]. convergence [LMNV20]. Converging [LBB20]. conversion [dMFOM^{+}20]. converting [ZA20]. Convolutional [JVGC21, PSG^{+}22]. cool [TDI^{+}20]. cool-temperate [TDI^{+}20]. cooperation [Gat21]. cooperative [GAC^{+}21, RBGB23]. coops [FC21]. coordinated [LYM^{+}23]. coordination [YD20, YJC^{+}24, ZSS21]. copepod [CSB22, KD20]. Copepoda [TBP^{+}23]. copper [GRAA22]. copper-induced [GRAA22]. coral [CN20, GGMR20, XHZH24]. cordatus [APGD^{+}21]. Cordyceps [CNSK21]. core [ALS^{+}21, TLH^{+}20]. Coro [CYP^{+}20]. Coronavirus [Gat20, CYP^{+}20, Cor20]. Correcting [BSCH23, RGD^{+}22]. corrections [HGBK23]. correlated [AP22, PDD^{+}22]. correlation [WZHP21]. correlative [HFSL^{+}24, SACEU^{+}21, TSdOCV21]. corridors [LHH^{+}24]. Corrigendum [AJG22b, AGS^{+}22, BdBG^{+}21b, BS20, CVTS20a, DZC^{+}23a, HES^{+}23, IBB^{+}22, KHH^{+}24, LhNA^{+}22, PYD^{+}21a, SE20a, SGTS23, WHE^{+}24a]. corroborate [FFM^{+}23]. coryi [CB20]. cost [CZ21, HCPG20, Lef23]. cost-effectiveness [HCPG20]. Costa [AHO^{+}20]. Costing [HPMB20]. costs [MJ24, NND^{+}22, RRP^{+}23]. cotton [GAMN^{+}23, WXH^{+}22]. cotton-wheat [GAMN^{+}23]. count [RCSS20]. countries [LMW^{+}23a]. counts [AP22, WWP20]. County [KT22, LYL^{+}24]. Coupled [GGE^{+}23, MTW^{+}22, WXT^{+}20, CXZ^{+}23, DVVH21, FMH^{+}20, KCS22, Mas23, MTT22, PG23, SINS22, WG21, WSW23, ZLL^{+}23]. Coupling [YHC^{+}21, BLA^{+}24, WLLW20, YXY^{+}20, YJC^{+}24, ZSS21, ZHWZ23, ZZG^{+}24]. course [LM23]. CoV [CYP^{+}20, Cor20]. cover [COS23, CPTD22, DCT^{+}21, DMdR24, HAZEF23, LSS^{+}21, MKB^{+}24, SCH^{+}21, WWC^{+}22, YCK20]. coverage [SGCFB22]. cowberry [ML20]. Coyote [MMMS23]. Cr [MTT22]. crab [CMFC23, OCT^{+}23, YFW^{+}21]. crane [NBMD21]. crassipes [XGT^{+}24]. created [SBS^{+}21]. Creating [HHFB23, EON23]. Creek [KFJS20]. Crimea [Vas23]. Criteria [USA21, KFM^{+}23, LAC^{+}24, MAN^{+}20, MSS24]. Critical [TA20, ABM20, WCG^{+}24]. critically [HMGK20]. croaker [RJY^{+}20]. Crop [ECC^{+}22, CSCT23, CJM^{+}23, CGVC21, DCT^{+}21, HKH^{+}23, HBC^{+}22, JPL21, LBT^{+}22, LDX^{+}23, LYM^{+}23, MVC^{+}22, OHH^{+}21, RGB^{+}23, SJA^{+}22, TGB^{+}23, WYJ^{+}22]. crop-weed [CGVC21, MVC^{+}22]. cropland [HCX^{+}21, ZYK22]. croplands [RMO22]. cropping [DDH^{+}22]. crops [CMSD^{+}22, CAFT22, DGVVP23, NAM22, dAWG23]. CROPSIM [PTF23].

CROPSIM-CERES [PTF23]. cross [JHL23, MBG^{+}22, RAC^{+}24, WHdBB21]. cross-border [JHL23]. cross-scale [MBG^{+}22]. cross-sectional [RAC^{+}24]. cross-validation [WHdBB21]. crossroad [BAR20]. crown [MAB23, dAPdBMT23]. crown-of-thorns [MAB23]. Crustaceae [BAAN24]. cruzi [EBMK20]. cryptic [BVR^{+}20]. Cryptolestes [AJJ24]. Cryptomeria [HKSM23]. CSCER [PTF23]. Cu [MTT22]. cucumber [TCG^{+}20]. cue [TPD^{+}24]. cue-based [TPD^{+}24]. Culex [FGM^{+}24]. culling [CAB^{+}21]. culls [Dav22]. cultivation [KBSS23, SSB22, Vas23]. culture [DMPLHGL20, FDSM20, SZL^{+}20]. cultures [PGB^{+}20]. Cumulative [TBP^{+}23, LPL22b, MMS^{+}24]. curated [SCACB23]. currency [CDB^{+}20]. Current [CFG^{+}22, JZS^{+}21, LHDBN20]. currents [PTD^{+}23]. Curricular [Por21]. curse [CBSB23]. Curtis [RP22]. curve [LMNV20, RGB22]. curves [BBSHC22, LGA^{+}21, SDP21]. cutaneous [ZMG^{+}21]. cutting [CSM23]. Cuyo [AZMD^{+}22]. cyanobacteria [GSRB22, LCLG20]. Cyanobacterial [LWZL23, JvVvHH24, ELS23]. cyanoFilter [OSH^{+}21]. cycle [AZAD^{+}23, BPL^{+}20, BP21, EL22, HKG^{+}24, KBSS23, LUT^{+}24, MMS^{+}24, NGA^{+}20, SLYL^{+}20, SGPSzC22, SGA22, SINS22, WZZF23]. cycle-based [EL22]. cycles [JMC^{+}22, Koo24a, MD23]. cycling [BKTM20, DMdR24, FWK^{+}20, KMAA23, SLB20b, SMP20]. cytometry [OSH^{+}21].

D [Des21, CGVC21, PMP^{+}21, PM22, VRRI^{+}24, WSW23]. D. [HKSM23]. DAESim [TBCV22]. daily [WHL^{+}23]. dam [SAdA^{+}22, WLZ^{+}23a]. damages [SAdA^{+}22]. damaskrose [NGA^{+}24]. damselfish [GGMR20]. Daphnia [PDH^{+}22]. Data [AVRC^{+}20, BTC^{+}22, AHH22, AADHA^{+}23, AMK21, AMP^{+}23, ATP20, AG22, ALK20, BSS^{+}24, BS20, BBSHC22, BVR^{+}20, CJM^{+}23, EMH^{+}22, GCSlR22, GCS^{+}21, GML^{+}20, ILLM20, IBB^{+}21, IBB^{+}22, Jag20, JVGC21, JMM^{+}20, KAMS22, KRM^{+}22, KKF^{+}23, KD20, KUK^{+}21, LDC^{+}20, LJP^{+}22, ŁKPK22, LAC^{+}24, MMM^{+}24, NGE^{+}21, NT23, OSH^{+}21, OW20, PK23, PSM22, PRdN^{+}20, PPH^{+}21, PM22, REGI^{+}24, RMA^{+}24, RGB22, RBVL^{+}20, SSM24, SNP^{+}23, TSdOCV21, VMH^{+}21, ZWL^{+}21]. data-based [AADHA^{+}23]. Data-driven [BTC^{+}22, CJM^{+}23]. data-poor [AMP^{+}23]. data-sparse [AMK21]. Database [TK20a]. dataset [LJRM^{+}20, ŁKPK22]. date [DIS23, JPL21, TBZ^{+}22]. DayCent [DDH^{+}22]. Dealing [BVR^{+}20, MVAD20]. death [NTHC23]. DEB [AVRC^{+}20, GSM^{+}20, JGG23, KK20, MPL^{+}20, OMC^{+}24, SLYL^{+}20]. DEB-structured [KK20]. DEB-tox [AVRC^{+}20]. DEBtox [Jag20]. Debugging [SLM24]. Decadal [HSL^{+}22, IBB^{+}21, IBB^{+}22]. decay [AB21]. December [Ano20c, Ano20o, Ano21c, Ano21o, Ano22c, Ano23c]. deciduous [SPPC21]. Decision [LZW20, NGE^{+}21, USA21, BVR^{+}20, CMSD^{+}22, For24, GZK^{+}20, JST^{+}22, MTW^{+}22, MSS24, PGP^{+}22, PBS21]. decision-making [BVR^{+}20, GZK^{+}20, JST^{+}22, MTW^{+}22, PBS21]. Decision-Support [LZW20, PGP^{+}22]. decisional [Pie23]. decisions [CDNN^{+}20, LE23, WDW^{+}22, YDLT24]. decline [LG20, RM20, XHZH24]. Decne [PCN^{+}23]. Decoding [HNC^{+}21]. decomposer [KKP^{+}21]. decomposition [CCH20, KAB^{+}23b, TGB^{+}23]. Deconstructing [WGK^{+}21]. decrease [dSBDH23]. deep [CCH20, EOÖ21, Lia23, WCSX22, WrJP^{+}22]. Deer [HCW^{+}22, BPF24, BLB^{+}24a, BGK^{+}20, KS22, TBM^{+}24, VRW^{+}21]. Defaunation [MMdP^{+}21]. defence [RTH^{+}21]. defence-reproduction [RTH^{+}21]. deficient [CAFT22, YXH20]. define [MLV24]. defines [ZGG22]. Defining [NBU^{+}23, RFHR24]. defoliator [PCL24]. Deforestation [dSBDH23, ZZW^{+}22]. Deforestation-induced [dSBDH23]. degradation [AMK21]. degraded [BWCD23, ESCA22]. degree [ZSS21]. degrees [XHZH24]. Delay [RCSS20, AWW24]. delayed [WWA^{+}22]. Delimitation [ZYK22]. Delineation [SD24]. Delta [CL21, SWP^{+}23, GMS^{+}23]. deltaic [BMAME^{+}23]. demand [KMP^{+}23, MRR^{+}20, MGdM^{+}20]. demanding [TSdOCV21]. demersal [BAAN24, DTD^{+}23, RSC^{+}21]. Demographic [BAZ22, PTBFP^{+}20, ABC^{+}22, CP20, MCMP^{+}24, PSM22, ZOJ^{+}23]. demography [JMC^{+}22, LVM^{+}22, PMB^{+}22, STW^{+}23]. demonstrates [JHKT^{+}23]. dengue [ABZ22, ZXK^{+}24]. denitrification [CL24]. denning [CLW23]. Dense [HHI^{+}22]. Density [HI22, CKC20, PSBE23, RFPRC23, VRW^{+}21, WRC21, dVDC20]. density-dependent [PSBE23, dVDC20]. dependence [WRC21]. dependency [FEPR24]. dependent [BHC22, CMFC23, CTL22, DGSM21, HML^{+}22, KRMP23, MGWG22, NGE^{+}21, PWN^{+}24, PSBE23, PGdAM^{+}20, RWF22, YB22, ZSS20, dVDC20]. depletion [BRAL23]. deposition [BPL^{+}20, RLdM20]. depredation [CDT^{+}23]. depredation-impacted [CDT^{+}23]. depth [CGRR^{+}24]. depth-resolved [CGRR^{+}24]. derived [OW20, SSJC23]. describe [HFP20, RCSS20, RRSG21]. Describing [BPN^{+}20, BWS^{+}20, CNNS^{+}21, PCN^{+}23, PAB^{+}22]. Description [RRSG21, TGB^{+}23, FSZ^{+}20]. descriptive [BSPT23a, BSPT23b]. desert [CWY^{+}24, CJM^{+}23, KWC24, LAK^{+}23, ZMS20]. desert-oasis [CJM^{+}23]. desertification [FMRMAS24]. deserts [BP21, WXX^{+}23]. design [BBCD23, CWH^{+}21, MAD^{+}20, NPS^{+}23, Tur20, WST^{+}23, ZJEG24]. Designing [SJA^{+}22]. designs [RAC^{+}24]. details [WTT^{+}21]. detect [LM23]. detecting [KKW^{+}21]. detection [APF21, DMM^{+}23, JJN22]. detections [NND^{+}22]. determinants [DDG^{+}21, LPD^{+}20]. determination [MGWG22]. determine [CSZY^{+}20, CDT^{+}23]. determined [HKH^{+}20]. determines [SBMF23]. Determining [ALS^{+}21, RPT21, USA21, SSM24]. deterministic [LK24b]. detritus [AWW24]. detritus-nutrient [AWW24]. Detroit [KWM^{+}24]. Developing [PFK^{+}22, SDP21]. Development [AIA^{+}23, DDG^{+}21, ESZ^{+}23, HL23, KLH^{+}24, MFF^{+}24, NO23, NT23, SSB22, ZZZD22, BSPT23a, nCLpW^{+}23, CPAT20, CMS^{+}21, DLZ^{+}20, FEPR24, GCMU23, HEL^{+}21, HBRK24, HCL^{+}22, HSD^{+}21, Jia21, JASW23, LPS^{+}21, ŁJ21, LYM^{+}23, RPB^{+}23, SGPSzC22, SYL^{+}24, SBWB24, TK20b, WRC21, XHZH24, ZSS21]. devices [NPK^{+}23, PDD^{+}22]. Dez [ZMK^{+}22]. DGVM [VK24]. diadromous [BOLC^{+}22]. diagnostic [KMH^{+}21]. diagrams [AIA^{+}23]. Dianchi [WCG^{+}24]. diapause [BVM^{+}21, PGdAM^{+}20]. DiaThor [GSJC22]. diatom [GSJC22]. diatoms [LCLG20, SBMF23]. Dicentrarchus [WBW^{+}20]. diel [CSB22, DZBB24]. Diels [ZWL^{+}21]. diet [LPA^{+}21, PTG^{+}20]. Dietary [MLP^{+}20, MPL^{+}20]. differences [AJJ24, CKK^{+}24a, GSM^{+}20, KWM^{+}24]. Different [XHZH24, ATP20, AJJ24, BRSW24, CHR^{+}23, DIS23, GVG^{+}23, Hop23, JL23, LYC^{+}21, MSLM^{+}24, Sak23, SBMG24, TLC^{+}22, TGL^{+}22, WSHZ24, WTT^{+}21, ZBH^{+}22, ZLZ^{+}24]. differential [COPM23, FGM^{+}24, VMM^{+}21, Wu24]. Differentially [RSC^{+}24]. Differentiating [BG22]. differentiation [GSRB22, HLZ^{+}22, ZSS21]. differentiations [LLZM20]. differently [RLdM20]. diffusion [CAH22, FMRMAS24]. Diffusivity [AJJ24]. digestion [AWW24]. digging [SLM24]. DigiBog Boreal [RPB^{+}23]. digital [KKW^{+}21, SG20]. dilated [PSG^{+}22]. Dilution [Oku21, CLHZ22]. dimensional [ABB23, Ano20-50, DBK20, LWE^{+}22, NBML21, TWL^{+}22, YLS^{+}21]. dimensions [KCS22, SZPW23]. dimictic [EL21]. DINAMICA [HPMB20]. Dingxi [LYL^{+}24]. dinoflagellate [GSJ^{+}22]. Dinosaurs [PR21]. dioecious [LBT^{+}22]. dioecy [MT23]. dioxide [BH20, BLB^{+}20]. diphtheria [LPSF24]. direct [BGA22, MTS^{+}20, TBM^{+}24, ZOJ^{+}23]. directional [STQ^{+}22]. discharge [Che21, YG21]. disciplines [BVR^{+}20]. Discrepancies [MYACC22]. discrete [mSzZK^{+}21, SE22]. discrete-time [SE22]. discrimination [SBFBDM24]. discus [DLZ^{+}21]. Disease [HCW^{+}22, BAPW20, BJKT21, BGK^{+}20, BRSW24, CBSB23, CLPZ20, CLHZ22, EBMK20, GSC^{+}23, KS22, NWJ^{+}22, PSG^{+}22, TBM^{+}24, WHE^{+}24a, WHE^{+}24b, WrJP^{+}22]. disease-induced [CBSB23]. diseases [GRGHM21, VSB22]. Disentangling [DAB^{+}20, QZSV21, WTT^{+}21]. disgust [PF23]. Dispersal [CGB^{+}23, LHH^{+}24, AE22, AM21, BOLC^{+}22, BSR21, CGV22, DZBB24, JPMA^{+}23, LK20, LHDBN20, LBB20, LSL20, VMM^{+}21, WSW23, dAdSCA20]. disperser [DAG^{+}21]. dispersion [IA23, MYH^{+}20]. displacement [CMA23]. disprove [LGK^{+}23]. dissimilarity [RP22]. dissolved [CFL^{+}22, FTH24, NKT^{+}20, YFW^{+}21]. distance [MVAD20, PPH^{+}21, RP22, RBEH21]. distance-independent [PPH^{+}21]. distinct [CTL22]. distinguishes [KWM^{+}24]. distort [ŁKPK22]. Distributed [RCSS20, NA22]. distribution [AHH22, AMP^{+}23, AFdSM23, ALJ^{+}22, BOLC^{+}22, BM22b, BM22a, BSCH23, BS20, BWSS21, BMS^{+}20, BG22, CSBS21, CLPZ20, CNSK21, CKC^{+}24, EWE20, FS22, GMM^{+}24, GWH20, HLZ^{+}22, HFSL^{+}24, Jia21, MGK^{+}23, MLJ^{+}23, MMRBLN20, MVAD20, Men22, MBD^{+}22, OW20, RBLS23, RSWS23, RSC^{+}24, RJY^{+}20, dSRDC^{+}23, RPT21, SLB^{+}20a, SVK20, SKW23, SBFBDM24, SC23, STRC24, SBMG24, TK20a, VRRI^{+}24, VMH^{+}21, VRS20, WSM24, WWP20, WSL^{+}22, YLS^{+}21, YCI20, ZZZD22, MLJ^{+}23]. distributional [BLA^{+}24]. distributions [HMDO20, JMM^{+}20, LK24b, OMP^{+}23, PEFW20, SRTS20]. District [MDP21]. disturbance [FdKO21, LVM^{+}22, LHL22, NL20, PZH^{+}21]. disturbances [YDQ^{+}23]. Disturbed [HHI^{+}22, KLNMG22, MHP^{+}23, MCB^{+}23, WLWL23]. diversicolor [GLA^{+}20]. diversion [BLB^{+}24a]. diversity [CLPZ20, Che22, HLW^{+}20, KAB^{+}23b, LPA^{+}21, PNGESS21, PSZ20, RFF^{+}23, SOP22, SG20, WZHP21, YXH20, ZL24]. DNDC [GAMN^{+}23, SMM^{+}22]. Do [WTT^{+}21, BAM23, CSCT23, RLdM20]. doab [AIA^{+}23]. DODM [STRC24]. Does [JJN22, CKC20, GZ23, KLNMG22, Pér24, PAS^{+}22, PNGESS21, SKH^{+}20, VSB22, XGT^{+}24]. doing [ASKR24]. dolomieu [LBI^{+}20]. dolphin [RBLS23]. domainal [Ano20-50]. domestic [GM23]. dominance [SIOM24]. dominant [AMP^{+}23, LHH^{+}24]. Don [HKSM23]. Dorosoma [PSBE23]. dorsalis [TBP^{+}23]. dose [YXH20, ZL24]. doses [MJdORPG23]. Douglas [BLB^{+}20]. Douglas-fir [BLB^{+}20]. down [EOÖ21]. downscaled [NLMM23]. Dr. [Des21]. drained [RGD20]. drift [LEC^{+}23]. drinking [ELS23, ZZZD22]. drip [GLK^{+}22, GZD^{+}23]. drip-irrigated [GLK^{+}22]. drive [FLG^{+}23, GZ23]. driven [ABR24, ABC^{+}22, BTC^{+}22, CMFC23, CJM^{+}23, nCLpW^{+}23, CGV22, ILLM20, JLLS23, KiAM^{+}20, Koo24a, LG20, NHGP23, PSM22, SJH^{+}23, SBR20, TK20b, WSW23]. driver [BA20, HLW^{+}20, MB22]. driver-response [BA20]. drivers [CFG^{+}22, DAB^{+}20, HBP^{+}21, MAGR^{+}22, MD23, WF24]. drives [ZFC^{+}23]. Droop [WGAH22]. Drosophila [RRSG21, SLM24]. drought [LPS^{+}21, SKPL24, SD24, ZZLF24]. drought-induced [LPS^{+}21]. droughts [HHBF24]. dry [DYMD22, DZC^{+}23a, DZC^{+}23b, YLLL23]. Dryland [CGV22]. dryness [JPL21]. DSSAT [PTF23]. dual [SZPW23]. due [AVP^{+}23, LDS^{+}23, MKR20, MYH^{+}20, ZOJ^{+}23]. DUET [MLJ^{+}23]. dugon [Lef23]. Dugong [Lef23]. dugongs [Lef23]. duration [LGA^{+}21]. during [CRZ^{+}23, CZ21, DZBB24, DZC^{+}23a, DZC^{+}23b, KFM^{+}23, OHH^{+}21, STQ^{+}22, Tsy22, dKS22]. Dutch [BAPW20]. Dwarakeswer [SD24]. dwelling [PSM^{+}21]. DynaGraM [MPCG21]. Dynamic [BM21, DMPLHGL20, FCP^{+}20, GLA^{+}20, HMGK20, LFN^{+}20, LZF^{+}24, PSM^{+}21, SC23, VLGG^{+}20, WWC^{+}22, ZLLZ24, ZLZ^{+}24, vdMvDSL20, ABKA24, AADHA^{+}23, BAAN24, CMGD^{+}22, CDB^{+}20, dSCdA21, DS23, DLZ^{+}21, ERP21, FdKO21, FSY22, GMM20, HBRK24, JB22, JM22, KWC24, KD20, Koo20b, LYC^{+}21, LAK20, LHL22, Mat23, RPG20, RJY^{+}20, RRCZ20, SVK20, STRC24, SRB^{+}20, SMS^{+}24, TBCV22, TSBE23, Tur20, VJK^{+}23, WCSX22, XWX^{+}24, YRK^{+}20, DPC20, KGBL24, RLH24]. dynamical [LPSF24]. Dynamically [SCG^{+}23, JPL21]. Dynamics [DBDR22, GWD^{+}23, MBR^{+}23, PG23, ADR21, AM22, AG22, APF21, APGD^{+}21, BS21, BOLC^{+}22, BGS^{+}20, BGR20, BGA^{+}23, BLP20, CR24, CLH23, CL24, CGV22, DKM^{+}21, DAQ20, DMGP20, DRZ^{+}24, DS24, EL21, EL22, FCFCM24, FDSM20, FSZ^{+}20, FZS^{+}20, GSR^{+}23, GAC^{+}21, GRAA22, GSJ^{+}20, HZ22, HSJ^{+}20, HSP^{+}22, Hop23, JLV^{+}21, JLLS23, KDH^{+}21, KCP^{+}23, KSJ22, KLH^{+}24, LE23, LRA^{+}20, LSL21, LQH24, LZW20, LSS^{+}21, LGL^{+}24, LCMB21, MLL20, MLMI23, MPSS24, MDB^{+}20, MSLMM22, MF20, MB22, MMC^{+}21, MPCG21, MKB^{+}24, NO23, NO24, NS20, NGE^{+}21, NLMM23, NK23, Oku21, OCT^{+}23, PDH^{+}22, PBH22, PSG21, PPH^{+}21, PKO^{+}21, PZH^{+}21, PMP^{+}21, PCK^{+}22, RCCLMGV22, RFJ^{+}22, RKN^{+}21, REGI^{+}24, RWS^{+}21, SCB^{+}20, SQDM21, SM22a, SMD^{+}20, SE21, SE22, SLB^{+}23, SJH^{+}23, SLC^{+}20, SBR20, TPD^{+}24, TK20b, Uch23, VSB22, WGAH22, Wu24]. dynamics [XYZR21, XFK21, ZMS20, ZLLZ24, ZYM^{+}21, GCSlR22].

E-PSR [RS23]. E2E [SHF24]. each [MJS24]. eagle [RM20]. early [LAK^{+}23, MFF^{+}24, WSFZ21]. earth [Kle23]. earthquake [BWS^{+}20]. earthworm [RPG20]. East [BWS^{+}20, COS23, MTW^{+}22]. eastern [BWS^{+}20, LAK^{+}23, LLC^{+}24, YJC^{+}24, GGMR20, MDB^{+}20, OCT^{+}23, SCG^{+}23]. easy [CXZ^{+}23]. easy-to-apply [CXZ^{+}23]. EasyCA [AXMJ22]. Echinococcus [MMMS23]. eco [Abe20, CN20, DPC20, GSC^{+}23, LLC^{+}24, LQH24, LW23, MF20, NWO21, NO23]. eco-efficiency [LW23]. eco-epidemiological [GSC^{+}23]. eco-evolutionary [Abe20, CN20, LQH24, MF20]. eco-hydrology [NWO21, NO23]. eco-physiological [DPC20, LLC^{+}24]. ecode [Wu24]. Ecodynamics [Ano20-49]. ecogeomorphic [PR22]. Ecohydraulic [Yao21]. ecohydrodynamic [DVVH21]. ecohydrologic [BKTM20]. ecohydrological [HKY24, MLMI23]. Ecol [BS20]. Ecological [AAAR24, AGS^{+}22, ARSC22, BdBG^{+}21b, BMS^{+}20, CVTS20a, DS21, HES^{+}23, KHH^{+}24, LC20, LBI^{+}20, LZD^{+}24, LSG21, LLW^{+}21, MMRBLN20, MP23, MGC24, Pie23, SE20a, SGTS23, TWL^{+}22, TLH^{+}20, VBH20, WHE^{+}24a, WCF^{+}22, XWX^{+}24, ZYZ22, ZZQ23, AS22, Ano20-50, AFG24, BANG20, BRC^{+}21, BAM23, BG22, CB20, CLY^{+}23, CNSK21, CIL^{+}20, CYP^{+}20, Cor20, DYMD22, DK22, DWG22, EGAS23, ESZ^{+}23, For24, GGE^{+}23, Gon23, GPN^{+}23, GCMU23, GZK^{+}20, HCL^{+}24, HYY^{+}24, HNC^{+}21, KMAA23, KNP^{+}21, KLH^{+}24, KAT^{+}22, KBSS23, LLZM20, LCM24, ŁKPK22, LPD^{+}20, LYL^{+}24, MMM^{+}24, MFPIBGES23, MS20, MZLRSZ21, NGA^{+}24, NBU^{+}23, NL20, OBFH22, PBH22, PHL20, PGB^{+}20, PAB^{+}22, PPSS22, PTNR^{+}20, RS23, RBMRR20, RFHR23, RFHR24, RBEC21, SZPW23, SACEU^{+}21, SCACB23, SML^{+}20, VRS20, VSK^{+}20, WG21, WXX^{+}23, WLQ^{+}23, WZLC24, XHS^{+}24, ZMK^{+}22, ZHJQ23, ZFP21, ZHC23, ZZG^{+}24]. ecological [ZGMFESD23, dFT^{+}20, Pas21]. ecological-economic [DWG22]. ecological-risk [GZK^{+}20]. Ecologically [Por21, CWF^{+}21, CKC^{+}24, GCSlR22, MFF^{+}24, WAHA^{+}24, WCF^{+}22]. Ecologically-Based [Por21]. ecologists [WG21]. Ecology [Mea20, Por21, BWSS21, CBW22, Fou21, PSW22, RFHR24, UGSG24, XJW21, dMT23]. econ [CXZ^{+}23]. econometric [BVR^{+}20]. Economic [DMPLHGL20, CZ21, DWG22, JLLS23, LZL22, NAM22, WAHA^{+}24, WZLC24, ZSS21]. economics [LM21, SM20]. economies [WF24, YD20]. economy [Ano20-49, FMSR20, GCMU23, SAL^{+}22, SMP20, ZA20]. Ecopath [BWS^{+}20, PTG^{+}20]. Ecophysics [GRS^{+}20, Pas21]. ecophysiological [CAK^{+}23, CAK^{+}24, TSdOCV21]. ecoregion [LhNA^{+}21, LhNA^{+}22]. ecoregions [HH23]. Ecosim [BWS^{+}20, WA20]. Ecosystem [BM21, FWK^{+}20, GSA24, LB21, PKB^{+}23, RSB^{+}23, YZF20, ZZŠ^{+}22, ASHK22, ABM20, BMBB24, BNN20, BNAR^{+}21, BAR24, BRBT22, BLSR^{+}20, BDA^{+}20, BGC^{+}24, CBG^{+}22, CL21, CLY^{+}23, CWW^{+}23, CXZ^{+}23, CVTS20a, CVTS20b, CL24, CPAT20, DS24, EON23, FMH^{+}20, FDSM20, GDF22, GSJ^{+}20, GZW^{+}20, HKB^{+}20, HES^{+}20, HES^{+}23, JHL23, KPLH24, LUV^{+}23, LL21, LZL23, LCW^{+}21, LPD^{+}20, LGA20, MPC^{+}22, MEGA21, MFH^{+}22, MTZA23, MBS22, NMS24, OHR^{+}23, PBF^{+}23, PGP^{+}22, PRdN^{+}20, PDGB21, PZH^{+}21, PCK^{+}22, RUTY23, RLS^{+}24, SBS^{+}20, SBS^{+}21, SBM^{+}22, SMD^{+}20, SHF24, SINS22, SZL^{+}20, TBCV22, WHL^{+}23, WYBX23, WFW^{+}24, XCH^{+}21, XCZ^{+}20, YPS^{+}21, ZYZ22, ZCJ^{+}20, ZYK22, ZHC23, ERP21]. ecosystem-atmosphere [BGC^{+}24]. Ecosystem-based [GSA24]. Ecosystem-scale [FWK^{+}20]. ecosystem-wide [HKB^{+}20]. Ecosystems [BAR20, AM22, AAC20, BA20, BLL^{+}23, BP21, BARF20, CP22, CTL22, DK22, DRZ^{+}24, FSZ^{+}20, FZS^{+}20, GPN^{+}23, GZK^{+}20, HLW^{+}20, HLZ^{+}22, LSA^{+}21, LWL24, MBG^{+}22, OTR^{+}21, OKS^{+}21, WHW^{+}23, RFHR23, RGRMRR20, WZHP21, WWCT24, ZFT20, ZFC^{+}23]. ecotone [TBZ^{+}22]. ecotoxicity [Jag20]. ecotoxicology [PFK^{+}22]. ectotherms [RRSG21]. edge [LCW^{+}21, YJC^{+}24]. edible [GMM20]. editorial [LSG21, Ano20y, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano20-48, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j]. Edmonton [NBML21]. eel [BdBG^{+}21a, BdBG^{+}21b]. eelgrass [KNP^{+}21]. Effect [AFdSM23, JMC^{+}22, JRD^{+}22, KFM^{+}23, LDX^{+}23, RGRMRR20, ZBW^{+}24, ABH^{+}20, BLB^{+}24a, BLLL21, CHR^{+}23, CLHZ22, CMS^{+}21, EWE20, KS22, KUK^{+}21, KBAC20, LL21, LP22, MSG21a, MDT^{+}23, MKR20, NDNN22, OMP^{+}23, PDH^{+}22, PZH^{+}21, RCCLMGV22, RMA^{+}24, XLM^{+}23, dAdSCA20, RLM23]. effective [BCMV20, PMZ^{+}21, SSM24]. effectiveness [BJKT21, HCPG20, XWX^{+}24]. Effects [CWH^{+}21, FdKO21, KKH^{+}24, LHDBN20, LCMB21, MMC^{+}21, MJS24, NBMD21, VRS20, YSFX23, ADR21, BKTM20, BH21, BSCH23, Che21, CGV22, dSCdA21, CP20, CSB22, DMM^{+}23, DWH^{+}21, EVT^{+}21, HCL^{+}22, Hop23, Ino22, KMAA23, KAB^{+}23b, KLPJ24, KPLH24, LK20, LBGD21, LYZ^{+}24, MMdP^{+}21, MMS^{+}24, MPSS24, MEPJ21, NK23, Oku21, PTBFP^{+}20, RWF22, RSC^{+}23, SDD^{+}20, SHF24, SWL^{+}23, SBWB24, SML^{+}20, TLL^{+}23, VVH^{+}20, VSK^{+}20, VBH20, WLZ^{+}23a, YPS^{+}21, ZFC^{+}23]. efficiencies [PGP^{+}22]. efficiency [BAM23, HLW^{+}20, LPL22a, LNK^{+}21, LW23, WWCT24, Yu20]. efficient [RBLS23]. efforts [XHZH24]. egg [BGR20, LEC^{+}23]. eggs [DLY^{+}22, LWE^{+}22]. EGO [HPMB20]. ELA [DMdR24]. elasticity [TK20a]. electricity [LGA^{+}21]. electrofishing [RGD^{+}22]. elements [MFZ23, PAPŁ^{+}23]. elephant [HBC^{+}22, MDP21, NMC^{+}20]. elephants [DDG^{+}21]. elevated [ZFT20]. elevation [LhNA^{+}21, LhNA^{+}22, PR22]. Elliptical [MLJ^{+}23]. Ellis [MTL^{+}23]. elm [BAPW20]. Elucidating [SIOM24]. embedded [KLH^{+}24]. Embedding [GXZ^{+}24, MS20]. embodied [XLM^{+}23, ZLML20]. Embracing [FCP^{+}20]. emergence [BR22, LDS^{+}23]. emergent [DAB^{+}20, LYC^{+}21]. emerges [WLWL23]. emerging [VSB22]. Emergy [CAS^{+}24, DYW^{+}22, ESCA22, EGAS23, LB21, AAM^{+}22, AAC20, CMHZ23, ENARM^{+}23, LL21, LDZ21, LZL22, LYL^{+}24, SAdA^{+}22, TLC^{+}22, ZMK^{+}22]. Emergy-based [ESCA22]. emigration [WGK^{+}21]. Emilia [MEGA21]. emission [IA23, KMH^{+}21, WF24, XLM^{+}23, ZMF^{+}20]. emissions [BGA22, BWT^{+}24, CMHZ23, DGVVP23, GAMN^{+}23, HWS^{+}23, MTS^{+}20, PG23, SMM^{+}22, TQM^{+}24, YLLL23, ZPZ^{+}22, ZBH^{+}22]. Empirical [BdSG^{+}23, BM21, dMFOM^{+}20, JM22, KWP^{+}23, LMNV20, MAN^{+}20, RBGB23, RBEH21, Ste20, WGAH22]. Employing [ATP20]. empowerment [NBMD21]. enables [VK24]. encounters [dKS22]. encourages [GLH24]. end [BNN20, CGRR^{+}24, CFG^{+}22]. end-to-end [BNN20, CGRR^{+}24, CFG^{+}22]. endangered [BLLL21, HMGK20, HBP^{+}21, MCB^{+}23]. endemic [HMGK20, WLZ^{+}23a]. enemies [LRA^{+}20]. energetic [KLA^{+}20, Lef23, TMFV22, WLLW20]. Energetics [CDB^{+}20, DGSM21, Koo20a, Koo24b, WWF^{+}23, WWA^{+}22]. Energy [DPC20, DS24, FCP^{+}20, GLA^{+}20, KGBL24, LFN^{+}20, MTP^{+}20, PDG^{+}20, RLH24, VLGG^{+}20, ZLML20, BPK^{+}20, BGC^{+}24, CZ21, CDB^{+}20, DLZ^{+}20, DLZ^{+}21, FMSR20, HMGK20, HWS^{+}23, JMHH22, KD20, Koo20b, LAK20, LBS20, MKS21, MPMT23, Mat23, RPG20, RJY^{+}20, RRCZ20, SLY^{+}21, SML^{+}20, VVH^{+}20, YRK^{+}20, vdMvDSL20]. energy-based [SLY^{+}21]. energy-budget [JMHH22]. Energy-Landscape [MTP^{+}20]. Energy-systems [DS24]. engineers [BMBB24, SBS^{+}20, SBS^{+}21]. English [DHW^{+}20]. enhance [ATP20, FCP^{+}20, Oku21]. enhanced [ESZ^{+}23, HAZEF23, SJH^{+}23]. enhancement [YRK^{+}20]. enhancements [CSCT23]. enough [LJRM^{+}20]. enrichment [OHR^{+}23, SM22a]. ensemble [CWW^{+}23, HBP^{+}21, LLLZ21, ZFP21]. ENSO [GGMR20]. entrainment [HKH^{+}20, LYZ^{+}24]. Entropy [ZSS20, GCS^{+}21, HMDO20, dRMJdS23, WT20, XJW21]. envelopes [MYACC22]. environ [LLW^{+}21]. environment [BSE21, BWS^{+}20, HSF^{+}20, HLB^{+}20, JASW23, KN20, KWG^{+}22b, MCB^{+}23, OEC24, PAdS^{+}24, PNGESS21, SBS^{+}21, SBMF23, SMP20, SGRDA^{+}24, ZSS21, ZL24].

Environmental [LMNV20, MBB20, OZSU22, Pas21, AZBdAI23, APGD^{+}21, BLA^{+}24, BGA^{+}23, BSR21, BAM23, BARF20, BDA^{+}20, CZ21, DMPLHGL20, FAC22, HL23, HWS^{+}23, JRD^{+}22, JASW23, KWC24, KCHZ21, LBGD21, LCP^{+}20, LBI^{+}20, MYACC22, MRR^{+}20, MBFS21, MLP^{+}20, NAM22, OMP^{+}23, PG23, PDW23, PYD^{+}21a, PYD^{+}21b, PM22, SLYL^{+}20, SM22a, Vas20, Vas23, WG21, WAHA^{+}24, WBD^{+}22a, YEL^{+}21, ZJEG24]. environmentally [ENARM^{+}23]. environments [BVM^{+}21, BR22, EHS^{+}20, FSY22, Koo24b, MMRBLN20, SG20, YXH20]. Enzo [Ano20-49]. ephemeral [AGP20, BGR20, BSR21]. Ephemeroptera [KFJS20]. EPIC [WYJ^{+}22]. epidemics [BCMV20]. Epidemiological [MD23, GSC^{+}23, VSB22]. epidemiology [MMMS23]. epigaeic [ZBW^{+}24]. epipelagic [CGRR^{+}24]. Epizootic [MSLMM22, EBS23]. equalization [HCL^{+}22]. equation [BPF24, CMG20, COPM23, DAQ20, KAB23a, NBU^{+}23, RFHR23, RFHR24, UGSG24, Wu24]. equations [AZAD^{+}23, FGM^{+}24, MAN^{+}20, VMM^{+}21]. equilibrium [CL24, SBMG24]. equitable [DL21]. equity [ZOJ^{+}23]. equivalence [HFP22]. eradication [HKB^{+}20]. erosion [AJG22a, AJG22b, GGMR20, LSS^{+}21, YG21, ZLZ^{+}24]. Error [KWG^{+}23, JTH24]. errors [OHH^{+}21]. eruptive [Fer23]. establishment [AV22, JHKT^{+}23, KKH^{+}24, NND^{+}22, SG20]. esters [WLQ^{+}23]. estimate [ACFM22, GAMN^{+}23, HCL^{+}24, HLS^{+}23, KAMS22, MCMP^{+}24, PSM22, RFdL^{+}24, SLB^{+}23, TTPA21]. estimated [LCMB21]. estimates [ALK20, BHC22, CLW^{+}24, DBV^{+}22]. Estimating [BPK^{+}20, GAM20, JS22, JMM^{+}20, KD20, LB21, MLL20, SZL^{+}21, WHE^{+}24a, WHE^{+}24b, ZPZ^{+}22, BBIP24, GLK^{+}22, LAK^{+}23, LAK20, RMA^{+}24, SWF21, SDP23, WYBX23]. Estimation [AJ22, II24, CCH20, EJ20, GZD^{+}23, HI22, HCX^{+}21, KLS20, KWP^{+}23, MTT22, OMC^{+}24, RFPRC23, SWHZ22, XWZ^{+}22, Yu20]. estimations [YSIB24]. estimators [LK24b]. estuaries [FTH24, MLP^{+}20, SJH^{+}23]. estuarine [AM22, CMFC23]. estuary [Che21, HMC20, KDH^{+}21, MKEH20, MBB20, DZC^{+}23a, DZC^{+}23b]. ESV [ZZQ23]. ESV-FLUS [ZZQ23]. Ethiopia [TGRM23]. eucalypt [BdSG^{+}23]. Eucalyptus [EHS^{+}20]. Euclidean [RP22]. europaea [PAPŁ^{+}23]. Europe [ASL^{+}20, EVT^{+}21, HKG^{+}24, HK23, Pre22, SBR20]. European [CH22, DHW^{+}20, LG20, MKR20, RBVL^{+}20, SLB^{+}20a, SKW23, WBW^{+}20, WBD^{+}22a]. eutrophic [BS21, ELS23, ZWZ^{+}20]. eutrophication [BNAR^{+}21, CDJ^{+}22, DBK20, DK22, NBML21, SRK^{+}20]. eutrophication-based [DK22]. evaluate [BGA^{+}23, GLA^{+}20, KDH^{+}21, KLPJ24, MAD^{+}20, MPA^{+}24, OEC24, TDI^{+}20, WHdBB21]. evaluates [MGC24]. Evaluating [CAK^{+}24, ERP21, GWH20, GSJ^{+}20, HBRK24, HFSL^{+}24, IBB^{+}21, IBB^{+}22, LDZ21, NGA^{+}24, SZL^{+}20, WLZ^{+}23a, XCZ^{+}20, DL21, DWH^{+}21, MDB^{+}20, MS23, NO23]. Evaluation [AAC20, CSZY^{+}20, NWO21, WHW^{+}23, STW^{+}23, BSPT23b, BLL^{+}23, CLY^{+}23, HYY^{+}24, Ino22, LPS^{+}21, LZL22, MAN^{+}20, MP23, PAB^{+}22, SLY^{+}21, SPPC21, SBFBDM24, SDP23, TK20b, WSZ^{+}24, YZF20, RPARM20].

Evaluations [LEC^{+}23]. evaporation [HKY24]. evapotranspiration [WLZ^{+}23b, ZFC^{+}23]. event [GRGHM21]. events [FS22, RFCEL22, RFPRC23, SLC^{+}23, SSB22, TTPA21]. Everglades [RGD20, ZRMW^{+}20]. evergreen [CLW^{+}24]. Evidence [LW23, PK23, SGH^{+}23, COJ23, CDC^{+}22, MMRBLN20, ZA20, BAM23]. Evolution [Gon23, LZSQ20, Gat21, JOE23, LPSF24, LZF^{+}24, LM21, LYL^{+}24, WXH^{+}22, YDQ^{+}23, YD20, GCMU23]. Evolutionary [GHK23, Abe20, AARS23, BDC^{+}22, CP20, CN20, GRCSA^{+}20, LQH24, LBS20, MF20, MOS21, MJdORPG23, PDW23]. Exaiptasia [KRMP23]. examine [BWT^{+}24]. Examining [CLG20]. example [CDNN^{+}20, LPSF24, WA20]. exchange [ASHK22, BLB^{+}20, HLS^{+}23, RSB^{+}23]. exchanges [SMP20]. exciting [GSH^{+}22]. exclusion [MYH^{+}20]. Exergy [VPAV22, NGA^{+}24]. exhibit [HSL^{+}22, PWN^{+}24, RBGB23]. existence [dAdSCA20]. existing [JS22, KD20, Kos22]. expansion [GCB^{+}22, LK20, MKEH20]. expectations [BMS^{+}20]. expected [PTD^{+}23]. experienced [SBMF23]. experiencing [RCCLMGV22]. experiment [ABC^{+}22, BLB^{+}24a, PNGESS21]. experimental [MPL^{+}20, NS20, DMdR24]. experiments [FZS^{+}20, JLV^{+}21, PSM^{+}21, Tur20, VN23]. expert [WBD^{+}22b]. Explaining [CKK^{+}24a]. explains [SGV^{+}24, XJW21]. explanation [MSG23]. explicit [AE22, BSU^{+}23, CAB^{+}21, DDG^{+}21, GCB^{+}22, GRCSA^{+}20, KS22, KWG^{+}23, LQH24, SLHB^{+}23, SJA^{+}22, SOV21, TPD^{+}24, WBW^{+}20, WBD^{+}22a]. explicitly [SBS^{+}21]. Exploitation [CDT^{+}23, TCB21]. exploited [CMGD^{+}22]. exploration [KAB^{+}23b, YYS^{+}22]. Exploratory [BTC^{+}22]. explore [BGS^{+}20, MSLMM22]. Exploring [AAAR23, AAAR24, BNN20, CKK^{+}24b, KMH^{+}21, LPL22a, LZL23, SBM^{+}22, SHF24, WZLC24, dMT23, BRBT22, DMGP20, GRS^{+}20, NGA^{+}20]. export [NO24]. exports [YG21]. exposed [JST^{+}22, WBD^{+}22b]. exposure [DEMJ24, DWH^{+}21, JSD^{+}22, MMS^{+}24, WSZ^{+}24, ZOJ^{+}23]. extended [JSL^{+}20, Laz23, NGA^{+}24, WMA21]. Extending [SDSP^{+}21, KMP^{+}23]. extension [BP21]. extensions [MKS21]. extensive [BNN20]. extensively [HH23]. extent [AFdSM23, GHK23]. extinction [Dav22, TCB21, Uch23]. extinctions [GRCSA^{+}20]. extrapolation [RRFLV^{+}22]. extreme [Dav22, SLC^{+}23, SCH^{+}21, ZOJ^{+}23].

facilitates [CKC20]. Facilitation [Pre22]. facing [BLB^{+}24a]. factor [ST24]. factored [PSG21]. Factors [BLB^{+}20, SOP22, BLA^{+}24, CDN^{+}23, CKK^{+}24b, DPC20, HYY^{+}24, HCL^{+}22, LHJ^{+}24, LW23, Vas20, YG21, YYS^{+}22, ZSS21, ZYM^{+}21]. facts [MZLRSZ21]. facultative [PGdAM^{+}20]. Fagus [LhNA^{+}21, LhNA^{+}22, SKW23]. fails [AE22]. failure [SAdA^{+}22]. family [SBFBDM24]. Farm [WrJP^{+}22, TLC^{+}22]. farmer [ASKR24, BAM23, LE23, MGdM^{+}20]. farming [BGA^{+}23, ENARM^{+}23, HSP^{+}22, MDB^{+}20, WAHA^{+}24]. farmland [LE23, OEC24, WWCT24, YLLL23]. farms [LBN^{+}22, MM21]. fast [AKM22, LM21, LCM24, SdlP20]. fastidiosa [BCMV20, GSC^{+}23]. fate [BSPT23a, BSPT23b, MLV24, SSJC23]. fates [WLQ^{+}23]. Fath [Des21]. fauna [Yao21]. favorable [KDC^{+}21]. favourability [MDT^{+}23]. favourability-based [MDT^{+}23]. Feasibility [AS22]. feature [LSG21]. Features [CLG20, SDD^{+}20]. February [Ano20d, Ano20p, Ano21d, Ano21p, Ano22p, Ano23p, Ano24k]. fed [Che21]. feed [dMFOM^{+}20, vOVvdW23]. feedback [BPN^{+}20, CMKF22, MLMI23, RGRMRR20, TK20b, ZWZ^{+}20]. feedback-driven [TK20b]. feedbacks [TK20b]. feeding [CMKF22, HSF^{+}20, JMHH22, SPB^{+}24]. felids [RSC^{+}24]. female [WBD^{+}22b]. Fenghuangshan [ZZZD22]. Fenneropenaeus [YRK^{+}20]. ferrugineus [AJJ24]. fertility [CAB^{+}21, VPAV22]. Fertilization [DGVVP23]. fertilizer [JLV^{+}21, LYM^{+}23]. fever [ABZ22, MSLMM22, MSLM^{+}24, MD23]. few [AXMJ22]. fidelity [BR22]. Field [CFL^{+}22, SOV21, BGA22, GZD^{+}23, HGBK23, LAC^{+}24, MGC24, NGE^{+}21, PSM22, PDD^{+}22, SGH^{+}23]. field-based [PDD^{+}22]. fields [BMAME^{+}23, GSR^{+}23]. fieldwork [DIS23]. filamentous [KBAC20]. filter [CWW^{+}23]. filtering [EMH^{+}22, VMH^{+}21]. Fine [Fou21, JSD^{+}22, KTB^{+}23, YSIB24]. fine-scale [JSD^{+}22]. Fine-tuning [Fou21]. finer [KKF^{+}23]. finless [DZC^{+}23a, DZC^{+}23b, HSF^{+}20]. finmarchicus [BVM^{+}21]. Finnish [KRM^{+}22]. fir [BLB^{+}20, SKW23]. fire [ABRG21, AAAR23, AAAR24, BKTM20, ERP21, HH23, JLH^{+}21, JB22, LMW^{+}23b, MMdP^{+}21, OHP^{+}23, SLHB^{+}23]. fire-land [JLH^{+}21]. fire-prone [ABRG21]. fires [dRMJdS23, NA22]. first [BPN^{+}20]. Fish [AWW24, BPK^{+}20, BGCJ21, BSR21, CSBS21, CFG^{+}22, DLY^{+}22, FTH24, HML^{+}22, JMHH22, KTB^{+}23, LDC^{+}20, LBN^{+}22, LCP^{+}20, LBI^{+}20, LZD^{+}24, LYZ^{+}24, LGA^{+}21, LUT^{+}24, MFF^{+}24, NT23, NPK^{+}23, NPS^{+}23, PBH22, RUTY23, RGB22, SLB^{+}20a, SRTS20, VRRI^{+}24, WLZ^{+}23a, WLY^{+}23, dKS22, PDD^{+}22]. Fisher [RBMRR20]. fisheries [BWS^{+}20, CDT^{+}23, GSJ^{+}20, Lov24, PHK23, WBW^{+}20, XHZH24]. fisheries-targeted [GSJ^{+}20]. fishery [AKM22, BNAR^{+}21, CMGD^{+}22, DMGP20, LGL^{+}24, MZLRSZ21, RSC^{+}23, SM22a]. fishes [HMDO20, LLB^{+}23, SGLK20]. fishing [BAAN24, KiAM^{+}20, KKH^{+}24, RWF22, XHZH24]. fit [PAS^{+}22]. five [IBB^{+}21, IBB^{+}22, ZMF^{+}20]. fixation [CL24]. fixed [GSRB22, HLS^{+}23, JPL21, MLV24]. fixing [GSRB22]. fjordic [GMM^{+}24]. fjords [AM22]. flammulated [CWH^{+}21]. flexibility [AKY^{+}22]. flexible [JMC^{+}22]. flight [KWG^{+}22a, SGV^{+}24]. flights [STQ^{+}22]. floating [HLS^{+}23]. floodplain [LUT^{+}24, TSUY23]. Florida [PTG^{+}20, ZRMW^{+}20, CB20, CRZ^{+}23, SPB^{+}24]. flow [ARSC22, HLM^{+}20, JB22, LZD^{+}24, NPK^{+}23, OSH^{+}21, SBWB24, WDW^{+}22]. flowering [KKW^{+}21]. flows [FMSR20, HKH^{+}20, MFZ23, RLS^{+}24, ZZWF21, ZLML20, dPGDdFS24]. fluctuating [BSE21]. Fluctuation [SGH^{+}23]. fluctuations [FCFCM24, GZ23]. Fluid [GCSlR22]. flume [LEC^{+}23]. fluorescence [SPPC21, YYS^{+}22]. FLUS [ZZQ23]. flux [ASHK22, CSB22, SPPC21, ZWZ^{+}20]. fluxes [ACFM22, BGC^{+}24, CLW^{+}24, GXZ^{+}24, HKSM23, SNP^{+}23]. fly [DBN^{+}22, YKZ^{+}23]. focal [For20]. focused [Laz23]. folivore [STRC24]. follow [KK20]. following [LVM^{+}22, cNKF24]. Food [GHbS21, SRK^{+}20, BAAN24, BNN20, CGRR^{+}24, DvDA24, DS24, HML^{+}22, JL23, KRMP23, KMH^{+}21, KKH^{+}24, KDC^{+}21, LPL22a, LLB^{+}23, LPA^{+}21, MAD^{+}20, MJ24, MPPM23, MGdM^{+}20, MEPJ21, MJS24, NHGP23, NK23, OKS^{+}21, VVH^{+}20, WLLW20, WA20, WSN^{+}21, YZZF22, ZBW^{+}24, vdHNH^{+}20]. footed [JZJ^{+}20]. footprint [EGAS23, LYL^{+}24, RPARM20, SML^{+}20, TWL^{+}22, TBZ^{+}22, ZMK^{+}22]. forage [MPA^{+}24, PAdS^{+}24]. foragers [CNNS^{+}21, GLH24]. foraging [BSR^{+}23, CDNN^{+}20, JOE23, KN20, RBEH21, YDLT24]. forcing [CBG^{+}22]. Forebay [LYZ^{+}24]. forecast [HEL^{+}21, WCSX22]. Forecasting [BPW^{+}21, HWS^{+}23, KiAM^{+}20, AAAR23, AAAR24, CZ21, HAZEF23, NLMM23, Pie23, ZFP21]. forecasts [NLMM23, WF24]. foreign [SML^{+}20]. Forest [BAR20, ETAAH21, Mea20, MTZA23, RBEC21, ABH^{+}20, AHO^{+}20, BSS^{+}24, BAPW20, BPF24, BdSG^{+}23, BLL^{+}23, BGC^{+}24, CM21, CWW^{+}23, CLW^{+}24, DBV^{+}22, DS24, FFM^{+}23, FEPR24, FG21, FSZ^{+}20, FZS^{+}20, FHP^{+}22, GMS^{+}23, GHK23, GSMD24, HKY24, HBRK24, IBB^{+}21, IBB^{+}22, JB22, KWP^{+}23, LDR^{+}20, LhNA^{+}21, LhNA^{+}22, ML20, MSLMM22, MSLM^{+}24, MLV24, MKB^{+}24, NMS24, NA22, PPH^{+}21, PZH^{+}21, PNGESS21, PSW22, PCL24, Pre22, PGSL^{+}20, RFJ^{+}22, RFF^{+}23, SE20a, SE20b, SAL^{+}22, SMD^{+}20, SLLW21, SPPC21, SNP^{+}23, TDI^{+}20, VK24, WZHP21, WYBX23, WLT^{+}23, YA24, ZCJ^{+}20, ZLL^{+}23, CKC^{+}24, VSK^{+}20]. forested [BPL^{+}20, HH23]. Forests [HHI^{+}22, AMP^{+}23, BSS^{+}24, CAK^{+}23, HHFB23, HHBF24, HLH^{+}21, IBB^{+}21, JHKT^{+}23, LPS^{+}21, LMW^{+}23b, MLL20, MAGR^{+}22, MHP^{+}23, MSG^{+}21b, MKB^{+}24, PKO^{+}21, dOPdMdS^{+}21, RRP^{+}23, RLdM20, RBVL^{+}20, SLHB^{+}23, SKPL24, SFGB23, SKW23, XJW21, YSIB24, ZRMW^{+}20]. form [MAN^{+}20]. formation [FMRMAS24, KFM^{+}23, SLC^{+}20, dKS22]. former [LGK^{+}23]. forming [CR24]. forms [LZSQ20, WGAH22]. formulation [BRAL23, FSZ^{+}20]. Fort [Pas21]. fouling [PF23]. foundation [KAT^{+}22]. four [AAC20, MKR20, PDW23, SINS22]. four-party [PDW23]. foxes [TLL^{+}23]. fragmentation [JRD^{+}22, cNKF24]. fragmented [BWCD23, CJM^{+}23, LSA^{+}21, RMV^{+}24]. fragments [LTR^{+}24, MSLMM22]. framework [AMK21, AAR23, BSU^{+}23, BGK^{+}20, BAR24, CMSD^{+}22, CIL^{+}20, CDC^{+}22, DBDR22, DBH^{+}22, II24, JLH^{+}21, JM22, KDH^{+}21, LDC^{+}20, LMNV20, LJRM^{+}20, LLLZ21, MTW^{+}22, MHP^{+}23, PAS23, PDGB21, RAC^{+}24, SGPSzC22, SGA22, TLC^{+}22, TWL^{+}22, TDI^{+}20, ZLL^{+}23]. frameworks [Mur24]. Framing [SC24b]. France [BPL^{+}20]. franksi [LVM^{+}22]. free [DGSM21, Men22, TBM^{+}24]. free-ranging [DGSM21, TBM^{+}24]. freezing [LCW^{+}21]. freezing-thawing [LCW^{+}21]. French [RMV^{+}24]. frequency [MMdP^{+}21, PCL24]. frequently [KLNMG22]. fresh [GGE^{+}23]. freshwater [KWM^{+}24, LDC^{+}20, SC24a, SRTS20, TCS21, Tsy22, WSZ^{+}24]. friendly [ENARM^{+}23]. front [ASL^{+}20]. frozen [LMW^{+}23b]. frugiperda [BSU^{+}23, dAWG23]. fruit [DBN^{+}22, LBT^{+}22]. fugacity [WLQ^{+}23]. full [SLYL^{+}20]. FullCAM [FEPR24]. Fully [JVGC21, PPSS22]. FunBies [PBF^{+}23]. function [DMG22, PSW22, SZPW23, YXY^{+}20, YJC^{+}24]. Functional [OEC24, BMBB24, BBSHC22, CLW^{+}24, DLT^{+}22, HES^{+}20, HES^{+}23, JHKT^{+}23, KAB^{+}23b, MAN^{+}20, NHGP23, PBF^{+}23, Pér24, PNGESS21, RP22, RFF^{+}23, SOP22, SHT21, SQDM21, dAdSCA20]. functioning [MPC^{+}22, MBG^{+}22, RLS^{+}24]. functions [ALK20, CNSK21, LAK20, PCN^{+}23, VPV21]. fundamental [EL21, JS22]. Fundão [SAdA^{+}22]. fungal [MGHT20]. Fusarium [GGS^{+}22]. fusion [SWP^{+}23]. future [BPW^{+}21, CZ21, CLW23, CLG20, ELS23, FHP^{+}22, FS22, GHK23, HSD^{+}21, KNP^{+}21, LhNA^{+}21, LhNA^{+}22, MBD^{+}22, PR22, ZPZ^{+}22, ZXK^{+}24]. futures [KMH^{+}21]. Fuzzy [DSHB23, KLH^{+}24, ZWL^{+}21].

Gadus [JOE23]. Gaia [Gat20]. gain [BPK^{+}20]. Galatheidae [BAAN24]. Game [MB22, DS23, GHbS21, LBN^{+}22, PDW23]. games [AARS23]. Gammaridae [DZBB24]. Ganges [RBLS23]. Gansu [XHS^{+}24]. gap [DTR^{+}22, RP22]. Gaps [AVRC^{+}20, ŁKPK22]. garden [SSJC23]. gardens [RPARM20]. Gas [dKS22, BST^{+}20, BWT^{+}24, CFL^{+}22, CPAT20, HCP^{+}20, KMH^{+}21, MTS^{+}20, TQM^{+}24]. GCM [TKO^{+}22, ZFT20]. GCM-related [TKO^{+}22]. geese [JZJ^{+}20]. gene [FLG^{+}23]. general [PSG21, RRSG21, DLT^{+}22]. generalizable [RBEH21]. Generalized [SE21, CIL^{+}20, KWC24, SE22, WSL^{+}22]. generated [TK20a]. Generation [EHS^{+}20, LGA^{+}21, MKS21]. generic [CdCAC^{+}20, DWG22]. generically [CDJ^{+}22]. genetic [SG20]. genetically [NS20]. Geo [DIS23]. Geo-spatial [DIS23]. geochemical [ACFM22, MTT22]. Geographic [LB21, HMDO20, RJY^{+}20, SM22b]. geographically [BBIP24, CKC^{+}24]. geography [MYACC22, USA21]. geolocation [NT23]. Georgia [MPPM23]. geospatial [KKF^{+}23]. geostatistical [BMB24]. Gerais [MRR^{+}20, dPGDdFS24]. German [HHBF24]. germanica [CH22]. germinans [LK20]. GHG [HWS^{+}23, KMH^{+}21]. giant [DRZ^{+}24, XFK21, HYY^{+}24]. GIS [DTR^{+}22, DIS23, ILLM20, KFM^{+}23, Lov24, SD24, USA21]. GIS-based [KFM^{+}23, Lov24]. given [SE20a, SE20b]. gizzard [PSBE23]. GLCM [DLT^{+}22]. glider [ACZ^{+}21]. GLM [CKC^{+}24]. Global [HRW^{+}24, LB21, MBFS21, CBW22, COPM23, CYP^{+}20, Cor20, CKC^{+}24, HES^{+}20, HES^{+}23, KMH^{+}21, LC20, LhNA^{+}21, LhNA^{+}22, LMW^{+}23a, MLP^{+}20, MPL^{+}20, OHH^{+}21, PG23, PSG^{+}22, SQDM21, SMP20, TRMB20, XLM^{+}23]. global-scale [CYP^{+}20, Cor20]. globe [HHFB23]. globosa [LCMB21]. goal [NMV^{+}23]. goals [GCMU23, NAM22, HL23]. going [ARSC22]. golden [BDP^{+}23]. Gompertz [CMG20]. good [BCMV20, LJRM^{+}20]. Gorges [XWX^{+}24]. governance [PDW23]. GPLake [CDJ^{+}22]. GPLake-S [CDJ^{+}22]. GPP [SPPC21, YYS^{+}22]. Grabarnik [Mea20]. gradient [ALS^{+}21, NSS^{+}23, SGTS21, SGTS23, SGCFB22]. gradients [BMS^{+}20, LBGD21, OMP^{+}23, RLS^{+}24]. grain [COJ23, KFM^{+}23]. granularity [OSH^{+}21]. grape [EGAS23]. grass [HGBK23, PAdS^{+}24]. grasshopper [LBGD21]. Grasshoppers [HSL^{+}22]. grassland [HCL^{+}24, HCP^{+}20, LWL24, MTW^{+}22, MPA^{+}24, WHW^{+}23, SOP22, WTT^{+}21, Yu20, ZFT20]. grasslands [AG22, ACFM22, BAZ22, CSM23, DP20, HHT21, MPCG21, PSM^{+}21, SHT21]. grazer [KWG^{+}22b]. grazing [CSM23, JZS^{+}21, ZGG22]. Great [KWM^{+}24, ZFT20, AJ22, ABR24, BWS^{+}20, KAT^{+}22]. greater [OHP^{+}23, ERP21]. Greece [LPD^{+}20]. green [GZW^{+}20, MKR20, RRCZ20, SBF^{+}21, SSB22, SMV^{+}20, SRB^{+}20]. green-lipped [RRCZ20]. greenhouse [AAC20, BST^{+}20, BWT^{+}24, HCP^{+}20, KMH^{+}21, MTS^{+}20, TQM^{+}24, TCG^{+}20, ZBH^{+}22]. greening [WLZ^{+}23b]. greenspace [CLY^{+}23]. grey [CAB^{+}21, HKG^{+}24]. grizzly [GCS^{+}21]. gross [GLK^{+}22, GZD^{+}23, HCX^{+}21, KWP^{+}23, ZZLF24]. Ground [ABC^{+}22, EOÖ21, FSZ^{+}20, FZS^{+}20]. ground-level [EOÖ21]. Ground-truthing [ABC^{+}22]. grounds [KiAM^{+}20]. Groundwater [MSS24, Che21, ZSS20, BPN^{+}20]. groundwater-dependent [ZSS20]. group [AIA^{+}23, GAC^{+}21]. groups [BLP20, NHGP23, RBGB23, vWVH20]. grouse [OHP^{+}23]. Growth [KDC^{+}21, dSMPA^{+}22, dOPdMdS^{+}21, VLGG^{+}20, WXT^{+}20, AZMD^{+}22, AZAD^{+}23, AXMJ22, AG22, AWW24, CZ21, CMG20, CAK^{+}23, CMS^{+}21, DHW^{+}20, DMG22, DPC20, dMFOM^{+}20, HMVR21, HGBK23, HFP22, Jia21, JASW23, KLS20, KAB23a, KABL21, KGBL24, KLA^{+}20, KBAC20, LK24a, LM21, LZF21, LCMB21, MPSS24, MGHT20, MJ24, MFF^{+}24, MLP^{+}20, PCN^{+}23, PKO^{+}21, Pre22, RGB22, SE20a, SE20b, SVK20, SGH^{+}23, mSzZK^{+}21, SMV^{+}20, SRB^{+}20, SLC^{+}20, TMFV22, TKO^{+}22, WYJ^{+}22, YA24, YSFX23, ZYK22, ZS20]. growth-maintenance [AG22]. Guangxi [ZHJQ23]. guarantee [CKC20]. guaranteed [AZBdAI23]. GUESS [ERP21]. guianensis [dOPdMdS^{+}21]. guidance [Gri20]. guide [BWCD23, DWG22, Tur20]. guided [HSJ^{+}20]. guideline [SACEU^{+}21]. guiding [PM22]. guild [CKC20]. guilds [LLZM20]. Gulf [BHC22, CGRR^{+}24, MZLRSZ21, RWF22, WSFZ21, ZHJQ23]. gull [vdMvDSL20]. gut [LSL21]. GWR [ZSS21].

H [HKSM23]. Habitat [HSF^{+}20, HK23, KWM^{+}24, KLNMG22, YHC^{+}21, BOLC^{+}22, BHC22, BLLL21, BBSHC22, BG22, CWH^{+}21, CLW23, CKK^{+}24a, CKK^{+}24b, CKK^{+}24c, DL21, DSHB23, GVG^{+}23, HYY^{+}24, ILLM20, JLH^{+}21, JHKT^{+}23, JRD^{+}22, LZF^{+}24, LZD^{+}24, LGA^{+}21, LUT^{+}24, MS23, OEC24, OHP^{+}23, PWN^{+}24, PTNR^{+}20, SvV21, SWW^{+}23, SKH^{+}20, WLZ^{+}23a, WWF^{+}23, WKL^{+}21, WWCT24, Yao21, YXY^{+}20, ZWL^{+}21, cNKF24]. habitats [Che21, DTD^{+}23, GRCSA^{+}20, KE24, YDQ^{+}23]. Haliaeetus [RM20]. Haliotis [DLZ^{+}21]. Handling [ŁKPK22]. Hanjiang [XZZ^{+}22]. hannai [DLZ^{+}21]. harbour [WSL^{+}22]. hardcover [Mea20]. hardwood [JLH^{+}21, MLL20]. hare [LG20]. Harmful [KSJ22, CR24, CMKF22, CLH23, PTG^{+}20, RFPRC23, SIOM24]. harvest [dSCdA21, DAG^{+}21, DDL^{+}20, SLLW21]. harvested [MLV24]. harvesting [BSE21, KLPJ24, NK23, PBH22, PHK23, SvV21, SM22a]. harvesting-induced [NK23]. hatching [LAK^{+}23]. hatchling [LHDBN20]. hazard [AAAR23, AAAR24]. healing [LHL22]. health [PF23]. heat [KFM^{+}23, SCH^{+}21, ZOJ^{+}23]. heatwave [SMV^{+}20]. heatwaves [MKR20]. heavy [LLW^{+}21, WSZ^{+}24]. Hebei [LZF21, LW23, ZLML20]. Hediste [GLA^{+}20]. Height [SE20a, SE20b, HBRK24, HLS^{+}23, HFP22, dAPdBMT23]. heights [IA23]. Heihe [XHS^{+}24]. helper [GAC^{+}21]. hemimetabolous [SKH^{+}20]. Hemiptera [PTBFP^{+}20]. Henriot [TCG^{+}20]. herbaceous [MLJ^{+}23]. herbicide [XGT^{+}24]. herbivore [Hop23, WSN^{+}21, YL22]. herbivores [GJC^{+}20, RMK21]. herbivorous [PGdAM^{+}20]. HermaDEB [LBS20]. hermaphrodites [LBS20]. hermaphroditism [MT23]. hermaphroditism-dioecy [MT23]. heterocyst [GSRB22]. Heterogeneity [GLH24, DPA21, LGFT^{+}20, LP22, vDC20]. heterogeneous [BR22, PEFW20, PMP^{+}21, SG20]. Heterosigma [SIOM24]. Hexi [LZF^{+}24]. hierarchical [BRSW24, KWC24, PDGB21, SWF21, SKW23, WSL^{+}22, WBD^{+}22b]. High [BRC^{+}21, LBGD21, YHC^{+}21, AMP^{+}23, AE22, ACZ^{+}21, BGA^{+}23, CJM^{+}23, nCLpW^{+}23, CDB^{+}20, RSB^{+}23, RPB^{+}23, YXH20, vdHNH^{+}20]. high-boreal [RPB^{+}23]. High-precision [YHC^{+}21]. high-quality [nCLpW^{+}23]. High-resolution [LBGD21, AMP^{+}23, RSB^{+}23]. Higher [CLHZ22, RFF^{+}23, CKC20, GJC^{+}20, LDS^{+}23, MT23]. Higher-order [CLHZ22]. Highly [DMGP20, FSY22, WGMR23]. highly-biodiverse [WGMR23]. hilly [ZBW^{+}24]. Himalayan [BS20]. Himalayas [BWSS21]. Hirst [AFdSM23]. hirundinacea [ARSC22]. Historical [ZPZ^{+}22, AG22]. histories [RPG20]. history [BMB21, CTL22, CP20, JOE23, KD20, NK23, SLYL^{+}20, vDC20]. history-dependent [CTL22]. hitchhiker [DWG22]. hive [LBT^{+}22]. Hoeksche [PRdN^{+}20]. holes [BMB24]. holey [TTPA21]. Homarus [BHC22]. homeostasis [AKY^{+}22]. Honey [LBT^{+}22]. honeybee [MMC^{+}21]. Hopf [NDNN22]. host [BRAL23, CLPZ20, MSLMM22, MM21, NWJ^{+}22, Oku21, PTBFP^{+}20, RWS^{+}21, SE21, VSB22, ZMG^{+}21, dGMCM^{+}23]. hosts [Oku21]. hot [LLZM20, MZLRSZ21]. Hotan [CWY^{+}24]. hotspot [BRC^{+}21, ZZWF21]. hotspots [CXZ^{+}23]. hourly [BH20]. household [GM23]. housing [CLY^{+}23, JASW23]. HSPF [LKP^{+}20]. Huangguashan [JLLS23]. Hugo [Pas21]. human [CBW22, CXZ^{+}23, DYMD22, FMSR20, GMS^{+}23, GRCSA^{+}20, HL23, KCS22, MDP21, MPMT23, PMB^{+}22, RGRMRR20, ST24, WG21, WFD23, YDQ^{+}23, YPS^{+}21]. human-altered [GRCSA^{+}20]. human-ecosystems [RGRMRR20]. human-natural [KCS22]. human-transformed [MPMT23]. humans [AARS23, ZHC23]. Humboldt [CFG^{+}22]. humidity [WBD^{+}22b]. humpback [SWF21]. hunter [LP22]. hunting [CPAT20, KLPJ24, RRFLV^{+}22]. hurdles [GSMD24]. Hutchinson [MYH^{+}20]. hybrid [BOLC^{+}22, CKK^{+}24c, LNK^{+}21, PR22]. hybridization [SCB^{+}20]. hybrids [CKK^{+}24a]. HyDiaD [BOLC^{+}22]. hydra [dAdSCA20]. hydraulic [KSJ22]. hydrocarbons [ZYZ22]. Hydrodynamic [YHC^{+}21, Che21, HSF^{+}20, HLS^{+}23, PRM^{+}24, ZZG^{+}24]. hydrodynamically [SJH^{+}23]. hydrologic [LHJ^{+}24, TKO^{+}22]. hydrological [GSA24, HLZ^{+}22, LZD^{+}24, NA22, ZRMW^{+}20, dPGDdFS24]. hydrology [LMW^{+}23b, NWO21, NO23]. hydropower [DYW^{+}22]. hyperallometric [SGLK20]. hyperbolic [PSZ20]. hypereutrophic [DAQ20]. hypotheses [SGLK20]. hypothesis [Lef23]. hypoxia [FTH24]. Hyrcanian [LhNA^{+}22, LhNA^{+}21].

ibex [LGFT^{+}20]. IBM [LBS20]. ice [DMdR24]. ice-cover [DMdR24]. ideal [Men22, dKS22]. idealized [KPLH24]. ideas [Ano20-49]. identifiability [SC21]. Identification [KT22, DTD^{+}23, DBH^{+}22, MP23, SGCFB22]. identified [BLB^{+}20, KPLH24]. identify [MYACC22, OSH^{+}21, SWW^{+}23]. Identifying [BGCJ21, JASW23, KRM^{+}22, MKEH20, XMWL23, BSU^{+}23, KDC^{+}21]. Ignoring [ŁKPK22]. II [BSPT23b, FZS^{+}20]. III [ML20]. Illustration [RGB22]. images [CSZY^{+}20, Lia23]. imbalance [BSCH23]. immersed [For20]. Impact [ELS23, JL23, JASW23, MSL^{+}21, NO24, SLYL^{+}20, AMK21, AFG24, BSS^{+}24, BMB21, CdCAC^{+}20, COJ23, CDN^{+}23, CDJ^{+}22, CL24, DAG^{+}21, HHBF24, HMC20, HLS^{+}23, JB22, KE24, LPSF24, LDXJ22, MGHT20, MTBB22, NA22, PDW23, QS23, RCCLMGV22, RBVL^{+}20, SLB^{+}23, TKO^{+}22, VMH^{+}21, Vas23, WAHA^{+}24, YLLL23, ZA20, ZYZ22, ZCJ^{+}20, ZZLF24]. impacted [CDT^{+}23, ZRMW^{+}20]. Impacts [CP22, CPAT20, KWG^{+}22b, XLS20, ZMF^{+}20, BPW^{+}21, BGS^{+}20, BNAR^{+}21, BRBT22, CRZ^{+}23, COPM23, CNSK21, Dav22, EBS23, ETAAH21, HKB^{+}20, KBSS23, LCW^{+}21, MCMP^{+}24, MCB^{+}23, MGC24, MBD^{+}22, MPA^{+}24, PK23, PTG^{+}20, WHW^{+}23, SBS^{+}20, SBS^{+}21, SG20, SBWB24, TGRM23, TCS21, XCZ^{+}20, ZXK^{+}24]. imperfect [BR22, SOV21]. implementation [BGA^{+}23, HKG^{+}24]. implemented [BGA22]. Implications [BHC22, CR24, CNSK21, HKSM23, LRA^{+}20, SE21, FEPR24, FLG^{+}23, FTH24, RBEC21, Pas21]. implicit [AE22]. import [NO24]. Importance [FS22, ALS^{+}21, CKK^{+}24c, EJ20, GMS^{+}23, HHT21, JHKT^{+}23, NPK^{+}23, WRC21, vdHNH^{+}20]. important [MFF^{+}24]. imprecise [XCZ^{+}20]. improve [CSCT23, GZD^{+}23, MVAD20, dRMJdS23, PTG^{+}20]. Improved [HKY24, HCX^{+}21, KWP^{+}23, KBAC20, Ste20, WMA21, BA20, CWW^{+}23, LCW^{+}21, NYH20, dOPdMdS^{+}21, SPPC21, Yu20]. improvements [LKP^{+}20]. improves [BSCH23, CLW^{+}24, JST^{+}22, OW20, RFF^{+}23]. Improving [AMP^{+}23, ASKR24, BLSR^{+}20, KMP^{+}23, PGSL^{+}20, SZPW23, SNM^{+}23, WXX^{+}23, YCI20, BGK^{+}20, GXZ^{+}24, TCG^{+}20]. IMTA [HK23]. in-season [HGBK23]. in-silico [PNGESS21, FCFCM24]. in-water [WWP20]. inbreeding [RMA^{+}24]. incidence [CMSD^{+}22]. including [DS24, HLZ^{+}22, YEL^{+}21]. inconsistency [Sak23]. incorporates [dAPdBMT23]. Incorporating [APF21, HML^{+}22, WBD^{+}22a, CL21, PGdAM^{+}20, VRRI^{+}24, WBD^{+}22b, XYZR21, ZYK22, ZHC23]. Incorporation [ALJ^{+}22]. Increase [AKM22, CAFT22, KHH^{+}22, KHH^{+}24]. Increased [cNKF24, FS22, GLH24, GGMR20, HHBF24, TCB21]. increases [ALJ^{+}22, KKF^{+}23]. Increasing [dGZHCA24, FWK^{+}20, SQDM21, WGMR23, WLZ^{+}23b]. incubation [CC20]. independent [HI22, PPH^{+}21]. indeterminate [CMG20]. index [HKSM23, HCL^{+}24, JPL21, PSZ20, WAHA^{+}24, ZSS20]. India [ESCA22, GMM20, JLV^{+}21, MSG^{+}21b, RSB^{+}23, SD24, SPPC21]. Indian [LHH^{+}24, BGC^{+}24, HMVR21, KWP^{+}23, RSC^{+}24]. indica [AFdSM23]. indicate [ZZŠ^{+}22]. indicators [WXX^{+}23, ZMK^{+}22]. indices [CSZY^{+}20, GSJC22, XMWL23, ZMK^{+}22]. Indigo [BGA22]. Indigo-N [BGA22]. Indirect [SWL^{+}23, KMAA23, TBM^{+}24]. Individual [APGD^{+}21, EV22, MKS21, Mea20, PMP^{+}21, RPG20, AVRC^{+}20, AHO^{+}20, APF21, BGS^{+}20, CDB^{+}20, CGVC21, CKK^{+}24a, CKK^{+}24c, DAB^{+}20, DWH^{+}21, EWE20, GRCSA^{+}20, HBMR24, HKG^{+}24, KCP^{+}23, KD20, Koo24a, KWG^{+}23, LGFT^{+}20, LDC^{+}20, LBI^{+}20, LQH24, MSG21a, MSG23, dSMPA^{+}22, MCMP^{+}24, MJ24, MPL^{+}20, OMC^{+}24, RFJ^{+}22, TBM^{+}24, WHE^{+}24a, WHE^{+}24b, WBW^{+}20, WYBX23, WBD^{+}22a, XJW21, YLS^{+}21, YDLT24, ZMS20, ZFP21, Mea20]. individual- [RFJ^{+}22]. Individual-based [APGD^{+}21, EV22, MKS21, Mea20, PMP^{+}21, RPG20, AHO^{+}20, BGS^{+}20, CGVC21, DAB^{+}20, DWH^{+}21, EWE20, GRCSA^{+}20, HBMR24, HKG^{+}24, KCP^{+}23, Koo24a, KWG^{+}23, LGFT^{+}20, LBI^{+}20, LQH24, MSG21a, MSG23, MCMP^{+}24, TBM^{+}24, WHE^{+}24a, WHE^{+}24b, WBW^{+}20, WBD^{+}22a, YDLT24, ZMS20, Mea20]. individual-level [AVRC^{+}20]. individual-specific [OMC^{+}24]. individuals [MOS21]. Indonesia [JASW23]. induced [dSBDH23, CBSB23, DBK20, GRAA22, GZ23, LPS^{+}21, NK23, SPPC21, YYS^{+}22]. industrial [DS24, KHS^{+}23]. industrial-natural [DS24]. industry [nCLpW^{+}23, PAPŁ^{+}23]. inequality [LMW^{+}23a]. inertial [CGV22]. infection [CGB^{+}23, CYP^{+}20, Cor20]. infections [DEMJ24]. infectious [GMM^{+}24, VSB22]. infer [GSM^{+}20, JVGC21]. inference [ŁKPK22, TDI^{+}20]. inferences [ATP20]. inferring [PDH^{+}22]. infestans [FC21]. infestation [BDP^{+}23, MM21]. influence [BHC22, BAM23, CKK^{+}24c, FCFCM24, FDSM20, GAMN^{+}23, IA23, JPMA^{+}23, KN20, KCHZ21, LCM24, MLMI23, MSLMM22, NL20, QZSV21, RLS^{+}24, SLB20b, SMD^{+}20, SKH^{+}20, WSM24, dAdSCA20]. influenced [KSJ22, XZZ^{+}22]. Influences [CMFC23, SHT21, CBW22, CLW23, MFH^{+}22, OHR^{+}23, SC24a, SM22a, SZL^{+}20]. influencing [LW23, XMWL23, ZSS21]. influential [BGCJ21, LBT^{+}22]. inform [CSMP21, GML^{+}20, WT20, WDW^{+}22]. Information [LB21, SPV^{+}20, ZL24, BAM23, HKB^{+}20, HNC^{+}21, LLW^{+}21, MGK^{+}23, MPMT23, RBMRR20, TTPA21, vWVH20]. information-content [MGK^{+}23]. informed [HMGK20]. Informing [HCW^{+}22]. infrastructure [SBF^{+}21]. inhabiting [SBMF23]. Inhibition [MYH^{+}20, FMRMAS24]. Initial [nCLpW^{+}23]. Initiative [LMW^{+}23a]. inland [XHS^{+}24]. inner [KNP^{+}21]. innovations [CAH22]. Ino [DLZ^{+}21]. inoculum [MGHT20]. inorganic [NKT^{+}20]. input [BS20, HLM^{+}20, MDB^{+}20, WHL^{+}23, WAHA^{+}24, XLM^{+}23, ZZŠ^{+}22]. input-output [HLM^{+}20, XLM^{+}23]. inputs [HRW^{+}24]. Insect [DBN^{+}22, DMG22, DKM^{+}21, MJ24, PSM^{+}21, PGdAM^{+}20, REGI^{+}24, RTH^{+}21, RCSS20, SE22, SKH^{+}20, SM22b, dGMCM^{+}23]. insects [CH22]. Insights [AADHA^{+}23, dSCdA21, HSP^{+}22, HCPG20, MAGR^{+}22, PYD^{+}21b, PTNR^{+}20, RMV^{+}24, ZWZ^{+}20, HBMR24, JZJ^{+}20, PSM^{+}21, SLB^{+}20a, dFT^{+}20, LJRM^{+}20, PYD^{+}21a]. inspires [PSW22]. instream [KTB^{+}23, LKP^{+}20]. instruments [BAM23]. intake [vdMvDSL20]. integer [SRÁMH^{+}20]. integral [JPL21, PSBE23, PFK^{+}22]. integrate [MTBB22, Mur24]. Integrated [BNAR^{+}21, DKM^{+}21, ETAAH21, BdCCdSAdS23, BMM22, DIS23, FMH^{+}20, GLA^{+}20, HK23, JZJ^{+}20, KFM^{+}23, KWG^{+}23, LUV^{+}23, LCMB21, LYM^{+}23, MSS24, MJdORPG23, PBF^{+}23, STW^{+}23, SWW^{+}23, SZL^{+}20, WAHA^{+}24, WSL^{+}22, Yao21]. integratedly [ZYK22]. Integrating [ATP20, Ano20-50, BKTM20, CLW^{+}24, JPL21, MBFS21, PGP^{+}22, SVK20, TPD^{+}24, VRS20, WDW^{+}22, WFD23, XJW21, CDB^{+}20, LZW20, PTNR^{+}20, SNP^{+}23]. Integration [BLB24b, LUT^{+}24, OW20, PRdN^{+}20, BNN20, OCT^{+}23, SKW23]. integrity [Gon23]. Intelligent [Lia23]. intensity [CP22, PAS23, TMNA20, ZLZ^{+}24]. intensive [HLS^{+}23, ZCJ^{+}20]. intentionally [SM22b]. inter [CKK^{+}24a, EVHS21, KD20, MPL^{+}20, PSG21]. inter- [PSG21]. inter-individual [CKK^{+}24a, KD20, MPL^{+}20]. inter-site [EVHS21]. interact [KN20]. interacting [MAB23]. interaction [CWY^{+}24, DYMD22, GGS^{+}22, HFP22, JBF22, MVC^{+}22, MM21, PVS^{+}21, RUTY23, SGCFB22, SGRDA^{+}24, TRMB20]. interactions [ABM20, AM21, BANG20, BNN20, CLPZ20, CZ21, CLHZ22, CDNN^{+}20, FHP^{+}22, GGMR20, HFP20, KNP^{+}21, LRA^{+}20, LP22, MDP21, OMP^{+}23, PSG21, PGSL^{+}20, RBEC21, SCG^{+}23, SGP^{+}20, SLLW21, SPV^{+}20, SPV^{+}21, WPL^{+}23, WMA21, WTT^{+}21, YL22, YEL^{+}21]. Interactive [KAB^{+}23b, KPLH24, TSBE23, CVTS20a, CVTS20b, ZFC^{+}23]. interacts [BSR21]. interception [LWL24]. intercropping [DDH^{+}22, KKP^{+}21]. interdependent [RK20]. Interdisciplinary [NPS^{+}23, LWZL23, RFHR24]. interface [MD23, TGL^{+}22, WPL^{+}23]. internal [VJK^{+}23]. International [Mea20]. interplay [AM21]. interpolation [RRFLV^{+}22]. interpolators [RAB21]. interpretability [AZAD^{+}23]. interpretation [AHH22, BM22a, LK24a, RSWS23]. interpretations [LPL22b].

Interpreting [AKY^{+}22]. interregional [ZLML20]. intersection [LHDBN20]. intertidal [GSM^{+}20, PR22, vdHNH^{+}20]. interval [ALK20, NMV^{+}23, RUTY23, ZZZD22]. interval-valued [RUTY23]. intervals [JPL21, LK24b]. intra [CKC20, RLM23]. intra-annual [RLM23]. intra-guild [CKC20]. intraguild [SM22a]. intraspecific [PSG21]. Intrinsic [MSML21, YZF20]. intrinsically [GZ23]. intrinsically-induced [GZ23]. introduced [SM22b]. Inuit [DSHB23]. invadable [MOS21]. invader [YB22]. invaders [MJS24, SG20]. invading [BAZ22]. invasion [AV22, Fou21, JL23, LBL22, MJS24, NL20, NTHC23, WLWL23]. invasive [CH22, JL23, LGG^{+}22, LDXJ22, SLB20b, SWL^{+}23, SC23, SBR20, TBP^{+}23, dGMCM^{+}23]. invasiveness [PBH22]. inventory [BSS^{+}24, IBB^{+}21, IBB^{+}22, PPH^{+}21]. inversion [TDI^{+}20]. investigate [DHW^{+}20, DPC20, MSG21a, MSG23, RWF22, Wu24]. Investigating [AAM^{+}22, CZ21, HZ22, MGHT20, PDH^{+}22, RSC^{+}23]. Investigation [GCSlR22, RKN^{+}21, WRC21]. investigations [CM21]. Ionian [RSC^{+}21, RSC^{+}23, LPD^{+}20]. ionising [VSK^{+}20]. ionizing [YXH20]. IOP [Pas21]. IPH [dPGDdFS24]. Iran [DYMD22, AAM^{+}22, CMHZ23, ETAAH21, EGAS23, ENARM^{+}23, KT22, NAM22, PK23, RSASP22, ZMK^{+}22]. irradiance [SDP21]. Irregular [CMKF22, BMB24, JM22]. irrigated [GLK^{+}22, JMPB24]. irrigation [GZD^{+}23, HKH^{+}23, LE23]. irrigation-related [LE23]. irruptive [STRC24]. ISBN [Mea20, Pas21]. ISBN-13 [Pas21]. Island [LUV^{+}23, XHZH24, SBS^{+}20, WF24, CLW23, GCSlR22]. Islands [LPD^{+}20, SOV21]. issues [PBH22]. Italian [GCMU23, NSS^{+}23]. Italy [GMS^{+}23, GSC^{+}23, MEGA21, SBF^{+}21, BBCD23, BARF20, CMHZ23, DSC^{+}21, LPA^{+}21]. iterative [FFM^{+}23]. iteroparous [KK20]. Ixodes [WBD^{+}22b].

January [Ano20e, Ano20q, Ano21e, Ano21q, Ano22q, Ano23q, Ano24l]. Japan [BWS^{+}20, II24, KiAM^{+}20, NO23, TDI^{+}20]. Japanese [HKSM23]. japonica [HKSM23]. jellyfish [EWE20, XFK21]. Jiangsu [KHS^{+}23]. joint [MJS24]. Jørgensen [Fat21]. JRR [Fat21]. July [Ano20f, Ano20r, Ano21f, Ano21r, Ano22r, Ano23r, Ano24m]. June [Ano20g, Ano20s, Ano21g, Ano21s, Ano22s, Ano23s, Ano24n]. Jurassic [PR21]. Just [PMZ^{+}21]. juvenile [BdBG^{+}21a, BdBG^{+}21b, Koo24a, LLB^{+}23, MBB20, MLP^{+}20, SPB^{+}24, SJH^{+}23, WRC21]. juvenile-driven [Koo24a]. JZB [XCZ^{+}20].

Kalman [CWW^{+}23]. Kamchatka [PA20]. Kansas [AAR23]. Karenia [GSJ^{+}22]. Karst [ZHJQ23]. Kashgar [WFW^{+}24]. Keeping [KAT^{+}22]. Kelp [VLGG^{+}20, SMS^{+}24]. Kemp [SPB^{+}24]. Kenya [GVG^{+}23]. Kerguelen [APGD^{+}21]. kernel [PSW22]. Key [YG21, CAK^{+}24, CNSK21, GAC^{+}21, GRCSA^{+}20, KDC^{+}21, XMWL23]. keystone [HNC^{+}21]. Kheyrud [ETAAH21]. Khuzestan [ENARM^{+}23].

Kinetics [GSJ^{+}22, AZAD^{+}23]. Kingdom [Pas21, KLA^{+}20]. kiwifruit [CMHZ23]. knowledge [CMGD^{+}22, DSHB23, JZS^{+}21, MS20, PRdN^{+}20, VRS20]. known [BGS^{+}20]. Korean [WLT^{+}23]. krige [BMB24]. krill [BMB21]. Kuznets [LMNV20].

L [LhNA^{+}22]. L. [EVT^{+}21, GRAA22, LhNA^{+}21, PAPŁ^{+}23, SWHZ22, SKW23, WSHZ24]. laboratories [Tur20]. laboratory [AZMD^{+}22, LEC^{+}23, SGH^{+}23, VN23]. labrax [WBW^{+}20]. lack [ZOJ^{+}23]. lactuca [LFN^{+}20]. Laemophloeidae [AJJ24]. lagoon [AACCIS20, CMGD^{+}22, PGB^{+}20]. Lagos [EON23]. Lagrangian [LWE^{+}22, LEC^{+}23, TMNA20, WSW23]. LAI [CSZY^{+}20]. Lake [PRM^{+}24, AJ22, BGCJ21, CDJ^{+}22, FWK^{+}20, GPN^{+}23, HSJ^{+}20, KUK^{+}21, TSUY23, Tsy22, WLLW20, WLQ^{+}23, WPL^{+}23, ZZŠ^{+}22, ZWZ^{+}20, DMdR24, DVVH21, HSF^{+}20, KKH^{+}24, MSL^{+}21, MFH^{+}22, SLB20b, TGRM23, WCG^{+}24, ZYZ22]. lake-floodplain [TSUY23]. lake-scale [AJ22]. Lakes [DMdR24, EL21, LCLG20, Ste20, YXY^{+}20, AJ22, KWM^{+}24, KAT^{+}22, LHJ^{+}24]. Lakeshore [SOV21]. Lamellidens [GMM20]. lampreys [AJ22]. Lancang [WLZ^{+}23a]. lance [SWF21]. Land [GXZ^{+}24, LZW20, MTP^{+}20, TLH^{+}20, ABC^{+}22, CP22, COS23, CPTD22, CLG20, DSC^{+}21, DWG22, Fou21, GZK^{+}20, HAZEF23, JLH^{+}21, JLLS23, JHL23, KMH^{+}21, KMP^{+}23, LBGD21, LZSQ20, Lia23, LSS^{+}21, LLW^{+}21, MTZA23, NMV^{+}23, NAM22, PHL20, RSASP22, SBM^{+}22, SWP^{+}23, TGRM23, TGL^{+}22, USA21, WWC^{+}22, WMA21, WFW^{+}24, YPS^{+}21, YKZ^{+}23, ZLLZ24, ZHWZ23]. land-atmosphere [GXZ^{+}24]. land-based [CLG20]. land-cover [WWC^{+}22]. Land-Use [TLH^{+}20, LZW20, CP22, COS23, JHL23, RSASP22, WWC^{+}22]. land-use-based [DWG22]. land-use/cover [COS23]. land-use/land-cover [WWC^{+}22]. lands [ESCA22, TSBE23]. Landscape [DLT^{+}22, HBP^{+}21, MTP^{+}20, dMT23, AJG22a, AJG22b, AGP20, BSR21, CWY^{+}24, FFM^{+}23, FHP^{+}22, GLH24, GSMD24, HMDO20, JSL^{+}20, KS22, Kos22, LDR^{+}20, LRA^{+}20, LP22, LZL23, MPMT23, MS23, MTZA23, OW20, PTRP23, SZPW23, SMS^{+}24, XMWL23, ZBW^{+}24, cNKF24]. Landscape-scale [HBP^{+}21]. landscapes [BWCD23, DYMD22, KLNMG22, LPD^{+}20, MSLM^{+}24, MGdM^{+}20, MMMS23, SJA^{+}22, VBH20, WGMR23]. LandSDS [LZW20]. Laplace [WSL^{+}22]. larch [GHK23]. large [BS21, BdCCdSAdS23, DYW^{+}22, GCB^{+}22, GJC^{+}20, GMM^{+}24, LBGD21, LEC^{+}23, PTRP23, PBS21, WPL^{+}23]. large-scale [GMM^{+}24]. Larger [CDC^{+}22]. Larimichthys [RJY^{+}20]. Larix [XLS20]. Larus [ILLM20]. larvae [GMM^{+}24, JPMA^{+}23]. Larval [PCL24, CMFC23, PYD^{+}21a, PYD^{+}21b]. last [BPL^{+}20]. late [HEL^{+}21, PR21]. Latent [Por21]. laterally [Che21]. latissima [VLGG^{+}20]. latitude [ALJ^{+}22]. latitude-adjusted [ALJ^{+}22]. latitudinal [NSS^{+}23]. Laurentian [KWM^{+}24]. law [KLS20, XJW21]. layer [FSZ^{+}20, FZS^{+}20, TMNA20]. layout [LBT^{+}22, ZZQ23]. lconnect [MS23]. leaching [BGA22]. lead [JJN22]. leads [CMA23, HKY24]. leaf [CLW^{+}24, HKSM23, KHH^{+}22, KHH^{+}24, Mas23, MLJ^{+}23, NYH20, PTF23, PCN^{+}23, PKO^{+}21, PNGESS21, SLHB^{+}23]. leaf-appearance [PTF23]. Learning [Ano20-51, ASKR24, ASHK22, BSCH23, CAK^{+}23, CAK^{+}24, EOÖ21, GZD^{+}23, HSJ^{+}20, KKW^{+}21, Lia23, MBFS21, PBS21, RSWS23, RFPRC23, SPV^{+}21, TK20b, WLY^{+}23, WCSX22, WKL^{+}21, WrJP^{+}22]. Learning-by-doing [ASKR24]. least [PCK^{+}22]. leatherback [CMG20]. Lei [ZCJ^{+}20]. leishmaniasis [ZMG^{+}21]. lemuru [PYD^{+}21b, PYD^{+}21a]. length [BPK^{+}20, LCMB21]. lens [DK22]. leopards [PWN^{+}24]. Leslie [SQDM21]. Lesotho [RMO22]. Less [XGT^{+}24, TSdOCV21]. lesser [BGS^{+}20, vdMvDSL20]. lesser-known [BGS^{+}20]. lesson [Fou21]. Lessons [For24, LBGD21]. Level [Por21, AVRC^{+}20, DvDA24, EOÖ21, LYL^{+}24, PPH^{+}21, PR22, STW^{+}23, TLC^{+}22, VSS^{+}22]. levels [CCH20, DBK20, Jor22, LYC^{+}21, PCK^{+}22, SKW23, SBMG24]. leveraged [YA24]. Lianas [dFT^{+}20]. library [LQH24]. lice [EVHS21, GMM^{+}24, MM21]. Liebl. [EVT^{+}21]. Liebl./Quercus [EVT^{+}21]. Life [JOE23, AZAD^{+}23, BMB21, CMGD^{+}22, CTL22, CP20, EL22, HKG^{+}24, KD20, KBSS23, LUT^{+}24, MMS^{+}24, MFF^{+}24, NGA^{+}20, NK23, RPG20, SLYL^{+}20, SINS22, vDC20]. life-histories [RPG20]. Life-history [JOE23, KD20, NK23, SLYL^{+}20]. lifespan [KN20]. Light [SBMF23, BWCD23, HMC20, PNGESS21, SMS^{+}24, TMNA20, Yu20]. light-climate [HMC20]. likelihood [NT23]. likely [WWA^{+}22]. limitation [LLB^{+}23, SGH^{+}23]. Limited [NK23, HKB^{+}20, HLW^{+}20, HLZ^{+}22, ZWL^{+}21]. limiting [CKK^{+}24b]. limnology [DBK20]. limpet [GSM^{+}20]. linear [SDD^{+}20]. lines [CDC^{+}22]. link [BPF24, MGWG22]. linkages [MFPIBGES23]. linked [PRM^{+}24]. Linking [HKH^{+}23, JTH24, JLH^{+}21, KWG^{+}22a, AB21, Jor22, REGI^{+}24]. links [AS22, MPMT23]. Liolaemus [NMGFOB21]. lipped [RRCZ20]. list [SCACB23]. literature [MMRBLN20]. Litopenaeus [dMFOM^{+}20]. litter [LDXJ22, MLJ^{+}23, SLHB^{+}23]. Liulin [ZZLF24]. lived [LGK^{+}23]. livelihoods [CNSK21]. lives [SCH^{+}21]. livestock [GJC^{+}20, JST^{+}22, MTS^{+}20, MD23]. Living [BLP20, TLH^{+}20, ZZQ23]. LMfireCF [ERP21]. loading [CR24]. loadings [WCG^{+}24]. loads [NA22]. Lobesia [AZMD^{+}22, AZAD^{+}23, CdCAC^{+}20, PSM22]. lobster [BHC22, HBMR24]. local [CMGD^{+}22, LSL21, MSL^{+}21, MAGR^{+}22, MKEH20, MM21, OZSU22, PDW23, PRdN^{+}20, PDGB21, RLS^{+}24]. localized [VRW^{+}21]. location [LDC^{+}20]. locations [MGK^{+}23]. lock [EOÖ21]. lock-down [EOÖ21]. locust [LAK^{+}23, SGV^{+}24]. Loess [SZL^{+}21, TGL^{+}22, WXT^{+}20, YLLL23]. Loggerhead [MZLRSZ21, CRZ^{+}23, II24, PTD^{+}23]. logging [BSS^{+}24]. LoggingLab [BSS^{+}24]. logic [DSHB23, KLH^{+}24]. logic-embedded [KLH^{+}24]. Logistic [SLC^{+}20, KAB23a, PSG21, YXY^{+}20]. Lombard [RCCLMGV22]. London [CCH20]. Long [RSASP22, ABH^{+}20, BLB^{+}20, DYMD22, FMRMAS24, JLV^{+}21, JSD^{+}22, PKO^{+}21, SZ20, YSIB24]. long-range [FMRMAS24]. Long-term [RSASP22, ABH^{+}20, BLB^{+}20, DYMD22, JLV^{+}21, JSD^{+}22, PKO^{+}21, YSIB24]. longiceps [HMVR21]. longleaf [JLH^{+}21, SLHB^{+}23]. look [BBSHC22]. loop [AIA^{+}23]. loops [RGRMRR20]. Lorestan [AAM^{+}22]. loses [MKN23]. loss [ALK20, JRD^{+}22, LSL20, LAK20, MMdP^{+}21, cNKF24]. losses [LLLZ21, SAdA^{+}22]. Lotka [NMGFOB21, SIOM24]. Low [HP20, MTL^{+}23, RM20, Lef23, LCM24, MDB^{+}20, WAHA^{+}24, YXH20, ZBW^{+}24, ZL24]. low-dose [YXH20, ZL24]. low-input [MDB^{+}20, WAHA^{+}24]. lower [GZ23, MJdORPG23, LUT^{+}24, XZZ^{+}22]. lowland [HCPG20]. Loxodonta [DDG^{+}21]. LPJ [ERP21]. LPJ-GUESS-LMfireCF [ERP21]. LSTM [ZHWZ23]. Ltd [Pas21]. LUCIA [WMA21]. lupus [HKG^{+}24]. Lycium [SWHZ22, WSHZ24]. lysis [KPLH24].

Macau [XWZ^{+}22]. Machine [Ano20-51, CAK^{+}23, ASHK22, BSCH23, CAK^{+}24, GZD^{+}23, HSJ^{+}20, KKW^{+}21, MBFS21, RSWS23, RFPRC23, WKL^{+}21]. Machine-learning [CAK^{+}23, BSCH23]. Macro [RTH^{+}21, NO23]. macro- [NO23]. Macro-symbiosis [RTH^{+}21]. macroalga [LFN^{+}20]. macroalgae [VN23]. macroecology [GRS^{+}20]. macroscopic [ZZWF21]. magnitude [CGB^{+}23]. main [BNAR^{+}21, HLW^{+}20, RJY^{+}20]. Maine [BHC22]. mainland [ZXK^{+}24]. maintained [SLHB^{+}23]. maintaining [MSG21a]. maintenance [AG22, MJ24]. maize [AMK21, dSBDH23, GLK^{+}22, GZD^{+}23, VSS^{+}22, WXH^{+}22, KFM^{+}23]. Making [CSMP21, For20, BVR^{+}20, GZK^{+}20, JST^{+}22, MTW^{+}22, PBS21]. Malaba [COS23]. Malakisi [COS23]. malaria [KE24]. Malayer [KT22]. males [HLB^{+}20]. mallards [WWF^{+}23]. mammal [Dav22, JSD^{+}22]. mammals [TSdOCV21]. Managed [Dav22, ACFM22, BLL^{+}23, DS24, MPCG21, SLB^{+}23]. Management [Ano20-51, Mea20, ABH^{+}20, AZBdAI23, BH21, BAM23, BM21, CSMP21, CAB^{+}21, DBN^{+}22, GSC^{+}23, HKH^{+}23, HAZEF23, JLH^{+}21, LWZL23, LCMB21, MLV24, MVC^{+}22, MJdORPG23, NGE^{+}21, OEC24, PGP^{+}22, PPH^{+}21, dOPdMdS^{+}21, PSM^{+}21, Pre22, RS23, SRÁMH^{+}20, SAL^{+}22, SFGB23, STW^{+}23, TSBE23, WBW^{+}20, Yao21, YKZ^{+}23, ZCJ^{+}20, ZBH^{+}22, Mea20]. Management1 [LL22]. Managing [BSU^{+}23, SYL^{+}24, FTH24, Lov24, MAB23]. manatee [Che21]. MANGA [BPN^{+}20]. mangrove [LK20, ZRMW^{+}20, BPN^{+}20]. Manning [HLS^{+}23]. map [DLT^{+}22, PTRP23, WHdBB21]. Mapping [EON23, GCS^{+}21, DK22, PDGB21]. maps [LCP^{+}20]. March [Ano20h, Ano20t, Ano21h, Ano21t, Ano22t, Ano23t, Ano24o]. margaritifera [SLYL^{+}20]. margin [LWL24]. Marginal [Men22, ESCA22]. marginalis [GMM20]. marginalized [KCS22]. marina [SIOM24]. Marine [BDA^{+}20, BWS^{+}20, BARF20, CMGD^{+}22, CBG^{+}22, CL24, CNNS^{+}21, DS23, Dav22, FSY22, GRGHM21, GDC24, HZ22, HES^{+}20, HES^{+}23, JPMA^{+}23, JSD^{+}22, KKPR20, LLB^{+}23, MS20, MBFS21, MSML21, MMRBLN20, MZLRSZ21, MFF^{+}24, MBD^{+}22, RUTY23, RAB21, SBS^{+}21, SMV^{+}20, WA20, YD20].

Marion [GCSlR22]. maritima [HBP^{+}21]. maritime [YD20]. marketability [CLY^{+}23]. Markov [LM23, SOV21]. marsh [LBGD21, QZSV21, ZZG^{+}24]. marten [SWL^{+}23]. Maryland [BRBT22]. mass [GRGHM21, LGA20, RWF22]. MASTIFF [MBG^{+}22]. masting [VK24]. Mate [WZWF20]. material [LZF21]. maternal [CLW23, CC20]. math [KAT^{+}22]. Mathematical [GGS^{+}22, LJP^{+}22, WGAH22, BDP^{+}23, BCMV20, EBS23, GAC^{+}21, GRAA22, Ino22, RCCLMGV22, RWF22, SE20a, SE20b, SvV21, ST24, WPL^{+}23, YXY^{+}20]. mathematics [ZWL^{+}21]. matrices [TK20a, TTPA21]. Matrix [TK20a, CCJK^{+}20, CP20, LKO20, LSG21, PTG^{+}20, SCB^{+}20, SdlP20, XYZR21, dVDC20]. Matt [EVT^{+}21]. matter [AB21, FMSR20, WTT^{+}21]. matters [BS20, Fou21, RWS^{+}21, VVH^{+}20]. Mauve [EWE20]. MaxEnt [HPMB20, SRTS20, SM22b, ATP20, SC24a]. maximize [DGVVP23]. Maximum [dSCdA21, AKM22, GCS^{+}21, HMDO20, dRMJdS23, dAPdBMT23, RLM23, WT20, XJW21]. May [Ano20i, Ano20u, Ano21i, Ano21u, MBD^{+}22, Ano22u, Ano23u, Ano24p]. mayfly [KFJS20]. MCDA [USA21]. MCDM [ŁJ21]. MCR [HYY^{+}24]. meadow [ALS^{+}21, PR22]. mean [Dav22]. measure [RP22]. measured [AAR23, GCSlR22]. measurements [ASHK22]. measures [JTH24]. Measuring [DPA21, GPN^{+}23]. mechanical [MLMI23]. mechanical-ecohydrological [MLMI23]. mechanics [YEL^{+}21]. mechanism [BGA^{+}23, CR24, DDH^{+}22, SIOM24, SZL^{+}21, YXH20]. mechanisms [SINS22, ZXK^{+}24]. Mechanistic [SMS^{+}24, BPW^{+}21, BGA22, CSBS21, EJ20, HMGK20, HZ22, HFSL^{+}24, KN20, KBAC20, LLC^{+}24, MK21, MJ24, MLJ^{+}23, MBG^{+}22, MOS21, RBEH21, dAWG23]. mechanized [NGA^{+}24]. media [PDW23, PAS23]. media-based [PAS23]. mediate [SM22a]. mediated [VVH^{+}20]. Mediterranean [RSC^{+}21, RSC^{+}23, AACCIS20, BRC^{+}21, BCMV20, DAQ20, LPD^{+}20, MTL^{+}23, PGSL^{+}20, SBF^{+}21, SMD^{+}20]. Mediterranean-type [AACCIS20]. megafauna [RAC^{+}24]. meiofauna [vdHNH^{+}20]. Mekong [LUT^{+}24]. melanogaster [SLM24]. melitensis [LGFT^{+}20]. mellifera [JSL^{+}20, LBT^{+}22]. Membracidae [PTBFP^{+}20]. memorized [Sak23]. memory [RBGB23, SZ20, TPD^{+}24]. Mercury [CB20]. Merging [RBEH21]. Meriones [ZMG^{+}21]. mesic [ZFT20]. Mesocarnivores [SOV21]. mesopelagic [CGRR^{+}24]. mesotidal [CMGD^{+}22]. Messina [SGP^{+}20]. meta [NMV^{+}23, OKS^{+}21, RLS^{+}24]. meta-ecosystem [RLS^{+}24]. meta-ecosystems [OKS^{+}21]. meta-goal [NMV^{+}23]. metabolic [CMA23, GSM^{+}20, KN20, LK24a, MTS^{+}20, SGLK20, VVH^{+}20, XJW21]. metabolism [BMB21, BCW^{+}23, CLG20, DvDA24, EL22, LZF21, MKS21, WZZF23]. metacommunities [AE22, LQH24]. Metacommunity [BACS22, LBL22]. MetaIBM [LQH24]. metal [LLW^{+}21]. metallophyte [BMS^{+}20]. metamodelling [CGVC21]. metamorphosis [PDG^{+}20]. metaphor [SGPSzC22]. metapopulation [FdKO21, FAC22, NTHC23, RGB^{+}23, SdlP20, BGC^{+}23]. metapopulations [BSR21, TA20]. meteorology [JPL21]. methane [SMM^{+}22]. method [AAM^{+}22, DL21, For20, FG21, KFM^{+}23, KK20, ML20, MSS24, PMZ^{+}21, SSM24, SZ20, SDH21]. Methodological [SLB^{+}20a, DBDR22, dMT23]. methodology [OMC^{+}24]. Methods [LAC^{+}24, Mea20, AIA^{+}23, GZD^{+}23, JL23, LGK^{+}23, MDT^{+}23, WT20]. metric [GAM20, OW20]. metrics [GSJC22, GWH20, SBFBDM24, SDSP^{+}21]. metrix [CC23]. Metropolitan [MTP^{+}20, AXMJ22, BBCD23, NSS^{+}23, WWC^{+}22, SBF^{+}21]. Mexico [CGRR^{+}24, MZLRSZ21, RWF22, WSFZ21]. MFA [GM23]. MGB [dPGDdFS24]. MGB-IPH [dPGDdFS24]. Michigan [PRM^{+}24, SLB20b]. micro [KKP^{+}21, NO23, PCK^{+}22]. micro-organisms [KKP^{+}21]. micro-plastics [NO23]. micro-scale [PCK^{+}22]. Microbial [MSL^{+}21, AKY^{+}22, KAB^{+}23b, KKPR20, MKN23, YXH20]. microbiome [LLZM20]. microbiota [LSL21]. microclimate [MK21]. microclimates [SVK20]. Microclimc [MK21]. Microcystis [DAQ20]. microphytobenthos [vdHNH^{+}20]. microphytobenthos-meiofauna [vdHNH^{+}20]. microplastics [BRBT22]. Micropterus [LBI^{+}20]. microsimulation [MKR20]. mid [KKW^{+}21, WWF^{+}23]. mid-continent [WWF^{+}23]. middle [XMWL23, XZZ^{+}22]. Middle-lower [XZZ^{+}22]. migrating [HBMR24, KTB^{+}23]. Migration [WWA^{+}22, AKM22, ASL^{+}20, BdBG^{+}21a, BdBG^{+}21b, CSB22, DZBB24, DLY^{+}22, Fer23, KiAM^{+}20, MBB20, NPS^{+}23, PTD^{+}23, PMB^{+}22, SdlP20, SJH^{+}23, WGK^{+}21, ZLL^{+}23]. migratory [ARSC22, DAB^{+}20, SGA22, SGV^{+}24, XCH^{+}21]. mikimotoi [GSJ^{+}22]. milk [OZSU22]. Mill. [SKW23]. Minas [MRR^{+}20, dPGDdFS24]. MinBAR [RPT21]. Mini [OKS^{+}21]. Mini-review [OKS^{+}21]. minimal [RPT21]. minimalistic [DDDC21]. Minimizing [MTS^{+}20]. mining [LLW^{+}21]. mirabilis [HBP^{+}21]. mismatch [AAR23, PCL24]. missing [JM22, ŁKPK22, MGWG22, RBVL^{+}20, SPV^{+}20]. Mississippi [WWF^{+}23]. Missouri [BGK^{+}20]. mite [MMC^{+}21]. mites [TCG^{+}20]. Mitigating [LG20]. Mitigation [CDN^{+}23, CLG20, MLV24, NMC^{+}20]. mixed [DMGP20, HMC20, LhNA^{+}21, LhNA^{+}22, MTZA23, SRÁMH^{+}20, SMD^{+}20, TMNA20, WLT^{+}23, YSIB24]. mixing [EVT^{+}21, KUK^{+}21]. mixoplankton [SGTS21, SGTS23]. mixotroph [CR24]. mixotrophic [LBPL20]. mixotrophs [CMA23]. mixotrophy [CR24]. mixtures [Pre22]. mobile [FdKO21, FSY22]. mode [KS22]. Model [BWT^{+}24, DLT^{+}22, KUK^{+}21, KAT^{+}22, LWL24, PKB^{+}23, RSB^{+}23, RCSS20, TLH^{+}20, Tur20, YHC^{+}21, AJG22a, ABKA24, Abe20, AB21, AZMD^{+}22, AZAD^{+}23, AXMJ22, AE22, AS22, AM22, APGD^{+}21, ABC^{+}22, ACZ^{+}21, AACCIS20, BSR^{+}23, BSPT23a, BSPT23b, BOLC^{+}22, BDP^{+}23, BLA^{+}24, BAAN24, BMB24, BKTM20, BPN^{+}20, BGS^{+}20, BANG20, BGR20, BGA^{+}23, BGCJ21, BA20, BGC^{+}23, BMBB24, BGA22, BST^{+}20, BWS^{+}20, BNN20, BAM23, BRBT22, BLSR^{+}20, BAZ22, BCMV20, BSM23, BRAL23, BARF20, BDA^{+}20, BGC^{+}24, BTC^{+}22, CMGD^{+}22, CGRR^{+}24, CM21, CBG^{+}22, CdCAC^{+}20, CDJ^{+}22, CHR^{+}23, COS23, Che21, CJM^{+}23, CWW^{+}23, CXZ^{+}23, CVTS20a, CVTS20b, COPM23, CAK^{+}24, CL24, CNNS^{+}21, CGVC21, CMS^{+}21, CYP^{+}20, Cor20, CSB22, CAB^{+}21, CDC^{+}22, CFG^{+}22, DS23, DHW^{+}20, DBK20, DAG^{+}21, DDH^{+}22, DGSM21, DGVVP23, DDG^{+}21, DP20, DBN^{+}22, DDDC21, DLZ^{+}21]. model [DVVH21, DWH^{+}21, EL21, EWE20, EBMK20, EHS^{+}20, EL22, EBS23, ERP21, ESZ^{+}23, FMH^{+}20, FC21, FFM^{+}23, FEPR24, FMRMAS24, FGM^{+}24, FAC22, FSZ^{+}20, FZS^{+}20, GSR^{+}23, GAMN^{+}23, GMM20, GSC^{+}23, GGMR20, GRAA22, GRCSA^{+}20, GSRB22, GXZ^{+}24, GWD^{+}23, HMGK20, HMVR21, HLZ^{+}22, HBMR24, HRW^{+}24, HBC^{+}22, HCL^{+}24, HYY^{+}24, HML^{+}22, HL23, HES^{+}20, HES^{+}23, HH23, HEL^{+}21, HGBK23, HCP^{+}20, HFSL^{+}24, HLM^{+}20, HCX^{+}21, HLS^{+}23, HFP20, HFP22, JvVvHH24, JTH24, JLH^{+}21, JMHH22, JGG23, JLV^{+}21, JPL21, Jia21, JZHT23, JHKT^{+}23, JSL^{+}20, JSD^{+}22, KiAM^{+}20, KAMS22, KMH^{+}21, KAB23a, KN20, KCP^{+}23, KABL21, KWG^{+}22a, KE24, KKF^{+}23, KLH^{+}24, KMP^{+}23, KS22, KD20, KKH^{+}24, Koo20b, KWG^{+}23, KTB^{+}21, KWP^{+}23, KPLH24, KKPR20, LTR^{+}24, LGFT^{+}20, LE23, LDR^{+}20, LFN^{+}20, LUV^{+}23, LLC^{+}24, LPSF24, LKP^{+}20, LBGD21, LBI^{+}20, LYC^{+}21, LJP^{+}22, LWE^{+}22, Lia23, LDX^{+}23, LZW20, LCW^{+}21, LPS^{+}21, LWZL23]. model [LNK^{+}21, LYL^{+}24, LGA20, LCMB21, MK21, MKR20, MT23, MS20, Mat23, MHP^{+}23, MJ24, MSML21, MLJ^{+}23, MSLMM22, MBG^{+}22, MTBB22, MVAD20, MOS21, MLV24, MZLRSZ21, MFF^{+}24, MJS24, MMMS23, MPCG21, MPL^{+}20, MPA^{+}24, NWO21, NO23, NHGP23, NGA^{+}20, NDNN22, NT23, NYH20, NTHC23, NMGFOB21, OFH^{+}20, OHR^{+}23, PBF^{+}23, PG23, PDH^{+}22, PFM^{+}21, PVS^{+}21, PSM22, PSG21, PSBE23, PDD^{+}22, PPH^{+}21, PTG^{+}20, PGB^{+}20, PMB^{+}22, PAdS^{+}24, PMP^{+}21, PGdAM^{+}20, PR22, dAPdBMT23, PM22, QZSV21, RSWS23, RCCLMGV22, RWF22, RFCEL22, RFJ^{+}22, RBMRR20, RJY^{+}20, RUTY23, RBEH21, RMA^{+}24, RRSG21, RGB^{+}23, RGB22, RBVL^{+}20, SBS^{+}21, SLHB^{+}23, STQ^{+}22, SJA^{+}22, SSM24, SFGB23, SvV21, SBM^{+}22, SMM^{+}22, SWHZ22, SIOM24, SLM24, mSzZK^{+}21, SACEU^{+}21, SHF24, SWW^{+}23, SWP^{+}23, SDH21, SJH^{+}23, SNP^{+}23, STRC24, SBWB24, SRB^{+}20, SDP21, SBR20]. model [SM22b, TBCV22, TBP^{+}23, TWL^{+}22, TLL^{+}23, TSUY23, TBM^{+}24, TDI^{+}20, TKO^{+}22, TSBE23, TK20b, VRRI^{+}24, VMH^{+}21, VVH^{+}20, VJK^{+}23, VRS20, VSS^{+}22, WBW^{+}20, WRC21, WYJ^{+}22, WHL^{+}23, WLQ^{+}23, WPL^{+}23, WSW23, WSZ^{+}24, WSHZ24, WYBX23, WMA21, WBD^{+}22a, WWF^{+}23, WCSX22, WWA^{+}22, WKL^{+}21, XYZR21, XLS20, XCZ^{+}20, XLM^{+}23, YDQ^{+}23, YRK^{+}20, YXY^{+}20, YLS^{+}21, Yu20, YCI20, ZMG^{+}21, ZYZ22, ZZZD22, ZPZ^{+}22, ZZQ23, ZLLZ24, ZL24, ZHWZ23, ZHC23, ZZG^{+}24, ZGMFESD23, ZS20, dGZHCA24, dPGDdFS24, dVDC20, dFT^{+}20, BSPT23a, BSPT23b, DBDR22, GXZ^{+}24, dKS22, AJG22b]. Model-based [KUK^{+}21, JvVvHH24, WRC21]. modeled [AAR23]. Modeling [AJG22a, ASHK22, AV22, ABM20, ABRG21, AVP^{+}23, BMAME^{+}23, BLLL21, BP21, BWCD23, CRZ^{+}23, DTR^{+}22, DCT^{+}21, DAG^{+}21, GSR^{+}23, GM23, HHI^{+}22, Hop23, JLV^{+}21, JMPB24, KRMP23, KCS22, LSA^{+}21, LSL21, LCLG20, LSS^{+}21, LGL^{+}24, LCM24, LDXJ22, MMS^{+}24, MPSS24, MEGA21, MCB^{+}23, NKT^{+}20, NAM22, PK23, PBH22, PRM^{+}24, PZH^{+}21, QS23, RBLS23, RRP^{+}23, RKN^{+}21, RGD20, SKPL24, SGTS21, SPB^{+}24, SBF^{+}21, SCH^{+}21, SINS22, SWP^{+}23, SDP21, TCB21, Tsy22, VRW^{+}21, Vas23, VLGG^{+}20, WXH^{+}22, WLY^{+}23, WSW23, YFW^{+}21, YEL^{+}21, ZRMW^{+}20, AGS^{+}21, AGS^{+}22, AHO^{+}20, BZZT22, BdCCdSAdS23, BPL^{+}20, BAR24, CAK^{+}23, CAH22, CKC^{+}24, DAQ20, EOÖ21, EVHS21, ELS23, EV22, ERP21, dMFOM^{+}20, For24, GCB^{+}22, GSA24, GXZ^{+}24, HBP^{+}21, HZ22, HPMB20, HMDO20, HBRK24, HSD^{+}21, HCPG20, JZS^{+}21, JMC^{+}22, JM22, KKP^{+}21, KGBL24, KFJS20, KLPJ24, KBAC20, LJRM^{+}20]. modeling [LBB20, LJP^{+}22, LGA20, LUT^{+}24, MGC24, MVC^{+}22, NBML21, NA22, PGP^{+}22, dOPdMdS^{+}21, Pre22, PCK^{+}22, RMV^{+}24, RK20, RFHR24, RFF^{+}23, RBEC21, RBEH21, SAL^{+}22, SLYL^{+}20, SLC^{+}23, SHT21, SOV21, SC21, SKH^{+}20, TSdOCV21, WDW^{+}22, YDLT24, ZOJ^{+}23, ZL24, ZWZ^{+}20, vdHNH^{+}20, vdMvDSL20, AJG22b, SGTS23]. Modell [BS20]. modelled [CLW^{+}24, GHK23, JOE23, KHH^{+}22, KHH^{+}24, MPPM23]. Modelling [ABH^{+}20, AAAR24, AGP20, ABR24, BJKT21, BH21, BLL^{+}23, BAR20, BGA22, CH22, CBW22, COJ23, CDN^{+}23, COS23, CVTS20a, CMG20, CIL^{+}20, CSB22, DYMD22, DMM^{+}23, EWE20, FMSR20, GMM^{+}24, GRGHM21, HLB^{+}20, HES^{+}23, KWC24, KTB^{+}23, KHH^{+}24, LE23, LK20, LRA^{+}20, LBPL20, MLMI23, MAN^{+}20, MSLM^{+}24, MMMS23, NSS^{+}23, NHGP23, NS20, NPK^{+}23, NTHC23, Pas21, PEFW20, PKO^{+}21, RPB^{+}23, RSC^{+}21, SBS^{+}20, SE20a, SGP^{+}20, SLB^{+}20a, SGTS23, SLB20b, SPPC21, SML^{+}20, TGRM23, VMM^{+}21, VSK^{+}20, WHE^{+}24a, WWP20, WXT^{+}20, YLLL23, ZSS20, dAWG23, vWVH20, AADHA^{+}23, AMK21, Ano20-50, BdSG^{+}23, BPW^{+}21, BRC^{+}21, BdBG^{+}21a, BdBG^{+}21b, BS20, CFL^{+}22, CDNN^{+}20, CSMP21, CN20, DTR^{+}22, DIS23, DK22, DZBB24, DKM^{+}21, DWG22, DSHB23, FSY22, FLG^{+}23, GDF22, GGS^{+}22, GCS^{+}21, HKB^{+}20, HSP^{+}22, HHT21, HK23, KFM^{+}23]. modelling [KWM^{+}24, KWG^{+}22b, Kos22, LG20, LQH24, LSG21, MSG21a, MSG23, MSL^{+}21, dSMPA^{+}22, MAD^{+}20, MDP21, MDB^{+}20, MTW^{+}22, MCMP^{+}24, MFPIBGES23, MTT22, MSS24, MM21, NMC^{+}20, OBFH22, PTRP23, PBS21, PMP^{+}21, PGSL^{+}20, RS23, RPG20, RFPRC23, dSRDC^{+}23, RLdM20, RAC^{+}24, SCB^{+}20, SRTS20, SVK20, SACEU^{+}21, SCACB23, SWF21, SGA22, SC23, TMNA20, WG21, WT20, YKZ^{+}23, ZMS20, ZJEG24, ZLL^{+}23, vOVvdW23]. Models [ECC^{+}22, AHH22, AJ22, AMP^{+}23, ABB23, AS22, ATP20, AFdSM23, AWW24, ALJ^{+}22, ASKR24, ACFM22, BPF24, BM22b, BM22a, BLB24b, BSCH23, BMM22, BWSS21, BRSW24, BLSR^{+}20, BG22, BVR^{+}20, CSZY^{+}20, CSBS21, CCJK^{+}20, CWH^{+}21, CDB^{+}20, CIL^{+}20, CPTD22, dSCdA21, CKC^{+}24, CKK^{+}24c, DMG22, DAB^{+}20, DRZ^{+}24, EJ20, FDSM20, Fou21, FHP^{+}22, GVG^{+}23, GSH^{+}22, GWH20, GML^{+}20, GSM^{+}20, GLK^{+}22, GSMD24, HKH^{+}23, HFSL^{+}24, Ino22, IBB^{+}21, IBB^{+}22, Jag20, JJN22, JZJ^{+}20, KWC24, Kle23, Koo24a, LL22, Laz23, LEC^{+}23, LK24b, LPHO^{+}20, LKO20, LAC^{+}24, MLD^{+}23, MKS21, MGK^{+}23, MYACC22, Mas23, MBFS21, MMRBLN20, MVAD20, MBS22, MBD^{+}22, NGE^{+}21, OTR^{+}21, OHH^{+}21, OMC^{+}24, OCT^{+}23, OHP^{+}23, OW20, OKS^{+}21, PGP^{+}22, PRM^{+}24, Pie23, PAB^{+}22, PFK^{+}22, PTNR^{+}20, RSWS23, RSC^{+}24, RRCZ20, REGI^{+}24, RPT21, RRFLV^{+}22, SdlP20, SVK20]. models [SOV21, SKW23, SE22, STW^{+}23, SBFBDM24, SSB22, SGRDA^{+}24, SDP23, SBMG24, Ste20, TMFV22, TQM^{+}24, TCG^{+}20, TGL^{+}22, Tur20, VRS20, WHE^{+}24a, WHE^{+}24b, WSM24, WSL^{+}22, WTT^{+}21, WBD^{+}22b, WFD23, ZZŠ^{+}22, ZFP21, dAdSCA20]. modern [Ano20-49, SACE^{+}20]. modes [BDC^{+}22]. Modified [TLC^{+}22, VK24, BPB^{+}20, CMG20, LWZL23, NS20, YXY^{+}20]. modulate [TRMB20, VBH20]. module [TGB^{+}23]. modules [SWW^{+}23, ZMF^{+}20]. moist [KWP^{+}23]. moisture [AJJ24, HKH^{+}23]. moment [BLP20]. Mongolia [NWO21]. Mongolian [ZYM^{+}21]. monitoring [JZJ^{+}20, LL22, LM23, LGK^{+}23, PSZ20, REGI^{+}24, WDW^{+}22, WrJP^{+}22, Yao21]. Monod [WGAH22]. monodon [BAAN24]. monoxide [KKPR20]. monsoon [NDNN22]. montane [ZLL^{+}23]. montdorensis [TCG^{+}20]. Monte [PM22, SDH21]. Monte-Carlo [PM22, SDH21]. monthly [WHL^{+}23]. moose [EBS23, KAMS22]. morhua [JOE23]. Moroccan [RKN^{+}21]. Morocco [JB22, ECC^{+}22]. morphotypes [GSM^{+}20]. Morris [PMZ^{+}21, dGZHCA24]. Mortality [PSM22, BAAN24, CAK^{+}23, CAK^{+}24, CDC^{+}22, GRGHM21, HHBF24, LPS^{+}21, MCMP^{+}24, MKR20, NTHC23, PCL24, SKPL24, LPS^{+}21]. mosquito [FGM^{+}24, VMM^{+}21, WRC21, WKL^{+}21]. mosquitoes [BM22b, HLB^{+}20, KWG^{+}22a, NS20]. most [KDC^{+}21]. moth [LGG^{+}22, PVS^{+}21, SBR20]. Moths [RBGB23]. Motility [SBMF23]. motion [RGB22, WLY^{+}23]. motivated [PF23]. mountain [ACFM22, BGA^{+}23, IBB^{+}21, IBB^{+}22, SFGB23, SKW23, SC23]. mountainous [ZLLZ24]. Mountains [WYBX23, ZZLF24, CWH^{+}21]. Movement [MKS21, Pér24, ABB23, ABR24, Ano20-50, AJJ24, BLB24b, BTC^{+}22, CDNN^{+}20, CNNS^{+}21, DDG^{+}21, JSD^{+}22, PBS21, TLL^{+}23, TPD^{+}24, WWF^{+}23, dKS22]. Movement-based [Pér24]. movements [BLP20, GMM^{+}24, Sak23, STQ^{+}22]. Moving [LBN^{+}22]. MRIO [XLM^{+}23]. MSY [RLM23]. Multi [BCW^{+}23, MTBB22, NMS24, TRMB20, Ano20-50, AHO^{+}20, BOLC^{+}22, BGCJ21, BM21, CWH^{+}21, EL21, GLA^{+}20, Gat21, HHBF24, HLM^{+}20, HK23, IBB^{+}21, IBB^{+}22, JVGC21, KFM^{+}23, MP23, MTZA23, MPCG21, MSS24, NWJ^{+}22, NAM22, OMC^{+}24, OSPZ22, PBS21, RMA^{+}24, SRÁMH^{+}20, SWP^{+}23, VRS20, VSS^{+}22, WLY^{+}23, WFW^{+}24, XLM^{+}23, USA21]. multi-action [SRÁMH^{+}20]. multi-agent [WLY^{+}23]. multi-armed [Gat21]. multi-criteria [KFM^{+}23, MSS24]. multi-decadal [IBB^{+}21, IBB^{+}22]. multi-dimensional [Ano20-50]. multi-domainal [Ano20-50]. multi-habitat [BOLC^{+}22]. multi-host [NWJ^{+}22]. multi-level [VSS^{+}22]. Multi-model [MTBB22, SWP^{+}23]. Multi-objective [NMS24, NAM22]. multi-parcel [OSPZ22]. multi-population [EL21]. multi-region [XLM^{+}23]. Multi-scale [TRMB20, CWH^{+}21, HLM^{+}20, PBS21]. multi-scaled [AHO^{+}20]. multi-scenario [MP23, MTZA23, WFW^{+}24]. Multi-sectoral [BCW^{+}23]. multi-sensor [JVGC21]. multi-species [BGCJ21, BM21, MPCG21, VRS20]. multi-state [RMA^{+}24]. multi-tier [OMC^{+}24]. multi-trophic [GLA^{+}20, HK23]. multi-year [HHBF24]. multiannual [PMP^{+}21]. Multidimensional [MFH^{+}22, SAL^{+}22]. multilayer [HNC^{+}21]. multilocularis [MMMS23]. Multimedia [WLQ^{+}23]. multimodal [FCP^{+}20]. multimodel [LLLZ21]. Multiple [CDC^{+}22, SGLK20, ABM20, AM21, AZBdAI23, CM21, DDL^{+}20, GSJ^{+}20, HCL^{+}22, JHL23, LPD^{+}20, MBG^{+}22, MFF^{+}24, MM21, WGMR23]. multiple-objective [WGMR23]. multipurpose [CXZ^{+}23]. multispecies [For20, XYZR21]. multistate [SDP23]. multitrait [CP20]. Multivariate [ZFC^{+}23]. municipal [CAS^{+}24]. mussel [FDSM20, RRCZ20, SSB22, Vas23]. mussels [BDP^{+}23, KWM^{+}24, SLB20b, WSZ^{+}24]. mutualism [AM21]. mutualisms [DAG^{+}21]. mutualistic [BDC^{+}22, LBL22, YL22]. my [LJRM^{+}20]. mycorrhizal [MGHT20]. mydas [SMV^{+}20]. mystery [GDC24].

N [AAR23, BGA22, GSR^{+}23, GSRB22, ZBH^{+}22]. N-fixing [GSRB22]. Nacella [GSM^{+}20]. Nærstad [HEL^{+}21]. Nanjing [LDZ21]. nanoflagellates [LBPL20]. Naples [SBF^{+}21]. Narragansett [HML^{+}22, LUV^{+}23]. Narran [LHJ^{+}24]. Nash [SACE^{+}20]. natal [vdMvDSL20]. National [ECC^{+}22, SOV21, BST^{+}20, PK23, PWN^{+}24, PPH^{+}21]. native [DP20, Ino22, KWM^{+}24, KKF^{+}23, MRR^{+}20, dOPdMdS^{+}21, SWL^{+}23, SC23]. Natural [Ano20-51, SMD^{+}20, AVP^{+}23, BWCD23, BARF20, BDA^{+}20, CDJ^{+}22, CPAT20, DS24, GPN^{+}23, KCS22, LRA^{+}20, LGG^{+}22, MSG^{+}21b, MPL^{+}20, RFCEL22, RFPRC23, TBCV22, WG21, ZPZ^{+}22]. naturalisation [RPARM20]. nature [Ano20-50, BWCD23, CXZ^{+}23, DYMD22, HPMB20, RLS^{+}24, WZLC24, HPMB20]. nature-based [WZLC24]. nearest [PSZ20]. nearest-neighbour [PSZ20]. nearshore [PRM^{+}24]. necked [KLPJ24]. need [AAM^{+}22, For24, MTS^{+}20]. neighborhood [QS23]. neighbour [BLP20, PSZ20]. neighbour-biased [BLP20]. Nemopilema [XFK21]. Neocloeon [KFJS20]. NEON [MMM^{+}24]. neonate [LHH^{+}24]. neonPlantEcology [MMM^{+}24]. Neophocaena [DZC^{+}23a, DZC^{+}23b]. Neotropical [KGBL24, PTBFP^{+}20]. Nepal [AAAR23, AAAR24]. nest [VJK^{+}23]. nested [WSL^{+}22]. nesting [CRZ^{+}23, GDC24]. net [ASHK22, ACFM22, DP20, RFdL^{+}24, SMD^{+}20, Yu20, ZCJ^{+}20, RSB^{+}23]. Netherlands [HSP^{+}22, PRdN^{+}20]. Network [JVGC21, LLZM20, Abe20, BLB^{+}20, CB20, CDNN^{+}20, DPA21, FDSM20, GCS^{+}21, GZK^{+}20, Jor22, LSL20, LMW^{+}23a, LLW^{+}21, LW23, MPC^{+}22, MS20, MP23, PSG^{+}22, PPSS22, dAPdBMT23, SRTS20, SZPW23, SZ20, SPV^{+}20, XHS^{+}24, ZWL^{+}21, KKH^{+}24]. networks [BDC^{+}22, BANG20, CCH20, DPA21, EOÖ21, FdKO21, Gon23, GML^{+}20, GCMU23, HMDO20, HNC^{+}21, KMAA23, LBL22, SNM^{+}23, SWW^{+}23, WLWL23]. Neural [JVGC21, BLB^{+}20, CCH20, MS20, PSG^{+}22, SNM^{+}23, SZ20]. neutral [EON23]. newly [BGA22]. Newman [RFHR24]. Next [EHS^{+}20, MKS21]. next-generation [MKS21]. nexus [DLZ^{+}20, NMV^{+}23]. NFI [PPH^{+}21]. Niche [Kos22, ARSC22, BMS^{+}20, BG22, CIL^{+}20, CYP^{+}20, Cor20, FSY22, Fou21, HLZ^{+}22, JS22, LLZM20, MYACC22, MMRBLN20, SBMF23, SACEU^{+}21, SCACB23, TSdOCV21, WCF^{+}22, ZFP21]. Nile [TGRM23, BGC^{+}23, KWG^{+}22a]. Niño [HSL^{+}22]. Niño/Southern [HSL^{+}22]. Nitocra [KD20]. nitrate [MKN23]. nitrogen [AKY^{+}22, BCW^{+}23, BH20, BGA22, CL24, DGVVP23, FSZ^{+}20, FZS^{+}20, GSR^{+}23, KKP^{+}21, LLLZ21, PGSL^{+}20, RLdM20, SMM^{+}22, Ste20, WZZF23, YG21, YZZF22, ZMF^{+}20, ZZWF21]. Nitrous [AAR23, ZMF^{+}20, ZPZ^{+}22]. Nitrous-oxide [AAR23]. no [DS23, Koo20b, CCH20]. no-take [DS23]. NoBa [SHF24]. nobilis [GRGHM21, HMGK20]. noctiluca [EWE20]. noise [JSD^{+}22, RCCLMGV22, SHF24]. nominal [PPSS22]. nomurai [XFK21]. non [AVP^{+}23, CWH^{+}21, Fer23, HI22, HSL^{+}22, KKF^{+}23, MOS21, PWN^{+}24, RSC^{+}24, vWVH20]. non-adaptive [AVP^{+}23]. non-attentional [vWVH20]. non-eruptive [Fer23]. non-independent [HI22]. non-invadable [MOS21]. non-native [KKF^{+}23]. non-representative [CWH^{+}21]. non-stationarity [HSL^{+}22, PWN^{+}24, RSC^{+}24]. Nonlinear [Vas20, ZXK^{+}24, KWC24, MLMI23]. nonparametric [RFPRC23]. nonstationary [CKK^{+}24b, HYY^{+}24]. Nordic [SHF24]. Norfolk [ZOJ^{+}23]. norm [MGWG22]. NORNE [HGBK23]. North [SGTS21, SGTS23, BBCD23, NT23, PMB^{+}22, PTNR^{+}20, TBZ^{+}22, ZBH^{+}22]. North-Western [BBCD23]. northeast [CBG^{+}22, LWL24, XLS20, DTD^{+}23, JOE23, MSG^{+}21b]. Northeastern [HLH^{+}21, WLT^{+}23]. northern [BGCJ21, SLLW21, STW^{+}23, ZBW^{+}24, CFG^{+}22, RSC^{+}21, RSC^{+}23]. Northwest [HCL^{+}24]. Norway [HEL^{+}21]. note [AAAR24]. Novel [LSG21, LGK^{+}23, SSM24, BANG20, BMBB24, CAS^{+}24, DDH^{+}22, GXZ^{+}24, ILLM20, II24, Jia21, Lef23, SBS^{+}21, SM22b, TWL^{+}22, WT20, ZHWZ23]. November [Ano20j, Ano20v, Ano21j, Ano21v, Ano22v, Ano23v]. NPP [XMWL23]. Number [CWF^{+}21, HLS^{+}23, WHE^{+}24a, WHE^{+}24b, WSM24]. numbers [CAB^{+}21]. Numerical [DZBB24, IA23, AB21, CFL^{+}22, SZL^{+}20]. Nunavik [DSHB23]. nungara [Fou21]. nurseries [LLB^{+}23]. nutrient [AWW24, CR24, CSM23, FWK^{+}20, GWD^{+}23, HCPG20, RLS^{+}24, TGRM23, Vas20, WCG^{+}24, YXH20, ZZŠ^{+}22, ZWZ^{+}20]. nutrient-deficient [YXH20]. nutrient-removal [HCPG20]. nutrients [BH21, RKN^{+}21, SSJC23]. nutritional [LBPL20]. Nutritionally [CAFT22].

O [AAR23, HKSM23, ZBH^{+}22]. Oak [BLB^{+}24a, EVT^{+}21, NYH20, PGSL^{+}20]. oasis [CWY^{+}24, CJM^{+}23]. Obama [Fou21]. OBEC [RFHR23]. objective [NMS24, NAM22, WGMR23, YA24]. objectives [DDL^{+}20]. observation [CFL^{+}22, SOV21]. observations [ACZ^{+}21, LDX^{+}23]. observed [BVR^{+}20]. obtained [AAM^{+}22]. occupancy [BBIP24, DL21, JJN22, SOV21, SDP23]. occurrence [BRSW24, RFCEL22, SSM24]. occurrences [BM22a]. Ocean [APGD^{+}21, RRCZ20, ACZ^{+}21, CBG^{+}22, CSB22, KiAM^{+}20, KPLH24, LHDBN20, MTL^{+}23, NPK^{+}23, LHH^{+}24, NT23, dSRDC^{+}23]. oceanic [CGRR^{+}24, KNP^{+}21, WSFZ21]. oceanography [LHDBN20]. oceans [ZGG22]. October [Ano20k, Ano20w, Ano21k, Ano21w, Ano22w, Ano23w]. ODD [Gri20]. ODE [RRSG21]. ODE-based [RRSG21]. Odocoileus [KS22]. odontocetes [RSC^{+}23]. off [BAAN24, BBIP24, CSCT23, CP20, KiAM^{+}20, NO24, Pér24]. offs [DSC^{+}21, JHL23, PF23]. offset [BLB^{+}24a]. offsetting [BSM23]. offshore [vOVvdW23]. oil [HMVR21]. oiled [WWA^{+}22]. Okhotsk [PA20]. old [BWCD23, KCP^{+}23, LDZ21, YA24]. older [JHKT^{+}23]. oleae [PVS^{+}21]. olgensis [XLS20]. oligotrophic [ZL24]. Oliv [ZWL^{+}21]. olive [PVS^{+}21]. omnivorous [JL23]. Oncorhynchus [BLLL21]. One [PAS^{+}22, DBK20, TCB21]. one-dimensional [DBK20]. only [GWH20, JJN22, KRM^{+}22, MBD^{+}22]. Ontario [DMdR24]. ontogenetic [mSzZK^{+}21]. open [LGA20]. operating [MM21]. operational [DLT^{+}22, MVC^{+}22, PM22]. Operationalising [DBDR22]. opinions [WBD^{+}22b]. opportunistic [EMH^{+}22, TCB21, VMH^{+}21]. opportunities [AGS^{+}21, AGS^{+}22, HL23, LSG21, MKEH20, WGMR23]. Optimal [BPB^{+}20, CC20, Fer23, BSU^{+}23, DMPLHGL20, ŁJ21, RBMRR20, VBH20, YDLT24]. Optimisation [TLH^{+}20]. Optimising [DDL^{+}20, NND^{+}22]. Optimization [MTP^{+}20, MTZA23, WCSX22, ZZQ23, CWW^{+}23, DGVVP23, FCP^{+}20, HBC^{+}22, HYY^{+}24, LWZL23, NMS24, RK20, WGMR23, WDW^{+}22, XHS^{+}24]. optimizing [ESZ^{+}23, NGE^{+}21]. option [KMH^{+}21]. Orbicella [LVM^{+}22]. orchard [LBT^{+}22]. order [CLHZ22]. ordinal [PPSS22]. ordinary [Wu24]. organic [AB21, CMSD^{+}22, JLV^{+}21, KBSS23, MAD^{+}20, MLP^{+}20, MPL^{+}20, RFHR23]. organisms [FLG^{+}23, Jia21, KKP^{+}21]. organization [YEL^{+}21]. organizational [Jor22]. organizing [EL21, LCP^{+}20]. oriental [LhNA^{+}21, LhNA^{+}22]. orientalis [LhNA^{+}21, LhNA^{+}22]. oriented [CM21, PTNR^{+}20, TPD^{+}24]. origin [DLZ^{+}20, SGRDA^{+}24]. ORYZA [GSR^{+}23, LDX^{+}23]. ORYZA-N [GSR^{+}23]. Oscillations [HSL^{+}22]. OSMOSE [XCZ^{+}20]. OSMOSE-JZB [XCZ^{+}20]. other [BAAN24, MJS24, RBGB23]. Oulujärvi [KKH^{+}24]. our [GCMU23]. Outbreak [HCW^{+}22, BJKT21, DMG22, Gat20, GSC^{+}23, PCL24, STRC24]. outbreaks [LDS^{+}23, MAB23, cNKF24]. outcome [vDC20]. outcomes [MKS21]. outdoor [BPF24, CCH20]. output [COJ23, HFSL^{+}24, HLM^{+}20, XLM^{+}23]. outputs [AVRC^{+}20, CHR^{+}23]. over-harvesting [PBH22]. overall [OMP^{+}23]. Overcoming [GSMD24]. overfitting [RSWS23]. overhead [LZSQ20]. overprediction [MVAD20]. Overview [HSD^{+}21]. ovigerous [HBMR24]. oviposition [dAWG23]. owls [CWH^{+}21]. oxidation [KKPR20]. oxide [AAR23, ZMF^{+}20, ZPZ^{+}22]. oxygen [FTH24, YFW^{+}21]. oyster [HLS^{+}23, SBWB24, YEL^{+}21]. oysters [LLC^{+}24]. ozone [EOÖ21, WHW^{+}23].

P [Mea20, Pas21, SDP21]. Pacific [HSL^{+}22, NT23, BWS^{+}20, dMFOM^{+}20, GGMR20, KiAM^{+}20]. package [BSS^{+}24, CC23, GSJC22, MMM^{+}24, MS23, OSH^{+}21, RPT21, SGPSzC22, Wu24]. packages [SCACB23]. paddy [GSR^{+}23, HCX^{+}21, SMM^{+}22]. page [AGS^{+}22]. PAHs [XWZ^{+}22]. paired [dPGDdFS24]. Pairwise [SPV^{+}21, MEPJ21]. Pakistan [SLY^{+}21]. palaeoclimate [PMB^{+}22]. pallida [KRMP23]. palm [DIS23]. Pan [VRRI^{+}24]. Pan-Atlantic [VRRI^{+}24]. panda [DRZ^{+}24, YDQ^{+}23, HYY^{+}24]. pandemic [EOÖ21]. pandemics [Laz23]. Panopea [LCMB21]. panther [CB20]. paradigm [AG22]. paradox [MYH^{+}20]. Parameter [COPM23, CWW^{+}23, EJ20, Mat23, MTT22, TDI^{+}20]. Parameter-varying [COPM23]. Parameterisation [RJY^{+}20, RRCZ20, JPMA^{+}23, MPL^{+}20]. parameterization [FCP^{+}20, KKPR20, NT23, PTG^{+}20, SNM^{+}23, FZS^{+}20]. parameterized [CDJ^{+}22, LAC^{+}24]. parameters [AS22, BGCJ21, CAK^{+}23, CAK^{+}24, HRW^{+}24, HCX^{+}21, JPL21, KAB23a, KD20, LAK20, LK24b, MSLMM22, MLP^{+}20, OMC^{+}24, PDH^{+}22, SSM24, SC21, VN23, XCZ^{+}20, YSIB24, dGZHCA24]. Parametric [AZAD^{+}23, RFPRC23, RP22, WBD^{+}22b]. parasite [EBMK20, GRGHM21, KE24, TA20]. parasite-produced [GRGHM21]. parasites [MMMS23]. parasitic [BPW^{+}21]. parasitism [Hop23, MMC^{+}21, Oku21, dFT^{+}20]. parasitoid [BMRF21, Oku21, SE21, YKZ^{+}23]. parasitoids [BRAL23, Hop23, MBR^{+}23]. parcel [OSPZ22]. park [BBCD23]. Part [MLP^{+}20, MPL^{+}20]. parthenogenetic [KFJS20]. partial [COPM23, FGM^{+}24, VMM^{+}21]. partially [SGV^{+}24]. particle [LWE^{+}22, RGB22, SJH^{+}23, WSW23, WSZ^{+}24]. partitioning [UGSG24, ZHWZ23]. party [PDW23]. passive [BVR^{+}20, PDD^{+}22]. past [BPL^{+}20, CLW23]. pastoral [TBZ^{+}22]. Patagonia [AM22]. patch [CDNN^{+}20, FdKO21, LSL20]. patches [AKM22, BACS22, CAFT22, MS23]. pathogen [AV22, MB22, MGC24, SWL^{+}23]. pathway [vdHNH^{+}20]. pathways [CDN^{+}23, GM23, MBB20]. pattern [ABZ22, CLHZ22, CGV22, FdKO21, FMRMAS24, HYY^{+}24, HLM^{+}20, SLC^{+}20, TPD^{+}24, WZLC24, XMWL23, ZHJQ23, ZBW^{+}24]. pattern-formation [FMRMAS24]. pattern-oriented [TPD^{+}24]. patterning [LCP^{+}20]. patterns [AAAR23, AAAR24, CLPZ20, CDB^{+}20, FHP^{+}22, GCSlR22, LDR^{+}20, LZF21, PKO^{+}21, RMV^{+}24, WZHP21, WWC^{+}22, ZPZ^{+}22, ZZW^{+}22, ZLZ^{+}24]. payment [BGA^{+}23]. PCA [LNK^{+}21]. PCA-SEM-ANN [LNK^{+}21]. peak [WF24]. Pearl [CL21, SWP^{+}23]. peatland [BST^{+}20, RPB^{+}23, RGD20]. Peatlands [BWT^{+}24, BST^{+}20, MLMI23]. peculiarities [SGRDA^{+}24]. pedestrianization [ST24]. pedology [CM21]. PEG [EL21]. Pelagia [EWE20]. pelagic [ESZ^{+}23, NT23, SLB^{+}20a, SINS22, SZL^{+}20, WLLW20]. pelagicus [RM20]. Pelotas [dMT23]. Penaeus [LGL^{+}24]. penguins [Koo20a]. per-capita [SGA22]. perceptive [GLH24]. percolation [HFP20, HFP22]. percolation-based [HFP22]. perennial [LGK^{+}23]. Performance [KTB^{+}21, LZL22, AHH22, BSCH23, CMHZ23, CKC^{+}24, ESZ^{+}23, KMP^{+}23, PTBFP^{+}20, RM20, RRFLV^{+}22, SC24a, SM22b, VMH^{+}21, VRS20, XCZ^{+}20]. peri [RBEC21, USA21]. peri-urban [RBEC21, USA21]. period [EOÖ21, LM23]. periodic [KK20]. Periphyton [HKH^{+}20]. permafrost [LCW^{+}21]. persistence [BACS22, FdKO21, LSA^{+}21, LJP^{+}22, dGMCM^{+}23]. Persistent [CCJK^{+}20, LKO20, MLP^{+}20, MPL^{+}20]. perspective [CBW22, CN20, Jor22, LSL21, MJdORPG23, SE20a, SE20b, SAdA^{+}22, WZLC24, ZZQ23]. perturbations [CTL22, WA20, dGZHCA24]. pest [BPW^{+}21, BJKT21, JBF22, LRA^{+}20, MJdORPG23, NGE^{+}21, NLMM23, PVS^{+}21, PEFW20, PSM^{+}21, RCSS20, SSM24, YKZ^{+}23, MTT22]. pesticide [BSPT23a, BSPT23b, BSU^{+}23, MJdORPG23, PEFW20, dVDC20]. pesticides [PDH^{+}22]. PESTIPOND [BSPT23a, BSPT23b]. pests [PSM^{+}21, WGK^{+}21]. petraea [EVT^{+}21, NYH20]. petrels [Koo20a]. PEWI [CVTS20a, CVTS20b]. PG [NYH20, SNM^{+}23, XLS20]. Phasianus [KLPJ24]. pheasant [KLPJ24]. phenModel [PGdAM^{+}20]. Phenological [KFJS20, CdCAC^{+}20, HCX^{+}21, PCL24]. phenology [GXZ^{+}24, LPHO^{+}20, NYH20, OFH^{+}20, PGdAM^{+}20, REGI^{+}24, VSS^{+}22]. phenology/voltinism [PGdAM^{+}20]. phenomenological [HFP20]. phenomenology [HFP22]. philippinarum [CMGD^{+}22]. philopatry [BR22]. phosphate [GSJ^{+}22]. Phosphorus [ADR21, AKY^{+}22, HSJ^{+}20, PRM^{+}24, RRP^{+}23, SLB20b, SGH^{+}23, WPL^{+}23, YG21, ZWZ^{+}20]. photography [KKW^{+}21]. photoperiod [SBR20]. photoperiod-driven [SBR20]. Photosynthesis [RSB^{+}23, Mas23, Yu20]. photosynthetic [HKSM23, SGH^{+}23, VN23]. Phragmites [YXY^{+}20]. PHREEQC [MTT22]. phthalic [WLQ^{+}23]. phylogeny [SPV^{+}21]. physical [AZMD^{+}22, FMH^{+}20, GGE^{+}23]. physical-biogeochemical [GGE^{+}23]. physics [GRS^{+}20]. physiognomies [DDDC21]. physiological [DPC20, LLC^{+}24, RRSG21]. Phytomass [CSM23]. Phytoplankton [PA20, SLC^{+}23, TSUY23, TMNA20, ACZ^{+}21, CDJ^{+}22, CMA23, HZ22, HMC20, MYH^{+}20, MS20, MBFS21, OSH^{+}21, RFCEL22, RKN^{+}21, SINS22, Vas20, ZGG22]. phytosanitary [BAR24]. Picea [BG22]. pigmentation [OSH^{+}21]. Pinctada [SLYL^{+}20]. pine [BdSG^{+}23, BAZ22, EVT^{+}21, JLH^{+}21, ML20, SLHB^{+}23, SWL^{+}23, SC23, SGRDA^{+}24, WLT^{+}23]. pine-hardwood [JLH^{+}21]. pink [JZJ^{+}20]. pink-footed [JZJ^{+}20]. Pinna [GRGHM21, HMGK20]. pinniped [WWP20]. Pinus [EVT^{+}21]. pioneering [Ano20-49]. pisha [ZLZ^{+}24]. Pityogenes [OFH^{+}20]. place [CNNS^{+}21]. placement [LBT^{+}22]. plague [RWS^{+}21]. Plains [ZFT20]. planarian [Fou21]. plankton [EL21, GWD^{+}23, MSML21, PCK^{+}22, RUTY23]. plankton-fish [RUTY23]. planktonic [MPPM23, MBS22]. Planning [MTP^{+}20, CL21, DLT^{+}22, MTZA23, NMV^{+}23, RBLS23, SRÁMH^{+}20, SACE^{+}20]. Plant [LM21, RMK21, TK20a, WXT^{+}20, dGMCM^{+}23, AADHA^{+}23, CGB^{+}23, CWF^{+}21, CSM23, DAG^{+}21, ERP21, HLW^{+}20, HHT21, Hop23, JBF22, KKP^{+}21, LSA^{+}21, MMM^{+}24, MGHT20, MGC24, MPCG21, MPA^{+}24, PSG^{+}22, PAPŁ^{+}23, PMP^{+}21, QZSV21, RRP^{+}23, SVK20, TGL^{+}22, VPV21, WLWL23, WMA21, WTT^{+}21, WCF^{+}22, YL22, YLS^{+}21, ZMS20]. plant-animal [WMA21]. Plant-host [dGMCM^{+}23]. plant-pest [JBF22]. plant-pollinator [YL22]. Plant-Water [WXT^{+}20]. plantation [HKSM23, LYM^{+}23, MSG^{+}21b]. plantations [BdSG^{+}23, SGRDA^{+}24, XLS20]. planted [WXH^{+}22]. planthopper [NDNN22]. planting [MRR^{+}20, TBZ^{+}22]. plants [KKF^{+}23, LC20, LZL22, MT23, Mas23]. Plastic [NO24]. plasticity [JRD^{+}22]. plastics [NO23]. Plateau [APGD^{+}21, SZL^{+}21, TGL^{+}22, YLLL23, WCG^{+}24, YJC^{+}24, ALS^{+}21, HCL^{+}24, LWL24, WXT^{+}20, ZPZ^{+}22, ZYM^{+}21]. platform [GDF22]. plausible [Koo20b]. play [AARS23]. plays [MSG21a]. Pleuroncodes [BAAN24]. plume [CMFC23]. pluralistic [MSG23]. PM [NSS^{+}23, SBF^{+}21]. poaching [NMC^{+}20]. point [GSH^{+}22, LJRM^{+}20, MYACC22, SBMG24]. points [WSM24]. Poland [ZA20]. polar [CLW23]. polder [YG21]. poleward [LK20]. policies [CL21, DWG22, Fer23]. pollination [CAFT22, CSCT23, JSL^{+}20, PTRP23, PDGB21, SJA^{+}22, SPV^{+}20, SPV^{+}21]. pollinator [WLWL23, YL22]. pollinator-plant-consumer [WLWL23]. pollinators [RMK21]. pollutants [MLP^{+}20, MPL^{+}20]. polluting [DGVVP23]. pollution [BH21, DBK20, KHS^{+}23, ZZZD22]. polyactis [RJY^{+}20]. polychaete [GLA^{+}20]. polycyclic [ZYZ22]. Polynesia [RMV^{+}24]. Pommerening [Mea20]. pond [YFW^{+}21]. ponds [BSPT23a, BSPT23b, BGR20, NBML21]. Pontederia [XGT^{+}24]. pool [CM21]. pooling [PSG^{+}22]. pools [SGP^{+}20]. poor [AMP^{+}23]. Pop [HBMR24]. Pop-up [HBMR24]. poplar [CSZY^{+}20]. Population [MMC^{+}21, TTPA21, VVH^{+}20, XFK21, YKZ^{+}23, AVP^{+}23, APGD^{+}21, BOLC^{+}22, BMM22, CMGD^{+}22, CCJK^{+}20, CLH23, CP20, CAB^{+}21, Dav22, DKM^{+}21, DRZ^{+}24, EL21, EVHS21, FCFCM24, Fer23, FC21, GCB^{+}22, GZ23, HKG^{+}24, Hop23, JLLS23, JZJ^{+}20, JASW23, KAMS22, KRM^{+}22, KLA^{+}20, Koo24a, KLNMG22, LP22, LHL22, LGA^{+}21, LKO20, LSG21, LGK^{+}23, LCMB21, MCB^{+}23, NS20, NGA^{+}20, NGE^{+}21, NLMM23, NBMD21, Oku21, OCT^{+}23, PDH^{+}22, PSM22, PSG21, RCCLMGV22, RUTY23, REGI^{+}24, RMA^{+}24, SE21, SE22, STW^{+}23, SDP23, SLC^{+}20, SBR20, TK20a, TRMB20, Uch23, VSK^{+}20, WZWF20, WGAH22, WLZ^{+}23a, WFD23, Yao21, dGMCM^{+}23]. populations [APF21, BGS^{+}20, CRZ^{+}23, COPM23, FSZ^{+}20, FZS^{+}20, GSJ^{+}20, KS22, KK20, KHH^{+}22, KHH^{+}24, KLPJ24, LSA^{+}21, LBI^{+}20, LGA^{+}21, MGC24, OSH^{+}21, PEFW20, RM20, RCSS20, SDD^{+}20, SLB^{+}23, SMS^{+}24, dVDC20]. porewater [ZRMW^{+}20]. porosity [IA23]. porpoise [DZC^{+}23a, DZC^{+}23b, HSF^{+}20, WSL^{+}22]. portfolio [SACE^{+}20]. portraying [RGB22]. position [LM23]. positions [Sak23]. positive [DMG22, WLWL23, ZWZ^{+}20]. possible [CSMP21, LLZM20, MSG21a, PR21]. post [DWH^{+}21, OHP^{+}23, vdMvDSL20, GGMR20]. post-ENSO [GGMR20]. post-exposure [DWH^{+}21]. post-fire [OHP^{+}23]. post-natal [vdMvDSL20]. postfire [LM23]. potato [DTR^{+}22, DGVVP23, HEL^{+}21, KBSS23, TBZ^{+}22, OHH^{+}21]. Potential [DTR^{+}22, HCW^{+}22, LhNA^{+}21, MRR^{+}20, AFdSM23, BGC^{+}23, BS20, BWCD23, CNSK21, CAS^{+}24, CN20, DCT^{+}21, DTD^{+}23, DL21, GLA^{+}20, LUV^{+}23, LLC^{+}24, LDXJ22, MEGA21, MTBB22, MGC24, MLV24, NA22, PK23, PDGB21, RMO22, SD24, SBFBDM24, YCK20, vOVvdW23, LhNA^{+}22]. potentially [PAPŁ^{+}23]. Power [KLS20, UGSG24]. Power-law [KLS20]. Poyang [HSF^{+}20, MFH^{+}22]. Practice [Ano20-51, BM22b, LSG21, ZJEG24]. practices [BAR24, BCMV20, TLC^{+}22, TCS21, YLLL23]. Prays [PVS^{+}21]. pre [LDC^{+}20]. pre-processing [LDC^{+}20]. Precipitation [WXT^{+}20, LM21, LWL24]. precision [KKF^{+}23, YHC^{+}21]. precocial [CC20]. Predation [dKS22, BACS22, BAAN24, CKC20, LK20, MAN^{+}20, SM22a, VVH^{+}20, vWVH20]. predator [dSCdA21, KKH^{+}24, MAN^{+}20, MF20, MMMS23, PHK23, SGP^{+}20, SQDM21, TCG^{+}20, dAdSCA20]. predator-prey [MMMS23]. predators [BSE21, MCMP^{+}24, PBH22]. predatory [TCG^{+}20]. predict [AE22, BdCCdSAdS23, CdCAC^{+}20, CYP^{+}20, Cor20, DVVH21, FEPR24, GCB^{+}22, JPL21, Jia21, KLNMG22, MAGR^{+}22, RFCEL22]. Predictability [BH20]. predictable [DEMJ24]. predicted [FS22]. Predicting [AADHA^{+}23, FHP^{+}22, HKB^{+}20, HSJ^{+}20, HMDO20, LLC^{+}24, LVM^{+}22, dSMPA^{+}22, NA22, REGI^{+}24, SDD^{+}20, SMM^{+}22, VVH^{+}20, WSN^{+}21, vdMvDSL20, FG21, KTB^{+}21, LWE^{+}22, MKS21, STQ^{+}22, Ste20, SPV^{+}21, VJK^{+}23, ZLL^{+}23]. prediction [CMSD^{+}22, GWH20, LJP^{+}22, LLLZ21, LDX^{+}23, LNK^{+}21, dRMJdS23, SZ20, SWP^{+}23, WrJP^{+}22, XZZ^{+}22]. predictions [ABC^{+}22, EVHS21, For20, HKY24, HRW^{+}24, JJN22]. predictive [CKC^{+}24, EJ20, JLH^{+}21, KMP^{+}23, RRFLV^{+}22, SC24a, TDI^{+}20, VRS20, YCI20]. predicts [WWA^{+}22]. predominantly [KHR23, PR23]. preference [DZC^{+}23a, DZC^{+}23b]. Preferences [BBCD23, BAM23, DDL^{+}20, PTNR^{+}20, VRS20]. preferential [AFG24]. PRELES [KWP^{+}23]. Preparing [MAB23, MMM^{+}24]. Presence [GWH20, CIL^{+}20, HBP^{+}21, JJN22, KRM^{+}22, WSM24]. presence-background [WSM24]. Presence-only [GWH20, JJN22, KRM^{+}22]. present [XGT^{+}24]. Preservation [DS21]. preserving [AM22]. pressure [DSC^{+}21, GJC^{+}20]. prevalence [GWH20, MSG23, MDT^{+}23]. prevent [GRCSA^{+}20]. prevention [BJKT21, GZK^{+}20]. prey [BSE21, CLH23, dSCdA21, MAN^{+}20, MF20, MMMS23, NPK^{+}23, PHK23, SGP^{+}20, SQDM21, SWL^{+}23, TCG^{+}20, dKS22, dAdSCA20]. prey-predator [SGP^{+}20]. price [CLY^{+}23, UGSG24]. primary [DP20, GAC^{+}21, GLK^{+}22, GZD^{+}23, HMC20, HCX^{+}21, KUK^{+}21, KWP^{+}23, MS20, MBFS21, RFdL^{+}24, SDP21, TSUY23, Yu20, ZCJ^{+}20, ZZLF24]. principle [HCL^{+}22]. principles [BPN^{+}20]. prior [BWS^{+}20, TTPA21, VRS20]. prioritization [HPMB20]. prioritizing [MS23, RBLS23]. Priority [PAS^{+}22]. probabilistic [DBH^{+}22]. probabilities [DMM^{+}23]. probability [LDS^{+}23, MDT^{+}23]. probability- [MDT^{+}23]. problem [LKO20, FSZ^{+}20]. problems [CCJK^{+}20, LCLG20, PAS^{+}22, TK20b]. procedure [EJ20, PEFW20]. procedures [FHP^{+}22]. Process [Por21, AZAD^{+}23, BdSG^{+}23, CAK^{+}24, GSH^{+}22, HSJ^{+}20, HEL^{+}21, HGBK23, HBRK24, HCPG20, KTB^{+}21, LJRM^{+}20, LCW^{+}21, MYACC22, MPCG21, NWO21, NO23, OKS^{+}21, PTNR^{+}20, TQM^{+}24, TSUY23, TPD^{+}24, VSS^{+}22, WGK^{+}21, ZPZ^{+}22]. process-based [BdSG^{+}23, CAK^{+}24, HEL^{+}21, HGBK23, HBRK24, HCPG20, KTB^{+}21, LCW^{+}21, MPCG21, NWO21, NO23, OKS^{+}21, TSUY23, VSS^{+}22, ZPZ^{+}22]. process-explicit [TPD^{+}24]. process-guided [HSJ^{+}20]. process-oriented [PTNR^{+}20]. Processes [YHC^{+}21, AJG22a, AJG22b, EL21, FFM^{+}23, FHP^{+}22, HKY24, LCW^{+}21, MTW^{+}22, MM21, NTHC23, QZSV21, SHT21, TRMB20, Tsy22, WTT^{+}21, ZZG^{+}24]. processing [LDC^{+}20]. produce [NAM22]. produced [GRGHM21]. producer [KWG^{+}22b]. producer-grazer [KWG^{+}22b]. produces [Sak23]. Production [LUV^{+}23, TLH^{+}20, ZZQ23, AAM^{+}22, AAC20, BLLL21, CMKF22, CDN^{+}23, CMHZ23, DTR^{+}22, DP20, ESCA22, EGAS23, GSR^{+}23, GLK^{+}22, HMC20, HCX^{+}21, KUK^{+}21, LYM^{+}23, MTS^{+}20, MS20, MBFS21, MEGA21, MGdM^{+}20, NGA^{+}24, NBMD21, OZSU22, PAdS^{+}24, SLY^{+}21, SMD^{+}20, SDP21, TBP^{+}23, TSUY23, TKO^{+}22, WGMR23, XLS20, ZBH^{+}22, vOVvdW23]. production-crop [LYM^{+}23]. Production-Living-Ecological [ZZQ23]. Productivity [LB21, SGRDA^{+}24, CGRR^{+}24, CSM23, GZD^{+}23, HCP^{+}20, KDH^{+}21, KWP^{+}23, MPA^{+}24, NYH20, QS23, WHW^{+}23, RFdL^{+}24, SE20a, SE20b, SHT21, TSUY23, WLT^{+}23, YSIB24, Yu20, ZCJ^{+}20, ZZLF24]. Productivity-environment [SGRDA^{+}24]. products [ACZ^{+}21, MLV24]. profiles [EMH^{+}22]. profiling [MGK^{+}23]. ProForM [SFGB23]. program [HSD^{+}21]. Programming [BM21, NMV^{+}23, NAM22, SRÁMH^{+}20]. programs [JZJ^{+}20]. progress [WZZF23, ZZWF21]. project [DBV^{+}22, XZZ^{+}22]. Projected [ZFT20]. Projecting [SLLW21, YPS^{+}21, APF21, MKB^{+}24]. projection [CP20, KK20, PSBE23, PFK^{+}22, TK20a, ZLLZ24, ZXK^{+}24]. Projections [KNP^{+}21, TTPA21]. projects [DYW^{+}22]. prolifera [SRB^{+}20]. promote [LGG^{+}22]. promotes [AM21, ZL24]. Promoting [LYM^{+}23, HL23, Yao21, BCMV20]. prone [ABRG21, KT22, SD24]. Proof [BRBT22]. Proof-of-concept [BRBT22]. propagation [KWG^{+}23]. propagules [LBB20]. properties [IA23, JRD^{+}22]. proposal [RFHR23]. Prorocentrum [SGH^{+}23]. prospect [DLZ^{+}20]. prospective [BH21]. protected [AADHA^{+}23, GMS^{+}23, PWN^{+}24, SvV21]. protection [HWS^{+}23, SFGB23, ZYK22]. protocol [Gri20, PAB^{+}22]. protozoa [JZS^{+}21]. provide [HBMR24]. Province [AAM^{+}22, KHS^{+}23, LZSQ20, MDB^{+}20, ENARM^{+}23, ZSS21, ZLML20]. provincial [PK23, SLY^{+}21]. provincial-scale [SLY^{+}21]. provision [MEGA21]. provisioning [PTRP23]. pseudoharengus [NGA^{+}20]. PSR [RS23]. pteropods [EV22]. public [CVTS20a, CVTS20b]. Publishing [Mea20, Pas21]. pullout [JZHT23]. pulses [MB22]. Puma [CB20]. PVA [LBGD21]. Python [LQH24]. Python-based [LQH24].

Q [MBMV20]. Qazvin [DYMD22]. Qinghai [ALS^{+}21, LWL24, Yu20, YJC^{+}24, ZPZ^{+}22]. Qinling [WYBX23]. Qinzhou [HLS^{+}23]. QtAC [SGPSzC22]. qualitative [FDSM20, HKB^{+}20, RBEC21, ZBW^{+}24]. qualitatively [CLPZ20]. quality [BZZT22, BM22b, BLLL21, BAM23, CC23, nCLpW^{+}23, ELS23, EMH^{+}22, HCP^{+}20, KLNMG22, LZF^{+}24, MSL^{+}21, MPA^{+}24, PRM^{+}24, Pie23, SBF^{+}21, SC21, VMH^{+}21]. Quantification [HLW^{+}20]. quantify [DYW^{+}22, Jia21, TBP^{+}23, WWF^{+}23]. Quantifying [AG22, HCPG20, KBSS23, LGA^{+}21, OMP^{+}23, RSWS23, SMP20, SGA22, TQM^{+}24, TKO^{+}22, PAS23, PRdN^{+}20, SRB^{+}20]. Quantitative [Ino22, vdHNH^{+}20, CAH22, For20, PSW22]. quantitatively [CLPZ20]. quantities [LCMB21]. quantization [MBS22]. Quanzhou [LZL23]. Quebec [MTBB22, RPB^{+}23]. Quercus [EVT^{+}21, NYH20]. questions [JTH24]. quick [Tur20]. quick-start [Tur20]. QWASI [WLQ^{+}23].

R [BSS^{+}24, CC23, GSJC22, MMM^{+}24, MS23, OSH^{+}21, SGPSzC22, SCACB23, Wu24]. R-package [SGPSzC22]. Rabies [TLL^{+}23]. radiation [DWH^{+}21, VSK^{+}20, YXH20, ZL24]. rafts [HLS^{+}23]. rainfall [dSBDH23, FS22]. rainfed [TKO^{+}22]. rainforest [AHO^{+}20]. ramp [NPS^{+}23]. Random [KAB23a, PDD^{+}22, SOV21, ABB23, AP22, BMB24, DMG22, DMM^{+}23, MLL20, CKC^{+}24]. randomly [TK20a]. range [FMRMAS24, KKW^{+}21, LK20, RBLS23]. range-wide [RBLS23]. rangeland [CPAT20, FS22, RS23]. rangelands [AJG22a, AJG22b]. ranging [DGSM21, TBM^{+}24]. ranked [PPSS22]. ranking [EJ20, ILLM20, PMZ^{+}21]. Raoiella [AFdSM23]. raphidophytes [SIOM24]. rapid [YZZF22, ZGG22]. raptor [STQ^{+}22]. rare [DTD^{+}23, TTPA21]. rate [CYP^{+}20, Cor20, DGVVP23, DP20, JLLS23, KN20, LK24a, PTF23, SE20a, SE20b, XJW21]. rates [APF21, CDT^{+}23, HFP22, II24, JST^{+}22, KABL21, RWF22]. ratio [MGWG22, WZWF20, ZFC^{+}23]. ration [DMPLHGL20]. rational [GSC^{+}23]. ratios [CDJ^{+}22]. ray [ZL24]. Re [SdlP20, AG22, KKH^{+}24]. re-analysis [AG22]. re-establishment [KKH^{+}24]. Re-scaling [SdlP20]. reaches [XMWL23]. reaction [FMRMAS24, MGWG22, RBEH21]. reactive [ZZWF21]. reading [RFHR24]. real [LJP^{+}22, PAS^{+}22, RGB22, SJA^{+}22]. real-time [LJP^{+}22]. realism [KHH^{+}22, KHH^{+}24]. realistic [PCK^{+}22]. Realities [Por21]. really [KHR23]. reanalysis [CBG^{+}22]. recapture [DMM^{+}23]. recharged [WLQ^{+}23]. Rechna [AIA^{+}23]. reclaimed [WLQ^{+}23]. recommendations [BM22b, EMH^{+}22, HSD^{+}21]. Reconciling [MGdM^{+}20]. Reconsiderations [GSRB22]. Reconsidering [UGSG24]. Reconstructing [SLB^{+}23, ZZW^{+}22]. reconversion [BGA^{+}23]. Recovery [HHI^{+}22, GCB^{+}22, GCS^{+}21, MHP^{+}23, SWL^{+}23]. recreation [BPF24, GCS^{+}21]. recreational [WBW^{+}20]. recruitment [AJ22, PYD^{+}21a, PYD^{+}21b]. recycling [AWW24]. Red [MSG21a, VSK^{+}20, BG22, MVC^{+}22]. reduce [CAB^{+}21, CDC^{+}22, Koo24a]. reduced [BSS^{+}24, VRW^{+}21]. reduced-impact [BSS^{+}24]. Reducing [TBZ^{+}22, DGVVP23]. reduction [Che21, CAK^{+}23, HWS^{+}23, MKR20, Pre22]. reed [YXY^{+}20]. reef [GGMR20, SBWB24, XHZH24, YEL^{+}21, ABR24]. reefs [CN20, DLY^{+}22, MAB23]. reference [CSM23]. refining [CB20]. reflect [TLC^{+}22]. Reflections [WG21]. reforestation [SAL^{+}22]. Reformulation [RCSS20]. regarding [RFHR23, VRS20]. regeneration [BLB^{+}24a, HBRK24, QZSV21]. regime [GWD^{+}23, LDXJ22, WCG^{+}24]. regimes [AADHA^{+}23, AACCIS20, LM21]. region [AZMD^{+}22, BS20, EGAS23, LZF^{+}24, LZF21, LW23, MEGA21, dRMJdS23, NO24, PZH^{+}21, SD24, SM22b, TLH^{+}20, WFW^{+}24, XLM^{+}23, Yu20, ZLLZ24, COS23, SLLW21]. regional [BLL^{+}23, ERP21, KMH^{+}21, KTB^{+}21, OZSU22, RMV^{+}24, RFCEL22, WYBX23, WKL^{+}21, ZA20]. regions [AMK21, BMB24, BCMV20, CDN^{+}23, GVG^{+}23, GSR^{+}23, GLK^{+}22, LZF^{+}24, LMW^{+}23a, MPSS24, NWO21, WFW^{+}24]. regression [JPL21, YCI20]. Regular [SLC^{+}20]. regulated [BS21]. regulates [SLC^{+}20]. regulating [GMS^{+}23]. regulation [BMB21, RGB22, SBF^{+}21]. regulations [PGB^{+}20]. reinforcement [SZL^{+}21, WLY^{+}23]. relascope [PSW22]. related [Kos22, LE23, MKB^{+}24, TKO^{+}22, WTT^{+}21, YSIB24, ZZŠ^{+}22]. relationship [CLPZ20, CLY^{+}23, LZL23, LK24b, MZLRSZ21, WLWL23, dMT23]. relationships [CKK^{+}24a, CKK^{+}24b, Vas20]. Relative [EL21, QZSV21, SDP23, WBD^{+}22b]. relay [DDH^{+}22]. release [HLB^{+}20, Lov24, ZWZ^{+}20]. relevance [Fer23]. relevant [GCSlR22]. reliability [dGZHCA24]. relictus [ILLM20]. reload [GRS^{+}20]. remnant [GMS^{+}23]. remote [AMP^{+}23, ACFM22, DIS23, HGBK23, HMDO20, ILLM20, KCHZ21, Lia23, SNP^{+}23]. removal [HKH^{+}20, HCPG20, NSS^{+}23, RGD^{+}22, SBF^{+}21]. renewable [HWS^{+}23]. renewal [LDZ21]. reoligotrophication [KUK^{+}21]. repeat [KKW^{+}21]. replace [AAM^{+}22]. replacement [SIOM24]. Report [PKB^{+}23]. reporting [BST^{+}20, DLT^{+}22]. represent [GSMD24]. representation [ASKR24, KBAC20, SDSP^{+}21, WGK^{+}21]. representative [CWH^{+}21]. Representing [BA20]. reproducible [LGA20]. reproduction [APF21, DWH^{+}21, KGBL24, KK20, MSG21a, MSG23, RTH^{+}21, SGLK20, SMV^{+}20, TBP^{+}23, WHE^{+}24a, WHE^{+}24b]. require [BM22a, Pér24]. requires [PCK^{+}22, YL22]. Research [WZZF23, ZZWF21, BZZT22, DK22, Fat21, LWZL23, MMM^{+}24, MS23, ZJEG24]. reserve [CTL22, GMS^{+}23, WST^{+}23]. reserves [DS23]. Reservoir [XWX^{+}24, BdCCdSAdS23, DAQ20, MKN23, ZMG^{+}21, LYZ^{+}24]. reservoirs [SC21, SZ20, Ste20]. residencies [BVR^{+}20]. resident [MGHT20]. residential [IA23]. residents [YZZF22]. residue [TGB^{+}23]. residues [DCT^{+}21, MEGA21]. Resilience [KMAA23, DBDR22, HSP^{+}22, JST^{+}22, JRD^{+}22, Mur24, RS23, RLM23, RGRMRR20, SLC^{+}23, SKPL24, SBM^{+}22]. resistance [BSU^{+}23, Hop23, WXH^{+}22, dVDC20]. resolution [AMP^{+}23, ACZ^{+}21, BRC^{+}21, BBIP24, CJM^{+}23, CDB^{+}20, ESZ^{+}23, LBGD21, RSB^{+}23]. resolutions [ATP20]. resolved [CGRR^{+}24, DMGP20]. resolves [HES^{+}20, HES^{+}23]. Resource [AWW24, LL22, YA24, BAAN24, CBSB23, GLH24, HLW^{+}20, KN20, Lia23, PF23, RK20, RLS^{+}24, TPD^{+}24, TK20b, WGAH22]. Resource-based [AWW24, WGAH22]. Resources [Ano20-51, BNAR^{+}21, CSM23, CDNN^{+}20, RK20, XHZH24, ZA20]. Respiration [RSB^{+}23, AG22, TDI^{+}20]. respond [RLdM20]. responding [LCP^{+}20]. Response [BPL^{+}20, NBU^{+}23, Por21, RFHR24, YCK20, BS21, BA20, BMBB24, FTH24, GSJ^{+}22, GGE^{+}23, HSL^{+}22, JOE23, LAK^{+}23, LE23, OHR^{+}23, PR22, RGB22, SQDM21, TRMB20, ZFT20, ZZŠ^{+}22, dAdSCA20]. responses [Che22, FWK^{+}20, KWC24, LJP^{+}22, OEC24, SRK^{+}20, WZLC24, WSN^{+}21]. resprout [ABRG21]. restoration [HYY^{+}24, LLC^{+}24, LVM^{+}22, ZRMW^{+}20]. restore [YA24]. restored [CWF^{+}21, WCF^{+}22]. restoring [OHP^{+}23, SDD^{+}20]. RESTREND [ZYM^{+}21]. RESTREND-based [ZYM^{+}21]. result [RFHR24, SWL^{+}23]. resulting [SAdA^{+}22, ZOJ^{+}23]. Results [MKR20, KCHZ21, ZFT20]. retention [Ste20, TGRM23]. Retraction [AAAR24]. retrospective [LM23]. reveal [CTL22, PR21, RSC^{+}24]. Revealed [BBCD23, EL21, TSUY23]. reveals [CP20, SPV^{+}20, SBR20, TLL^{+}23, WGMR23]. reverse [CLHZ22]. Review [FLG^{+}23, WYJ^{+}22, ABM20, BZZT22, DRZ^{+}24, KAB23a, Laz23, MMRBLN20, OKS^{+}21, dSRDC^{+}23, SDH21, TGL^{+}22, WGAH22]. reviews [Fat21]. revised [PTG^{+}20]. Revisiting [Jag20, MPL^{+}20, RRFLV^{+}22]. Revolutionising [ABKA24]. rheophilic [LWE^{+}22]. rheophilic-spawning [LWE^{+}22]. Rhode [LUV^{+}23]. Rican [AHO^{+}20]. Rice [PSG^{+}22, BMAME^{+}23, ENARM^{+}23, GSR^{+}23, HCX^{+}21, MVC^{+}22, NDNN22, SMM^{+}22, TKO^{+}22]. rich [CAFT22, SM22a]. richness [MAGR^{+}22, VPV21, YB22]. ricinus [WBD^{+}22b]. RIDL [NS20]. ridley [SPB^{+}24]. Riera [NBU^{+}23]. right [WHdBB21]. rights [nCLpW^{+}23]. Rim [YD20]. ring [KLPJ24]. ring-necked [KLPJ24]. rise [PR22]. risk [AAAR23, AAAR24, BRSW24, CB20, CDC^{+}22, For24, FLG^{+}23, GZK^{+}20, HP20, HEL^{+}21, JOE23, LBI^{+}20, LYZ^{+}24, LDS^{+}23, LLW^{+}21, Oku21, TLL^{+}23, TCB21, WLQ^{+}23]. River [CWY^{+}24, CL21, DZC^{+}23a, DZC^{+}23b, JASW23, MP23, SWP^{+}23, TGRM23, WLZ^{+}23a, ZGMFESD23, BS21, CFL^{+}22, KWM^{+}24, KSJ22, LDC^{+}20, RBLS23, SRTS20, SD24, XHS^{+}24, Yao21, BH21, nCLpW^{+}23, GJC^{+}20, JHL23, MBB20, TLH^{+}20, XZZ^{+}22, XMWL23]. riverine [CSBS21, JASW23, LBI^{+}20, MBG^{+}22]. rivers [KWM^{+}24, LWE^{+}22, LEC^{+}23, SJH^{+}23]. Road [LMW^{+}23a]. roadmap [RFHR24]. robur [EVT^{+}21]. Robust [DTD^{+}23]. robusta [KTB^{+}21]. Robustness [LSL20, LBL22, MP23]. rocky [SBS^{+}20, CWH^{+}21]. rocky-shore [SBS^{+}20]. rodent [KWC24]. rodents [Fer23]. role [BDC^{+}22, DvDA24, FSY22, JSL^{+}20, LTR^{+}24, MSG21a, PSM^{+}21, SRTS20, TA20]. roles [EL21, RSC^{+}21]. Romagna [MEGA21]. roofs [MKR20]. rookeries [LHH^{+}24]. root [AG22, HFP20, HFP22, JZHT23, JMC^{+}22, PMP^{+}21, SZL^{+}21, YLS^{+}21]. root-soil [HFP20, HFP22]. roots [YSIB24]. rotation [GAMN^{+}23, XLS20]. Rothamsted [JLV^{+}21]. roughness [HLS^{+}23]. route [KTB^{+}23]. routes [NPS^{+}23, PTD^{+}23, SGV^{+}24]. Rpath [LGA20, WA20]. RS [SD24]. rubens [BG22]. Ruditapes [CMGD^{+}22]. rule [RLH24]. ruminant [MDB^{+}20]. run [BLLL21]. runoff [NA22]. rural [CNSK21, HCPG20, JLLS23, MFPIBGES23, YZZF22]. rust [CGB^{+}23]. Rwanda [GVG^{+}23].

S [CDJ^{+}22, FSZ^{+}20, FZS^{+}20, KMP^{+}23, COS23]. Sable [HBP^{+}21]. Saccharina [VLGG^{+}20]. Sacramento [BLLL21]. sacrorum [SZL^{+}21]. safe [DP20]. safeguarding [XCH^{+}21]. sage [OHP^{+}23]. sage-grouse [OHP^{+}23]. saira [KiAM^{+}20]. sal [KWP^{+}23]. Salicornia [PAPŁ^{+}23]. salinity [LYC^{+}21, TSBE23, YSFX23, ZRMW^{+}20]. Salish [KNP^{+}21]. salmon [BLLL21, HP20, MBB20, WDW^{+}22]. salmonid [MM21]. salmonids [MM21, SJH^{+}23]. salt [ZZG^{+}24]. sample [GWH20, MDT^{+}23]. sampled [MGK^{+}23]. samples [BBIP24, JM22, RGD^{+}22]. Sampling [PSW22, AFG24, CWH^{+}21, MDT^{+}23, RSC^{+}24, RAC^{+}24, VPV21]. sanctioning [LL22]. sand [SWF21]. sandstone [ZLZ^{+}24]. Sanggou [FMH^{+}20]. sapidus [CMFC23]. saplings [BLB^{+}24a]. sardine [HMVR21, PYD^{+}21a, PYD^{+}21b]. Sardinella [HMVR21, PYD^{+}21a, PYD^{+}21b]. SARS [CYP^{+}20, Cor20]. SARS-CoV-2 [CYP^{+}20, Cor20]. satellite [CSZY^{+}20, HBMR24]. saturation [dSCdA21]. saury [KiAM^{+}20]. savanna [ABRG21, DDG^{+}21, DDDC21, MMdP^{+}21, dRMJdS23, YCK20]. savannahs [MTW^{+}22]. savannas [SKPL24, WMA21]. saved [SCH^{+}21]. Scale [BHC22, YB22, AJ22, BLL^{+}23, CBW22, CLPZ20, CSM23, CXZ^{+}23, CWH^{+}21, CYP^{+}20, Cor20, CKK^{+}24c, FWK^{+}20, FFM^{+}23, GMM^{+}24, HBP^{+}21, HAZEF23, HLM^{+}20, JSD^{+}22, KTB^{+}23, KKF^{+}23, Kos22, KTB^{+}21, MAGR^{+}22, MBG^{+}22, PWN^{+}24, PDGB21, PBS21, PMP^{+}21, PCK^{+}22, STQ^{+}22, SdlP20, SLY^{+}21, SKW23, TLL^{+}23, TRMB20, VPV21]. Scale-dependent [BHC22, YB22, PWN^{+}24]. scaled [AHO^{+}20]. scales [ATP20, BAR20, DKM^{+}21, KMH^{+}21, MKS21, MM21, OZSU22, PTRP23, RLM23]. scaling [SdlP20, XJW21]. scarce [GSRB22, MJ24]. scarcity [DYMD22, RSASP22]. scavenger [KHR23, PR23]. Scavengers [PR21]. scenario [DSC^{+}21, MP23, MTZA23, WFW^{+}24]. scenario-based [DSC^{+}21]. scenarios [BOLC^{+}22, JHL23, MTBB22, MSS24, SBM^{+}22, SMV^{+}20]. scene [TWL^{+}22]. scheme [VPV21]. schemes [BJKT21, OEC24]. Schichaa [TCG^{+}20]. school [dKS22]. schooling [WLY^{+}23, dKS22]. science [DK22, EMH^{+}22, JMM^{+}20, KRM^{+}22, LJRM^{+}20, VMH^{+}21, ZJEG24]. scope [OHP^{+}23, YYS^{+}22]. Scotland [CKK^{+}24a]. Scots [ML20, SGRDA^{+}24]. Scottish [CKC^{+}24, CKK^{+}24c]. screening [PMZ^{+}21, dGZHCA24]. scrofa [KT22]. SDM [LJRM^{+}20, MGK^{+}23]. SDMs [MDT^{+}23]. Sea [RSC^{+}21, RSC^{+}23, SRB^{+}20, AJ22, APGD^{+}21, CRZ^{+}23, EVHS21, GMM^{+}24, GGMR20, LHH^{+}24, II24, LHDBN20, LBB20, MM21, PTD^{+}23, PR22, RM20, SPB^{+}24, WBW^{+}20, WBD^{+}22a, BNN20, EWE20, GSJ^{+}20, KNP^{+}21, NHGP23, OCT^{+}23, PA20, PTNR^{+}20, SGTS21, SGTS23, SRK^{+}20]. seabass [DHW^{+}20]. seabird [ARSC22]. seabream [DTD^{+}23]. seagrass [LTR^{+}24, PR22]. seagrasses [VN23]. Seamlessly [LZW20]. Searching [BWSS21, CBSB23]. Seas [SHF24]. SEAS5 [NLMM23]. Seaside [HBP^{+}21]. season [DZC^{+}23a, DZC^{+}23b, HGBK23, LZD^{+}24]. Seasonal [NLMM23, ZWZ^{+}20, BMB21, DMdR24, HMVR21, HKSM23, KABL21, SSM24, SIOM24, STW^{+}23, WHL^{+}23, XYZR21, ZS20]. Seasonality [MPPM23, NWJ^{+}22, MMC^{+}21]. seasons [CRZ^{+}23, GDC24]. seaweed [FMH^{+}20, LUV^{+}23, WT20, vOVvdW23]. seaweeds [vOVvdW23]. secondary [CGV22]. section [LDC^{+}20, XHS^{+}24]. sectional [RAC^{+}24]. Sector [XLM^{+}23]. sectoral [BCW^{+}23]. security [GHbS21, HYY^{+}24, LPA^{+}21, LYL^{+}24, WZLC24, ZHJQ23, ZHC23]. sediment [HKH^{+}20, LCLG20, NA22, SBMF23, TGRM23, WPL^{+}23, WSZ^{+}24, ZWZ^{+}20]. sediment-inhabiting [SBMF23]. sediment-water [WPL^{+}23]. sedimentation [WSZ^{+}24]. Seed [LKO20, CGV22, WSW23]. SEEMS [CXZ^{+}23]. segregation [Kos22]. Segura [ZGMFESD23]. SEIB [VK24]. SEIB-DGVM [VK24]. SEIR [BGC^{+}23]. SEIR-Metapopulation [BGC^{+}23]. select [VPV21]. Selecting [BMM22]. Selection [BS20, ŁJ21, OEC24, PWN^{+}24, SC24a, SM20, SBMG24, WWF^{+}23, WT20, dVDC20]. Selective [ZGG22, BSS^{+}24, BACS22, BdBG^{+}21a, BdBG^{+}21b, NK23]. Self [LHL22, AVP^{+}23, EL21, GSH^{+}22, KBAC20, LCP^{+}20, PF23, YEL^{+}21]. self-compatibility [AVP^{+}23]. self-exciting [GSH^{+}22]. self-fouling [PF23]. Self-healing [LHL22]. self-organization [YEL^{+}21]. self-organizing [EL21, LCP^{+}20]. self-shading [KBAC20]. SEM [BPF24, LNK^{+}21]. semantic [PTRP23]. semelparous [TK20a]. Semi [AZMD^{+}22, BGA22, FS22, KWP^{+}23, LHJ^{+}24, MPSS24, NWO21, PAdS^{+}24, WFW^{+}24]. semi-arid [FS22, LHJ^{+}24, MPSS24, NWO21, PAdS^{+}24, WFW^{+}24]. semi-mechanistic [BGA22]. semiarid [PGSL^{+}20]. sensing [AMP^{+}23, ACFM22, DIS23, HGBK23, ILLM20, LDR^{+}20, Lia23, SNP^{+}23]. sensitive [GWH20, PMB^{+}22, SBMG24]. Sensitivity [CGVC21, FEPR24, HLH^{+}21, MLP^{+}20, OFH^{+}20, PMZ^{+}21, RGB^{+}23, SC21, WSHZ24, WXT^{+}20, WLT^{+}23, YSIB24, vDC20, BGCJ21, BLSR^{+}20, DMdR24, HRW^{+}24, HEL^{+}21, Mat23, PFM^{+}21, RS23, TGB^{+}23, TK20b, WA20, XHS^{+}24, dGZHCA24]. sensor [JVGC21]. sensors [HKH^{+}23]. Sentinel [CSZY^{+}20]. Sentinel-2 [CSZY^{+}20]. Separating [GSH^{+}22]. September [Ano20l, Ano20x, Ano21l, Ano21x, Ano22x, Ano23x]. Sequential [SDH21, WST^{+}23]. sequestration [MTBB22, RMO22, WZLC24]. series [DMG22, LCP^{+}20]. service [CVTS20a, CVTS20b, GZW^{+}20, JSL^{+}20, MEGA21, MTZA23, PTRP23, PDGB21]. services [ABM20, BGA^{+}23, BDA^{+}20, CAFT22, CL21, CLY^{+}23, GMS^{+}23, GSA24, JHL23, LL21, LZL23, LPD^{+}20, NMS24, PGP^{+}22, PRdN^{+}20, SJA^{+}22, WFW^{+}24, XCH^{+}21, YPS^{+}21, ZHC23]. sessile [NYH20]. set [LBT^{+}22]. sets [BS20]. settlement [SC24b]. several [GLK^{+}22]. severity [AB21, PCL24]. sewage [LZL22]. sex [MGWG22, WZWF20]. sexual [MSG21a, MSG23]. sexualization [MGWG22]. Shaanxi [LZSQ20]. shad [PSBE23]. shading [GZW^{+}20, KBAC20]. shallow [WPL^{+}23, WCG^{+}24, YXY^{+}20]. Shandong [LGL^{+}24]. shape [CIL^{+}20]. shape-constrained [CIL^{+}20]. shaped [LPL22a, dKS22]. shared [SWL^{+}23]. sharing [ZL24]. shark [KCHZ21]. shawi [ZMG^{+}21]. sheep [DGSM21]. shelf [SPB^{+}24, PTG^{+}20]. shelled [EV22]. shellfish [GMM20, LUV^{+}23, PGB^{+}20]. shift [LhNA^{+}21, LhNA^{+}22, LDXJ22, WCG^{+}24, dGMCM^{+}23]. shifting [HBRK24]. shifts [LVM^{+}22, NHGP23]. shikokuense [SGH^{+}23]. shocks [JST^{+}22]. shore [SBS^{+}20]. shoreline [MSL^{+}21]. short [BP21, EVHS21, JvVvHH24, LCP^{+}20, LGK^{+}23, SZ20]. short-lived [LGK^{+}23]. short-term [EVHS21, JvVvHH24]. short-time [LCP^{+}20]. should [RAB21]. shows [GGMR20, SG20]. shrimp [dMFOM^{+}20, LGL^{+}24, RWF22]. shrub [ALS^{+}21]. shrub-meadow [ALS^{+}21]. Siberia [GHK23]. Sichuan [HCL^{+}24]. sickness [Gat20]. SIF [YYS^{+}22]. sightings [WWP20]. sigmoidal [PCN^{+}23]. sign [WSFZ21]. signal [GSH^{+}22]. significantly [CAFT22]. silico [PNGESS21, dFT^{+}20, FCFCM24, PSM^{+}21]. silver [SKW23]. similar [ML20]. simple [JMHH22, MM21, TDI^{+}20]. simplification [AIA^{+}23]. simplified [Jag20]. Simplifying [CGVC21]. SIMREG [PPH^{+}21]. simulate [BSS^{+}24, DDH^{+}22, DBN^{+}22, EHS^{+}20, KLH^{+}24, LPHO^{+}20, MPCG21, PPH^{+}21, TGB^{+}23, TPD^{+}24]. Simulated [SMV^{+}20, AP22, KKH^{+}24, LVM^{+}22, LCW^{+}21, WHL^{+}23, WSHZ24, YYS^{+}22]. Simulating [ADR21, BGC^{+}24, CGRR^{+}24, CWW^{+}23, EVT^{+}21, FTH24, KDH^{+}21, KS22, LMW^{+}23b, PBS21, SSJC23, SJH^{+}23, WGK^{+}21, WCG^{+}24, YLS^{+}21, AMK21, CXZ^{+}23, EON23, LPS^{+}21, SCG^{+}23, TGL^{+}22, ZZG^{+}24]. Simulation [ASL^{+}20, JLLS23, LZW20, SNP^{+}23, ZBH^{+}22, ZZLF24, AXMJ22, BLB^{+}24a, BSR21, CM21, CLG20, DPA21, DPC20, FZS^{+}20, GDF22, JHKT^{+}23, LKP^{+}20, LP22, LZF^{+}24, LDX^{+}23, LWL24, MP23, MTZA23, MTT22, NGA^{+}20, OHH^{+}21, OHP^{+}23, SFGB23, SG20, TBCV22, TSBE23, Tur20, VK24, WWC^{+}22, WXX^{+}23, WMA21, WFW^{+}24, ZLLZ24, ZL24, ZHWZ23, ZLZ^{+}24, dPGDdFS24]. simulation-based [DPA21].

Simulations [LPL22b, PMB^{+}22, DMGP20, GGE^{+}23, KHH^{+}22, KHH^{+}24, PR21, RSB^{+}23, RRCZ20, SMS^{+}24, WA20]. simulator [CXZ^{+}23, KWG^{+}22a]. simultaneous [MTT22, RMK21]. simultaneously [GSJ^{+}20]. sinensis [ZWL^{+}21]. single [GAM20, RCCLMGV22]. sink [BWT^{+}24]. sinusoidal [LPL22a]. sinusoidal-shaped [LPL22a]. Sio [COS23]. Sio-Malaba-Malakisi [COS23]. SIR [Laz23]. Sistan [EGAS23]. site [EVHS21, FFM^{+}23, NND^{+}22, STW^{+}23, WT20]. site-level [STW^{+}23]. site-scale [FFM^{+}23]. sites [PAPŁ^{+}23, WLZ^{+}23a, WWP20]. six [CDN^{+}23]. size [BGCJ21, DMPLHGL20, HES^{+}20, HES^{+}23, KAMS22, LP22, MBS22, NK23, PAS^{+}22, PSZ20, RLH24, RGD^{+}22, RMA^{+}24, WZHP21, XYZR21, ZZŠ^{+}22]. size-selective [NK23]. size-spectrum [HES^{+}20, HES^{+}23]. size-structured [MBS22, XYZR21]. sizes [HH23]. sketches [LLZM20]. skill [HFSL^{+}24]. Skudai [BH21]. slope [ETAAH21, SZL^{+}21]. slopes [ZLZ^{+}24]. slow [LM21]. small [APF21, CSM23, CXZ^{+}23, CSMP21, GMM^{+}24, RSC^{+}24, RJY^{+}20, SLB^{+}20a]. small-scale [CSM23, CXZ^{+}23, GMM^{+}24]. smallholder [ASKR24]. smallmouth [LBI^{+}20]. smart [CMSD^{+}22]. smarter [BCW^{+}23]. smolts [WDW^{+}22]. snail [YKZ^{+}23]. Snow [PWN^{+}24, OCT^{+}23]. Sobol [LAC^{+}24]. Sociability [BSR21]. Social [MFPIBGES23, ASKR24, AAC20, BGS^{+}20, BAM23, GCS^{+}21, KS22, LCM24, PAS23, PHL20, RTH^{+}21, WAHA^{+}24]. Social-ecological [MFPIBGES23, BAM23, PHL20]. Socio [CXZ^{+}23, DS21, LPD^{+}20, OBFH22, BAR24, GM23, KLH^{+}24, LCM24, PG23, RBEC21, WG21, ZJEG24, ZGMFESD23]. Socio-Ecological [DS21, LPD^{+}20, OBFH22, KLH^{+}24, LCM24, RBEC21, WG21, ZGMFESD23]. Socio-econ-ecosystem [CXZ^{+}23]. socio-ecosystem [BAR24]. socio-environmental [PG23, WG21, ZJEG24]. socio-MFA [GM23]. soda [PAPŁ^{+}23]. sodicity [TSBE23]. soft [vdHNH^{+}20]. soft-bottom [vdHNH^{+}20]. Soil [TK20b, AMK21, AB21, AKY^{+}22, BMAME^{+}23, BMS^{+}20, CM21, DCT^{+}21, DPA21, ETAAH21, HKH^{+}23, HGBK23, HFP20, HFP22, JMPB24, KKP^{+}21, LLLZ21, LCW^{+}21, LSS^{+}21, LLW^{+}21, LMW^{+}23b, MSG^{+}21b, MTT22, PAPŁ^{+}23, PSM^{+}21, RGD20, TLC^{+}22, TDI^{+}20, TGL^{+}22, YG21, YLLL23, ZRMW^{+}20, ZLZ^{+}24]. soil-dwelling [PSM^{+}21]. soil-plant [PAPŁ^{+}23, TGL^{+}22]. soils [SMM^{+}22, XWZ^{+}22]. Solander [MTL^{+}23]. solar [YYS^{+}22]. solar-induced [YYS^{+}22]. sole [MLP^{+}20, MPL^{+}20]. Solea [MLP^{+}20, MPL^{+}20]. solid [CAS^{+}24, IA23]. solution [JVGC21, JASW23, Mas23, NND^{+}22, SACE^{+}20]. solutions [MTT22, WZLC24]. solving [MYH^{+}20]. somatic [KABL21]. sophisticated [KHH^{+}22, KHH^{+}24]. source [ELS23, GJC^{+}20, LGA20, ZZZD22]. sources [DBK20, HWS^{+}23, PSM^{+}21]. souring [MKN23]. South [MDB^{+}20, ARSC22, CRZ^{+}23]. southeastern [SLHB^{+}23]. Southern [APGD^{+}21, HSL^{+}22, LCW^{+}21, PTNR^{+}20, BARF20, GSA24, SGTS21, SGTS23]. Southwest [LZL22, ZHJQ23]. southwestern [YA24, dSRDC^{+}23]. soybean [dSBDH23, Hop23, RFdL^{+}24]. soybean-maize [dSBDH23]. sp [KDC^{+}21]. Space [RWS^{+}21, TLH^{+}20, ABB23, DDG^{+}21, EVHS21, HHFB23, KMH^{+}21, MM21, ZS20]. spaces [GZW^{+}20]. Spain [ZGMFESD23]. Sparrow [HBP^{+}21]. sparse [AMK21]. Spatial [BLP20, CCH20, LGG^{+}22, LW23, RS23, Sak23, SACE^{+}20, WHdBB21, WZHP21, AAAR23, AAAR24, ATP20, BRC^{+}21, BRSW24, BBIP24, BSM23, CLY^{+}23, CJM^{+}23, CSM23, CLG20, DIS23, DMM^{+}23, EON23, ESZ^{+}23, FTH24, GSA24, GWH20, GZK^{+}20, HSL^{+}22, LGFT^{+}20, LDR^{+}20, LZL23, LQH24, MKR20, MLJ^{+}23, MSLM^{+}24, MMRBLN20, NHGP23, NL20, OCT^{+}23, PEFW20, PSZ20, RAB21, RRFLV^{+}22, RWS^{+}21, SRTS20, SC24b, TWL^{+}22, ZZQ23, dMT23, dGMCM^{+}23, dAdSCA20]. spatial-explicit [LQH24]. Spatial-moment [BLP20]. spatial-temporal [EON23]. Spatially [ABZ22, GCB^{+}22, SOV21, AE22, BSU^{+}23, BSR21, BR22, CAB^{+}21, DMM^{+}23, DDG^{+}21, GRCSA^{+}20, JMM^{+}20, JSL^{+}20, KS22, KWG^{+}23, PRdN^{+}20, SLHB^{+}23, SJA^{+}22, WBW^{+}20, WBD^{+}22a]. Spatially-explicit [GCB^{+}22, DDG^{+}21, KWG^{+}23, WBD^{+}22a]. spatio [GSJ^{+}20, LRA^{+}20, NWO21, NO23, ZHWZ23]. spatio-temporal [GSJ^{+}20, LRA^{+}20, NWO21, NO23, ZHWZ23]. Spatiotemporal [CL21, LYL^{+}24, ZSS21, DMGP20, SBS^{+}21, WWC^{+}22, YD20, ZPZ^{+}22]. spawning [DSHB23, LWE^{+}22, WLZ^{+}23a]. special [CSM23, LSG21, Men22]. Speciation [FAC22]. specie [ILLM20, ZLL^{+}23]. Species [BM22b, BM22a, CLPZ20, EMH^{+}22, HHI^{+}22, HPMB20, dSRDC^{+}23, AHH22, AMP^{+}23, AM21, AFdSM23, ALJ^{+}22, AARS23, BOLC^{+}22, BAAN24, BSCH23, BGCJ21, BS20, BWSS21, BBIP24, BM21, CSBS21, CWF^{+}21, CNSK21, CIL^{+}20, CPTD22, dSCdA21, CKC^{+}24, CKK^{+}24b, DTD^{+}23, EVT^{+}21, FSY22, For20, FG21, GGS^{+}22, GWH20, HLZ^{+}22, HBRK24, HFSL^{+}24, HLH^{+}21, HNC^{+}21, Ino22, JLH^{+}21, JL23, JMM^{+}20, JHKT^{+}23, KWC24, KABL21, LYZ^{+}24, LDXJ22, MSG23, MGK^{+}23, MRR^{+}20, MAGR^{+}22, MMRBLN20, MVAD20, MPCG21, MBD^{+}22, NL20, NTHC23, NMGFOB21, OMP^{+}23, OW20, Pre22, RSWS23, RCCLMGV22, RBGB23, RSC^{+}24, RPT21, RAB21, SVK20, SvV21, SACEU^{+}21, SKW23, SWL^{+}23, SC23, SBMG24, TK20a, TBP^{+}23, TRMB20, TCB21, VMH^{+}21, VPV21, VRS20, WZHP21, WSM24, WT20, WTT^{+}21, XCH^{+}21, YXH20, YB22, YCI20, ZWL^{+}21, dAdSCA20]. species-area [VPV21]. species-predictive [JLH^{+}21]. Species-specific [HPMB20]. specific [HPMB20, KWM^{+}24, OMC^{+}24, TK20a]. spectral [CSZY^{+}20]. spectrum [BGCJ21, HES^{+}20, HES^{+}23, LM21]. speeds [Gat21]. spillover [MB22]. spinipes [KD20]. Spodoptera [BSU^{+}23, dAWG23]. spot [MZLRSZ21]. spp [GGS^{+}22]. Spread [BAPW20, AVP^{+}23, BGC^{+}23, CAH22, EVHS21, JB22, LDXJ22, MSLM^{+}24, WHE^{+}24a, WHE^{+}24b]. spreading [ABZ22]. spring [Che21, LLZM20, Tsy22]. spring-fed [Che21]. Springer [Mea20]. spruce [BG22, SCG^{+}23, STRC24, cNKF24]. squid [DvDA24]. squirrels [CAB^{+}21]. St [MSL^{+}21]. St. [KWM^{+}24]. Stability [NDNN22, BACS22, CP22, ETAAH21, GWH20, RLM23, SE21, SZL^{+}21, WLWL23]. stabilize [Oku21]. stabilizing [YA24]. Stable [YL22, JvVvHH24]. stage [HKG^{+}24, PSM22, QZSV21, dVDC20]. stage-based [HKG^{+}24]. stage-classified [dVDC20]. stage-structured [PSM22, QZSV21]. stages [DIS23, JMHH22, Jia21, MFF^{+}24]. stakeholder [DDL^{+}20, NBMD21]. stamp [AXMJ22]. stand [BLL^{+}23, BLB^{+}20]. Standard [Mat23, JGG23, Koo20b, PAB^{+}22]. standardization [For24]. stands [CSZY^{+}20, CWW^{+}23, EVT^{+}21]. stanza [DPC20]. starfish [MAB23]. start [Tur20]. starting [RFHR23]. starvation [WWA^{+}22]. State [DGSM21, EVHS21, MRR^{+}20, BdCCdSAdS23, FWK^{+}20, HHFB23, JZS^{+}21, NBU^{+}23, OHP^{+}23, RFHR23, RFHR24, RMA^{+}24, ZS20]. State- [DGSM21]. state-and-transition [OHP^{+}23]. State-space [EVHS21, ZS20]. Stated [BBCD23]. States [CBG^{+}22, JvVvHH24, KNP^{+}21, YA24]. stationarity [HSL^{+}22, PWN^{+}24, RSC^{+}24]. statistical [CSBS21, RSWS23, VPV21]. statistics [CDNN^{+}20]. status [AFG24, BS21, PGB^{+}20, RSC^{+}23, WLZ^{+}23a, Yao21]. steady [AXMJ22]. Steller [RM20]. step [EON23, OCT^{+}23, SACEU^{+}21]. step-by-step [EON23, SACEU^{+}21]. sterile [HLB^{+}20, DBN^{+}22]. Sterna [ARSC22]. stinger [EWE20]. stink [KGBL24]. Stochastic [STQ^{+}22, SdlP20, BVM^{+}21, FCFCM24, GWD^{+}23, KAB23a, KWG^{+}22b, LCMB21, NTHC23, PDH^{+}22]. stochasticity [vDC20]. stock [AJ22, DIS23, YRK^{+}20]. stock-recruitment [AJ22]. stocking [DP20, JST^{+}22]. stocks [FEPR24, MSG^{+}21b, YA24]. Stocktake [ECC^{+}22]. stoichiometric [AKY^{+}22, CDJ^{+}22, KWG^{+}22b]. stomatal [Mas23]. storage [PZH^{+}21, RFF^{+}23, TLC^{+}22, WYBX23, XLS20]. storms [SMS^{+}24]. stormwater [NBML21]. story [PSW22]. Strait [MPPM23]. strategic [GHbS21, RFHR24]. strategies [BSU^{+}23, BSE21, CTL22, CC20, DDL^{+}20, HCPG20, LBPL20, MPA^{+}24, MD23, PSM^{+}21, RSC^{+}24, SBMG24, TLL^{+}23, VBH20, WZWF20, WLQ^{+}23, WWCT24]. strategy [AXMJ22, BPB^{+}20]. streak [SYL^{+}24]. stream [BdBG^{+}21a, BdBG^{+}21b, HMDO20]. Strebel [GSM^{+}20]. street [GZW^{+}20]. Strengbach [BPL^{+}20]. strength [BWT^{+}24]. strengths [CDNN^{+}20]. stress [GSJ^{+}22, GRAA22, KFM^{+}23, LCP^{+}20, LBI^{+}20, RFdL^{+}24, RGB22, VVH^{+}20]. stressed [MBG^{+}22]. Stressor [SGCFB22]. stressors [MAB23]. strong [RBGB23]. structural [BPF24, DAQ20, MFZ23, SBS^{+}20, SBS^{+}21, YZZF22, YEL^{+}21, dFT^{+}20]. Structure [LB21, AE22, AM22, AHO^{+}20, BMM22, BSR21, CSB22, DvDA24, JSL^{+}20, KTB^{+}23, KHS^{+}23, LRA^{+}20, LGG^{+}22, LSL20, LW23, MPMT23, MLJ^{+}23, MPPM23, MMC^{+}21, NL20, PDH^{+}22, RRSG21, RWS^{+}21, SZPW23, VSB22, WSFZ21, XFK21, YA24]. structure-preserving [AM22]. structured [FGM^{+}24, KK20, MBS22, PSM22, QZSV21, WZWF20, WFD23, XYZR21]. structures [KSJ22]. studies [CdCAC^{+}20, ŁKPK22]. Study [DBK20, PSG21, SZ20, YJC^{+}24, AFdSM23, AIA^{+}23, BCW^{+}23, BH21, BGK^{+}20, BAR24, BG22, BARF20, Che21, CPTD22, ETAAH21, ENARM^{+}23, GAC^{+}21, Gon23, GSJ^{+}20, HHT21, HCL^{+}22, JLLS23, LPSF24, LP22, LZSQ20, LDZ21, LZF^{+}24, LZL23, LDX^{+}23, LPA^{+}21, LCMB21, MMS^{+}24, MS20, MTT22, MVC^{+}22, OHH^{+}21, OBFH22, OCT^{+}23, PHK23, PAPŁ^{+}23, PMP^{+}21, PTNR^{+}20, RBGB23, RKN^{+}21, RAB21, SLC^{+}23, SHT21, SG20, SPPC21, SWP^{+}23, SZL^{+}20, TCG^{+}20, TLH^{+}20, TMNA20, WBD^{+}22b, XWZ^{+}22, XHS^{+}24, XMWL23, YKZ^{+}23, ZMK^{+}22, ZHJQ23, ZZQ23, ZLML20, ZLZ^{+}24]. studying [DRZ^{+}24]. subcatchment [SRTS20]. subject [AWW24]. submarine [Che21]. submerged [NKT^{+}20]. submodules [PTF23]. Subnational [HLM^{+}20]. subpinnata [MTL^{+}23]. subsidence [RGD20]. subsidized [STQ^{+}22]. substitutable [RK20]. substitution [QS23]. subtidal [GSM^{+}20]. subtropical [BGC^{+}24, CLW^{+}24, PZH^{+}21, RGD20, SC21]. success [KN20, MJS24, OHP^{+}23, SG20, SKH^{+}20, WZWF20]. Succession [YXH20, HSF^{+}20, LM23, SLLW21, ZGG22]. successional [FFM^{+}23]. Sugar [VLGG^{+}20, AAM^{+}22]. sugarcane [BMRF21, MBR^{+}23]. suggestions [LCLG20]. suggests [MJ24]. Suitability [YHC^{+}21, BOLC^{+}22, BHC22, BBSHC22, HK23, ILLM20, JLH^{+}21, KKF^{+}23, LZD^{+}24, ZWL^{+}21]. Suitable [HYY^{+}24, YXY^{+}20, BG22, USA21]. sulfide [CWW^{+}23]. sum [LBN^{+}22]. Summer [KFM^{+}23]. Summit [PKB^{+}23]. sums [DMG22]. sun [SPPC21]. sun-induced [SPPC21]. Sungo [SZL^{+}20]. sunlight [ZOJ^{+}23]. superiority [GSRB22]. supersaturated [CFL^{+}22]. supply [GSA24]. Support [LZW20, BRC^{+}21, EMH^{+}22, For24, Gri20, MVC^{+}22, NGE^{+}21, PGP^{+}22, WBW^{+}20, ZWZ^{+}20]. supporting [GSC^{+}23]. supports [RSWS23]. suppression [SWL^{+}23]. surface [ABC^{+}22, BZZT22, DCT^{+}21, HSJ^{+}20, LCLG20, LCW^{+}21, PFM^{+}21, SBM^{+}22, TGB^{+}23]. surface-based [PFM^{+}21]. surfaces [OW20]. Surveillance [HCW^{+}22, BGK^{+}20, BRSW24, SLB^{+}23]. survey [KCHZ21]. surveys [PM22]. survival [AWW24, APF21, ALK20, dSCdA21, DHW^{+}20, II24, PSBE23, RMA^{+}24, SdlP20, WRC21, WBD^{+}22b]. survivorship [RAC^{+}24]. Sus [KT22]. Sustainability [SLY^{+}21, ZMK^{+}22, ABKA24, AAC20, DYW^{+}22, ESCA22, HL23, LL21, LGL^{+}24, MDB^{+}20, MSS24, NGA^{+}24, PHL20, RGRMRR20, RBEC21, TWL^{+}22, WAHA^{+}24]. Sustainable [BSE21, LPA^{+}21, ŁJ21, Ano20-49, AKM22, dSCdA21, EON23, GCMU23, HAZEF23, HWS^{+}23, MTZA23, NMV^{+}23, PGB^{+}20, RLM23, SvV21, SSB22, WrJP^{+}22, XHZH24, HL23]. Suzhou [BCW^{+}23]. suzukii [RRSG21]. swarming [LAK^{+}23]. swarms [SGV^{+}24]. Sweden [BLL^{+}23]. swimming [LHDBN20]. swine [MD23]. swirskii [TCG^{+}20]. switching [GWD^{+}23]. Switzerland [Mea20]. sylvatic [EBMK20]. sylvatica [SKW23]. sylvestris [EVT^{+}21]. symbionts [KKP^{+}21, MGC24]. symbiosis [KRMP23, RTH^{+}21]. sympatric [Kos22, NMGFOB21]. symptom [Gat20]. synchronized [KK20]. synchrony [GZ23]. syncronization [ABZ22]. synergistic [MMdP^{+}21]. synthesis [TLC^{+}22]. System [EL22, LB21, YDQ^{+}23, ABKA24, AM21, BGA^{+}23, CMGD^{+}22, CBSB23, ENARM^{+}23, GLA^{+}20, GM23, HSP^{+}22, HCL^{+}22, HSD^{+}21, HWS^{+}23, HK23, JZHT23, JST^{+}22, KMH^{+}21, KKP^{+}21, KCS22, KLH^{+}24, KWG^{+}22b, Kle23, LZW20, LYM^{+}23, MPSS24, MDB^{+}20, MF20, MMMS23, OSPZ22, PHK23, PAPŁ^{+}23, SQDM21, SM22a, ST24, TSUY23, TK20b, ZLLZ24, CFG^{+}22, DBDR22]. Systematic [DK22, ECC^{+}22, dSRDC^{+}23]. Systems [DS21, VLGG^{+}20, ADR21, AGS^{+}21, AGS^{+}22, CXZ^{+}23, DS24, FS22, GMM^{+}24, HSP^{+}22, JPMA^{+}23, JMPB24, KLH^{+}24, LHDBN20, LZSQ20, LCM24, MAD^{+}20, MAN^{+}20, MDB^{+}20, MTW^{+}22, MFPIBGES23, MSG^{+}21b, MBMV20, NGA^{+}24, NWJ^{+}22, OBFH22, PHL20, PMP^{+}21, SC24a, SGPSzC22, SLY^{+}21, SC24b, TQM^{+}24, Tur20, WG21, Wu24, ZJEG24, ZZWF21, ZBH^{+}22].

T. [EBMK20]. tackle [DSC^{+}21]. tadpoles [CMS^{+}21, PDG^{+}20]. tagging [PDD^{+}22]. tags [HBMR24]. Taihang [ZZLF24]. tail [ŁJ21]. Tailed [HCW^{+}22, BGK^{+}20, KS22, TBM^{+}24, VRW^{+}21]. take [DS23]. Tana [TGRM23]. tangent [PSZ20]. Tanzania [MDP21]. targeted [GSJ^{+}20, PEFW20]. targeting [MVC^{+}22, TCB21]. tasks [AGS^{+}21, AGS^{+}22]. taxa [CGRR^{+}24]. taxes [SML^{+}20]. taxonomic [RBGB23]. Taylor [XJW21]. Teaching [Por21]. technique [CCH20, DBN^{+}22]. techniques [ASHK22, BAR20, GM23, HLB^{+}20, SD24, SWP^{+}23, USA21]. technogenic [PAPŁ^{+}23]. technologies [COJ23]. teleconnected [LL21]. telemetry [BVR^{+}20, LDC^{+}20]. temperate [BGCJ21, BWT^{+}24, CAK^{+}23, KABL21, TDI^{+}20, VK24]. Temperature [MBMV20, AJJ24, CRZ^{+}23, DMdR24, GSJ^{+}22, HML^{+}22, JPL21, MGWG22, NGE^{+}21, PGdAM^{+}20, RLH24, SSM24, SBR20, SMS^{+}24, TA20, VJK^{+}23, WBD^{+}22b, YSFX23]. temperature- [SBR20]. temperature-dependent [HML^{+}22, MGWG22, NGE^{+}21, PGdAM^{+}20]. temperature-size [RLH24]. temperatures [ABC^{+}22, FWK^{+}20]. Temporal [CPTD22, AAAR23, AAAR24, BSR21, DKM^{+}21, EON23, FS22, GSJ^{+}20, HMC20, LRA^{+}20, LCP^{+}20, LSL21, NWO21, NO23, NK23, PAS23, RAC^{+}24, SBR20, WYBX23, ZHWZ23]. temporally [BR22]. tempos [Ano20-49]. tendency [LJP^{+}22]. tendency-prediction [LJP^{+}22]. Tenuipalpidae [AFdSM23]. term [ABH^{+}20, BLB^{+}20, DYMD22, EVHS21, JvVvHH24, JLV^{+}21, JSD^{+}22, PKO^{+}21, RSASP22, SZ20, YSIB24]. termite [LPL22a]. Tern [ARSC22]. terrestrial [ASHK22, LC20, OTR^{+}21, OKS^{+}21, QS23, WHL^{+}23, WWP20, WLZ^{+}23b, ZS20]. territorial [BAR24, ŁJ21, MCMP^{+}24]. territoriality [TPD^{+}24]. territories [CDC^{+}22, KRM^{+}22, MOS21]. territory [SM20]. Test [HFP22, RFJ^{+}22, SBMG24]. tested [FG21]. Testing [JMHH22, MDT^{+}23, SM22b, CdCAC^{+}20]. Tetragastris [dOPdMdS^{+}21]. Texas [YCK20]. Thanks [Des21]. Thau [PGB^{+}20]. thawing [LCW^{+}21]. their [BCMV20, CZ21, DZBB24, FSZ^{+}20, FZS^{+}20, FHP^{+}22, GMM^{+}24, HHFB23, Jia21, LZL22, OKS^{+}21, RRFLV^{+}22, SINS22, YD20, ZFC^{+}23]. theorem [Men22]. Theoretical [MAN^{+}20, BMS^{+}20, GRS^{+}20, KAB^{+}23b, LAC^{+}24]. Theory [Ano20-51, VLGG^{+}20, CIL^{+}20, DYW^{+}22, GLA^{+}20, HFP20, LPL22b, LSG21, MTS^{+}20, MBMV20, PSW22, RLH24, SGLK20, SACE^{+}20, XJW21, XHS^{+}24, ZSS20]. Thermal [Che22, MGWG22, Che21, Tsy22, VJK^{+}23]. Thermodynamic [CWF^{+}21, WCF^{+}22]. thermodynamics [MPMT23]. thermoregulation [HBMR24]. thinning [EVT^{+}21]. Thirteen [LGK^{+}23]. thorns [MAB23]. threat [JASW23, MAB23, SRÁMH^{+}20]. threatened [HP20]. threats [GHbS21]. Three [NBML21, ABB23, LWE^{+}22, LPHO^{+}20, NSS^{+}23, PCN^{+}23, PCK^{+}22, TWL^{+}22, WA20, YLS^{+}21, XWX^{+}24]. Three-dimensional [NBML21, ABB23, LWE^{+}22, TWL^{+}22, YLS^{+}21]. threshold [HH23]. thresholds [HKH^{+}20, TA20]. thrips [CMSD^{+}22]. Thunnus [DPC20]. Tianjin [LZF21, LW23]. Tianshan [ZLLZ24]. Tibet [ALS^{+}21, LWL24, WLZ^{+}23a, Yu20, YJC^{+}24]. Tibetan [ZPZ^{+}22]. tick [AV22, EBS23]. tick-borne [AV22]. ticks [WBD^{+}22b]. tidal [BdBG^{+}21a, BdBG^{+}21b, KSJ22]. tides [SRB^{+}20]. tier [OMC^{+}24]. Tiezzi [Ano20-49]. tiger [AAAR23, AAAR24]. tilapia [DMPLHGL20]. tillage [GAMN^{+}23, YLLL23]. timber [XLS20]. time [BAAN24, DMG22, HMC20, Jia21, LCP^{+}20, LJP^{+}22, LHL22, MLV24, MM21, RLM23, SE22]. time-averaged [HMC20]. time-dynamic [BAAN24]. time-scales [RLM23]. timescale [AAR23]. timescales [MSML21]. Timing [BVM^{+}21, CGB^{+}23, FS22, LAK^{+}23, LBGD21]. Titicaca [DVVH21]. tool [GZK^{+}20, MGK^{+}23, MPC^{+}22, MS23, MM21, NMC^{+}20, OEC24]. tools [Yao21]. Tooth [MSG21a]. top [CDN^{+}23, KKH^{+}24]. Topological [LDR^{+}20]. Topology [MEPJ21]. topsoil [VPAV22]. total [PRM^{+}24]. tourism [GCS^{+}21, LL21]. tower [ASHK22]. tox [AVRC^{+}20]. toxic [PAPŁ^{+}23]. toxicants [MMS^{+}24]. toxicity [BMS^{+}20, CBSB23]. toxicodynamic [JGG23]. toxicokinetic [JGG23, MLP^{+}20]. toxicokinetic-toxicodynamic [JGG23]. toxin [CMKF22]. toxoid [LPSF24]. tracking [LWE^{+}22, SJH^{+}23, WSW23, ZWZ^{+}20]. Trade [JHL23, PF23, BBIP24, CSCT23, CP20, DSC^{+}21, LMW^{+}23a, NO24, Pér24, SML^{+}20, ZLML20]. trade-off [BBIP24, CSCT23, CP20, NO24, Pér24]. Trade-offs [JHL23, PF23, DSC^{+}21]. tradeoff [RTH^{+}21]. tradeoffs [LGA^{+}21]. traditional [NGA^{+}24]. trained [HFSL^{+}24]. training [MS20]. trait [BDC^{+}22, HHT21, Mur24, RFJ^{+}22, ZMS20]. trait- [ZMS20]. trait-based [BDC^{+}22, Mur24, RFJ^{+}22]. traits [AVP^{+}23, CMGD^{+}22, CLW^{+}24, EHS^{+}20, GRCSA^{+}20, HKSM23, PKO^{+}21, SLYL^{+}20, SHT21, SPV^{+}21, Vas20]. trajectories [AADHA^{+}23, BAR24, COS23, FG21]. trajectory [SDSP^{+}21]. trans [KMP^{+}23]. trans-CLUE-S [KMP^{+}23]. transboundary [COS23]. transcriptomic [GML^{+}20]. Transdisciplinary [Por21]. Transeius [TCG^{+}20]. Transfer [XZZ^{+}22, vdHNH^{+}20]. Transferability [GVG^{+}23, RFJ^{+}22, CPTD22]. transfers [KKP^{+}21, SML^{+}20]. transformed [GCMU23, MPMT23]. Transient [CTL22, ZZZD22]. transit [MLV24]. Transition [MT23, OHP^{+}23, TTPA21]. Transitioning [Ano20-51]. transitions [HCX^{+}21, KMP^{+}23]. transmission [CLHZ22, EBMK20, KE24, KS22, LGFT^{+}20, MSLM^{+}24, MD23, RWS^{+}21, TLL^{+}23, TBM^{+}24, ZXK^{+}24, vWVH20]. transmitted [MMMS23]. TransparC2U [CM21]. transpiration [HFP22, ZFC^{+}23]. transport [BdBG^{+}21a, BdBG^{+}21b, CMFC23, LWE^{+}22, SPB^{+}24, WPL^{+}23, MLJ^{+}23]. transportation [LPL22a]. trap [AP22]. traps [HI22]. travel [Lef23]. treatment [BSU^{+}23, LZL22, MKN23]. treatments [GAMN^{+}23]. Tree [HHI^{+}22, dSMPA^{+}22, ABRG21, CAK^{+}23, CAK^{+}24, CNSK21, DMG22, HHBF24, HBRK24, HFP22, JMC^{+}22, LGG^{+}22, LPS^{+}21, MRR^{+}20, MLJ^{+}23, PVS^{+}21, PPH^{+}21, PKO^{+}21, Pre22, SLHB^{+}23, SKW23, SCH^{+}21, SBR20, VJK^{+}23, WHE^{+}24a, WHE^{+}24b, WHL^{+}23, WYBX23, ZOJ^{+}23, ZLL^{+}23]. tree-based [WYBX23]. tree-level [PPH^{+}21]. tree-species-shifting [HBRK24]. treehopper [PTBFP^{+}20]. treeline [BS20]. trees [AMP^{+}23, FG21, GHK23, GSMD24, IA23, LDR^{+}20, LMW^{+}23b, YCI20]. trend [AFG24]. trends [CZ21, dSRDC^{+}23, RAC^{+}24, SDP23, ZPZ^{+}22]. trials [NS20]. triangulifer [KFJS20]. Triatoma [FC21]. Tribolium [dVDC20]. tributaries [MSL^{+}21]. Trichodesmium [ABR24]. triggered [CMKF22, CLH23, RTH^{+}21]. tripartite [WLWL23]. triplet [ABKA24]. TRIPLEX [LPS^{+}21]. TRIPLEX-Mortality [LPS^{+}21]. Triticum [GGS^{+}22]. tritrophic [BMBB24, PVS^{+}21]. Trojan [BPB^{+}20]. Trophic [KKH^{+}24, OHR^{+}23, BdCCdSAdS23, BNN20, FWK^{+}20, GLA^{+}20, HK23, LPL22b, MZLRSZ21, PCK^{+}22, RSC^{+}21, SBS^{+}20, WSFZ21, vdHNH^{+}20]. tropical [BSS^{+}24, BH21, DBV^{+}22, FG21, HLB^{+}20, MAGR^{+}22, MHP^{+}23, PYD^{+}21a, PYD^{+}21b, PKO^{+}21, SKPL24, SPPC21, TSUY23, WGMR23, XJW21]. tropics [AMK21, RFJ^{+}22]. truthing [ABC^{+}22]. tsetse [GVG^{+}23]. tshawytscha [BLLL21]. Tumen [JHL23, TLH^{+}20]. tuna [DPC20, NPK^{+}23, PDD^{+}22]. tuning [Fou21]. tunnel [LPL22a]. turbidity [SZ20]. turbine [PM22, STQ^{+}22]. turbine-scale [STQ^{+}22]. turbulence [WSW23]. turbulent [LWE^{+}22, TMNA20]. turning [Ano20-49]. turnoff [BACS22]. turnover [AG22]. turtle [CRZ^{+}23, KABL21, LHDBN20, MZLRSZ21]. turtles [GDC24, LHH^{+}24, II24, PTD^{+}23, SPB^{+}24, SMV^{+}20]. Tuscany [LPA^{+}21]. TVDI [RFdL^{+}24]. TVDI-based [RFdL^{+}24]. two [AKM22, CM21, CWW^{+}23, DKM^{+}21, DPC20, KABL21, Kos22, LHDBN20, LYZ^{+}24, MBR^{+}23, NMGFOB21, PWN^{+}24, TCG^{+}20, WG21, WTT^{+}21, ZHC23]. two-pool [CM21]. two-stanza [DPC20]. type [AM21, AACCIS20, ETAAH21, Fat21, LLW^{+}21, SQDM21]. types [BRSW24, Sak23, SINS22, YYS^{+}22, dAdSCA20]. typhoon [PZH^{+}21]. Typical [WXT^{+}20, MFZ23].

U.S [LUV^{+}23]. Ubatuba [OBFH22]. Ulloa [MZLRSZ21]. Ulva [LFN^{+}20, SRB^{+}20]. uncertain [VRS20, YSIB24]. uncertainties [CM21]. Uncertainty [HAZEF23, AAR23, KKF^{+}23, LDX^{+}23, LKO20, MLP^{+}20, MKB^{+}24, NMS24, SACE^{+}20, TKO^{+}22]. underestimate [MBD^{+}22]. Undergraduate [Por21]. Underlying [ZXK^{+}24, CLH23, CL24, GSH^{+}22]. underpinning [EL21]. understand [CDNN^{+}20, DAQ20, MFPIBGES23, MPMT23, ZMS20]. Understanding [CFG^{+}22, HSP^{+}22, Por21, PCK^{+}22, ZOJ^{+}23, PTG^{+}20, PGSL^{+}20, SBWB24]. Understory [HHI^{+}22, MLJ^{+}23]. underwater [KCHZ21, SHF24]. uneven [JS22]. Unfolding [AS22, LMW^{+}23a]. ungulate [HBRK24]. unified [FMRMAS24, LMNV20]. unimodal [CLHZ22]. United [Pas21, CBG^{+}22, YA24]. units [ŁJ21, SRTS20]. universal [GAM20]. Unlocking [CAS^{+}24]. unmarked [HI22]. unpredictability [cNKF24]. Unraveling [BMBB24]. unravels [vdHNH^{+}20]. unsampled [MGK^{+}23]. update [Gri20]. updraft [STQ^{+}22]. updraft-subsidized [STQ^{+}22]. upgradation [LZL22]. upland [MKEH20]. upon [FAC22]. Upper [TGRM23, DYW^{+}22, SD24, XMWL23]. upregulated [MJ24]. upscaling [VPV21]. upstream [KTB^{+}23]. uptake [CWW^{+}23, GSJ^{+}22, JMC^{+}22]. upwelling [AACCIS20, GGE^{+}23]. Upwind [SGV^{+}24]. urban [AXMJ22, BAPW20, BCW^{+}23, BBCD23, CLY^{+}23, DBDR22, EL22, EON23, GZW^{+}20, JHL23, LL21, MFZ23, MPSS24, MKEH20, MTZA23, MMMS23, OBFH22, RBEC21, SBF^{+}21, SCH^{+}21, SSJC23, ST24, USA21, WZZF23, XWZ^{+}22, YZZF22, ZYK22, ZHWZ23, dMT23]. urbanization [BH21, QS23, YZZF22]. urbanizing [SZPW23]. urchin [APGD^{+}21, GGMR20]. Uruguayan [DP20]. USA [AAR23, HH23, ZFT20, ZRMW^{+}20]. usages [AZBdAI23]. Use [ACFM22, BM22b, GML^{+}20, LDX^{+}23, MTP^{+}20, dRMJdS23, TLH^{+}20, ACZ^{+}21, CP22, DSC^{+}21, DWG22, GCS^{+}21, Gri20, GZD^{+}23, HAZEF23, JLLS23, JHL23, KMH^{+}21, LBGD21, Lia23, LAK20, LZW20, LNK^{+}21, MMM^{+}24, MTZA23, NMV^{+}23, PHL20, PEFW20, PTBFP^{+}20, RSASP22, RAB21, SWP^{+}23, TGRM23, WMA21, WFW^{+}24, YPS^{+}21, Yu20, ZLLZ24]. use/cover [COS23]. use/land [WWC^{+}22]. user [CM21]. uses [TGL^{+}22]. Using [BPF24, BAAN24, BWT^{+}24, CSMP21, DAQ20, FDSM20, GLA^{+}20, HCL^{+}22, HLS^{+}23, JGG23, JZJ^{+}20, KAMS22, KFM^{+}23, LPSF24, MSG21a, MSG23, MCMP^{+}24, MPMT23, MBD^{+}22, NGE^{+}21, OHP^{+}23, TTPA21, AJG22a, ABKA24, ASHK22, AJG22b, AMP^{+}23, AHO^{+}20, BSS^{+}24, BSR^{+}23, BA20, BGA22, BRSW24, BNN20, BM21, BLB^{+}20, BGC^{+}24, CCH20, CMGD^{+}22, CGRR^{+}24, CSZY^{+}20, CHR^{+}23, COS23, Che21, CWW^{+}23, CMG20, CIL^{+}20, CSB22, CAB^{+}21, CKC^{+}24, DTR^{+}22, DDH^{+}22, DMM^{+}23, DAB^{+}20, EWE20, GM23, GCSlR22, GCS^{+}21, GRCSA^{+}20, GWH20, HBP^{+}21, HKB^{+}20, HSJ^{+}20, HBC^{+}22, HGBK23, HMDO20, HBRK24, HLM^{+}20, ILLM20, Ino22, II24, IBB^{+}21, IBB^{+}22, JB22, JLV^{+}21, JPL21, JLLS23, KiAM^{+}20, KRM^{+}22, KS22, KWP^{+}23, LE23, LUV^{+}23, LLB^{+}23, Lia23, LZL22, LGA20, LAC^{+}24, MLL20, MMS^{+}24, MPSS24, MLJ^{+}23, MSS24, NWO21, NGA^{+}24, NHGP23, NA22]. using [NTHC23, NPS^{+}23, OHH^{+}21, OSH^{+}21, PMZ^{+}21, PSG^{+}22, PRM^{+}24, PBS21, PM22, RSB^{+}23, RBEH21, RCSS20, SSM24, SRTS20, SD24, SMM^{+}22, SOV21, SIOM24, SLM24, SACE^{+}20, SNM^{+}23, STW^{+}23, SHF24, SWP^{+}23, SJH^{+}23, SC23, SM22b, TK20a, TMFV22, USA21, VLGG^{+}20, VMM^{+}21, WLQ^{+}23, WSW23, WA20, WSL^{+}22, XLS20, YKZ^{+}23, Yu20, YCI20, YDLT24, ZOJ^{+}23, ZSS20, ZFT20, ZSS21, ZPZ^{+}22, ZLLZ24, dGZHCA24, dKS22]. Utilization [GAMN^{+}23, DDG^{+}21, HLW^{+}20, Vas20]. Utilizing [KKW^{+}21, PSBE23].

UV [DWH^{+}21]. UV-B [DWH^{+}21].

v3 [LDX^{+}23, BGA22]. VA [ZOJ^{+}23]. vaccine [LPSF24]. Validation [dPGDdFS24, CDC^{+}22, DDG^{+}21, FZS^{+}20, FHP^{+}22, dRMJdS23, OFH^{+}20, PFM^{+}21, PAdS^{+}24, WHdBB21]. Valley [BLLL21, WWF^{+}23]. Valuation [SAdA^{+}22]. Value [DS21, LB21, BARF20, CL21, GPN^{+}23, LC20, Men22, WFW^{+}24, YZF20]. valued [RUTY23]. values [CLY^{+}23]. vannamei [dMFOM^{+}20]. Variability [VN23, CFG^{+}22, FTH24, FS22, HHT21, HMC20, KD20, MSML21, MLP^{+}20, MPL^{+}20, NK23, PYD^{+}21a, PYD^{+}21b, PCK^{+}22, SLYL^{+}20, WBD^{+}22a]. Variable [CKK^{+}24c, NTHC23]. variables [AXMJ22, ALJ^{+}22, BRC^{+}21, DMG22, MRR^{+}20, PPSS22, RCSS20]. Variance [vDC20]. variation [APF21, BSR21, CMFC23, CSM23, HGBK23, KAB23a, KABL21, LDX^{+}23, Oku21, XJW21]. variations [NWO21]. variety [LZD^{+}24]. various [DBK20, GAMN^{+}23, MFZ23, SBMG24, WBD^{+}22b]. Varroa [MMC^{+}21]. vary [FWK^{+}20]. varying [COPM23]. vector [KWG^{+}22a, MSLMM22, OW20, TK20a]. vectored [RLS^{+}24]. vectoring [MGC24]. vegetable [AAC20, ZBH^{+}22]. Vegetation [WXT^{+}20, BKTM20, BWCD23, CWY^{+}24, CGV22, DDDC21, ETAAH21, ERP21, FMRMAS24, FSZ^{+}20, FZS^{+}20, JPL21, LYC^{+}21, LCW^{+}21, LHJ^{+}24, LWL24, LMW^{+}23b, MKEH20, NKT^{+}20, QS23, WHW^{+}23, YYS^{+}22, Yu20, ZSS20, ZYM^{+}21, ZLZ^{+}24, RSB^{+}23]. vegetation-desertification [FMRMAS24]. vegetations [SOP22]. vegetative [LTR^{+}24]. vehicle [BLB24b]. velocity [AJJ24, ZLL^{+}23]. versatile [MS23]. versus [RFPRC23, Pie23]. vertebrates [ZS20]. Vertical [Jor22, ABR24, CSB22, GMM^{+}24]. Very [RSB^{+}23]. Vespula [CH22]. via [BdBG^{+}21a, BdBG^{+}21b, HZ22, KBAC20, LGA^{+}21, PF23, WLY^{+}23, ZWZ^{+}20, GML^{+}20]. Viability [OSPZ22, AZBdAI23, CDT^{+}23, HKG^{+}24, LGK^{+}23, dGMCM^{+}23]. Vibrio [DEMJ24]. video [KCHZ21]. Vietnam [KTB^{+}21]. view [EBMK20, KE24]. viewing [PAS23]. Viewpoint [Kle23]. village [JLLS23]. vineyards [KT22]. violation [BP21]. viral [KPLH24]. virginianus [KS22]. virtual [HHFB23]. virulence [Hop23]. virus [BGC^{+}23, KWG^{+}22a, SYL^{+}24]. viscous [DLY^{+}22]. visual [RBEH21, vWVH20]. Visualizing [YDLT24]. volatile [Koo24b]. vole [VSK^{+}20]. Volterra [NMGFOB21, SIOM24]. voltinism [CdCAC^{+}20, PGdAM^{+}20]. Vorticity [DZC^{+}23b, DZC^{+}23a]. VPRM [RSB^{+}23]. vs [GHbS21, ML20, YL22, PA20]. Vulnerability [FS22, HCL^{+}24, MTL^{+}23]. vulnificus [DEMJ24]. vulture [PR21].

Waard [PRdN^{+}20]. wading [YDLT24]. walk [ABB23, PDD^{+}22]. walks [AP22, BMB24, HBC^{+}22]. Wanggu [JASW23]. Wanhe [DZC^{+}23a, DZC^{+}23b]. Want [SACEU^{+}21]. war [MKN23]. Warm [Des21, PA20]. warming [Che22, KPLH24, LhNA^{+}21, MTL^{+}23, TRMB20, ZFT20, ZZŠ^{+}22, ZRMW^{+}20, LhNA^{+}22]. warning [WSFZ21]. Was [KHR23, PR23]. Wasp [CH22, BPW^{+}21]. waste [CAS^{+}24, NO24, ZA20]. wastewater [GM23]. Wasting [HCW^{+}22, BGK^{+}20, BRSW24, KS22, TBM^{+}24]. Water [DLZ^{+}20, ELS23, WXT^{+}20, BZZT22, BP21, BAM23, BGC^{+}24, CC23, nCLpW^{+}23, CLW^{+}24, DEMJ24, DCT^{+}21, DMdR24, GGE^{+}23, HLW^{+}20, HLZ^{+}22, HSJ^{+}20, HLS^{+}23, JMC^{+}22, KHS^{+}23, Lef23, LYC^{+}21, LNK^{+}21, MSL^{+}21, MFPIBGES23, NMV^{+}23, NWO21, PRM^{+}24, PGSL^{+}20, RSASP22, RFdL^{+}24, RPARM20, SC21, TBZ^{+}22, VRRI^{+}24, WLQ^{+}23, WPL^{+}23, WWP20, YJC^{+}24, ZSS21, ZZZD22, XZZ^{+}22]. Water-energy [DLZ^{+}20]. water-land [NMV^{+}23]. water-limited [HLW^{+}20, HLZ^{+}22]. waterbirds [WWA^{+}22]. waterbodies [RFCEL22]. waters [RJY^{+}20]. watershed [BH21, GSA24, HAZEF23, HCPG20, JB22, LSS^{+}21, MTZA23, PGP^{+}22, ZZLF24, dPGDdFS24]. wave [SBWB24]. wavelet [CCH20]. way [DLT^{+}22, WHdBB21]. Ways [Koo24a]. Weak [CKC20, ABZ22]. weaken [CLHZ22]. web [BAAN24, BNN20, CGRR^{+}24, CVTS20a, CVTS20b, DvDA24, DS24, GHbS21, HML^{+}22, KKH^{+}24, MAD^{+}20, MPPM23, NHGP23, OKS^{+}21, SRK^{+}20, ZBW^{+}24, vdHNH^{+}20]. web-based [CVTS20a, CVTS20b]. webs [JL23, MEPJ21, MJS24, NK23, WLLW20, WA20, vdHNH^{+}20]. weed [ABC^{+}22, CGVC21, MVC^{+}22, PBF^{+}23, PMP^{+}21]. weeds [MVC^{+}22]. weight [BPK^{+}20]. weighted [CKC^{+}24]. weighting [ZFP21]. well [GMS^{+}23, HL23, HMC20, YPS^{+}21]. well-being [GMS^{+}23, HL23, YPS^{+}21]. well-mixed [HMC20]. wellbeing [KCS22]. west [HMVR21, BGC^{+}23, KWG^{+}22a, PTG^{+}20, SPB^{+}24, SD24]. western [AM22, HH23, KT22, STRC24, TDI^{+}20, ASL^{+}20, BBCD23, LHH^{+}24, PA20]. wet [BPK^{+}20]. Wetland [MLD^{+}23, EON23, OTR^{+}21]. wetlands [AGP20, DBK20, LHJ^{+}24]. whale [MCB^{+}23, SWF21]. wheat [CDN^{+}23, GAMN^{+}23, GGS^{+}22, HRW^{+}24, Hop23, LPHO^{+}20, PTF23, PK23]. Which [ENARM^{+}23, RAB21]. while [DGVVP23, XCH^{+}21, YA24]. White [HCW^{+}22, BGK^{+}20, dMFOM^{+}20, KS22, TBM^{+}24, VRW^{+}21, KFJS20]. White-Tailed [HCW^{+}22, BGK^{+}20, KS22, TBM^{+}24, VRW^{+}21]. whitefly [SYL^{+}24]. whole [DLT^{+}22, NO23, PKO^{+}21]. whole-tree [PKO^{+}21]. wide [HKB^{+}20, RBLS23]. wider [Gri20]. wild [DAG^{+}21, GJC^{+}20, KT22, LP22, LBN^{+}22, MM21, RRFLV^{+}22]. wildcat [CKK^{+}24a, CKK^{+}24c]. wildcats [CKC^{+}24]. wildfire [AB21, ABH^{+}20, SCG^{+}23, YA24]. wildflower [CAFT22, CSCT23]. wildflower/crop [CSCT23]. wildlife [BGK^{+}20, BLB24b, COPM23, CSMP21, MD23, PAS23]. wildlife-livestock [MD23]. Williston [LYZ^{+}24]. Wind [HHI^{+}22, AJG22a, AJG22b, CMFC23, GCSlR22, JB22, WGK^{+}21, WSW23]. wind-borne [WGK^{+}21]. Wind-Disturbed [HHI^{+}22]. wind-driven [CMFC23, WSW23]. wins [MKN23]. winter [BLLL21, EBS23, GGS^{+}22, KUK^{+}21, SMS^{+}24]. winter-run [BLLL21]. within [BKTM20, CDNN^{+}20, FFM^{+}23, HGBK23, KDH^{+}21, MK21, PHK23]. within-canopy [MK21]. within-field [HGBK23]. within-patch [CDNN^{+}20]. WOFOST [SWHZ22, WSHZ24]. wolf [BGS^{+}20, KRM^{+}22]. wolves [CDC^{+}22, HKG^{+}24]. wood [MRR^{+}20, MLV24]. woodland [WHW^{+}23, SDD^{+}20]. woody [dAPdBMT23, YCK20]. working [SGLK20]. workshop [PKB^{+}23]. world [GCMU23, PAS^{+}22, vOVvdW23]. worldwide [CAK^{+}24]. Wrangel [CLW23]. Wuhan [WWC^{+}22, ZZQ23].

X [ZL24]. X-ray [ZL24]. Xiamen [LZL23]. Xinjiang [CWY^{+}24, WFW^{+}24]. Xylella [BCMV20, GSC^{+}23].

Y-chromosome [BPB^{+}20]. Yaghooti [EGAS23]. Yangtze [DZC^{+}23a, DZC^{+}23b, DYW^{+}22, HSF^{+}20]. year [HHBF24]. Yearning [Por21]. years [BPL^{+}20, LGK^{+}23, PA20]. yellow [MSLMM22, MSLM^{+}24, RJY^{+}20, nCLpW^{+}23, GJC^{+}20, MP23, SRB^{+}20, XMWL23]. yellowfin [DPC20]. Yellowstone [ERP21]. Yield [SWHZ22, AMK21, AKM22, CSCT23, CJM^{+}23, dSCdA21, DTR^{+}22, DGVVP23, For20, GAMN^{+}23, HRW^{+}24, JPL21, KTB^{+}21, LPHO^{+}20, PK23, RWF22, WSHZ24]. yields [dSBDH23, RLM23, RRFLV^{+}22]. yolk [JMHH22]. yolk-feeding [JMHH22]. Yongxing [XHZH24].

zero [LBN^{+}22]. zero-sum [LBN^{+}22]. Zhangzhou [LZL23]. zone [ALS^{+}21, Kos22, RWF22, Vas23, WXX^{+}23]. zones [AAAR23, AAAR24, TWL^{+}22]. zoning [XWX^{+}24]. zoonotic [MMMS23]. zooplankton [CMKF22, DWH^{+}21, HES^{+}20, HES^{+}23, KNP^{+}21, ZZŠ^{+}22, ZGG22].

References

AlFaisal:2023:EFS

[AAAR23] Abdullah Al Faisal, Abdulla Al Kafy, Farzana Afroz, and Zullyadini A. Rahaman. Exploring and forecasting spatial and temporal patterns of fire hazard risk in Nepal’s tiger conservation zones. Ecological Modelling, 476(??): ??, February 2023. CODEN ECMODT. ISSN ????

               URL http://www.sciencedirect.com/science/article/

pii/S0304380022003428. See retraction notice [AAAR24].

AlFaisal:2024:RNE

[AAAR24] Abdullah Al Faisal, Abdulla Al Kafy, Farzana Afroz, and Zullyadini A. Rahaman. Retraction note to “Exploring and forecasting spatial and temporal patterns of fire hazard risk in Nepal’s tiger conservation zones” [Ecological Modelling 476 (2023) 1–12/110244]. Ecological Modelling, 494(??):??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001534.

Asgharipour:2020:ESF

[AAC20] Mohammad Reza Asgharipour, Zahra Amiri, and Daniel E. Campbell. Evaluation of the sustainability of four greenhouse vegetable production ecosystems based on an analysis of emergy and social characteristics”. Ecological Modelling, 424(??):??, May 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300934.

Aveytua-Alcazar:2020:CUR

[AACCIS20] L. Aveytua-Alcazar, D. Melaku Canu, V. F. Camacho-Ibar, and C. Solidoro. Changes in upwelling regimes in a Mediterranean-type lagoon: a model application. Ecological Modelling, 418(??):??, February 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019304168.

Acosta-Arreola:2023:PDT

[AADHA^{+}23] Jaime Acosta-Arreola, Elisa Domínguez-Hüttinger, Pablo Aguirre, Nicolás González, and Jorge A. Meave. Predicting dynamic trajectories of a protected plant community under contrasting conservation regimes: Insights from databased modelling. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001801.

Amiri:2022:INR

[AAM^{+}22] Zahra Amiri, Mohammad Reza Asgharipour, Esfandiar Hassani Moghadam, Ebrahim Kakolvand, and Daniel E. Campbell. Investigating the need to replace the conventional method of sugar beet production in Lorestan Province, Iran based on the arguments obtained from emergy calculations. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002200196X.

Arango:2023:CFA

[AAR23] Miguel A. Arango, Aavudai Anandhi, and Charles W. Rice. Conceptual framework addressing timescale mismatch uncertainty: Nitrous-oxide (N_{2}O) modeled and measured, Kansas, USA. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002661.

Antoci:2023:WHP

[AARS23] Angelo Antoci, Marco Apollonio, Paolo Russu, and Massimo Scandura. When humans play evolutionary games with animal species. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT. ISSN ???? URL http://www.

sciencedirect.com/science/article/pii/S0304380022003192.

Aedo:2021:NML

[AB21] Sebastián A. Aedo and Carlos A. Bonilla. A numerical model for linking soil organic matter decay and wildfire severity. Ecological Modelling, 447(??):??, May 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000776.

Ahmed:2023:RWM

[ABB23] Danish A. Ahmed, Joseph D. Bailey, and Michael B. Bonsall. On random walk models as a baseline for animal movement in three-dimensional space. Ecological Modelling, 475(??):??, January 2023. CODEN ???? ISSN

????  URL http://www.sciencedirect.com/science/
article/pii/S0304380022002708.

Augustinus:2022:GTP

[ABC^{+}22] Benno A. Augustinus, Moshe Blum, Sandra Citterio, Rodolfo Gentili, David Helman, David Nestel, Urs Schaffner, Heinz Müller-Schärer, and Itamar M. Lensky. Ground-truthing predictions of a demographic model driven by land surface temperatures with a weed biocontrol cage experiment. Ecological Modelling, 466(??):??, April 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000242.

Abernethy:2020:AEE

[Abe20] Gavin M. Abernethy. Allometry in an eco-evolutionary network model. Ecological Modelling, 427(??):??, July 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301629.

Ager:2020:MEA

[ABH^{+}20] Alan A. Ager, Ana M. G. Barros, Rachel Houtman, Rob Seli, and Michelle A. Day. Modelling the effect of accelerated forest management on long-term wildfire activity. Ecological Modelling, 421(??):??, April 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002030034X.

Abbas:2024:RSU

[ABKA24] Shahbaz Abbas, Abdur Rehman Bin Nasir Rao, Farrukh Khattak, and Yasir Ahmad. Revolutionising sustainability using a new triplet: a system dynamic model. Ecological Modelling, 492(??):??, June 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002400108X.

Agudelo:2020:MIA

[ABM20] César Augusto Ruiz Agudelo, Sandra Liliana Hurtado Bustos, and Carmen Alicia Parrado Moreno. Modeling interactions among multiple ecosystem services. A critical review. Ecological Modelling, 429(??):??, August 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301757.

Ani:2024:MBD

[ABR24] Chinenye J. Ani, Mark Baird, and Barbara Robson. Modelling buoyancy-driven vertical movement of Trichodesmium application in the Great Barrier Reef. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380023002971.

Akpoue:2021:MBA

[ABRG21] Kouadio Jean-Philippe Akpoué, Sébastien Barot, Xavier Raynaud, and Jacques Gignoux. Modeling the biomass allocation of tree resprout in a fire-prone savanna. Ecological Modelling, 448(??):??, May 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000983.

Abdalgader:2022:SWS

[ABZ22] Tarteel Abdalgader, Malay Banerjee, and Lai Zhang. Spatially weak syncronization of spreading pattern between Aedes albopictus and dengue fever. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002241.

Argenti:2022:URS

[ACFM22] G. Argenti, M. Chiesi, L. Fibbi, and F. Maselli. Use of remote sensing and bio-geochemical models to estimate the net carbon fluxes of managed mountain grasslands. Ecological Modelling, 474(??):??, December 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002538.

Aulicino:2021:UAH

[ACZ^{+}21] Giuseppe Aulicino, Cinzia Cesarano, Mohamed Zerrouki, Simon Ruiz, Giorgio Budillon, and Yuri Cotroneo. On the use of ABACUS high resolution glider observations for the assessment of phytoplankton ocean biomass from CMEMS model products. Ecological Modelling, 455(??):??, September 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021001812.

Adhurya:2021:SEA

[ADR21] Sagar Adhurya, Suvendu Das, and Santanu Ray. Simulating the effects of aquatic avifauna on the phosphorus dynamics of aquatic systems. Ecological Modelling, 445(??):??, April 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021000661.

Ai:2022:SIM

[AE22] Dexiecuo Ai and M. D. Farnon Ellwood. A spatially implicit model fails to predict the structure of spatially explicit metacommunities under high dispersal. Ecological Modelling, 474(??):??, December 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002526.

Amaro:2023:ESA

[AFdSM23] George Amaro, Elisangela Gomes Fidelis, Ricardo Siqueira da Silva, and Cesar Augusto Marchioro. Effect of study area extent on the potential distribution of species: a case study with models for Raoiella indica Hirst (Acari: Tenuipalpidae). Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001850.

Aubry:2024:IPS

[AFG24] Philippe Aubry, Charlotte Francesiaz, and Matthieu Guillemain. On the impact of preferential sampling on ecological status and trend assessment. Ecological Modelling, 492(??): ??, June 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000954.

Amato:2022:QRT

[AG22] Matthew T. Amato and Daniel Giménez. Quantifying root turnover in grasslands from biomass dynamics: Application of the growth-maintenance respiration paradigm and re-analysis of historical data. Ecological Modelling, 467(??):??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002200062X.

Allen:2020:MCE

[AGP20] Corrie Allen, Rodolphe Gonzales, and Lael Parrott. Modelling the contribution of ephemeral wetlands to landscape connectivity. Ecological Modelling, 419(??):??, March 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300156.

An:2021:CTO

[AGS^{+}21] Li An, Volker Grimm, Abigail Sullivan, B. L. Turner II, Nicolas Malleson, Alison Heppenstall, Christian Vincenot, Derek Robinson, Xinyue Ye, Jianguo Liu, Emilie Lindkvist, and Wenwu Tang. Challenges, tasks, and opportunities in modeling agent-based complex systems. Ecological Modelling, 457(??):??, October 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S030438002100243X. See corrigendum [AGS^{+}22].

An:2022:CCT

[AGS^{+}22] Li An, Volker Grimm, Abigail Sullivan, B. L. Turner, Nicolas Malleson, Alison Heppenstall, Christian Vincenot, Derek Robinson, Xinyue Ye, Jianguo Liu, Emilie Lindkvist, and Wenwu Tang. Corrigendum to “Challenges, tasks, and opportunities in modeling agent-based complex systems”, Ecological Modeling, 2021, 457: 109,685, page 1– 15. Ecological Modelling, 471(??):??, September 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001727. See [AGS^{+}21].

Abdulwahab:2022:CCD

[AHH22] Umarfarooq A. Abdulwahab, Edd Hammill, and Charles P. Hawkins. Choice of climate data affects the performance and interpretation of species distribution models. Ecological Modelling, 471(??):??, September 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001521.

Armstrong:2020:MSA

[AHO^{+}20] A. H. Armstrong, A. Huth, B. Osmanoglu, G. Sun, K. J. Ranson, and R. Fischer. A multi-scaled analysis of forest structure using individual-based modeling in a Costa Rican rainforest. Ecological Modelling, 433(??):??, October 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020302969.

Asif:2023:DMS

[AIA^{+}23] Muhammad Asif, Azhar Inam, Jan Adamowski, Muhammad Shoaib, Hisham Tariq, Shakil Ahmad, Mohammad Reza Alizadeh, and Aftab Nazeer. Development of methods for the simplification of complex group built causal loop diagrams: a case study of the Rechna doab. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022002903.

Adams:2022:ELS

[AJ22] Jean V. Adams and Michael L. Jones. Estimation of lakescale stock-recruitment models for Great Lakes sea lampreys. Ecological Modelling, 467(??):??, May 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000394.

A:2022:MLW

[AJG22a] Tadesse A., Jaehak Jeong, and Colleen H. M. Green. Modeling landscape wind erosion processes on rangelands using the APEX model. Ecological Modelling, 467(??):??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022000485. See [AJG22b].

Abitew:2022:CML

[AJG22b] Tadesse A. Abitew, Jaehak Jeong, and Colleen H. M. Green. Corrigendum to Modeling landscape wind erosion processes on rangelands using the APEX model. Ecological Modelling, 472(??):??, October 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001879. See [AJG22a].

Anukiruthika:2024:DBM

[AJJ24] T. Anukiruthika, Digvir S. Jayas, and Fuji Jian. Diffusivity and bias movement velocity of Cryptolestes ferrugineus (Coleoptera: Laemophloeidae) adults under different temperature and moisture differences. Ecological Modelling, 491(??): ??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000747.

Auger:2022:IMS

[AKM22] Pierre Auger, Bob Kooi, and Ali Moussaoui. Increase of maximum sustainable yield for fishery in two patches with fast migration. Ecological Modelling, 467(??):??, May 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000254.

Asada:2022:ISH

[AKY^{+}22] K. Asada, T. Kanda, N. Yamashita, M. Asano, and S. Eguchi. Interpreting stoichiometric homeostasis and flexibility of soil microbial biomass carbon, nitrogen, and phosphorus. Ecological Modelling, 470(??):??, August 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001296.

Andersen:2022:ILA

[ALJ^{+}22] Desiree Andersen, Spartak N. Litvinchuk, Hoan Jin Jang, Jianping Jiang, Kyo Soung Koo, Irina Maslova, Daemin Kim, Yikweon Jang, and Amaël Borzée. Incorporation of latitudeadjusted bioclimatic variables increases accuracy in species distribution models. Ecological Modelling, 469(??):??, July 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022001004.

Augustine:2020:CLF

[ALK20] Starrlight Augustine, Konstadia Lika, and Sebastiaan A. L. M. Kooijman. Comparing loss functions and interval estimates for survival data. Ecological Modelling, 430(??):??, August 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020301496.

An:2021:DIC

[ALS^{+}21] Yi An, Shiliang Liu, Yongxiu Sun, Fangning Shi, Yixuan Liu, and Robert Beazley. Determining the importance of core areas in the alpine shrub-meadow gradient zone of the Qinghai– Tibet Plateau. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304567.

Aliyu:2021:IBM

[AM21] Murtala Bello Aliyu and Mohd Hafiz Mohd. The interplay between mutualism, competition and dispersal promotes species coexistence in a multiple interactions type system. Ecological Modelling, 452(??):??, July 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001575.

Almonacid:2022:SPM

[AM22] Pablo Mata Almonacid and Carolina Medel. A structurepreserving model for the dynamics of estuarine ecosystems and its application in western Patagonia fjords. Ecological Modelling, 466(??):??, April 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021004087.

Adiku:2021:CMF

[AMK21] Samuel G. K. Adiku, Dilys S. MacCarthy, and Samuel K. Kumahor. A conceptual modelling framework for simulating the impact of soil degradation on maize yield in datasparse regions of the tropics. Ecological Modelling, 448(??):??, May 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002100096X.

Ahmadi:2023:ISD

[AMP^{+}23] Kourosh Ahmadi, Shirin Mahmoodi, Subodh Chandra Pal, Asish Saha, Indrajit Chowdhuri, Trinh Trong Nguyen, Scott Jarvie, Marta Szostak, Jaroslaw Socha, and Van Nam Thai. Improving species distribution models for dominant trees in climate data-poor forests using high-resolution remote sensing. Ecological Modelling, 475(??):??, January 2023. CODEN

???? ISSN ???? URL http://www.sciencedirect.com/
science/article/pii/S0304380022002885.

Anonymous:2020:Aa

[Ano20a] Anonymous. 1 April 2020. Ecological Modelling, 421(??):??, April 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Ac

[Ano20b] Anonymous. 1 August 2020. Ecological Modelling, 429(??):??, August 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Da

[Ano20c] Anonymous. 1 December 2020. Ecological Modelling, 437(??): ??, December 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Fa

[Ano20d] Anonymous. 1 February 2020. Ecological Modelling, 417(??): ??, February 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Ja

[Ano20e] Anonymous. 1 January 2020. Ecological Modelling, 415(??): ??, January 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Je

[Ano20f] Anonymous. 1 July 2020. Ecological Modelling, 427(??):??, July 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Jc

[Ano20g] Anonymous. 1 June 2020. Ecological Modelling, 425(??):??, June 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Ma

[Ano20h] Anonymous. 1 March 2020. Ecological Modelling, 419(??):??, March 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Mc

[Ano20i] Anonymous. 1 May 2020. Ecological Modelling, 423(??):??, May 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Na

[Ano20j] Anonymous. 1 November 2020. Ecological Modelling, 435(??): ??, November 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Oa

[Ano20k] Anonymous. 1 October 2020. Ecological Modelling, 433(??): ??, October 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Sa

[Ano20l] Anonymous. 1 September 2020. Ecological Modelling, 431(??): ??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Ab

[Ano20m] Anonymous. 15 April 2020. Ecological Modelling, 422(??):??, April 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Ad

[Ano20n] Anonymous. 15 August 2020. Ecological Modelling, 430(??): ??, August 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Db

[Ano20o] Anonymous. 15 December 2020. Ecological Modelling, 438 (??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Fb

[Ano20p] Anonymous. 15 February 2020. Ecological Modelling, 418(??): ??, February 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Jb

[Ano20q] Anonymous. 15 January 2020. Ecological Modelling, 416(??): ??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Jf

[Ano20r] Anonymous. 15 July 2020. Ecological Modelling, 428(??):??, July 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Jd

[Ano20s] Anonymous. 15 June 2020. Ecological Modelling, 426(??):??, June 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Mb

[Ano20t] Anonymous. 15 March 2020. Ecological Modelling, 420(??):??, March 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Md

[Ano20u] Anonymous. 15 May 2020. Ecological Modelling, 424(??):??, May 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Nb

[Ano20v] Anonymous. 15 November 2020. Ecological Modelling, 436 (??):??, November 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Ob

[Ano20w] Anonymous. 15 October 2020. Ecological Modelling, 434(??): ??, October 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:Sb

[Ano20x] Anonymous. 15 September 2020. Ecological Modelling, 432 (??):??, September 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2020:EBa

[Ano20y] Anonymous. Editorial Board. Ecological Modelling, 415 (??):??, January 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019304041.

Anonymous:2020:EBb

[Ano20z] Anonymous. Editorial Board. Ecological Modelling, 416(??): ??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020300053.

Anonymous:2020:EBc

[Ano20-27] Anonymous. Editorial Board. Ecological Modelling, 417(??): ??, February 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020300259.

Anonymous:2020:EBd

[Ano20-28] Anonymous. Editorial Board. Ecological Modelling, 418 (??):??, February 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020300429.

Anonymous:2020:EBe

[Ano20-29] Anonymous. Editorial Board. Ecological Modelling, 419 (??):??, March 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020300533.

Anonymous:2020:EBf

[Ano20-30] Anonymous. Editorial Board. Ecological Modelling, 420 (??):??, March 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020300636.

Anonymous:2020:EBg

[Ano20-31] Anonymous. Editorial Board. Ecological Modelling, 421(??): ??, April 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020300855.

Anonymous:2020:EBh

[Ano20-32] Anonymous. Editorial Board. Ecological Modelling, 422 (??):??, April 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020300971.

Anonymous:2020:EBi

[Ano20-33] Anonymous. Editorial Board. Ecological Modelling, 423(??): ??, May 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301149.

Anonymous:2020:EBj

[Ano20-34] Anonymous. Editorial Board. Ecological Modelling, 424 (??):??, May 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020301381.

Anonymous:2020:EBk

[Ano20-35] Anonymous. Editorial Board. Ecological Modelling, 425(??): ??, June 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301538.

Anonymous:2020:EBl

[Ano20-36] Anonymous. Editorial Board. Ecological Modelling, 426 (??):??, June 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020301708.

Anonymous:2020:EBm

[Ano20-37] Anonymous. Editorial Board. Ecological Modelling, 427(??): ??, July 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301861.

Anonymous:2020:EBn

[Ano20-38] Anonymous. Editorial Board. Ecological Modelling, 428 (??):??, July 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020302143.

Anonymous:2020:EBo

[Ano20-39] Anonymous. Editorial Board. Ecological Modelling, 429 (??):??, August 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020302271.

Anonymous:2020:EBp

[Ano20-40] Anonymous. Editorial Board. Ecological Modelling, 430 (??):??, August 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020302416.

Anonymous:2020:EBq

[Ano20-41] Anonymous. Editorial Board. Ecological Modelling, 431 (??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020302854.

Anonymous:2020:EBr

[Ano20-42] Anonymous. Editorial Board. Ecological Modelling, 432 (??):??, September 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020303070.

Anonymous:2020:EBs

[Ano20-43] Anonymous. Editorial Board. Ecological Modelling, 433 (??):??, October 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020303240.

Anonymous:2020:EBt

[Ano20-44] Anonymous. Editorial Board. Ecological Modelling, 434(??): ??, October 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020303458.

Anonymous:2020:EBu

[Ano20-45] Anonymous. Editorial Board. Ecological Modelling, 435(??): ??, November 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020303732.

Anonymous:2020:EBv

[Ano20-46] Anonymous. Editorial Board. Ecological Modelling, 436 (??):??, November 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020303860.

Anonymous:2020:EBw

[Ano20-47] Anonymous. Editorial Board. Ecological Modelling, 437(??): ??, December 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438002030404X.

Anonymous:2020:EBx

[Ano20-48] Anonymous. Editorial Board. Ecological Modelling, 438 (??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020304294.

Anonymous:2020:ETT

[Ano20-49] Anonymous. Enzo tiezzi, turning pioneering into modern ideas: tempos, ecodynamics and sustainable economy. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302337.

Anonymous:2020:IMD

[Ano20-50] Anonymous. Integrating the multi-domainal and multidimensional nature of animal movement into ecological modelling. Ecological Modelling, 436(??):??, November 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302908.

Anonymous:2020:TML

[Ano20-51] Anonymous. Transitioning machine learning from theory to practice in natural resources management. Ecological Modelling, 435(??):??, November 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303276.

Anonymous:2021:Aa

[Ano21a] Anonymous. 1 April 2021. Ecological Modelling, 445(??):??, April 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Ac

[Ano21b] Anonymous. 1 August 2021. Ecological Modelling, 453(??):??, August 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Da

[Ano21c] Anonymous. 1 December 2021. Ecological Modelling, 461(??): ??, December 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Fa

[Ano21d] Anonymous. 1 February 2021. Ecological Modelling, 441(??): ??, February 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Ja

[Ano21e] Anonymous. 1 January 2021. Ecological Modelling, 439(??): ??, January 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Je

[Ano21f] Anonymous. 1 July 2021. Ecological Modelling, 451(??):??, July 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Jc

[Ano21g] Anonymous. 1 June 2021. Ecological Modelling, 449(??):??, June 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Ma

[Ano21h] Anonymous. 1 March 2021. Ecological Modelling, 443(??):??, March 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Mc

[Ano21i] Anonymous. 1 May 2021. Ecological Modelling, 447(??):??, May 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Na

[Ano21j] Anonymous. 1 November 2021. Ecological Modelling, 459(??): ??, November 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Oa

[Ano21k] Anonymous. 1 October 2021. Ecological Modelling, 457(??): ??, October 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Sa

[Ano21l] Anonymous. 1 September 2021. Ecological Modelling, 455(??): ??, September 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Ab

[Ano21m] Anonymous. 15 April 2021. Ecological Modelling, 446(??):??, April 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Ad

[Ano21n] Anonymous. 15 August 2021. Ecological Modelling, 454(??): ??, August 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Db

[Ano21o] Anonymous. 15 December 2021. Ecological Modelling, 462 (??):??, December 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Fb

[Ano21p] Anonymous. 15 February 2021. Ecological Modelling, 442(??): ??, February 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Jb

[Ano21q] Anonymous. 15 January 2021. Ecological Modelling, 440(??): ??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Jf

[Ano21r] Anonymous. 15 July 2021. Ecological Modelling, 452(??):??, July 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Jd

[Ano21s] Anonymous. 15 June 2021. Ecological Modelling, 450(??):??, June 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Mb

[Ano21t] Anonymous. 15 March 2021. Ecological Modelling, 444(??):??, March 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Md

[Ano21u] Anonymous. 15 May 2021. Ecological Modelling, 448(??):??, May 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Nb

[Ano21v] Anonymous. 15 November 2021. Ecological Modelling, 460 (??):??, November 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Ob

[Ano21w] Anonymous. 15 October 2021. Ecological Modelling, 458(??): ??, October 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:Sb

[Ano21x] Anonymous. 15 September 2021. Ecological Modelling, 456 (??):??, September 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:EBa

[Ano21y] Anonymous. Editorial Board. Ecological Modelling, 439 (??):??, January 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438002030435X.

Anonymous:2021:EBb

[Ano21z] Anonymous. Editorial Board. Ecological Modelling, 440(??): ??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021000065.

Anonymous:2021:EBc

[Ano21-27] Anonymous. Editorial Board. Ecological Modelling, 441(??): ??, February 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021000132.

Anonymous:2021:EBd

[Ano21-28] Anonymous. Editorial Board. Ecological Modelling, 442 (??):??, February 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021000363.

Anonymous:2021:EBe

[Ano21-29] Anonymous. Editorial Board. Ecological Modelling, 443 (??):??, March 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021000508.

Anonymous:2021:EBf

[Ano21-30] Anonymous. Editorial Board. Ecological Modelling, 444 (??):??, March 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021000594.

Anonymous:2021:EBg

[Ano21-31] Anonymous. Editorial Board. Ecological Modelling, 445(??): ??, April 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021000867.

Anonymous:2021:EBh

[Ano21-32] Anonymous. Editorial Board. Ecological Modelling, 446 (??):??, April 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021000934.

Anonymous:2021:EBi

[Ano21-33] Anonymous. Editorial Board. Ecological Modelling, 447(??): ??, May 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021001071.

Anonymous:2021:EBj

[Ano21-34] Anonymous. Editorial Board. Ecological Modelling, 448 (??):??, May 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021001174.

Anonymous:2021:EBk

[Ano21-35] Anonymous. Editorial Board. Ecological Modelling, 449(??): ??, June 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002100140X.

Anonymous:2021:EBl

[Ano21-36] Anonymous. Editorial Board. Ecological Modelling, 450 (??):??, June 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021001472.

Anonymous:2021:EBm

[Ano21-37] Anonymous. Editorial Board. Ecological Modelling, 451(??): ??, July 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021001617.

Anonymous:2021:EBn

[Ano21-38] Anonymous. Editorial Board. Ecological Modelling, 452 (??):??, July 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021001782.

Anonymous:2021:EBo

[Ano21-39] Anonymous. Editorial Board. Ecological Modelling, 453 (??):??, August 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438002100185X.

Anonymous:2021:EBp

[Ano21-40] Anonymous. Editorial Board. Ecological Modelling, 454 (??):??, August 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021002039.

Anonymous:2021:EBq

[Ano21-41] Anonymous. Editorial Board. Ecological Modelling, 455 (??):??, September 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021002271.

Anonymous:2021:EBr

[Ano21-42] Anonymous. Editorial Board. Ecological Modelling, 456 (??):??, September 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021002556.

Anonymous:2021:EBs

[Ano21-43] Anonymous. Editorial Board. Ecological Modelling, 457 (??):??, October 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021002702.

Anonymous:2021:EBt

[Ano21-44] Anonymous. Editorial Board. Ecological Modelling, 458(??): ??, October 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021002969.

Anonymous:2021:EBu

[Ano21-45] Anonymous. Editorial Board. Ecological Modelling, 459(??): ??, November 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021003082.

Anonymous:2021:EBv

[Ano21-46] Anonymous. Editorial Board. Ecological Modelling, 460 (??):??, November 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021003185.

Anonymous:2021:EBw

[Ano21-47] Anonymous. Editorial Board. Ecological Modelling, 461(??): ??, December 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438002100332X.

Anonymous:2021:EBx

[Ano21-48] Anonymous. Editorial Board. Ecological Modelling, 462 (??):??, December 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021003537.

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Ecological Modelling, 466(??):??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. Ecological Modelling, 470(??):??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:D

[Ano22c] Anonymous. December 2022. Ecological Modelling, 474(??): ??, December 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:EBa

[Ano22d] Anonymous. Editorial Board. Ecological Modelling, 463 (??):??, January 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021003732.

Anonymous:2022:EBb

[Ano22e] Anonymous. Editorial Board. Ecological Modelling, 464 (??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021004038.

Anonymous:2022:EBc

[Ano22f] Anonymous. Editorial Board. Ecological Modelling, 465(??): ??, March 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000199.

Anonymous:2022:EBd

[Ano22g] Anonymous. Editorial Board. Ecological Modelling, 466(??): ??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000448.

Anonymous:2022:EBe

[Ano22h] Anonymous. Editorial Board. Ecological Modelling, 467(??): ??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000941.

Anonymous:2022:EBf

[Ano22i] Anonymous. Editorial Board. Ecological Modelling, 468(??): ??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001041.

Anonymous:2022:EBg

[Ano22j] Anonymous. Editorial Board. Ecological Modelling, 469(??): ??, July 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001363.

Anonymous:2022:EBh

[Ano22k] Anonymous. Editorial Board. Ecological Modelling, 470(??): ??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001570.

Anonymous:2022:EBi

[Ano22l] Anonymous. Editorial Board. Ecological Modelling, 471 (??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380022001910.

Anonymous:2022:EBj

[Ano22m] Anonymous. Editorial Board. Ecological Modelling, 472 (??):??, October 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380022002289.

Anonymous:2022:EBk

[Ano22n] Anonymous. Editorial Board. Ecological Modelling, 473 (??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380022002629.

Anonymous:2022:EBl

[Ano22o] Anonymous. Editorial Board. Ecological Modelling, 474 (??):??, December 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380022002952.

Anonymous:2022:F

[Ano22p] Anonymous. February 2022. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:Ja

[Ano22q] Anonymous. January 2022. Ecological Modelling, 463(??):??, January 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:Jc

[Ano22r] Anonymous. July 2022. Ecological Modelling, 469(??):??, July 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:Jb

[Ano22s] Anonymous. June 2022. Ecological Modelling, 468(??):??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:Ma

[Ano22t] Anonymous. March 2022. Ecological Modelling, 465(??):??, March 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:Mb

[Ano22u] Anonymous. May 2022. Ecological Modelling, 467(??):??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:N

[Ano22v] Anonymous. November 2022. Ecological Modelling, 473(??): ??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:O

[Ano22w] Anonymous. October 2022. Ecological Modelling, 472(??):??, October 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2022:S

[Ano22x] Anonymous. September 2022. Ecological Modelling, 471(??): ??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023. Ecological Modelling, 478(??):??, April 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023. Ecological Modelling, 482(??):??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. Ecological Modelling, 486(??): ??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. Ecological Modelling, 475 (??):??, January 2023. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003374.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. Ecological Modelling, 476 (??):??, February 2023. CODEN ECMODT. ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003647.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. Ecological Modelling, 477(??): ??, March 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000182.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. Ecological Modelling, 478(??): ??, April 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002300042X.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. Ecological Modelling, 479(??): ??, May 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000650.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. Ecological Modelling, 480(??): ??, June 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000789.

Anonymous:2023:EBg

[Ano23j] Anonymous. Editorial Board. Ecological Modelling, 481(??): ??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001199.

Anonymous:2023:EBh

[Ano23k] Anonymous. Editorial Board. Ecological Modelling, 482(??): ??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001618.

Anonymous:2023:EBi

[Ano23l] Anonymous. Editorial Board. Ecological Modelling, 483 (??):??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380023001965.

Anonymous:2023:EBj

[Ano23m] Anonymous. Editorial Board. Ecological Modelling, 484 (??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380023002259.

Anonymous:2023:EBk

[Ano23n] Anonymous. Editorial Board. Ecological Modelling, 485 (??):??, November 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438002300251X.

Anonymous:2023:EBl

[Ano23o] Anonymous. Editorial Board. Ecological Modelling, 486 (??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380023002855.

Anonymous:2023:F

[Ano23p] Anonymous. February 2023. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT. ISSN ????

Anonymous:2023:Ja

[Ano23q] Anonymous. January 2023. Ecological Modelling, 475(??):??, January 2023. CODEN ???? ISSN ????

Anonymous:2023:Jc

[Ano23r] Anonymous. July 2023. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2023:Jb

[Ano23s] Anonymous. June 2023. Ecological Modelling, 480(??):??, June 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2023:Ma

[Ano23t] Anonymous. March 2023. Ecological Modelling, 477(??):??, March 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2023:Mb

[Ano23u] Anonymous. May 2023. Ecological Modelling, 479(??):??, May 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2023:N

[Ano23v] Anonymous. November 2023. Ecological Modelling, 485(??): ??, November 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2023:O

[Ano23w] Anonymous. October 2023. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2023:S

[Ano23x] Anonymous. September 2023. Ecological Modelling, 483(??): ??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Ecological Modelling, 490(??):??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. August 2024. Ecological Modelling, 494(??):??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2024:EBa

[Ano24c] Anonymous. Editorial Board. Ecological Modelling, 487 (??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380023003137.

Anonymous:2024:EBb

[Ano24d] Anonymous. Editorial Board. Ecological Modelling, 488 (??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380024000048.

Anonymous:2024:EBc

[Ano24e] Anonymous. Editorial Board. Ecological Modelling, 489(??): ??, March 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000309.

Anonymous:2024:EBd

[Ano24f] Anonymous. Editorial Board. Ecological Modelling, 490(??): ??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000693.

Anonymous:2024:EBe

[Ano24g] Anonymous. Editorial Board. Ecological Modelling, 491(??): ??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000991.

Anonymous:2024:EBf

[Ano24h] Anonymous. Editorial Board. Ecological Modelling, 492(??): ??, June 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001339.

Anonymous:2024:EBg

[Ano24i] Anonymous. Editorial Board. Ecological Modelling, 493(??): ??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001601.

Anonymous:2024:EBh

[Ano24j] Anonymous. Editorial Board. Ecological Modelling, 494(??): ??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001820.

Anonymous:2024:F

[Ano24k] Anonymous. February 2024. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2024:Ja

[Ano24l] Anonymous. January 2024. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2024:Jc

[Ano24m] Anonymous. July 2024. Ecological Modelling, 493(??):??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2024:Jb

[Ano24n] Anonymous. June 2024. Ecological Modelling, 492(??):??, June 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2024:Ma

[Ano24o] Anonymous. March 2024. Ecological Modelling, 489(??):??, March 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2024:Mb

[Ano24p] Anonymous. May 2024. Ecological Modelling, 491(??):??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Alqubori:2022:AST

[AP22] Omar Alqubori and Sergei Petrovskii. Analysis of simulated trap counts arising from correlated and biased random walks. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001272.

Armstrong:2021:IIV

[APF21] Doug P. Armstrong, Elizabeth H. Parlato, and Peter G. H. Frost. Incorporating individual variation in survival, reproduction and detection rates when projecting dynamics of small populations. Ecological Modelling, 455(??):??, September 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002076.

Arnould-Petre:2021:IBM

[APGD^{+}21] Margot Arnould-Pétré, Charlène Guillaumot, Bruno Danis, Jean-Pierre Féral, and Thomas Saucède. Individualbased model of population dynamics in a sea urchin of the Kerguelen Plateau (Southern Ocean), Abatus cordatus, under changing environmental conditions. Ecological Modelling, 440(??):??, January 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002030418X.

Azuaje-Rodriguez:2022:GFE

[ARSC22] Roxiris A. Azuaje-Rodríguez, Sofia Marques Silva, and Caio J. Carlos. Not going with the flow: Ecological niche of a migratory seabird, the South American tern Sterna hirundinacea. Ecological Modelling, 463(??):??, January 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003495.

AlAdwani:2022:FCE

[AS22] Mohammad AlAdwani and Serguei Saavedra. Feasibility conditions of ecological models: Unfolding links between model parameters. Ecological Modelling, 466(??):??, April 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000266.

Abbasian:2022:MTN

[ASHK22] Hassan Abbasian, Eisa Solgi, Seyed Mohsen Hosseini, and Seyed Hossein Kia. Modeling terrestrial net ecosystem exchange using machine learning techniques based on flux tower measurements. Ecological Modelling, 466(??):??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000278.

Apetrei:2024:IRS

[ASKR24] Cristina I. Apetrei, Nikita Strelkovskii, Nikolay Khabarov, and Valeria Javalera Rincón. Improving the representation of smallholder farmers’ adaptive behaviour in agentbased models: Learning-by-doing and social learning. Ecological Modelling, 489(??):??, March 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023003393.

Aurbach:2020:SBF

[ASL^{+}20] Annika Aurbach, Baptiste Schmid, Felix Liechti, Ndaona Chokani, and Reza Abhari. Simulation of broad front bird migration across Western Europe. Ecological Modelling, 415 (??):??, January 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019303874.

Alsamadisi:2020:EIA

[ATP20] Adam G. Alsamadisi, Liem T. Tran, and Monica Papeş. Employing inferences across scales: Integrating spatial data with different resolutions to enhance Maxent models. Ecological Modelling, 415(??):??, January 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019303655.

Ackleh:2022:MIE

[AV22] Azmy S. Ackleh and Amy Veprauskas. Modeling the invasion and establishment of a tick-borne pathogen. Ecological Modelling, 467(??):??, May 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000382.

Andrianova:2023:MSN

[AVP^{+}23] Elena Andrianova, Ivan Vladimirov, Olga Pavlova, Eugene Skrebnikov, and Denis Bogomaz. Modeling of the spread of non-adaptive traits in a natural population due to selfcompatibility. Ecological Modelling, 482(??):??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002300128X.

Accolla:2020:DTD

[AVRC^{+}20] Chiara Accolla, Maxime Vaugeois, Pamela Rueda-Cediel, Adrian Moore, Gonçalo M. Marques, Purvaja Marella, and Valery E. Forbes. DEB-tox and data gaps: Consequences for individual-level outputs. Ecological Modelling, 431(??):??, September 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301794.

An:2024:FSS

[AWW24] Qi An, Hao Wang, and Xiunan Wang. Fish survival subject to algal bloom: Resource-based growth models with algal digestion delay and detritus-nutrient recycling delay. Ecological Modelling, 491(??):??, May 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000607.

Ai:2022:EMF

[AXMJ22] Bin Ai, Dixiang Xie, Shifa Ma, and Haiyan Jiang. An EasyCA model with few steady variables and clone stamp strategy for simulation of urban growth in metropolitan areas. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000709.

Aguirre-Zapata:2023:PIG

[AZAD^{+}23] Estefania Aguirre-Zapata, Hernan Alvarez, Carla Vanina Dagatti, Fernando di Sciascio, and Adriana N. Amicarelli. Parametric interpretability of growth kinetics equations in a process model for the life cycle of Lobesia botrana. Ecological Modelling, 482(??):??, August 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001382.

Alvarez:2023:CME

[AZBdAI23] Isabelle Alvarez, Laetitia Zaleski, Jean-Pierre Briot, and Marta de A. Irving. Collective management of environmental commons with multiple usages: a guaranteed viability approach. Ecological Modelling, 475(??):??, January 2023. CO-

DEN ???? ISSN ???? URL http://www.sciencedirect.
com/science/article/pii/S0304380022002848.

Aguirre-Zapata:2022:SPG

[AZMD^{+}22] Estefania Aguirre-Zapata, Humberto Morales, Carla V. Dagatti, Fernando di Sciascio, and Adriana N. Amicarelli. Semi physical growth model of Lobesia botrana under laboratory conditions for Argentina’s Cuyo region. Ecological Modelling, 464(??):??, February 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003483.

Bentley:2020:RDR

[BA20] Chance Bentley and Aavudai Anandhi. Representing driverresponse complexity in ecosystems using an improved conceptual model. Ecological Modelling, 437(??):??, December 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020303902.

Barros:2024:UTD

[BAAN24] Mónica E. Barros, Ana Arriagada, Hugo Arancibia, and Sergio Neira. Using a time-dynamic food web model to compare predation and fishing mortality in Pleuroncodes monodon (Galatheidae: Crustaceae) and other benthic and demersal resource species off central Chile. Ecological Modelling, 487(??):??, January 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002764.

Bagchi:2022:MSP

[BACS22] Dweepabiswa Bagchi, Ramesh Arumugam, V. K. Chandrasekar, and D. V. Senthilkumar. Metacommunity stability and persistence for predation turnoff in selective patches. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001259.

Bourceret:2023:HDF

[BAM23] Amélie Bourceret, Laurence Amblard, and Jean-Denis Mathias. How do farmers’ environmental preferences influence the efficiency of information instruments for water quality management? Evidence from a social-ecological agent-based model. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000285.

Baumgartner:2020:NCM

[BANG20] Matheus T. Baumgartner, Mário Almeida-Neto, and Luiz C. Gomes. A novel coextinction model considering compensation and new interactions in ecological networks. Ecological Modelling, 416(??):??, January 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303849.

Bajeux:2020:SDE

[BAPW20] Nicolas Bajeux, Julien Arino, Stéphanie Portet, and Richard Westwood. Spread of Dutch elm disease in an urban forest. Ecological Modelling, 438(??):??, December 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002030363X.

Blanco:2020:MFE

[BAR20] Juan A. Blanco, Aitor Ameztegui, and Francisco Rodríguez. Modelling forest ecosystems: a crossroad between scales, techniques and applications. Ecological Modelling, 425(??):??, June 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301022.

Bourceret:2024:MFT

[BAR24] Amélie Bourceret, Francesco Accatino, and Corinne Robert. A modeling framework of a territorial socio-ecosystem to study the trajectories of change in agricultural phytosanitary practices. Ecological Modelling, 494(??):??, August 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001157.

Buonocore:2020:ANC

[BARF20] Elvira Buonocore, Luca Appolloni, Giovanni F. Russo, and Pier Paolo Franzese. Assessing natural capital value in marine ecosystems through an environmental accounting model: a case study in Southern Italy. Ecological Modelling, 419(??):??, March 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020300302.

Brancatelli:2022:DMA

[BAZ22] Gabriela I. E. Brancatelli, Martín R. Amodeo, and Sergio M. Zalba. Demographic model for Aleppo pine invading Argentinean grasslands. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022002447.

Bottero:2023:CRS

[BBCD23] Marta Bottero, Marina Bravi, Caterina Caprioli, and Federico Dell’Anna. Combining revealed and stated preferences to design a new urban park in a metropolitan area of North-Western Italy. Ecological Modelling, 483(??):??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001679.

Boyd:2024:TBA

[BBIP24] Robin J. Boyd, Diana E. Bowler, Nick J. B. Isaac, and Oliver L. Pescott. On the trade-off between accuracy and spatial resolution when estimating species occupancy from geographically biased samples. Ecological Modelling, 493(??): ??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001273.

Boudreault:2022:NLH

[BBSHC22] Jeremie Boudreault, Normand E. Bergeron, Andre St-Hilaire, and Fateh Chebana. A new look at habitat suitability curves through functional data analysis. Ecological Modelling, 467(??):??, May 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002200031X.

Brunetti:2020:MMX

[BCMV20] Matteo Brunetti, Vincenzo Capasso, Matteo Montagna, and Ezio Venturino. A mathematical model for Xylella fastidiosa epidemics in the Mediterranean regions. Promoting good agronomic practices for their effective control. Ecological Modelling, 432(??):??, September 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002030274X.

Beck:2023:MSA

[BCW^{+}23] M. Bruce Beck, Chen Chen, Rodrigo Villarroel Walker, Zongguo Wen, and Jiangxue Han. Multi-sectoral analysis of smarter urban nitrogen metabolism: a case study of Suzhou, China. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000145.

Buonocore:2020:MNC

[BDA^{+}20] Elvira Buonocore, Luigia Donnarumma, Luca Appolloni, Antonino Miccio, Giovanni F. Russo, and Pier Paolo Franzese. Marine natural capital and ecosystem services: an environmental accounting model. Ecological Modelling, 424(??):??, May 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301010.

Benson:2021:ABM

[BdBG^{+}21a] Thomas Benson, Jasper de Bie, Jennifer Gaskell, Paolo Vezza, James R. Kerr, Darren Lumbroso, Markus R. Owen, and Paul S. Kemp. Agent-based modelling of juvenile eel migration via selective tidal stream transport. Ecological Modelling, 443(??):??, March 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000211.

Benson:2021:CAB

[BdBG^{+}21b] Thomas Benson, Jasper de Bie, Jennifer Gaskell, Paolo Vezza, James R. Kerr, Darren Lumbroso, Markus R. Owen, and Paul S. Kemp. Corrigendum to ‘Agent-based modelling of juvenile eel migration via selective tidal stream transport’ [Ecological Modelling 443 (2021) 109448]. Ecological Modelling, 453(??):??, August 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021001678.

Bastazini:2022:REM

[BDC^{+}22] Vinicius Augusto Galvão Bastazini, Vanderlei Debastiani, Laura Cappelatti, Paulo Guimarães, and Valério D. Pillar. The role of evolutionary modes for trait-based cascades in mutualistic networks. Ecological Modelling, 470(??):??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000989.

Barbosa:2023:IMA

[BdCCdSAdS23] Carolina Cerqueira Barbosa, Maria do Carmo Calijuri, Phelipe da Silva Anjinho, and André Cordeiro Alves dos Santos. An integrated modeling approach to predict trophic state changes in a large Brazilian reservoir. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003258.

Barbosa:2023:AMM

[BDP^{+}23] Charles H. X. B. Barbosa, Claudia M. Dias, Dayse H. Pastore, José C. R. Silva, Anna R. C. Costa, Isaac P. Santos, Ramoni Z. S. Azevedo, Raquel M. A. Figueira, and Humberto F. M. Fortunato. Analysis of a mathematical model for golden mussels infestation. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002326.

Barbosa:2023:CFP

[BdSG^{+}23] Lorena Oliveira Barbosa, Juscelina Arcanjo dos Santos, Anny Francielly Ataide Gonçalves, Otávio Camargo Campoe, José Roberto Soares Scolforo, and Henrique Ferraço Scolforo. Competition in forest plantations: Empirical and process-based modelling in pine and eucalypt plantations. Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001412.

Brown:2022:DBD

[BG22] Christian H. Brown and Heather P. Griscom. Differentiating between distribution and suitable habitat in ecological niche models: a red spruce (Picea rubens) case study. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002058.

Bockstaller:2022:MDF

[BGA22] Christian Bockstaller, Victor Galland, and Angel Avadí. Modelling direct field nitrogen emissions using a semimechanistic leaching model newly implemented in Indigo-N v3. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002125.

Benavides:2023:ASD

[BGA^{+}23] Raúl Andrés Molina Benavides, Rómulo Campos Gaona, Alberto Stanislao Atzori, Luisa Fernanda Sánchez, and Hugo Sánchez Guerrero. Application of a system dynamics model to evaluate the implementation of payment for environmental services as a reconversion mechanism in high mountain farming. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001990.

Bhowmick:2023:SMM

[BGC^{+}23] Suman Bhowmick, Jörn Gethmann, Franz J. Conraths, Igor M. Sokolov, and Hartmut H. K. Lentz. SEIR-Metapopulation model of potential spread of West Nile virus. Ecological Modelling, 476(??):??, February 2023. CODEN EC-

MODT. ISSN ???? URL http://www.sciencedirect.com/
science/article/pii/S0304380022003118.

Burman:2024:SEA

[BGC^{+}24] Pramit Kumar Deb Burman, Prajeesh A. G., Supriyo Chakraborty, Yogesh K. Tiwari, Dipankar Sarma, and Nirmali Gogoi. Simulating the ecosystem-atmosphere carbon, water and energy fluxes at a subtropical Indian forest using an ecosystem model. Ecological Modelling, 490(??):??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000267.

Benoit:2021:IIP

[BGCJ21] David M. Benoit, Henrique C. Giacomini, Cindy Chu, and Donald A. Jackson. Identifying influential parameters of a multi-species fish size spectrum model for a northern temperate lake through sensitivity analyses. Ecological Modelling, 460(??):??, November 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002908.

Belsare:2020:ABF

[BGK^{+}20] Aniruddha V. Belsare, Matthew E. Gompper, Barbara Keller, Jason Sumners, Lonnie Hansen, and Joshua J. Millspaugh. An agent-based framework for improving wildlife disease surveillance: a case study of chronic wasting disease in Missouri white-tailed deer. Ecological Modelling, 417(??):??, February 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019304272.

Bellin:2020:MEB

[BGR20] Nicolò Bellin, Maria Groppi, and Valeria Rossi. A model of egg bank dynamics in ephemeral ponds. Ecological Modelling, 430(??):??, August 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301988.

Bauduin:2020:IBM

[BGS^{+}20] Sarah Bauduin, Oksana Grente, Nina Luisa Santostasi, Paolo Ciucci, Christophe Duchamp, and Olivier Gimenez. An individual-based model to explore the impacts of lesserknown social dynamics on wolf populations. Ecological Modelling, 433(??):??, October 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020302799.

Behm:2020:PHN

[BH20] Svenia Behm and Harry Haupt. Predictability of hourly nitrogen dioxide concentration. Ecological Modelling, 428(??):??, July 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301484.

Bello:2021:MEU

[BH21] Al-Amin Danladi Bello and Mohd Ridza Mohd Haniffah. Modelling the effects of urbanization on nutrients pollution for prospective management of a tropical watershed: a case study of Skudai River watershed. Ecological Modelling, 459(??):??, November 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002751.

Behan:2022:SDA

[BHC22] Jamie T. Behan, Cameron T. Hodgdon, and Yong Chen. Scale-dependent assumptions influence habitat suitability estimates for the American lobster (Homarus americanus): Implications for a changing Gulf of Maine. Ecological Modelling, 470(??):??, August 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001405.

Bate:2021:MEC

[BJKT21] Andrew M. Bate, Glyn Jones, Adam Kleczkowski, and Julia Touza. Modelling the effectiveness of collaborative schemes for disease and pest outbreak prevention. Ecological Modelling, 442(??):??, February 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304701.

Bart:2020:IFE

[BKTM20] Ryan R. Bart, Maureen C. Kennedy, Christina L. Tague, and Donald McKenzie. Integrating fire effects on vegetation carbon cycling within an ecohydrologic model. Ecological Modelling, 416(??):??, January 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303886.

Barlett:2024:NDM

[BLA^{+}24] Trinidad Ruiz Barlett, María Fabiana Laguna, Guillermo Abramson, Adrian Monjeau, and Gabriel Martin. A new distributional model coupling environmental and biotic factors. Ecological Modelling, 489(??):??, March 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002300340X.

Briegel:2020:FCL

[BLB^{+}20] Ferdinand Briegel, Sung Ching Lee, T. Andrew Black, Rachhpal S. Jassal, and Andreas Christen. Factors controlling longterm carbon dioxide exchange between a Douglas-fir stand and the atmosphere identified using an artificial neural network approach. Ecological Modelling, 435(??):??, November 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020303367.

Barrere:2024:ORF

[BLB^{+}24a] Julien Barrere, Gauthier Ligot, Vincent Boulanger, Catherine Collet, Benoît Courbaud, François de Coligny, Anders Mårell, Sonia Saïd, and Philippe Balandier. Oak regeneration facing deer browsing: Can competition between saplings offset the diversion effect? A simulation experiment. Ecological Modelling, 489(??):??, March 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003381.

Benard:2024:IAM

[BLB24b] Annaëlle Bénard, Thierry Lengagne, and Christophe Bonenfant. Integration of animal movement into wildlife–vehicle collision models. Ecological Modelling, 492(??):??, June 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000784.

Bergkvist:2023:MMF

[BLL^{+}23] John Bergkvist, Fredrik Lagergren, Maj-Lena Finnander Linderson, Paul Miller, Mats Lindeskog, and Anna Maria Jönsson. Modelling managed forest ecosystems in Sweden: an evaluation from the stand to the regional scale. Ecological Modelling, 477(??):??, March 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003519.

Bellido-Leiva:2021:MEH

[BLLL21] F. J. Bellido-Leiva, Robert A. Lusardi, and Jay R. Lund. Modeling the effect of habitat availability and quality on endangered winter-run Chinook salmon (Oncorhynchus tshawytscha) production in the Sacramento Valley. Ecological Modelling, 447(??):??, May 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002100082X.

Binny:2020:LGS

[BLP20] Rachelle N. Binny, Richard Law, and Michael J. Plank. Living in groups: Spatial-moment dynamics with neighbourbiased movements. Ecological Modelling, 415(??):??, January 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019303333.

Bracis:2020:ICC

[BLSR^{+}20] Chloe Bracis, Sigrid Lehuta, Marie Savina-Rolland, Morgane Travers-Trolet, and Raphaël Girardin. Improving confidence in complex ecosystem models: the sensitivity analysis of an Atlantis ecosystem model. Ecological Modelling, 431(??):??, September 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020302052.

Brias:2021:EBM

[BM21] Antoine Brias and Stephan B. Munch. Ecosystem based multi-species management using empirical dynamic programming. Ecological Modelling, 441(??):??, February 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304804.

Barker:2022:SDMb

[BM22a] Justin R. Barker and Hugh J. MacIsaac. Species distribution models: Administrative boundary centroid occurrences require careful interpretation. Ecological Modelling, 472(??):??, October 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002101.

Barker:2022:SDMa

[BM22b] Justin R. Barker and Hugh J. MacIsaac. Species distribution models applied to mosquitoes: Use, quality assessment, and recommendations for best practice. Ecological Modelling, 472(??):??, October 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001806.

Belenguer-Manzanedo:2023:MSA

[BMAME^{+}23] María Belenguer-Manzanedo, Carles Alcaraz, Maite Martínez-Eixarch, Antonio Camacho, James T. Morris, and Carles Ibáñez. Modeling soil accretion and carbon accumulation in deltaic rice fields. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001862.

Bahlburg:2021:ISR

[BMB21] Dominik Bahlburg, Bettina Meyer, and Uta Berger. The impact of seasonal regulation of metabolism on the life history of Antarctic krill. Ecological Modelling, 442(??):??, February 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021000028.

Barry:2024:GMB

[BMB24] Ronald P. Barry, Julie McIntyre, and Jordan Bernard. A geostatistical model based on random walks to krige regions with irregular boundaries and holes. Ecological Modelling, 491(??): ??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000541.

Biswas:2024:BEE

[BMBB24] Sukdev Biswas, Sk Golam Mortoja, Ritesh Kumar Bera, and Sabyasachi Bhattacharya. Bacteria as ecosystem engineers: Unraveling clues through a novel functional response and tritrophic model. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002910.

Besbeas:2022:SAS

[BMM22] P. T. Besbeas, R. S. McCrea, and B. J. T. Morgan. Selecting age structure in integrated population models. Ecological Modelling, 473(??):??, November 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002137.

Bezerra:2021:BCC

[BMRF21] João I. M. Bezerra, Alexandre Molter, Marat Rafikov, and Daiane F. Frighetto. Biological control of the chaotic sugarcane borer-parasitoid agroecosystem. Ecological Modelling, 450(??):??, June 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021001332.

Boisson:2020:END

[BMS^{+}20] Sylvain Boisson, Arnaud Monty, Maxime Séleck, Mylor Ngoy Shutcha, Michel-Pierre Faucon, and Grégory Mahy. Ecological niche distribution along soil toxicity gradients: Bridging theoretical expectations and metallophyte conservation. Ecological Modelling, 415(??):??, January 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303692.

Bossier:2021:IEI

[BNAR^{+}21] Sieme Bossier, J. Rasmus Nielsen, Elin Almroth-Rosell, Anders Höglund, Francois Bastardie, Stefan Neuenfeldt, Iréne Wåhlström, and Asbjørn Christensen. Integrated ecosystem impacts of climate change and eutrophication on main Baltic fishery resources. Ecological Modelling, 453(??):??, August 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002100171X.

Bossier:2020:ETI

[BNN20] Sieme Bossier, J. Rasmus Nielsen, and Stefan Neuenfeldt. Exploring trophic interactions and cascades in the Baltic Sea using a complex end-to-end ecosystem model with extensive food web integration. Ecological Modelling, 436(??):??, November 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303513.

Barber-OMalley:2022:HHS

[BOLC^{+}22] Betsy Barber-O’Malley, Géraldine Lassalle, Guillem Chust, Estibaliz Diaz, Andrew O’Malley, César Paradinas Blázquez, Javier Pórtoles Marquina, and Patrick Lambert. Hy-DiaD: a hybrid species distribution model combining dispersal, multi-habitat suitability, and population dynamics for diadromous species under climate change scenarios. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001119.

Bila:2021:MCE

[BP21] Jiri Bila and Jan Pokorny. Modeling of complex ecosystems, extension of deserts and violation of the short water cycle. Ecological Modelling, 461(??):??, December 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003112.

Beauregard:2020:OCA

[BPB^{+}20] Matthew A. Beauregard, Rana D. Parshad, Sarah Boon, Harley Conaway, Thomas Griffin, and Jingjing Lyu. Optimal control and analysis of a modified Trojan Y-chromosome strategy. Ecological Modelling, 416(??):??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001930362X.

Bakshi:2024:USE

[BPF24] Baishali Bakshi, Stephen Polasky, and Lee E. Frelich. Using structural equation models (SEM) to link climate change, forest composition, deer, and outdoor recreation. Ecological Modelling, 493(??):??, July 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024001194.

Bavcevic:2020:EFE

[BPK^{+}20] Lav Bavčević, Siniša Petrović, Vatroslav Karamarko, Umberto Luzzana, and Tin Klanjšček. Estimating fish energy content and gain from length and wet weight. Ecological Modelling, 436(??):??, November 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303501.

Beaulieu:2020:RFC

[BPL^{+}20] Emilie Beaulieu, Marie-Claire Pierret, Arnaud Legout, François Chabaux, Yves Goddéris, Daniel Viville, and Agnès Herrmann. Response of a forested catchment over the last 25 years to past acid deposition assessed by biogeochemical cycle modeling (Strengbach, France). Ecological Modelling, 430(??):??, August 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301964.

Bathmann:2020:MGF

[BPN^{+}20] Jasper Bathmann, Ronny Peters, Dmitri Naumov, Thomas Fischer, Uta Berger, and Marc Walther. The MANgrove– GroundwAter feedback model (MANGA) — describing belowground competition based on first principles. Ecological Modelling, 420(??):??, March 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300454.

Barton:2021:FIB

[BPW^{+}21] Madeleine Barton, Hazel Parry, Samantha Ward, Ary A. Hoffmann, Paul A. Umina, Maarten van Helden, and Sarina Macfadyen. Forecasting impacts of biological control under future climates: mechanistic modelling of an aphid pest and a parasitic wasp. Ecological Modelling, 457(??):??, October 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002374.

Byer:2022:EIP

[BR22] Nathan W. Byer and Brendan N. Reid. The emergence of imperfect philopatry and fidelity in spatially and temporally heterogeneous environments. Ecological Modelling, 468(??): ??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000850.

Bruzzone:2023:NMF

[BRAL23] Octavio A. Bruzzone, Luca Rossini, M. Belén Aguirre, and Guillermo Logarzo. A new model formulation for host depletion in parasitoids. Ecological Modelling, 475 (??):??, January 2023. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S030438002200312X.

Boyer:2022:PCM

[BRBT22] Jessica Boyer, Kasondra Rubalcava, Shawn Booth, and Howard Townsend. Proof-of-concept model for exploring the impacts of microplastics accumulation in the Maryland coastal bays ecosystem. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003872.

Bazzato:2021:HSR

[BRC^{+}21] Erika Bazzato, Leonardo Rosati, Simona Canu, Michele Fiori, Emmanuele Farris, and Michela Marignani. High spatial resolution bioclimatic variables to support ecological modelling in a Mediterranean biodiversity hotspot. Ecological Modelling, 441(??):??, February 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304208.

Bondo:2024:CRC

[BRSW24] Kristin J. Bondo, Christopher S. Rosenberry, David Stainbrook, and W. David Walter. Comparing risk of chronic wasting disease occurrence using Bayesian hierarchical spatial models and different surveillance types. Ecological Modelling, 493(??):??, July 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024001443.

Bobrowski:2020:CWI

[BS20] Maria Bobrowski and Udo Schickhoff. Corrigendum to “Why input matters: Selection of climate data sets for modelling the potential distribution of a treeline species in the Himalayan region” [Ecol. Modell. 359 (2017) 92–102]. Ecological Modelling, 418(??):??, February 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002030017X.

Bae:2021:CAB

[BS21] Sunim Bae and Dongil Seo. Changes in algal bloom dynamics in a regulated large river in response to eutrophic status. Ecological Modelling, 454(??):??, August 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001551.

Benkendorf:2023:CEC

[BSCH23] Donald J. Benkendorf, Samuel D. Schwartz, D. Richard Cutler, and Charles P. Hawkins. Correcting for the effects of class imbalance improves the performance of machine-learning based species distribution models. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002300145X.

Bellier:2021:SSH

[BSE21] Edwige Bellier, Bernt-Erik Sæther, and Steinar Engen. Sustainable strategies for harvesting predators and prey in a fluctuating environment. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020304166.

Brunetti:2023:SMB

[BSM23] Ilaria Brunetti, Rodolphe Sabatier, and Lauriane Mouysset. A spatial model for biodiversity offsetting. Ecological Modelling, 481(??):??, July 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000923.

Bahi:2023:PDMa

[BSPT23a] Aya Bahi, Sabine Sauvage, Sylvain Payraudeau, and Julien Tournebize. PESTIPOND: a descriptive model of pesticide fate in artificial ponds: I. Model development. Ecological Modelling, 485(??):??, November 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001849.

Bahi:2023:PDMb

[BSPT23b] Aya Bahi, Sabine Sauvage, Sylvain Payraudeau, and Julien Tournebize. PESTIPOND: a descriptive model of pesticide fate in artificial ponds: II. Model application and evaluation. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002028.

Blanchard:2021:SIT

[BSR21] Jesse R. Blanchard, Rolando O. Santos, and Jennifer S. Rehage. Sociability interacts with temporal environmental variation to spatially structure metapopulations: a fish dispersal simulation in an ephemeral landscape. Ecological Modelling, 443(??):??, March 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000302.

Baey:2023:CBB

[BSR^{+}23] Charlotte Baey, Henrik G. Smith, Maj Rundlöf, Ola Olsson, Yann Clough, and Ullrika Sahlin. Calibration of a bumble bee foraging model using Approximate Bayesian Computation. Ecological Modelling, 477(??):??, March 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022003490.

Badouard:2024:LPS

[BSS^{+}24] Vincyane Badouard, Sylvain Schmitt, Guillaume Salzet, Thomas Gaquiere, Margaux Rojat, Caroline Bedeau, Olivier Brunaux, and Géraldine Derroire. LoggingLab: an R package to simulate reduced-impact selective logging in tropical forests using forest inventory data. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002697.

Bona:2020:CMP

[BST^{+}20] Kelly Ann Bona, Cindy Shaw, Dan K. Thompson, Oleksandra Hararuk, Kara Webster, Gary Zhang, Mihai Voicu, and Werner A. Kurz. The Canadian model for peatlands (CaMP): a peatland carbon model for national greenhouse gas reporting. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020302350.

Baudrot:2023:MPR

[BSU^{+}23] Virgile Baudrot, Rafael Schouten, Paul A. Umina, Ary A. Hoffmann, Lisa Bird, Melina Miles, and James L. Maino. Managing pesticide resistance in Spodoptera frugiperda: a spatially explicit framework for identifying optimal treatment strategies. Ecological Modelling, 483(??):??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001473.

Butts:2022:DDA

[BTC^{+}22] David J. Butts, Noelle E. Thompson, Sonja A. Christensen, David M. Williams, and Michael S. Murillo. Data-driven agent-based model building for animal movement through Exploratory Data Analysis. Ecological Modelling, 470(??):??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001132.

Bandara:2021:TCF

[BVM^{+}21] Kanchana Bandara, Øystein Varpe, Frédéric Maps, Rubao Ji, Ketil Eiane, and Vigdis Tverberg. Timing of Calanus finmarchicus diapause in stochastic environments. Ecological Modelling, 460(??):??, November 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002891.

Bruneel:2020:CDD

[BVR^{+}20] Stijn Bruneel, Pieterjan Verhelst, Jan Reubens, Stijn Luca, Johan Coeck, Tom Moens, and Peter Goethals. Combining disciplines: Dealing with observed and cryptic animal residencies in passive telemetry data by applying econometric decision-making models. Ecological Modelling, 438(??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020304075.

Bourdouxhe:2023:MPN

[BWCD23] Axel Bourdouxhe, Lionel Wibail, Hugues Claessens, and Marc Dufrêne. Modeling potential natural vegetation: a new light on an old concept to guide nature conservation in fragmented and degraded landscapes. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001138.

Booth:2020:PEP

[BWS^{+}20] Shawn Booth, William J. Walters, Jeroen Steenbeek, Villy Christensen, and Sabine Charmasson. An Ecopath with Ecosim model for the Pacific coast of eastern Japan: Describing the marine environment and its fisheries prior to the Great East Japan earthquake. Ecological Modelling, 428(??):??, July 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020301599.

Bobrowski:2021:SES

[BWSS21] Maria Bobrowski, Johannes Weidinger, Niels Schwab, and Udo Schickhoff. Searching for ecology in species distribution models in the Himalayas. Ecological Modelling, 458(??):??, October 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021002490.

Bona:2024:UCM

[BWT^{+}24] Kelly A. Bona, Kara L. Webster, Dan K. Thompson, Oleksandra Hararuk, Gary Zhang, and Werner A. Kurz. Using the Canadian Model for Peatlands (CaMP) to examine greenhouse gas emissions and carbon sink strength in Canada’s boreal and temperate peatlands. Ecological Modelling, 490(??): ??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002400022X.

Bai:2022:ARR

[BZZT22] Jing Bai, Jian Zhao, Zhenyu Zhang, and Ziqiang Tian. Assessment and a review of research on surface water quality modeling. Ecological Modelling, 466(??):??, April 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000151.

Croft:2021:SEP

[CAB^{+}21] S. Croft, J. N. Aegerter, S. Beatham, J. Coats, and G. Massei. A spatially explicit population model to compare management using culling and fertility control to reduce numbers of grey squirrels. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304506.

Capera-Aragones:2022:NRW

[CAFT22] Pau Capera-Aragones, Eric Foxall, and Rebecca C. Tyson. Nutritionally rich wildflower patches adjacent to nutritionally deficient crops significantly increase pollination services. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000588.

Clark:2022:QAD

[CAH22] Matt Clark, Jeffrey Andrews, and Vicken Hillis. A quantitative application of diffusion of innovations for modeling the spread of conservation behaviors. Ecological Modelling, 473(??):??, November 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002460.

Cho:2023:MLM

[CAK^{+}23] Nanghyun Cho, Casimir Agossou, Eunsook Kim, Jong-Hwan Lim, Jeong-Wook Seo, and Sinkyu Kang. Machine-learning modeling on tree mortality and growth reduction of temperate forests with climatic and ecophysiological parameters. Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001874.

Cho:2024:EKC

[CAK^{+}24] Nanghyun Cho, Casimir Agossou, Eunsook Kim, Jong-Hwan Lim, Taehee Hwang, and Sinkyu Kang. Evaluating key climatic and ecophysiological parameters of worldwide tree mortality with a process-based BGC model and machine learning algorithms. Ecological Modelling, 491(??):??, May 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000760.

Clasen:2024:UPM

[CAS^{+}24] Arno P. Clasen, Feni Agostinho, Federico Sulis, Cecília M. V. B. Almeida, and Biagio F. Giannetti. Unlocking the potential of municipal solid waste: Emergy accounting applied in a novel biorefinery. Ecological Modelling, 492(??):??, June 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001133.

Carriger:2020:BNA

[CB20] John F. Carriger and Mace G. Barron. A Bayesian network approach to refining ecological risk assessments: Mercury and the Florida panther (Puma concolor coryi). Ecological Modelling, 418(??):??, February 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019304193.

Caracappa:2022:NUS

[CBG^{+}22] Joseph C. Caracappa, Andrew Beet, Sarah Gaichas, Robert J. Gamble, Kimberly J. W. Hyde, Scott I. Large, Ryan E. Morse, Charles A. Stock, and Vincent S. Saba. A northeast United States Atlantis marine ecosystem model with ocean reanalysis and ocean color forcing. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002200148X.

Chattopadhyay:2023:TCB

[CBSB23] Arnab Chattopadhyay, Swarnendu Banerjee, Amit Samadder, and Sabyasachi Bhattacharya. Is toxicity a curse or blessing, or both? — searching answer from a disease-induced consumer-resource system. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002648.

Cepic:2022:MHI

[CBW22] Michael Cepic, Ulrike Bechtold, and Harald Wilfing. Modelling human influences on biodiversity at a global scale — a human ecology perspective. Ecological Modelling, 465(??):??, March 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021003926.

Cones:2020:OMI

[CC20] Alexandra G. Cones and Philip H. Crowley. Optimal maternal incubation strategies for altricial and precocial birds. Ecological Modelling, 436(??):??, November 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303604.

Cabrera:2023:PMR

[CC23] Juan M. Cabrera and Julieta Capeletti. metrix: an R package for assessing the biological quality of water. Ecological Modelling, 484(??):??, October 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002300203X.

Cabaneros:2020:SEO

[CCH20] Sheen Mclean Cabaneros, John Kaiser Calautit, and Ben Hughes. Spatial estimation of outdoor NO_{2} levels in Central London using deep neural networks and a wavelet decomposition technique. Ecological Modelling, 424(??):??, May 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020300892.

Che-Castaldo:2020:CPP

[CCJK^{+}20] Judy Che-Castaldo, Owen R. Jones, Bruce E. Kendall, Jean H. Burns, Dylan Z. Childs, Thomas H. G. Ezard, Haydee Hernandez-Yanez, David J. Hodgson, Eelke Jongejans, Tiffany Knight, Cory Merow, Satu Ramula, Iain Stott, Yngvild Vindenes, Hiroyuki Yokomizo, and Roberto Salguero-Gómez. Comments to “Persistent problems in the construction of matrix population models”. Ecological Modelling, 416(??):??, January 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019304211.

Chimienti:2020:ECC

[CDB^{+}20] Marianna Chimienti, Jean-Pierre Desforges, Larissa T. Beumer, Jacob Nabe-Nielsen, Floris M. van Beest, and Niels Martin Schmidt. Energetics as common currency for integrating high resolution activity patterns into dynamic energy budget-individual based models. Ecological Modelling, 434(??):??, October 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020303203.

Crouse:2022:LTR

[CDC^{+}22] Kristin N. Crouse, Nisarg P. Desai, Kira A. Cassidy, Erin E. Stahler, Clarence L. Lehman, and Michael L. Wilson. Larger territories reduce mortality risk for chimpanzees, wolves, and agents: Multiple lines of evidence in a model validation framework. Ecological Modelling, 471(??):??, September 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001715.

Castex:2020:ATG

[CdCAC^{+}20] V. Castex, I. García de Cortázar-Atauri, P. Calanca, M. Beniston, and J. Moreau. Assembling and testing a generic phenological model to predict Lobesia botrana voltinism for impact studies. Ecological Modelling, 420(??):??, March 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020300181.

Chang:2022:GPM

[CDJ^{+}22] Manqi Chang, Donald L. DeAngelis, Jan H. Janse, Annette B. G. Janssen, Tineke A. Troost, Dianneke van Wijk, Wolf M. Mooij, and Sven Teurlincx. A generically parameterized model of lake eutrophication: the impact of stoichiometric ratios and constraints on the abundance of natural phytoplankton communities (GPLake-S). Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002435.

Chandio:2023:MPT

[CDN^{+}23] Abbas Ali Chandio, Devi Prasad Dash, Solomon Prince Nathaniel, Ghulam Raza Sargani, and Yuansheng Jiang. Mitigation pathways towards climate change: Modelling the impact of climatological factors on wheat production in top six regions of China. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001126.

Chudzinska:2020:CSA

[CDNN^{+}20] Magda Chudzinska, Yoko L. Dupont, Jacob Nabe-Nielsen, Kate P. Maia, Marie V. Henriksen, Claus Rasmussen, W. Daniel Kissling, Melanie Hagen, and Kristian Trøjelsgaard. Combining the strengths of agent-based modelling and network statistics to understand animal movement and interactions with resources: example from within-patch foraging decisions of bumblebees. Ecological Modelling, 430(??):??, August 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301915.

Clavareau:2023:EDR

[CDT^{+}23] Lyndsay Clavareau, Jeffrey M. Dambacher, Verena M. Trenkel, Sophie Gourguet, Paul Tixier, and Martin P. Marzloff. Exploitation and depredation rates determine viability of depredation-impacted fisheries. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001758.

Cruz:2022:UDF

[CFG^{+}22] Mariana Hill Cruz, Ivy Frenger, Julia Getzlaff, Iris Kriest, Tianfei Xue, and Yunne-Jai Shin. Understanding the drivers of fish variability in an end-to-end model of the Northern Humboldt Current System. Ecological Modelling, 472(??):??, October 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002009.

Chen:2022:FON

[CFL^{+}22] Zhuo Chen, Jingjie Feng, Ran Li, Yuanming Wang, Fangjun Peng, and Kefeng Li. Field observation and numerical modelling of supersaturated dissolved gas at river confluence. Ecological Modelling, 471(??):??, September 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001284.

Cauwelaert:2023:DPA

[CGB^{+}23] Emilio Mora Van Cauwelaert, Cecilia González González, Denis Boyer, Zachary Hajian-Forooshani, John Vandermeer, and Mariana Benítez. Dispersal and plant arrangement condition the timing and magnitude of coffee rust infection. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003040.

Calhoun-Grosch:2024:SPC

[CGRR^{+}24] Stacy Calhoun-Grosch, Jim J. Ruzicka, Kelly L. Robinson, Verena H. Wang, Tracey Sutton, Cameron Ainsworth, and Frank Hernandez. Simulating productivity changes of epipelagic, mesopelagic, and bathypelagic taxa using a depthresolved, end-to-end food web model for the oceanic Gulf of Mexico. Ecological Modelling, 489(??):??, March 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000127.

Consolo:2022:DVP

[CGV22] Giancarlo Consolo, Gabriele Grifó, and Giovanna Valenti. Dryland vegetation pattern dynamics driven by inertial effects and secondary seed dispersal. Ecological Modelling, 474 (??):??, December 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380022002721.

Colas:2021:SCC

[CGVC21] Floriane Colas, Jean-Pierre Gauchi, Jean Villerd, and Nathalie Colbach. Simplifying a complex computer model: Sensitivity analysis and metamodelling of an 3D individualbased crop-weed canopy model. Ecological Modelling, 454 (??):??, August 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021001691.

Cacho:2022:MBI

[CH22] Oscar J. Cacho and Susan M. Hester. Modelling biocontrol of invasive insects: an application to European wasp (Vespula germanica) in Australia. Ecological Modelling, 467(??):??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000618.

Chen:2021:SER

[Che21] XinJian Chen. A study of effects of reduction of submarine groundwater discharge on thermal habitats for manatee in a spring-fed estuary using a laterally averaged hydrodynamic model. Ecological Modelling, 456(??):??, September 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002131.

Chen:2022:TDA

[Che22] Bingzhang Chen. Thermal diversity affects community responses to warming. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003847.

Chapagain:2023:AEU

[CHR^{+}23] Ranju Chapagain, Neil Huth, Tomas A. Remenyi, Caroline L. Mohammed, and Jonathan J. Ojeda. Assessing the effect of using different APSIM model configurations on model outputs. Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001825.

Citores:2020:MSP

[CIL^{+}20] L. Citores, L. Ibaibarriaga, D.-J. Lee, M. J. Brewer, M. Santos, and G. Chust. Modelling species presence– absence in the ecological niche theory framework using shape-constrained generalized additive models. Ecological Modelling, 418(??):??, February 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438001930434X.

Chen:2023:DDH

[CJM^{+}23] Qiting Chen, Li Jia, Massimo Menenti, Guangcheng Hu, Kun Wang, Zhiwei Yi, Jie Zhou, Fei Peng, Shaoxiu Ma, Quangang You, Xiaojie Chen, and Xian Xue. A data-driven high spatial resolution model of biomass accumulation and crop yield: Application to a fragmented desert-oasis agroecosystem. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022002800.

Chang:2020:WIG

[CKC20] Feng-Hsun Chang, Po-Ju Ke, and Bradley Cardinale. Weak intra-guild predation facilitates consumer coexistence but does not guarantee higher consumer density. Ecological Modelling, 424(??):??, May 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300910.

Cushman:2024:CPG

[CKC^{+}24] S. A. Cushman, K. Kilshaw, R. D. Campbell, Z. Kaszta, M. Gaywood, and D. W. Macdonald. Comparing the performance of global, geographically weighted and ecologically weighted species distribution models for Scottish wildcats using GLM and Random Forest predictive modeling. Ecological Modelling, 492(??):??, June 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000796.

Cushman:2024:EII

[CKK^{+}24a] S. A. Cushman, K. Kilshaw, Z. Kaszta, R. D. Campbell, M. Gaywood, and D. W. Macdonald. Explaining interindividual differences in habitat relationships among wildcat hybrids in Scotland. Ecological Modelling, 491(??):??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024000450.

Cushman:2024:ENL

[CKK^{+}24b] S. A. Cushman, K. Kilshaw, Z. Kaszta, R. D. Campbell, M. Gaywood, and D. W. Macdonald. Exploring nonstationary limiting factors in species habitat relationships. Ecological Modelling, 490(??):??, April 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000516.

Cushman:2024:VIS

[CKK^{+}24c] S. A. Cushman, K. Kilshaw, Z. Kaszta, R. D. Campbell, M. Gaywood, and D. W. Macdonald. Variable importance and scale of influence across individual Scottish wildcat hybrid habitat models. Ecological Modelling, 491(??):??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024000863.

Chen:2021:SCE

[CL21] Chengjing Chen and Yihua Liu. Spatiotemporal changes of ecosystem services value by incorporating planning policies: a case of the Pearl River Delta, China. Ecological Modelling, 461(??):??, December 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003252.

Choi:2024:BDU

[CL24] Jang-Geun Choi and Thomas C. Lippmann. Biogeochemical dynamics underlying equilibrium between nitrogen fixation and denitrification and its impact on a coastal marine ecosystem model. Ecological Modelling, 494(??):??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001558.

Cui:2020:ESC

[CLG20] Xuezhu Cui, Shaoying Li, and Feng Gao. Examining spatial carbon metabolism: Features, future simulation, and landbased mitigation. Ecological Modelling, 438(??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002030394X.

Choi:2023:APD

[CLH23] Jang-Geun Choi, Thomas C. Lippmann, and Elizabeth L. Harvey. Analytical population dynamics underlying harmful algal blooms triggered by prey avoidance. Ecological Modelling, 481(??):??, July 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000947.

Chen:2022:HOI

[CLHZ22] Shiliang Chen, Xiang Liu, Qiang He, and Shurong Zhou. Higher-order interactions on disease transmission can reverse the dilution effect or weaken the amplification effect to unimodal pattern. Ecological Modelling, 474(??):??, December 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022002575.

Chen:2020:SDP

[CLPZ20] Lifan Chen, Xiang Liu, Zechen Peng, and Shurong Zhou. Species distribution patterns and the scale of host interactions quantitatively but not qualitatively affect the diversity– disease relationship. Ecological Modelling, 435(??):??, November 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303380.

Chinn:2023:APF

[CLW23] Sarah M. Chinn, Glen E. Liston, and Ryan R. Wilson. Assessing past and future climatic influences on the availability of polar bear maternal denning habitat on Wrangel Island. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002090.

Chen:2024:ILF

[CLW^{+}24] Bin Chen, Yue Li, Shaoqiang Wang, Jinghua Chen, Xuanze Zhang, Zhenhai Liu, and Holly Croft. Integrating leaf functional traits improves modelled estimates of carbon and water fluxes at a subtropical evergreen conifer forest. Ecological Modelling, 488(??):??, February 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002300323X.

Chen:2023:CEU

[CLY^{+}23] Yu Chen, Gengyuan Liu, Ningyu Yan, Qing Yang, He Gao, Liya Su, and Remo Santagata. Comprehensive evaluation of urban greenspace ecological values marketability through the spatial relationship between housing price and ecosystem services. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002120.

Callesen:2021:TTP

[CM21] I. Callesen and A. Magnussen. TransparC2U — a twopool, pedology oriented forest soil carbon simulation model aimed at user investigations of multiple uncertainties. Ecological Modelling, 453(??):??, August 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021001654.

Chu:2023:CBP

[CMA23] Tianyi Chu, Holly V. Moeller, and Kevin M. Archibald. Competition between phytoplankton and mixotrophs leads to metabolic character displacement. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380023000595.

Caracappa:2023:IBD

[CMFC23] Joseph C. Caracappa, Daphne M. Munroe, Heidi L. Fuchs, and Robert J. Chant. Influences of brood-dependent behavioral variation on blue crab (Callinectes sapidus) larval transport in a wind-driven estuarine plume. Ecological Modelling, 478(??):??, April 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000236.

Chevallier:2020:MLB

[CMG20] Damien Chevallier, Baptiste Mourrain, and Marc Girondot. Modelling leatherback biphasic indeterminate growth using a modified Gompertz equation. Ecological Modelling, 426 (??):??, June 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020301095.

Caill-Milly:2022:ADS

[CMGD^{+}22] N. Caill-Milly, J. Bald Garmendia, F. D’Amico, O. Guyader, C. Dang, and N. Bru. Adapting a dynamic system model using life traits and local fishery knowledge — application to a population of exploited marine bivalves (Ruditapes philippinarum) in a mesotidal coastal lagoon. Ecological Modelling, 470(??): ??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001442.

Chen:2023:CCP

[CMHZ23] Qiuwen Chen, Xiaohan Ma, Jiayu Hu, and Xiaohong Zhang. Comparison of comprehensive performance of kiwifruit production in China, Iran, and Italy based on emergy and carbon emissions. Ecological Modelling, 483(??):??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001709.

Chakraborty:2022:IHA

[CMKF22] Subhendu Chakraborty, Stefanie D. Moorthi, Rajat Karnatak, and Ulrike Feudel. Irregular harmful algal blooms triggered by feedback between toxin production and zooplankton feeding. Ecological Modelling, 473(??):??, November 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002216.

Colomer:2021:CMA

[CMS^{+}21] Ma. Àngels Colomer, Antoni Margalida, Isabel Sanuy, Gustavo A. Llorente, Delfí Sanuy, and Eudald Pujol-Buxó. A computational model approach to assess the effect of climate change on the growth and development of tadpoles. Ecological Modelling, 461(??):??, December 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003124.

Campos:2022:SDF

[CMSD^{+}22] Jean C. Campos, José Manrique-Silupú, Bogdan Dorneanu, William Ipanaqué, and Harvey Arellano-García. A smart decision framework for the prediction of thrips incidence in organic banana crops. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022002484.

Cropp:2020:PCR

[CN20] Roger Cropp and John Norbury. The potential for coral reefs to adapt to a changing climate — an eco-evolutionary modelling perspective. Ecological Modelling, 426(??):??, June 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020301101.

cNie:2024:IUS

[cNKF24] Pierce M. c Nie, Daniel Kneeshaw, and Élise Filotas. Increased unpredictability in spruce budworm outbreaks following habitat loss and landscape fragmentation. Ecological Modelling, 491(??):??, May 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000632.

Chudzinska:2021:AAB

[CNNS^{+}21] Magda Chudzinska, Jacob Nabe-Nielsen, Sophie Smout, Geert Aarts, Sophie Brasseur, Isla Graham, Paul Thompson, and Bernie McConnell. AgentSeal: Agent-based model describing movement of marine central-place foragers. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304610.

Choden:2021:PIC

[CNSK21] Kunzang Choden, Craig R. Nitschke, Stephen B. Stewart, and Rodney J. Keenan. The potential impacts of climate change on the distribution of key tree species and Cordyceps in Bhutan: Implications for ecological functions and rural livelihoods. Ecological Modelling, 455(??):??, September 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021002106.

Chandio:2023:MIC

[COJ23] Abbas Ali Chandio, Dicle Ozdemir, and Yuansheng Jiang. Modelling the impact of climate change and advanced agricultural technologies on grain output: Recent evidence from China. Ecological Modelling, 485(??):??, November 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002314.

Chhaytle:2023:PVP

[COPM23] Mohamad Chhaytle, Régis Ouvrard, Thierry Poinot, and Lauriane Mouysset. Parameter-varying partial differential equation to model the global change impacts on wildlife populations. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002466.

Coro:2020:GSE

[Cor20] Gianpaolo Coro. A global-scale ecological niche model to predict SARS-CoV-2 coronavirus infection rate. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302581. See comment [CYP^{+}20].

Chasia:2023:MLU

[COS23] Stanley Chasia, Luke O. Olang, and Lewis Sitoki. Modelling of land-use/cover change trajectories in a transboundary catchment of the Sio-Malaba-Malakisi Region in East Africa using the CLUE-s model. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT. ISSN

????  URL http://www.sciencedirect.com/science/
article/pii/S0304380022003544.

Coste:2020:AMP

[CP20] Christophe F. D. Coste and Samuel Pavard. Analysis of a multitrait population projection matrix reveals the evolutionary and demographic effects of a life history trade-off. Ecological Modelling, 418(??):??, February 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019304235.

Canelas:2022:ILU

[CP22] Joana Viana Canelas and Henrique Miguel Pereira. Impacts of land-use intensity on ecosystems stability. Ecological Modelling, 472(??):??, October 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001971.

Cisneros-Pineda:2020:ICH

[CPAT20] Alfredo Cisneros-Pineda, David Aadland, and John Tschirhart. Impacts of cattle, hunting, and natural gas development in a rangeland ecosystem. Ecological Modelling, 431(??): ??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020302453.

Coppee:2022:TTS

[CPTD22] Thomas Coppée, Jean-Yves Paquet, Nicolas Titeux, and Marc Dufrêne. Temporal transferability of species abundance models to study the changes of breeding bird species based on land cover changes. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002200237X.

Cagle:2024:CMD

[CR24] Sierra E. Cagle and Daniel L. Roelke. Chaotic mixotroph dynamics arise with nutrient loading: Implications for mixotrophy as a harmful bloom forming mechanism. Ecological Modelling, 492(??):??, June 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024001029.

Catron:2023:MIT

[CRZ^{+}23] Spencer Catron, Sarah Roth, Francesca Zumpano, Jason Bintz, James A. Fordyce, Suzanne Lenhart, Debra L. Miller, and Jeanette Wyneken. Modeling the impacts of temperature during nesting seasons on loggerhead (Caretta caretta) sea turtle populations in South Florida. Ecological Modelling, 481(??):??, July 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000911.

Countryman:2022:MEC

[CSB22] C. E. Countryman, D. K. Steinberg, and A. B. Burd. Modelling the effects of copepod diel vertical migration and community structure on ocean carbon flux using an agent-based model. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001156.

Caradima:2021:BMC

[CSBS21] Bogdan Caradima, Andreas Scheidegger, Jakob Brodersen, and Nele Schuwirth. Bridging mechanistic conceptual models and statistical species distribution models of riverine fish. Ecological Modelling, 457(??):??, October 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002386.

Carturan:2023:BBP

[CSCT23] Bruno S. Carturan, Nourridine Siewe, Christina A. Cobbold, and Rebecca C. Tyson. Bumble bee pollination and the wildflower/crop trade-off: When do wildflower enhancements improve crop yield? Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001783.

Chen:2023:PPC

[CSM23] Jun Chen, Masae Shiyomi, and Satoru Morita. Phytomass productivity of cutting and grazing grasslands with special reference to small-scale spatial variation in plant nutrient resources. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002417.

Crevier:2021:MBP

[CSMP21] Lucas Phillip Crevier, Joseph H. Salkeld, Jessa Marley, and Lael Parrott. Making the best possible choice: Using agentbased modelling to inform wildlife management in small communities. Ecological Modelling, 446(??):??, April 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000764.

Canete-Salinas:2020:EMD

[CSZY^{+}20] Paulo Cañete-Salinas, Francisco Zamudio, Marco Yáñez, John Gajardo, Héctor Valdés, Cristian Espinosa, Jaime Venegas, Luis Retamal, Samuel Ortega-Farias, and César Acevedo-Opazo. Evaluation of models to determine LAI on poplar stands using spectral indices from Sentinel-2 satellite images. Ecological Modelling, 428(??):??, July 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301307.

Chen:2022:TPR

[CTL22] Renfei Chen, Chengyi Tu, and Quan-Xing Liu. Transient perturbations reveal distinct strategies for reserve benefits in life history-dependent ecosystems. Ecological Modelling, 466(??): ??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000229.

Chennault:2020:CPI

[CVTS20a] Carrie M. Chennault, Robert M. Valek, John C. Tyndall, and Lisa A. Schulte. Corrigendum to PEWI: an interactive web-based ecosystem service model for a broad public audience: Ecological Modelling 2020 431: 109165. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302660. See [CVTS20b].

Chennault:2020:PIW

[CVTS20b] Carrie M. Chennault, Robert M. Valek, John C. Tyndall, and Lisa A. Schulte. PEWI: an interactive webbased ecosystem service model for a broad public audience. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020302362. See corrigen-

dum [CVTS20a].

Chen:2021:TAE

[CWF^{+}21] Mingli Chen, Zijian Wu, Xinxi Fu, Linnan Ouyang, and Xiaofu Wu. Thermodynamic analysis of an ecologically restored plant community:number of species. Ecological Modelling, 455(??):??, September 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021001940.

Chiaverini:2021:ENR

[CWH^{+}21] Luca Chiaverini, Ho Yi Wan, Beth Hahn, Amy Cilimburg, Tzeidle N. Wasserman, and Samuel A. Cushman. Effects of non-representative sampling design on multi-scale habitat models: flammulated owls in the Rocky Mountains. Ecological Modelling, 450(??):??, June 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001356.

Chen:2023:SCC

[CWW^{+}23] Bin Chen, Pengyuan Wang, Shaoqiang Wang, Weimin Ju, Zhenhai Liu, and Yinghui Zhang. Simulating canopy carbonyl sulfide uptake of two forest stands through an improved ecosystem model and parameter optimization using an ensemble Kalman filter. Ecological Modelling, 475(??):??, January 2023. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003106.

Cai:2024:LAI

[CWY^{+}24] Yimeng Cai, Jiaxin Wu, Tudi Yimiti, Zhouyuan Li, Xiuchun Yang, and Shikui Dong. The landscape altered the interaction between vegetation and climate in the desert oasis of Hotan River Basin, Xinjiang, China. Ecological Modelling, 491(??): ??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000759.

Chen:2023:SEE

[CXZ^{+}23] Yufeng Chen, Liyan Xu, Xiao Zhang, Zilin Wang, Hailong Li, Yansheng Yang, Hong You, and Dihua Li. Socio-econ-ecosystem multipurpose simulator (SEEMS): an easy-to-apply agent-based model for simulating small-scale coupled human and nature systems in biological conservation hotspots. Ecological Modelling, 476(??):??, February 2023. CODEN EC-

MODT. ISSN ???? URL http://www.sciencedirect.com/
science/article/pii/S0304380022003301.

Contina:2020:CGS

[CYP^{+}20] Andrea Contina, Scott W. Yanco, Allison K. Pierce, Michelle DePrenger-Levin, Michael B. Wunder, Andreas M. Neophytou, C. Phoebe Lostroh, Richard J. Telford, Blas M. Benito, Joseph Chipperfield, Robert B. O’Hara, and Colin J. Carlson. Comment on “A global-scale ecological niche model to predict SARS-CoV-2 coronavirus infection rate”, author Coro. Ecological Modelling, 436(??):??, November 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303586. See [Cor20].

Chen:2021:IIB

[CZ21] Yangfan Chen and Xiaohong Zhang. Investigating the interactions between Chinese economic growth, energy consumption and its air environmental cost during 1989–2016 and forecasting their future trends. Ecological Modelling, 461(??):??, December 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021003148.

Dodson:2020:DBA

[DAB^{+}20] Stephanie Dodson, Briana Abrahms, Steven J. Bograd, Jerome Fiechter, and Elliott L. Hazen. Disentangling the biotic and abiotic drivers of emergent migratory behavior using individual-based models. Ecological Modelling, 432(??):??, September 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020302957.

dosAnjos:2020:CEH

[dAdSCA20] Lucas dos Anjos, Michel Iskin da S. Costa, and Regina C. Almeida. Characterizing the existence of hydra effect in spatial predator–prey models and the influence of functional response types and species dispersal. Ecological Modelling, 428(??):??, July 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301812.

DeAngeli:2021:MIW

[DAG^{+}21] Kevin De Angeli, Eeman Abbasi, Alan Gan, Daniel J. Ingram, Xingli Giam, and Charlotte H. Chang. Modeling the impact of wild harvest on plant–disperser mutualisms: Plant and disperser co-harvest model. Ecological Modelling, 439(??):??, January 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303951.

Prado:2023:WCN

[dAPdBMT23] Carlos Henrique Britto de Assis Prado and Dilma Maria de Brito Melo Trovão. The woody crown network model incorporates maximum height. Ecological Modelling, 481(??): ??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002300073X.

Deutsch:2020:USE

[DAQ20] Eliza S. Deutsch, Ibrahim Alameddine, and Song S. Qian. Using structural equation modeling to better understand Microcystis biovolume dynamics in a Mediterranean hypereutrophic reservoir. Ecological Modelling, 435(??):??, November 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303525.

Davis:2022:MCM

[Dav22] Katrina J. Davis. Managed culls mean extinction for a marine mammal population when combined with extreme climate impacts. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002200223X.

dosAnjos:2023:MBS

[dAWG23] Lucas dos Anjos, Igor Daniel Weber, and Wesley Augusto Conde Godoy. Modelling the biocontrol of Spodoptera frugiperda: a mechanistic approach considering Bt crops and oviposition behaviour. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002300220X.

Datola:2022:ORM

[DBDR22] Giulia Datola, Marta Bottero, Elena De Angelis, and Francesco Romagnoli. Operationalising resilience: a methodological framework for assessing urban resilience through System Dynamics Model. Ecological Modelling, 465(??):??, March 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003896.

Dentinger:2022:PFB

[DBH^{+}22] Jane E. Dentinger, Luca Börger, Mark D. Holton, Ruholla Jafari-Marandi, Durham A. Norman, Brian K. Smith, Seth F. Oppenheimer, Bronson K. Strickland, Rory P. Wilson, and Garrett M. Street. A probabilistic framework for behavioral identification from animal-borne accelerometers. Ecological Modelling, 464(??):??, February 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003628.

Dash:2020:SLW

[DBK20] Siddhant Dash, Smitom Swapna Borah, and Ajay S. Kalamdhad. Study of the limnology of wetlands through a one-dimensional model for assessing the eutrophication levels induced by various pollution sources. Ecological Modelling, 416(??):??, January 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019304156.

Diouf:2022:ABM

[DBN^{+}22] Esther Gnilane Diouf, Thierry Brévault, Saliou Ndiaye, Emile Faye, Anaïs Chailleux, Paterne Diatta, and Cyril Piou. An agent-based model to simulate the boosted Sterile Insect Technique for fruit fly management. Ecological Modelling, 468(??): ??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000710.

David:2022:TBC

[DBV^{+}22] Hassan C. David, Reinaldo I. Barbosa, Alexander C. Vibrans, Luciano F. Watzlawick, Jonathan W. Trautenmuller, Rafaelo Balbinot, Sabina C. Ribeiro, Laércio A. G. Jacovine, Ana Paula D. Corte, Carlos R. Sanquetta, Alessandra Calegari da Silva, Joberto Veloso de Freitas, and David W. MacFarlane. The tropical biomass & carbon project — an application for forest biomass and carbon estimates. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001752.

Dann:2021:MWP

[DCT^{+}21] Carson E. Dann, M. L. Cabrera, R. Thapa, S. Mirsky, K. Tully, C. Reberg-Horton, R. Hitchcock, and F. Morari. Modeling water potential of cover crop residues on the soil surface. Ecological Modelling, 459(??):??, November 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002635.

Djeumen:2021:MMV

[DDDC21] I. V. Yatat Djeumen, Y. Dumont, A. Doizy, and P. Couteron. A minimalistic model of vegetation physiognomies in the savanna biome. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304452.

Diaz:2021:DVS

[DDG^{+}21] Stephanie G. Diaz, Donald L. DeAngelis, Michael S. Gaines, Andrew Purdon, Michael A. Mole, and Rudi J. van Aarde. Development and validation of a spatially-explicit agentbased model for space utilization by African savanna elephants (Loxodonta africana) based on determinants of movement. Ecological Modelling, 447(??):??, May 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000703.

DellaChiesa:2022:NMS

[DDH^{+}22] Tomas Della Chiesa, Stephen J. Del Grosso, Melannie D. Hartman, William J. Parton, Laura Echarte, Laura Yahdjian, and Gervasio Piñeiro. A novel mechanism to simulate intercropping and relay cropping using the DayCent model. Ecological Modelling, 465(??):??, March 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021004075.

Dowling:2020:OHS

[DDL^{+}20] Natalie A. Dowling, Catherine M. Dichmont, George M. Leigh, Sean Pascoe, Rachel J. Pears, Tom Roberts, Sian Breen, Toni Cannard, Aaron Mamula, and Marc Mangel. Optimising harvest strategies over multiple objectives and stakeholder preferences. Ecological Modelling, 435(??):??, November 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303136.

Daniels:2024:VVI

[DEMJ24] Robert Daniels, Ava Ellett, Katherine McCombs, and John Jacobs. Vibrio vulnificus infections from water exposure in the Chesapeake Bay: Are they predictable? Ecological Modelling, 488(??):??, February 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023003241.

Descoteaux:2021:WTD

[Des21] Danielle Descoteaux. Warm thanks to Dr. Brian D. Fath. Ecological Modelling, 445(??):??, April 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000326.

diPorciaeBrugnera:2020:LSE

[dFT^{+}20] Manfredo di Porcia e Brugnera, Rico Fischer, Franziska Taubert, Andreas Huth, and Hans Verbeeck. Lianas in silico, ecological insights from a model of structural parasitism. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302301.

deGodoy:2023:PHS

[dGMCM^{+}23] Isabelle Bueno Silva de Godoy, Blake McGrane-Corrigan, Oliver Mason, Rafael de Andrade Moral, and Wesley Augusto Conde Godoy. Plant-host shift, spatial persistence, and the viability of an invasive insect population. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022002733.

Denryter:2021:SCD

[DGSM21] Kristin Denryter, David W. German, Thomas R. Stephenson, and Kevin L. Monteith. State- and context-dependent applications of an energetics model in free-ranging bighorn sheep. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304154.

Diaz-Gonzalez:2023:FRO

[DGVVP23] Freddy A. Diaz-Gonzalez, Jose. Vuelvas, Victoria E. Vallejo, and D. Patino. Fertilization rate optimization model for potato crops to maximize yield while reducing polluting nitrogen emissions. Ecological Modelling, 485(??):??, November 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002284.

deGamiz-Zearra:2024:IRB

[dGZHCA24] A. Lopez de Gamiz-Zearra, C. Hansen, X. Corrales, and E. Andonegi. Increasing the reliability of the Bay of Biscay Atlantis model: a sensitivity analysis to parameters perturbations using a Morris screening approach. Ecological Modelling, 488(??):??, February 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023003290.

Dambrine:2020:CBM

[DHW^{+}20] Chloé Dambrine, Martin Huret, Mathieu Woillez, Laure Pecquerie, François Allal, Arianna Servili, and Hélène de Pontual. Contribution of a bioenergetics model to investigate the growth and survival of European seabass in the Bay of Biscay–English Channel area. Ecological Modelling, 423(??): ??, May 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002030079X.

Dahy:2023:GSM

[DIS23] Basam Dahy, Salem Issa, and Nazmi Saleous. Geo-spatial modelling of carbon stock assessment of date palm at different age stages: an integrated approach of fieldwork, remote sensing and GIS. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001084.

Dash:2022:SBR

[DK22] Siddhant Dash and Ajay S. Kalamdhad. Systematic bibliographic research on eutrophication-based ecological modelling of aquatic ecosystems through the lens of science mapping. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001855.

Dennis:2021:IMI

[DKM^{+}21] Emily B. Dennis, Marc Kéry, Byron J. T. Morgan, Armin Coray, Michael Schaub, and Bruno Baur. Integrated modelling of insect population dynamics at two temporal scales. Ecological Modelling, 441(??):??, February 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304695.

deKerckhove:2022:PSF

[dKS22] Derrick T. de Kerckhove and Brian J. Shuter. Predation on schooling fish is shaped by encounters between prey during school formation using an Ideal Gas Model of animal movement. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001193.

Drielsma:2021:EME

[DL21] Michael Drielsma and Jamie Love. An equitable method for evaluating habitat amount and potential occupancy. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002030452X.

Drielsma:2022:GLC

[DLT^{+}22] Michael J. Drielsma, Jamie Love, Subhashni Taylor, Rajesh Thapa, and Kristen J. Williams. General Landscape Connectivity Model (GLCM): a new way to map whole of landscape biodiversity functional connectivity for operational planning and reporting. Ecological Modelling, 465(??):??, March 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021003963.

Ding:2022:MVF

[DLY^{+}22] Yu Ding, Haifei Liu, Daode Yu, Jingjing Song, and Guosheng Duan. The migration of viscous fish eggs in artificial reefs. Ecological Modelling, 469(??):??, July 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000990.

Ding:2020:WEN

[DLZ^{+}20] Tao Ding, Liang Liang, Kaile Zhou, Min Yang, and Yuqi Wei. Water-energy nexus: the origin, development and prospect. Ecological Modelling, 419(??):??, March 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300144.

Duan:2021:DEB

[DLZ^{+}21] Jiaoyang Duan, Hui Liu, Jianxin Zhu, Longfei Lu, Siqing Chen, Fan Lin, and Lirong Chang. A dynamic energy budget model for abalone, Haliotis discus hannai Ino. Ecological Modelling, 451(??):??, July 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001277.

Dibike:2024:CSS

[DMdR24] Yonas Dibike, Rebecca Marshall, and Laurent de Rham. Climatic sensitivity of seasonal ice-cover, water temperature and biogeochemical cycling in Lake 239 of the Experimental Lakes Area (ELA), Ontario, Canada. Ecological Modelling, 489(??): ??, March 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000103.

Filho:2020:EMF

[dMFOM^{+}20] Marcos Estevão Santiago de Melo Filho, Marco Shizuo Owatari, José Luiz Pedreira Mouriño, Bruno Augusto Mattar Carciofi, and Hugo Moreira Soares. Empirical modeling of feed conversion in Pacific white shrimp (Litopenaeus vannamei) growth. Ecological Modelling, 437(??):??, December 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303616.

Debaly:2022:AMT

[DMG22] Zinsou Max Debaly, Philippe Marchand, and Miguel Montoro Girona. Autoregressive models for time series of random sums of positive variables: Application to tree growth as a function of climate and insect outbreak. Ecological Modelling, 471(??):??, September 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001636.

Dolder:2020:HRS

[DMGP20] Paul J. Dolder, Cóilín Minto, Jean-Marc Guarini, and Jan Jaap Poos. Highly resolved spatiotemporal simulations for exploring mixed fishery dynamics. Ecological Modelling, 424(??):??, May 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300727.

Dey:2023:MSA

[DMM^{+}23] Soumen Dey, Ehsan Moqanaki, Cyril Milleret, Pierre Dupont, Mahdieh Tourani, and Richard Bischof. Modelling spatially autocorrelated detection probabilities in spatial capturerecapture using random effects. Ecological Modelling, 479(??): ??, May 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000522.

Dominguez-May:2020:DOR

[DMPLHGL20] Roger Domínguez-May, Gaspar R. Poot-López, Juan Hernández, and Eucario Gasca-Leyva. Dynamic optimal ration size in tilapia culture: Economic and environmental considerations. Ecological Modelling, 420(??):??, March 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300016.

dAcampora:2023:LEU

[dMT23] Bárbara Heliodora Alves d’Acampora, Clarice Maraschin, and Cleiton Guollo Taufemback. Landscape ecology and urban spatial configuration: Exploring a methodological relationship. application in Pelotas, Brazil. Ecological Modelling, 486(??):??, December 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002600.

Pires:2021:GMC

[dOPdMdS^{+}21] Sandra Aguiar de Oliveira Pires, Adriano Ribeiro de Mendonça, Gilson Fernandes da Silva, Marcus Vinícius Neves d’Oliveira, Luís Claudio de Oliveira, Jeferson Pereira Martins Silva, and Evandro Ferreira da Silva. Growth modeling of Carapa guianensis and Tetragastris altissima for improved management in native forests in the Amazon. Ecological Modelling, 456(??):??, September 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002416.

Dieguez:2020:UNG

[DP20] Francisco José Dieguez and Marcelo Pereira. Uruguayan native grasslands net aerial primary production model and its application on safe stocking rate concept. Ecological Modelling, 430(??):??, August 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301320.

Davis:2021:MHS

[DPA21] Natalie Davis, J. Gareth Polhill, and M. J. Aitkenhead. Measuring heterogeneity in soil networks: a network analysis and simulation-based approach. Ecological Modelling, 439(??):??, January 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303781.

Dortel:2020:DEB

[DPC20] E. Dortel, L. Pecquerie, and E. Chassot. A Dynamic Energy Budget simulation approach to investigate the eco-physiological factors behind the two-stanza growth of yellowfin tuna (Thunnus albacares). Ecological Modelling, 437(??):??, December 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020303677.

dePaiva:2024:VMI

[dPGDdFS24] Marllus Henrique Ribeiro de Paiva, Paula Cristine Silva Gomes, Lívia Cristina Pinto Dias, and Aníbal da Fonseca Santiago. Validation of the MGB-IPH hydrological model for flows simulation in paired watershed in Minas Gerais, Brazil. Ecological Modelling, 491(??):??, May 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000875.

Miranda:2023:UME

[dRMJdS23] Jonathan da Rocha Miranda, Ronie Silva Juvanhol, and Rosane Gomes da Silva. Use of maximum entropy to improve validation and prediction of active fires in a Brazilian savanna region. Ecological Modelling, 475(??):??, January 2023. CO-

DEN ???? ISSN ???? URL http://www.sciencedirect.
com/science/article/pii/S0304380022003179.

Duan:2024:RCM

[DRZ^{+}24] Yingying Duan, Haina Rong, Gexiang Zhang, Sergey Gorbachev, Dunwu Qi, Luis Valencia-Cabrera, and Mario J. Pérez-Jiménez. A review of computing models for studying population dynamics of giant panda ecosystems. Ecological Modelling, 487(??):??, January 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002739.

Dragicevic:2021:PVS

[DS21] Arnaud Z. Dragicevic and Jason F. Shogren. Preservation value in socio-ecological systems. Ecological Modelling, 443 (??):??, March 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021000247.

Dahmouni:2023:DGM

[DS23] Ilyass Dahmouni and Rashid U. Sumaila. A dynamic game model for no-take marine reserves. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380023000881.

Dunlap:2024:ESA

[DS24] J. Dunlap and J. R. Schramski. Energy-systems accounting in industrial-natural systems; an energy analysis of a managed forest ecosystem including food web biomass dynamics. Ecological Modelling, 488(??):??, February 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003289.

Batista:2023:DIC

[dSBDH23] Fabiana de Souza Batista, Confidence Duku, and Lars Hein. Deforestation-induced changes in rainfall decrease soybeanmaize yields in Brazil. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002636.

DiPirro:2021:SBA

[DSC^{+}21] E. Di Pirro, L. Sallustio, G. Capotorti, M. Marchetti, and B. Lasserre. A scenario-based approach to tackle tradeoffs between biodiversity conservation and land use pressure in Central Italy. Ecological Modelling, 448(??):??, May 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021001034.

Costa:2021:MSY

[dSCdA21] Michel Iskin da S. Costa and Lucas dos Anjos. Maximum sustainable yield and species survival: Insights from effects of prey harvest saturation on dynamic predator–prey models. Ecological Modelling, 461(??):??, December 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003136.

Dubos:2023:FLM

[DSHB23] Véronique Dubos, André St-Hilaire, and Normand E. Bergeron. Fuzzy logic modelling of anadromous Arctic char spawning habitat from Nunavik Inuit knowledge. Ecological Modelling, 477(??):??, March 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002200360X.

Magalhaes:2022:PTG

[dSMPA^{+}22] Juliana G. de S. Magalhaes, Adam P. Polinko, Mariano M. Amoroso, Gursimran S. Kohli, and Bruce C. Larson. The predicting tree growth app: an algorithmic approach to modelling individual tree growth. Ecological Modelling, 467(??): ??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000552.

Rodrigues:2023:SDM

[dSRDC^{+}23] Lucas dos Santos Rodrigues, Nicholas Winterle Daudt, Luis Gustavo Cardoso, Paul Gerhard Kinas, David Conesa, and Maria Grazia Pennino. Species distribution modelling in the Southwestern Atlantic Ocean: a systematic review and trends. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002442.

DeCubber:2023:RIP

[DTD^{+}23] Lola De Cubber, Verena M. Trenkel, Guzman Diez, Juan Gil-Herrera, Ana Maria Novoa Pabon, David Eme, and Pascal Lorance. Robust identification of potential habitats of a rare demersal species (blackspot seabream) in the Northeast Atlantic. Ecological Modelling, 477(??):??, March 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022003532.

Dadrasi:2022:MPP

[DTR^{+}22] Amir Dadrasi, Benjamin Torabi, Asghar Rahimi, Afshin Soltani, and Ebrahim Zeinali. Modeling potential production and yield gap of potato using modelling and GIS approaches. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001600.

Denechere:2024:RSF

[DvDA24] Rémy Denéchère, P. Daniël van Denderen, and Ken H. Andersen. The role of squid for food web structure and communitylevel metabolism. Ecological Modelling, 493(??):??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001170.

deVries:2020:MMD

[dVDC20] Charlotte de Vries, Robert A. Desharnais, and Hal Caswell. A matrix model for density-dependent selection in stageclassified populations, with application to pesticide resistance in Tribolium. Ecological Modelling, 416(??):??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019303837.

Duquesne:2021:CEM

[DVVH21] François Duquesne, Valentin Vallaeys, Prem Jai Vidaurre, and Emmanuel Hanert. A coupled ecohydrodynamic model to predict algal blooms in Lake Titicaca. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304750.

Drechsler:2022:HGG

[DWG22] Martin Drechsler, Frank Wätzold, and Volker Grimm. The hitchhiker’s guide to generic ecological-economic modelling of land-use-based biodiversity conservation policies. Ecological Modelling, 465(??):??, March 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003999.

Dur:2021:IBM

[DWH^{+}21] Gaël Dur, Eun-Ji Won, Jeonghoon Han, Jae-Seong Lee, and Sami Souissi. An individual-based model for evaluating postexposure effects of UV-B radiation on zooplankton reproduction. Ecological Modelling, 441(??):??, February 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304440.

Darvishi:2022:MES

[DYMD22] Asef Darvishi, Maryam Yousefi, Joan Marull, and Naghmeh Mobarghaee Dinan. Modelling ecological scarcity considering the long-term interaction between human and nature in dry agricultural landscapes. Application in Qazvin (Iran). Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002095.

Du:2022:ETQ

[DYW^{+}22] Hailong Du, Liu Yang, Wenzhong Wang, Lunhui Lu, and Zhe Li. Emergy theory to quantify the sustainability of large cascade hydropower projects in the upper Yangtze. Ecological Modelling, 468(??):??, June 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000746.

Dauvin:2024:NMD

[DZBB24] Jean-Claude Dauvin, Assia Zarrouki, Martial Boutet, and Anne-Claire Bennis. Numerical modelling of dispersal of Ampelisca (Amphipoda Gammaridae) during their diel migration. Ecological Modelling, 493(??):??, July 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001431.

Ding:2023:CLV

[DZC^{+}23a] Rui Ding, Xiaoke Zhang, Minmin Chen, Yuxi Lian, and Daoping Yu. Corrigendum to ‘Vorticity preference of the Yangtze finless porpoise (Neophocaena asiaeorientalis) during the dry season at the Wanhe River Estuary confluence’. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002569. See [DZC^{+}23b].

Ding:2023:VPY

[DZC^{+}23b] Rui Ding, Xiaoke Zhang, Minmin Chen, Yuxi Lian, and Daoping Yu. Vorticity preference of the Yangtze finless porpoise (Neophocaena asiaeorientalis) during the dry season at the Wanhe River Estuary confluence. Ecological Modelling, 485(??):??, November 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S030438002300234X. See corrigendum [DZC^{+}23a].

ElSaadi:2020:STC

[EBMK20] N. El Saadi, A. Bah, T. Mahdjoub, and C. Kribs. On the sylvatic transmission of T. cruzi, the parasite causing Chagas disease: a view from an agent-based model. Ecological Modelling, 423(??):??, May 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300739.

Elzinga:2023:MMI

[EBS23] David C. Elzinga, Charlotte Beckford, and W. Christopher Strickland. A mathematical model of the impacts of climate change on the winter tick epizootic in moose. Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001527.

Epule:2022:SNS

[ECC^{+}22] Terence Epule Epule, Abdelghani Chehbouni, Tarik Chfadi, Victor Ongoma, Salah Er-Raki, Said Khabba, Daniel Etongo, Adán L. Martínez-Cruz, Ernest L. Molua, Soumia Achli, Wiam Salih, Clifford Chuwah, Martin Jemo, and Ikram Chairi. A systematic national stocktake of crop models in Morocco. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001466.

Enayat:2023:EEF

[EGAS23] Fatemeh Fartout Enayat, Seyed Ahmad Ghanbari, Mohammad Reza Asgharipour, and Esmaeel Seyedabadi. Emergy ecological footprint analysis of yaghooti grape production in the Sistan region of Iran. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000601.

Elli:2020:AAN

[EHS^{+}20] Elvis Felipe Elli, Neil Huth, Paulo Cesar Sentelhas, Rafaela Lorenzato Carneiro, and Clayton Alcarde Alvares. Ability of the APSIM Next Generation Eucalyptus model to simulate complex traits across contrasting environments. Ecological Modelling, 419(??):??, March 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300314.

Elevitch:2020:PRP

[EJ20] Craig R. Elevitch and C. Richard Johnson. A procedure for ranking parameter importance for estimation in predictive mechanistic models. Ecological Modelling, 419(??):??, March 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002030020X.

E:2021:RRF

[EL21] Cagle Sierra E. and Roelke Daniel L. Relative roles of fundamental processes underpinning PEG dynamics in dimictic lakes as revealed by a self-organizing, multi-population plankton model. Ecological Modelling, 462(??):??, December 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003380.

Elliot:2022:SDL

[EL22] Thomas Elliot and Annie Levasseur. System dynamics life cycle-based carbon model for consumption changes in urban metabolism. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022001211.

Elhabashy:2023:WQM

[ELS23] Ahmed Elhabashy, Jing Li, and Ekaterina Sokolova. Water quality modeling of a eutrophic drinking water source: Impact of future climate on Cyanobacterial blooms. Ecological Modelling, 477(??):??, March 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000030.

Eupen:2022:SPS

[EMH^{+}22] Camille Van Eupen, Dirk Maes, Marc Herremans, Kristijn R. R. Swinnen, Ben Somers, and Stijn Luca. Species profiles support recommendations for quality filtering of opportunistic citizen science data. Ecological Modelling, 467(??): ??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000345.

Eyni-Nargeseh:2023:WRF

[ENARM^{+}23] Hamed Eyni-Nargeseh, Mohammad Reza Asgharipour, Sajjad Rahimi-Moghaddam, Abdolali Gilani, Abdolmajid Mahdavi Damghani, and Khosro Azizi. Which rice farming system is more environmentally friendly in Khuzestan province, Iran? A study based on emergy analysis. Ecological Modelling, 481(??): ??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001011.

Enoh:2023:MSS

[EON23] Mfoniso Asuquo Enoh, Uzoma Chinenye Okeke, and Baywood Ceciliar Nkechi. Mapping and simulating the spatialtemporal changes in the Lagos wetland ecosystem: a step-by-step approach to creating a carbon-neutral, sustainable urban city. Ecological Modelling, 482(??):??, August 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001308.

Ekinci:2021:CAM

[EOÖ21] Ekin Ekinci, Sevinç İlhan Omurca, and Bilge Özbay. Comparative assessment of modeling deep learning networks for modeling ground-level ozone concentrations of pandemic lockdown period. Ecological Modelling, 457(??):??, October 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002349.

Emmett:2021:ADV

[ERP21] Kristen D. Emmett, Katherine M. Renwick, and Benjamin Poulter. Adapting a dynamic vegetation model for regional biomass, plant biogeography, and fire modeling in the Greater Yellowstone Ecosystem: Evaluating LPJ-GUESS-LMfireCF. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304762.

Edrisi:2022:EBS

[ESCA22] Sheikh Adil Edrisi, Sheikh Arshiya Sahiba, Bin Chen, and P. C. Abhilash. Emergy-based sustainability analysis of bioenergy production from marginal and degraded lands of India. Ecological Modelling, 466(??):??, April 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000291.

Erturk:2023:DPB

[ESZ^{+}23] Ali Ertürk, Ilona Sakurova, Mindaugas Zilius, Petras Zemlys, Georg Umgiesser, Burak Kaynaroglu, Renata Pilkaitytė, and Artūras Razinkovas-Baziukas. Development of a pelagic biogeochemical model with enhanced computational performance by optimizing ecological complexity and spatial resolution. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002399.

Emadi-Tafti:2021:IIV

[ETAAH21] Mohsen Emadi-Tafti, Behzad Ataie-Ashtiani, and Seiyed Mossa Hosseini. Integrated impacts of vegetation and soil type on slope stability: a case study of Kheyrud Forest, Iran. Ecological Modelling, 446(??):??, April 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000697.

Elizondo:2022:IBM

[EV22] Urs Hofmann Elizondo and Meike Vogt. Individual-based modeling of shelled pteropods. Ecological Modelling, 468(??): ??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000655.

Elghafghuf:2021:SSM

[EVHS21] Adel Elghafghuf, Raphael Vanderstichel, Larry Hammell, and Henrik Stryhn. State-space modeling for inter-site spread of sea lice with short-term population predictions. Ecological Modelling, 452(??):??, July 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001642.

Engel:2021:SET

[EVT^{+}21] Markus Engel, Sonja Vospernik, Maude Toïgo, Xavier Morin, Antonio Tomao, Carlo Trotta, Mathias Steckel, Anna Barbati, Arne Nothdurft, Hans Pretzsch, Miren del Rio, Jerzy Skrzyszewski, Quentin Ponette, Magnus Löf, Āris Jansons, and Gediminas Brazaitis. Simulating the effects of thinning and species mixing on stands of oak (Quercus petraea (Matt.) Liebl./Quercus robur L.) and pine (Pinus sylvestris L.) across Europe. Ecological Modelling, 442(??):??, February 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020304683.

ElRahi:2020:MEB

[EWE20] Joe El Rahi, Marc P. Weeber, and Ghada El Serafy. Modelling the effect of behavior on the distribution of the jellyfish Mauve stinger (Pelagia noctiluca) in the Balearic Sea using an individual-based model. Ecological Modelling, 433(??):??, October 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303008.

Freitas:2022:SMM

[FAC22] Osmar Freitas, Sabrina B. L. Araujo, and Paulo R. A. Campos. Speciation in a metapopulation model upon environmental changes. Ecological Modelling, 468(??):??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002200076X.

Fath:2021:ANA

[Fat21] Brian D. Fath. Announcing a new article type — Jørgensen research and reviews (JRR). Ecological Modelling, 460(??):??, November 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003045.

Ferrieres:2021:PMT

[FC21] S. A. Ferrieres and C. A. Condat. A population model for Triatoma infestans in chicken coops. Ecological Modelling, 460(??):??, November 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003021.

Falgueras-Cano:2024:ISF

[FCFCM24] Javier Falgueras-Cano, Juan-Antonio Falgueras-Cano, and Andrés Moya. The influence of stochastic fluctuations on population dynamics: an in-silico approach. Ecological Modelling, 490(??):??, April 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000565.

Filgueira:2020:EMO

[FCP^{+}20] Ramón Filgueira, Manuel Chica, Juan José Palacios, Tore Strohmeier, Romain Lavaud, Antonio Agüera, Sergio Damas, and Øivind Strand. Embracing multimodal optimization to enhance dynamic energy budget parameterization. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302106.

Fabritius:2021:EMD

[FdKO21] Henna Fabritius, Henrik de Knegt, and Otso Ovaskainen. Effects of a mobile disturbance pattern on dynamic patch networks and metapopulation persistence. Ecological Modelling, 460(??):??, November 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002100288X.

Forget:2020:UQN

[FDSM20] Nathalie L. Forget, Daniel E. Duplisea, Fany Sardenne, and Christopher W. McKindsey. Using qualitative network models to assess the influence of mussel culture on ecosystem dynamics. Ecological Modelling, 430(??):??, August 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301423.

Forrester:2024:SAF

[FEPR24] David I. Forrester, Jacqueline R. England, Keryn I. Paul, and Stephen H. Roxburgh. Sensitivity analysis of the FullCAM model: Context dependency and implications for model development to predict Australia’s forest carbon stocks. Ecological Modelling, 489(??):??, March 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000206.

Ferreira:2023:OCP

[Fer23] João Miguel Ferreira. Optimal control policies for a noneruptive population of rodents — the relevance of migration. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002077.

Fitts:2023:ISS

[FFM^{+}23] Lucia A. Fitts, Jacob S. Fraser, Brian R. Miranda, Grant M. Domke, Matthew B. Russell, and Brian R. Sturtevant. An iterative site-scale approach to calibrate and corroborate successional processes within a forest landscape model. Ecological Modelling, 477(??):??, March 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000029.

Fort:2021:MPS

[FG21] Hugo Fort and Tomás S. Grigera. A method for predicting species trajectories tested with trees in Barro Colorado tropical forest. Ecological Modelling, 446(??):??, April 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021000752.

Frantz:2024:ASP

[FGM^{+}24] Rachel M. Frantz, Humberto Godinez, Kaitlyn Martinez, William S. Cuello, and Carrie Manore. Age structured partial differential equations model for Culex mosquito abundance. Ecological Modelling, 494(??):??, August 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001522.

Furniss:2022:PFP

[FHP^{+}22] Tucker J. Furniss, Paul F. Hessburg, Nicholas A. Povak, R. Brion Salter, and Mark S. Wigmosta. Predicting future patterns, processes, and their interactions: Benchmark calibration and validation procedures for forest landscape models. Ecological Modelling, 473(??):??, November 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002022.

Friess:2023:RGD

[FLG^{+}23] Johannes L. Frieß, Carina R. Lalyer, Bernd Giese, Samson Simon, and Mathias Otto. Review of gene drive modelling and implications for risk assessment of gene drive organisms. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000133.

Fan:2020:PBC

[FMH^{+}20] L. I. N. Fan, D. U. Meirong, L. I. U. Hui, F. A. N. G. Jianguang, Asplin Lars, and J. I. A. N. G. Zengjie. A physical–biological coupled ecosystem model for integrated aquaculture of bivalve and seaweed in Sanggou Bay. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302520.

Franco:2024:NRD

[FMRMAS24] Rebeca Franco, Marco A. Morales, J. Isrrael Rodríguez-Mora, and Ricardo Agustín-Serrano. A new reaction– diffusion–advection model with long-range inhibition for vegetation-desertification pattern-formation as a unified approach. Ecological Modelling, 492(??):??, June 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001108.

Franzese:2020:MME

[FMSR20] Pier Paolo Franzese, Fausto Manes, Michele Scardi, and Angelo Riccio. Modelling matter and energy flows in the biosphere and human economy. Ecological Modelling, 422(??):??, April 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020300569.

Fort:2020:MQP

[For20] Hugo Fort. Making quantitative predictions on the yield of a species immersed in a multispecies community: the focal species method. Ecological Modelling, 430(??):??, August 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020301800.

Forbes:2024:NSE

[For24] Valery E. Forbes. The need for standardization in ecological modeling for decision support: Lessons from ecological risk assessment. Ecological Modelling, 492(??):??, June 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001248.

Fourcade:2021:FTN

[Fou21] Yoan Fourcade. Fine-tuning niche models matters in invasion ecology. A lesson from the land planarian Obama nungara. Ecological Modelling, 457(??):??, October 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002441.

Fust:2022:ITV

[FS22] Pascal Fust and Eva Schlecht. Importance of timing: Vulnerability of semi-arid rangeland systems to increased variability in temporal distribution of rainfall events as predicted by future climate change. Ecological Modelling, 468(??):??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000795.

Fernandez:2022:RAN

[FSY22] Marc Fernandez, Neftali Sillero, and Chris Yesson. To be or not to be: the role of absences in niche modelling for highly mobile species in dynamic marine environments. Ecological Modelling, 471(??):??, September 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001508.

Frolov:2020:CMGa

[FSZ^{+}20] Pavel Frolov, Vladimir Shanin, Elena Zubkova, Sergey Bykhovets, and Pavel Grabarnik. CAMPUS-S — the model of ground layer vegetation populations in forest ecosystems and their contribution to the dynamics of carbon and nitrogen. I. Problem formulation and description of the model. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302556.

Fulford:2024:SIF

[FTH24] Richard S. Fulford, Jessica L. Tolan, and James D. Hagy. Simulating implications of fish behavioral response for managing hypoxia in estuaries with spatial dissolved oxygen variability. Ecological Modelling, 490(??):??, April 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000243.

Farrell:2020:ESN

[FWK^{+}20] Kaitlin J. Farrell, Nicole K. Ward, Arianna I. Krinos, Paul C. Hanson, Vahid Daneshmand, Renato J. Figueiredo, and Cayelan C. Carey. Ecosystem-scale nutrient cycling responses to increasing air temperatures vary with lake trophic state. Ecological Modelling, 430(??):??, August 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302064.

Frolov:2020:CMGb

[FZS^{+}20] Pavel Frolov, Elena Zubkova, Vladimir Shanin, Sergey Bykhovets, Raisa Mäkipää, and Maija Salemaa. CAMPUS-S — the model of ground layer vegetation populations in forest ecosystems and their contribution to the dynamics of carbon and nitrogen. II. Parameterization, validation and simulation experiments. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302544.

Ghosh:2021:PHA

[GAC^{+}21] Sinchan Ghosh, Fahad Al Basir, Ganesh Chowdhury, Sabyasachi Bhattacharya, and Santanu Ray. Is the primary helper always a key group for the dynamics of cooperative birds? A mathematical study on cooperative breeding birds. Ecological Modelling, 459(??):??, November 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002805.

Gatti:2020:ECB

[GAM20] Roberto Cazzolla Gatti, Nicola Amoroso, and Alfonso Monaco. Estimating and comparing biodiversity with a single universal metric. Ecological Modelling, 424(??):??, May 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020300922.

Gheisari:2023:UDM

[GAMN^{+}23] Aboalghasem Gheisari, Mohammad Reza Asgharipour, Mohsen Mousavi-Nik, Ahmad Ghanbari, and Mohammad Ali Javaheri. Utilization of the DNDC model to estimate yield and CO_{2} and CH_{4} emissions in a cotton-wheat rotation under the influence of various tillage treatments. Ecological Modelling, 481(??):??, July 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000856.

Gatti:2020:COS

[Gat20] Roberto Cazzolla Gatti. Coronavirus outbreak is a symptom of Gaia’s sickness. Ecological Modelling, 426(??):??, June 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020301472.

Gatti:2021:MAB

[Gat21] Roberto Cazzolla Gatti. A multi-armed bandit algorithm speeds up the evolution of cooperation. Ecological Modelling, 439(??):??, January 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020304142.

Gantchoff:2022:SEP

[GCB^{+}22] Mariela G. Gantchoff, Laura Conlee, Melanie R. Boudreau, Raymond B. Iglay, Charles Anderson, and Jerrold L. Belant. Spatially-explicit population modeling to predict large carnivore recovery and expansion. Ecological Modelling, 470(??):??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001430.

Guarino:2023:HSD

[GCMU23] Raffaele Guarino, Giulio Corsi, and Enrique Muñoz-Ulecia. How sustainable development goals have transformed our world? Evolution of the ecological networks of the Italian economy. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002041.

Goodbody:2021:MRT

[GCS^{+}21] Tristan R. H. Goodbody, Nicholas C. Coops, Vivek Srivastava, Bethany Parsons, Sean P. Kearney, Gregory J. M. Rickbeil, and Gordon B. Stenhouse. Mapping recreation and tourism use across grizzly bear recovery areas using social network data and maximum entropy modelling. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304427.

Goddard:2022:IER

[GCSlR22] K. A. Goddard, K. J. Craig, J. Schoombie, and P. C. le Roux. Investigation of ecologically relevant wind patterns on Marion Island using Computational Fluid Dynamics and measured data. Ecological Modelling, 464(??):??, February 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003689.

Girondot:2024:MBN

[GDC24] Marc Girondot, Ambre Dejoie, and Michel Charpentier. The mystery of bimodal nesting seasons in marine turtles. Ecological Modelling, 490(??):??, April 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000425.

Gignoux:2022:SPE

[GDF22] Jacques Gignoux, Ian D. Davies, and Shayne R. Flint. 3Worlds, a simulation platform for ecosystem modelling. Ecological Modelling, 473(??):??, November 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002228.

Gera:2023:CPB

[GGE^{+}23] Anitha Gera, R. Gayathri, P. Ezhilarasan, V. Ranga Rao, and M. V. Ramana Murthy. Coupled physical-biogeochemical simulations of upwelling, ecological response to fresh water. Ecological Modelling, 476(??):??, February 2023. CODEN EC-

MODT. ISSN ???? URL http://www.sciencedirect.com/
science/article/pii/S0304380022003441.

Glynn:2020:ABM

[GGMR20] Peter J. Glynn, Peter W. Glynn, Juan Maté, and Bernhard Riegl. Agent-based model of Eastern Pacific damselfish and sea urchin interactions shows increased coral reef erosion under post-ENSO conditions. Ecological Modelling, 423(??):??, May 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020300715.

Goncharov:2022:MMI

[GGS^{+}22] Anton A. Goncharov, Anna S. Gorbatova, Alena A. Sidorova, Alexei V. Tiunov, and Gennady A. Bocharov. Mathematical modelling of the interaction of winter wheat (Triticum aestivum) and Fusarium species (Fusarium spp.). Ecological Modelling, 465(??):??, March 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002100394X.

Gatmiry:2021:FWC

[GHbS21] Zohreh S. Gatmiry, Ashkan Hafezalkotob, Morteza Khakzar bafruei, and Roya Soltani. Food web conservation vs. strategic threats: a security game approach. Ecological Modelling, 442(??):??, February 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000016.

Gloy:2023:EAT

[GHK23] Josias Gloy, Ulrike Herzschuh, and Stefan Kruse. Evolutionary adaptation of trees and modelled future larch forest extent in Siberia. Ecological Modelling, 478(??):??, April 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023000066.

Gao:2020:CPL

[GJC^{+}20] Hongmei Gao, Feng Jiang, Xiangwen Chi, Guangying Li, Zhenyuan Cai, Wen Qin, Jingjie Zhang, Tong Wu, and Tongzuo Zhang. The carrying pressure of livestock is higher than that of large wild herbivores in Yellow River source area, China. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020302349.

Galasso:2020:UDE

[GLA^{+}20] Helena Lopes Galasso, Sébastien Lefebvre, Catherine Aliaume, Bastien Sadoul, and Myriam D. Callier. Using the dynamic energy budget theory to evaluate the bioremediation potential of the polychaete Hediste diversicolor in an integrated multi-trophic aquaculture system. Ecological Modelling, 437(??):??, December 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303665.

Gibbs:2024:HRL

[GLH24] Richard Gibbs, Pietro Landi, and Cang Hui. Heterogeneity in the resource landscape encourages increased cognitive and perceptive capabilities in foragers. Ecological Modelling, 492(??): ??, June 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000814.

Guo:2022:CSM

[GLK^{+}22] Hui Guo, Sien Li, Shaozhong Kang, Taisheng Du, Wenfeng Liu, Ling Tong, Xinmei Hao, and Risheng Ding. Comparison of several models for estimating gross primary production of drip-irrigated maize in arid regions. Ecological Modelling, 468(??):??, June 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000515.

Ghani:2023:MDW

[GM23] Latifah Abdul Ghani and Noor Zalina Mahmood. Modeling domestic wastewater pathways on household system using the socio-MFA techniques. Ecological Modelling, 480(??): ??, June 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002300056X.

Guadagno:2020:UTD

[GML^{+}20] C. R. Guadagno, D. Millar, R. Lai, D. S. Mackay, J. R. Pleban, C. R. McClung, C. Weinig, D. R. Wang, and B. E. Ewers. Use of transcriptomic data to inform biophysical models via Bayesian networks. Ecological Modelling, 429(??):??, August 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301587.

Ghosh:2020:DMC

[GMM20] Santu Ghosh, Mukul Mal, and Sudipto Mandal. A dynamic model of cadmium bioaccumulation in Lamellidens marginalis, an edible shellfish in India. Ecological Modelling, 419(??):??, March 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300296.

Garnier:2024:MAH

[GMM^{+}24] Soizic Garnier, Meadhbh Moriarty, Rory O’Hara Murray, Alejandro Gallego, and Alexander G. Murray. Modelling assessment of how small-scale vertical movements of infectious sea lice larvae can affect their large-scale distribution in fjordic systems. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023003174.

Gaglio:2023:RRS

[GMS^{+}23] Mattias Gaglio, Alexandra Nicoleta Muresan, Alessandro Sebastiani, Davide Cavicchi, Elisa Anna Fano, and Giuseppe Castaldelli. A “reserve” of regulating services: the importance of a remnant protected forest for human well-being in the Po delta (Italy). Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002156.

Gonzalez:2023:ECE

[Gon23] Cecilia González. Evolution of the concept of ecological integrity and its study through networks. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003222.

Grande:2023:MNC

[GPN^{+}23] U. Grande, A. Piernik, A. Nienartowicz, E. Buonocore, and P. P. Franzese. Measuring natural capital value and ecological complexity of lake ecosystems. Ecological Modelling, 482(??): ??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001321.

Gonzalez-Ramirez:2022:MMC

[GRAA22] Laura R. González-Ramírez, Deniz Alaçam, and Aysegul Akpinar. A mathematical model of Chenopodium album L. dynamics under copper-induced stress. Ecological Modelling, 469(??):??, July 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000849.

Gracia:2020:AKE

[GRCSA^{+}20] Eva Graciá, Roberto C. Rodríguez-Caro, Ana Sanz-Aguilar, José D. Anadón, Francisco Botella, Angel Luis García-García, Thorsten Wiegand, and Andrés Giménez. Assessment of the key evolutionary traits that prevent extinctions in humanaltered habitats using a spatially explicit individual-based model. Ecological Modelling, 415(??):??, January 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001930331X.

Gimenez-Romero:2021:MPP

[GRGHM21] Àlex Giménez-Romero, Amalia Grau, Iris E. Hendriks, and Manuel A. Matias. Modelling parasite-produced marine diseases: the case of the mass mortality event of Pinna nobilis. Ecological Modelling, 459(??):??, November 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002100260X.

Grimm:2020:OPU

[Gri20] Volker Grimm. The ODD protocol: an update with guidance to support wider and more consistent use. Ecological Modelling, 428(??):??, July 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301770.

Gouveia:2020:ERE

[GRS^{+}20] Sidney F. Gouveia, Juan G. Rubalcaba, Vladislav Soukhovolsky, Olga Tarasova, A. Márcia Barbosa, and Raimundo Real. Ecophysics reload — exploring applications of theoretical physics in macroecology. Ecological Modelling, 424(??):??, May 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301046.

Gianuca:2024:EBS

[GSA24] Kahuam Gianuca, Tatiana Silva, and Milton Asmus. Ecosystembased spatial modeling: Assessing the supply of hydrological services in a watershed in Southern Brazil. Ecological Modelling, 492(??):??, June 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002400111X.

Gilioli:2023:EEM

[GSC^{+}23] Gianni Gilioli, Anna Simonetto, Michele Colturato, Noelia Bazarra, José R. Fernández, Maria Grazia Naso, Boscia Donato, Domenico Bosco, Crescenza Dongiovanni, Andrea Maiorano, Olaf Mosbach-Schulz, Juan A. Navas Cortés, and Maria Saponari. An eco-epidemiological model supporting rational disease management of Xylella fastidiosa. An application to the outbreak in Apulia (Italy). Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003246.

Grames:2022:SAS

[GSH^{+}22] Eliza M. Grames, Piper L. Stepule, Susan Z. Herrick, Benjamin T. Ranelli, and Chris S. Elphick. Separating acoustic signal into underlying behaviors with self-exciting point process models. Ecological Modelling, 468(??):??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000837.

Guan:2020:EST

[GSJ^{+}20] Lisha Guan, Xiujuan Shan, Xianshi Jin, Harry Gorfine, Tao Yang, and Zhongyi Li. Evaluating spatio-temporal dynamics of multiple fisheries-targeted populations simultaneously: a case study of the Bohai Sea ecosystem in China. Ecological Modelling, 422(??):??, April 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300594.

Gao:2022:KPU

[GSJ^{+}22] Shufei Gao, Anglu Shen, Jie Jiang, Hao Wang, and Sanling Yuan. Kinetics of phosphate uptake in the dinoflagellate Karenia mikimotoi in response to phosphate stress and temperature. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000333.

Gelis:2022:DPC

[GSJC22] M. M. Nicolosi Gelis, M. B. Sathicq, J. Jupke, and J. Cochero. DiaThor: R package for computing diatom metrics and biotic indices. Ecological Modelling, 465(??):??, March 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003975.

Guillaumot:2020:CDM

[GSM^{+}20] Charlène Guillaumot, Thomas Saucède, Simon A. Morley, Starrlight Augustine, Bruno Danis, and Sebastiaan Kooijman. Can DEB models infer metabolic differences between intertidal and subtidal morphotypes of the Antarctic limpet Nacella concinna (Strebel, 1908)? Ecological Modelling, 430(??):??, August 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301605.

Gustafson:2024:OCH

[GSMD24] Eric J. Gustafson, Brian R. Sturtevant, Brian R. Miranda, and Matthew J. Duveneck. Overcoming conceptual hurdles to accurately represent trees as cohorts in forest landscape models. Ecological Modelling, 490(??):??, April 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000462.

Gao:2023:MND

[GSR^{+}23] Ya Gao, Chen Sun, Tiago B. Ramos, Zailin Huo, Guanhua Huang, and Xu Xu. Modeling nitrogen dynamics and biomass production in rice paddy fields of cold regions with the ORYZA-N model. Ecological Modelling, 475(??):??, January 2023. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002824.

Grover:2022:CSF

[GSRB22] James P. Grover, J. Thad Scott, Daniel L. Roelke, and Bryan W. Brooks. Competitive superiority of N-fixing cyanobacteria when fixed N is scarce: Reconsiderations based on a model with heterocyst differentiation. Ecological Modelling, 466(??):??, April 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000308.

Gachoki:2023:TTH

[GVG^{+}23] Stella Gachoki, Anton Vrieling, Thomas A. Groen, Andrew Skidmore, and Daniel Masiga. Transferability of tsetse habitat models between different regions in Kenya and Rwanda. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002788.

Guo:2023:DSN

[GWD^{+}23] Qing Guo, Yi Wang, Chuanjun Dai, Lijun Wang, He Liu, Jianbing Li, Pankaj Kumar Tiwari, and Min Zhao. Dynamics of a stochastic nutrient–plankton model with regime switching. Ecological Modelling, 477(??):??, March 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022003477.

Grimmett:2020:POS

[GWH20] Liam Grimmett, Rachel Whitsed, and Ana Horta. Presenceonly species distribution models are sensitive to sample prevalence: Evaluating models using spatial prediction stability and accuracy metrics. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302659.

Gui:2024:ENP

[GXZ^{+}24] Hanliang Gui, Qinchuan Xin, Xuewen Zhou, Zhenhua Xiong, and Kun Xiao. Embedding a novel phenology model into the Common Land Model for improving the modeling of landatmosphere fluxes. Ecological Modelling, 494(??):??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001704.

Grosklos:2023:CDD

[GZ23] Guenchik Grosklos and Jia Zhao. Chaos does not drive lower synchrony for intrinsically-induced population fluctuations. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003015.

Guo:2023:UML

[GZD^{+}23] Hui Guo, Xiao Zhou, Yi Dong, Yahui Wang, and Sien Li. On the use of machine learning methods to improve the estimation of gross primary productivity of maize field with drip irrigation. Ecological Modelling, 476(??): ??, February 2023. CODEN ECMODT. ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003489.

Guo:2020:SBN

[GZK^{+}20] Kai Guo, Xinchang Zhang, Xi Kuai, Zhifeng Wu, Yiyun Chen, and Yi Liu. A spatial bayesian-network approach as a decision-making tool for ecological-risk prevention in land ecosystems. Ecological Modelling, 419(??):??, March 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019304375.

Guo:2020:BSA

[GZW^{+}20] Zhen Guo, Zhiwei Zhang, Xiaogang Wu, Jing Wang, Peidong Zhang, Dan Ma, and Yuan Liu. Building shading affects the ecosystem service of urban green spaces: Carbon capture in street canyons. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302490.

Harik:2023:UFL

[HAZEF23] G. Harik, I. Alameddine, R. Zurayk, and M. El-Fadel. Uncertainty in forecasting land cover land use at a watershed scale: Towards enhanced sustainable land management. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002454.

He:2022:ABM

[HBC^{+}22] Haosen He, Erin Buchholtz, Frederick Chen, Susanne Vogel, and Chu A. (Alex) Yu. An agent-based model of elephant crop consumption walks using combinatorial optimization. Ecological Modelling, 464(??):??, February 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003902.

Hanley:2024:PSA

[HBMR24] Patricia H. Hanley, Eric P. Bjorkstedt, Bryan L. Morse, and Rémy Rochette. Pop-up satellite archival tags and an individual-based model provide new insights on behavioural thermoregulation in migrating ovigerous American lobster. Ecological Modelling, 489(??):??, March 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002300337X.

Haider:2021:LSD

[HBP^{+}21] Saira M. Haider, Allison M. Benscoter, Leonard Pearlstine, Laura E. D’Acunto, and Stephanie S. Romañach. Landscapescale drivers of endangered Cape Sable Seaside Sparrow (Ammospiza maritima mirabilis) presence using an ensemble modeling approach. Ecological Modelling, 461(??):??, December 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003239.

Holzer:2024:EDT

[HBRK24] Dominik Holzer, Kai Bödeker, Werner Rammer, and Thomas Knoke. Evaluating dynamic tree-species-shifting and height development caused by ungulate browsing in forest regeneration using a process-based modeling approach. Ecological Modelling, 493(??):??, July 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024001297.

Hong:2022:USE

[HCL^{+}22] Wei Hong, Bingxue Chen, Yuntao Lu, Chuanqi Lu, and Shengping Liu. Using system equalization principle to study the effects of multiple factors to the development of bee colony. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001144.

He:2024:CCM

[HCL^{+}24] Huayun He, Chao Chen, Qiuyue Li, Bin Guo, Qi Hu, Lina Zhang, and Yanmei Pang. Combining CENTURY model and vulnerability assessment index to estimate aboveground biomass and ecological vulnerability of grassland: an analysis in Northwest Sichuan Plateau. Ecological Modelling, 493(??): ??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001212.

Hoglind:2020:PBM

[HCP^{+}20] Mats Höglind, David Cameron, Tomas Persson, Xiao Huang, and Marcel van Oijen. BASGRA N: a model for grassland productivity, quality and greenhouse gas balance. Ecological Modelling, 417(??):??, February 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019304338.

Huang:2020:QCE

[HCPG20] Jiacong Huang, Qiuwen Chen, Jian Peng, and Junfeng Gao. Quantifying the cost-effectiveness of nutrient-removal strategies for a lowland rural watershed: Insights from processbased modeling. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301952.

Hanley:2022:IST

[HCW^{+}22] Brenda J. Hanley, Michelle Carstensen, Daniel P. Walsh, Sonja A. Christensen, Daniel J. Storm, James G. Booth, Joseph Guinness, Cara E. Them, Md Sohel Ahmed, and Krysten L. Schuler. Informing surveillance through the characterization of outbreak potential of chronic wasting disease in white-tailed deer. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022001648.

Huang:2021:IEG

[HCX^{+}21] Duan Huang, Hong Chi, Fengfei Xin, Akira Miyata, Minseok Kang, Kaiwen Liu, Rendong Li, Haishan Dang, Yuanwei Qin, and Xiangming Xiao. Improved estimation of gross primary production of paddy rice cropland with changing model parameters over phenological transitions. Ecological Modelling, 445(??):??, April 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000636.

Hjelkrem:2021:PBM

[HEL^{+}21] Anne-Grete Roer Hjelkrem, Håvard Eikemo, Vinh Hong Le, Arne Hermansen, and Ragnhild Nærstad. A process-based model to forecast risk of potato late blight in Norway (the Nærstad model): model development, sensitivity analysis and Bayesian calibration. Ecological Modelling, 450(??):??, June 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021001344.

Heneghan:2020:FSS

[HES^{+}20] Ryan F. Heneghan, Jason D. Everett, Patrick Sykes, Sonia D. Batten, Martin Edwards, Kunio Takahashi, Iain M. Suthers, Julia L. Blanchard, and Anthony J. Richardson. A functional size-spectrum model of the global marine ecosystem that resolves zooplankton composition. Ecological Modelling, 435(??):??, November 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0304380020303355. See corrigendum [HES^{+}23].

Heneghan:2023:CFS

[HES^{+}23] Ryan F. Heneghan, Jason D. Everett, Patrick Sykes, Sonia D. Batten, Martin Edwards, Kunio Takahashi, Iain M. Suthers, Julia L. Blanchard, and Anthony J. Richardson. Corrigendum to “A functional size-spectrum model of the global marine ecosystem that resolves zooplankton composition”, Ecological Modelling, 2020, 435: 109265. Ecological Modelling, 480(??):??, June 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000376. See [HES^{+}20].

Hunt:2020:NPM

[HFP20] Allen G. Hunt, Boris Faybishenko, and Thomas L. Powell. A new phenomenological model to describe root-soil interactions based on percolation theory. Ecological Modelling, 433(??):??, October 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302751.

Hunt:2022:TME

[HFP22] A. G. Hunt, B. Faybishenko, and T. L. Powell. Test of model of equivalence of tree height growth and transpiration rates in percolation-based phenomenology for root-soil interaction. Ecological Modelling, 465(??):??, March 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003914.

Horemans:2024:ESC

[HFSL^{+}24] Dante M. L. Horemans, Marjorie A. M. Friedrichs, Pierre St-Laurent, Raleigh R. Hood, and Christopher W. Brown. Evaluating the skill of correlative species distribution models trained with mechanistic model output. Ecological Modelling, 491(??): ??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000802.

Hjelkrem:2023:NPB

[HGBK23] Anne-Grete Roer Hjelkrem, Jakob Geipel, Anne Kjersti Bakken, and Audun Korsaeth. NORNE, a process-based grass growth model accounting for within-field soil variation using remote sensing for in-season corrections. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001643.

Henne:2023:ATM

[HH23] Paul D. Henne and Todd J. Hawbaker. An aridity threshold model of fire sizes and annual area burned in extensively forested ecoregions of the western USA. Ecological Modelling, 477(??):??, March 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000054.

Holtmann:2024:AIM

[HHBF24] Anne Holtmann, Andreas Huth, Friedrich Bohn, and Rico Fischer. Assessing the impact of multi-year droughts on German forests in the context of increased tree mortality. Ecological Modelling, 492(??):??, June 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002400084X.

Henniger:2023:CVF

[HHFB23] Hans Henniger, Andreas Huth, Karin Frank, and Friedrich J. Bohn. Creating virtual forests around the globe and analysing their state space. Ecological Modelling, 483(??):??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001357.

Haga:2022:MTR

[HHI^{+}22] Chihiro Haga, Wataru Hotta, Takahiro Inoue, Takanori Matsui, Masahiro Aiba, Toshiaki Owari, Satoshi N. Suzuki, Hideaki Shibata, and Junko Morimoto. Modeling tree recovery in wind-disturbed forests with dense understory species under climate change. Ecological Modelling, 472(??):??, October 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002200179X.

Hetzer:2021:IPT

[HHT21] Jessica Hetzer, Andreas Huth, and Franziska Taubert. The importance of plant trait variability in grasslands: a modelling study. Ecological Modelling, 453(??):??, August 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002100168X.

Hayashi:2022:DEN

[HI22] Kohta Hayashi and Hayato Iijima. Density estimation of non-independent unmarked animals from camera traps. Ecological Modelling, 472(??):??, October 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002034.

Hughes:2023:HSM

[HK23] Conchúr Hughes and Jonathan W. King. Habitat suitability modelling for an integrated multi-trophic aquaculture (IMTA) system along Europe’s Atlantic coast. Ecological Modelling, 484(??):??, October 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001904.

Han:2020:PEW

[HKB^{+}20] Yi Han, Nadiah P. Kristensen, Yvonne M. Buckley, Dion J. Maple, Judith West, and Eve McDonald-Madden. Predicting the ecosystem-wide impacts of eradication with limited information using a qualitative modelling approach. Ecological Modelling, 430(??):??, August 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301940.

Hatlauf:2024:SBL

[HKG^{+}24] J. Hatlauf, F. Kunz, P. Griesberger, F. Sachser, and K. Hackländer. A stage-based life cycle implementation for individual-based population viability analyses of grey wolves (Canis lupus) in Europe. Ecological Modelling, 491(??):??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000887.

Haddadchi:2020:PRF

[HKH^{+}20] Arman Haddadchi, Anika Kuczynski, Joanna T. Hoyle, Cathy Kilroy, Doug J. Booker, and Murray Hicks. Periphyton removal flows determined by sediment entrainment thresholds. Ecological Modelling, 434(??):??, October 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303331.

Haddon:2023:LSM

[HKH^{+}23] Antoine Haddon, Loïc Kechichian, Jérôme Harmand, Cyril Dejean, and Nassim Ait-Mouheb. Linking soil moisture sensors and crop models for irrigation management. Ecological Modelling, 484(??):??, October 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002004.

Hata:2023:ISC

[HKSM23] Yoshiaki Hata, Tomo’omi Kumagai, Takanori Shimizu, and Yoshiyuki Miyazawa. Implications of seasonal changes in photosynthetic traits and leaf area index for canopy CO_{2} and H_{2}O fluxes in a Japanese cedar (Cryptomeria japonica D. Don) plantation. Ecological Modelling, 477(??):??, March 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022003696.

Haas:2024:IFC

[HKY24] Henrique Haas, Latif Kalin, and Haw Yen. Improved forest canopy evaporation leads to better predictions of ecohydrological processes. Ecological Modelling, 489(??):??, March 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024000097.

Henderson:2023:MSD

[HL23] Kirsten Henderson and Michel Loreau. A model of Sustainable Development Goals: Challenges and opportunities in promoting human well-being and environmental sustainability. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022002654.

Haramboure:2020:MCA

[HLB^{+}20] Marion Haramboure, Pierrick Labbé, Thierry Baldet, David Damiens, Louis Clément Gouagna, Jérémy Bouyer, and Annelise Tran. Modelling the control of Aedes albopictus mosquitoes based on sterile males release techniques in a tropical environment. Ecological Modelling, 424(??):??, May 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020300740.

Huang:2021:SAB

[HLH^{+}21] Chao Huang, Yu Liang, Hong S. He, Mia M. Wu, Bo Liu, and Tianxiao Ma. Sensitivity of aboveground biomass and species composition to climate change in boreal forests of Northeastern China. Ecological Modelling, 445(??):??, April 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021000442.

Hu:2020:SCF

[HLM^{+}20] Junmei Hu, Gengyuan Liu, Fanxin Meng, Yuanchao Hu, and Marco Casazza. Subnational carbon flow pattern analysis using multi-scale input-output model. Ecological Modelling, 431(??):??, September 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002030209X.

Huang:2023:URH

[HLS^{+}23] Yu Huang, Zenghui Li, Changyang Sun, Zhiwei Feng, Jingyao Li, Danyi Wei, Bin Wang, Shang Jiang, Keliang Chen, and Xiang Sun. Using the roughness height and Manning number in hydrodynamic model to estimate the impact of intensive oyster aquaculture by floating & fixed rafts on water exchange with an application in Qinzhou Bay, China. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003283.

Han:2020:QWR

[HLW^{+}20] Zhi-Quan Han, Tong Liu, Tao Wang, Hua-Feng Liu, Xiao-Ran Hao, Yi-Neng Ouyang, Bo Zheng, and Bai-Lian Li. Quantification of water resource utilization efficiency as the main driver of plant diversity in the water-limited ecosystems. Ecological Modelling, 429(??):??, August 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300466.

Han:2022:NSA

[HLZ^{+}22] Zhi-Quan Han, Tong Liu, Wen-Xuan Zhao, Han-Yue Wang, Qin-Ming Sun, Hui Sun, and Bai-Lian Li. A new species abundance distribution model including the hydrological niche differentiation in water-limited ecosystems. Ecological Modelling, 470(??):??, August 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002200120X.

Horemans:2020:ITV

[HMC20] Dante M. L. Horemans, Patrick Meire, and Tom J. S. Cox. The impact of temporal variability in light-climate on timeaveraged primary production and a phytoplankton bloom in a well-mixed estuary. Ecological Modelling, 436(??):??, November 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303574.

Holder:2020:PGD

[HMDO20] Anna M. Holder, Arev Markarian, Jessie M. Doyle, and John R. Olson. Predicting geographic distributions of fishes in remote stream networks using maximum entropy modeling and landscape characterizations. Ecological Modelling, 433(??):??, October 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002030301X.

Haberle:2020:DEB

[HMGK20] Ines Haberle, Nina Marn, Sunčana Geček, and Tin Klanjšček. Dynamic energy budget of endemic and critically endangered bivalve Pinna nobilis: a mechanistic model for informed conservation. Ecological Modelling, 434(??):??, October 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020302775.

Heinichen:2022:ITD

[HML^{+}22] Margaret Heinichen, M. Conor McManus, Sean M. Lucey, Kerim Aydin, Austin Humphries, Anne Innes-Gold, and Jeremy Collie. Incorporating temperature-dependent fish bioenergetics into a Narragansett Bay food web model. Ecological Modelling, 466(??):??, April 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000357.

Hamza:2021:BMS

[HMVR21] Faseela Hamza, Anju M., Vinu Valsala, and Smitha B. R. A bioenergetics model for seasonal growth of Indian oil sardine (Sardinella longiceps) in the Indian west coast. Ecological Modelling, 456(??):??, September 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002209.

Huaylla:2021:DIM

[HNC^{+}21] Claudia A. Huaylla, Marcos E. Nacif, Carolina Coulin, Marcelo N. Kuperman, and Lucas A. Garibaldi. Decoding information in multilayer ecological networks: the keystone species case. Ecological Modelling, 460(??):??, November 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002842.

Hopper:2023:MEP

[Hop23] Keith R. Hopper. Modeling the effects of plant resistance, herbivore virulence, and parasitism, on the population dynamics of aphids and parasitoids in wheat and soybean in different climates. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001072.

Hinrichsen:2020:LCC

[HP20] Richard A. Hinrichsen and Charles M. Paulsen. Low carrying capacity a risk for threatened Chinook salmon. Ecological Modelling, 432(??):??, September 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302933.

Hemati:2020:SSN

[HPMB20] Touraj Hemati, Sharareh Pourebrahim, Masoud Monavari, and Akbar Baghvand. Species-specific nature conservation prioritization (a combination of MaxEnt, Costing Nature and DINAMICA EGO modeling approaches). Ecological Modelling, 429(??):??, August 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301654.

Hao:2024:GSA

[HRW^{+}24] Shirui Hao, Dongryeol Ryu, Andrew W. Western, Eileen Perry, Heye Bogena, and Harrie Jan Hendricks Franssen. Global sensitivity analysis of APSIM-wheat yield predictions to model parameters and inputs. Ecological Modelling, 487 (??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380023002818.

Hood:2021:CBP

[HSD^{+}21] Raleigh R. Hood, Gary W. Shenk, Rachel L. Dixon, Sean M. C. Smith, William P. Ball, Jesse O. Bash, Rich Batiuk, Kathy Boomer, Damian C. Brady, Carl Cerco, Peter Claggett, Kim de Mutsert, Zachary M. Easton, Andrew J. Elmore, Marjorie A. M. Friedrichs, Lora A. Harris, Thomas F. Ihde, Lara Lacher, Li Li, Lewis C. Linker, Andrew Miller, Julia Moriarty, Gregory B. Noe, George E. Onyullo, Kenneth Rose, Katie Skalak, Richard Tian, Tamie L. Veith, Lisa Wainger, Donald Weller, and Yinglong Joseph Zhang. The Chesapeake Bay program modeling system: Overview and recommendations for future development. Ecological Modelling, 456(??):??, September 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021001964.

Han:2020:HSY

[HSF^{+}20] Yong Han, Zhiyu Sun, Hongwei Fang, Sen Bai, Lei Huang, and Guojian He. Habitat succession of the Yangtze finless porpoise in Poyang Lake under the changing hydrodynamic and feeding environment. Ecological Modelling, 424(??):??, May 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301071.

Hanson:2020:PLS

[HSJ^{+}20] Paul C. Hanson, Aviah B. Stillman, Xiaowei Jia, Anuj Karpatne, Hilary A. Dugan, Cayelan C. Carey, Jemma Stachelek, Nicole K. Ward, Yu Zhang, Jordan S. Read, and Vipin Kumar. Predicting lake surface water phosphorus dynamics using process-guided machine learning. Ecological Modelling, 430(??):??, August 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302076.

Humphreys:2022:GEA

[HSL^{+}22] John M. Humphreys, Robert B. Srygley, Douglas Lawton, Amy R. Hudson, and David H. Branson. Grasshoppers exhibit asynchrony and spatial non-stationarity in response to the El Niño/Southern and Pacific Decadal Oscillations. Ecological Modelling, 471(??):??, September 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001533.

Herrera:2022:URF

[HSP^{+}22] H. Herrera, L. Schütz, W. Paas, P. Reidsma, and B. Kopainsky. Understanding resilience of farming systems: Insights from system dynamics modelling for an arable farming system in The Netherlands. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003860.

Huang:2023:FGE

[HWS^{+}23] Jiaqing Huang, Linlin Wang, Abu Bakkar Siddik, Zulkiflee Abdul-Samad, Arpit Bhardwaj, and Bharat Singh. Forecasting GHG emissions for environmental protection with energy consumption reduction from renewable sources: a sustainable environmental system. Ecological Modelling, 475(??):??, January 2023. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002794.

He:2024:SHE

[HYY^{+}24] Hui He, Xiangke Yu, Hui Yu, Zhigang Ma, Yong Luo, Tao Liu, Ziwei Rong, Jinhong Xu, Dianpeng Chen, Pengshang Li, Fuxin Yuan, and Yongtao Zhao. Suitable habitat evaluation and ecological security pattern optimization for the ecological restoration of Giant Panda habitat based on nonstationary factors and MCR model. Ecological Modelling, 494(??):??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001480.

Han:2022:IBC

[HZ22] Yue Han and Yuntao Zhou. Investigating biophysical control of marine phytoplankton dynamics via Bayesian mechanistic modeling. Ecological Modelling, 474(??):??, December 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022002691.

Issakhov:2023:NAS

[IA23] Alibek Issakhov and Aizhan Abylkassymova. Numerical analysis of solid barrier heights and trees with porosity properties influence on the automobile’s emission dispersion in the residential area. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001266.

Irauschek:2021:EFF

[IBB^{+}21] Florian Irauschek, Ivan Barka, Harald Bugmann, Benoit Courbaud, Che Elkin, Tomáš Hlásny, Matija Klopcic, Marco Mina, Werner Rammer, and Manfred J. Lexer. Evaluating five forest models using multi-decadal inventory data from mountain forests. Ecological Modelling, 445(??):??, April 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021000648. See corrigen-

dum [IBB^{+}22].

Irauschek:2022:CEF

[IBB^{+}22] Florian Irauschek, Ivan Barka, Harald Bugmann, Benoit Courbaud, Che Elkin, Tomáš Hlásny, Matija Klopcic, Marco Mina, Werner Rammer, and Manfred J. Lexer. Corrigendum to evaluating five forest models using multidecadal inventory data from mountain forests’. Ecological Modelling, 470(??):??, August 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001235. See [IBB^{+}21].

Inoue:2024:ESR

[II24] Naoto K. Inoue and Takashi Ishihara. Estimation of survival rates of loggerhead sea turtles (Caretta caretta) in Japan using a novel framework. Ecological Modelling, 493(??):??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001509.

Ikhumhen:2020:AND

[ILLM20] Harrison Odion Ikhumhen, Tianxin Li, Shanlong Lu, and Nametso Matomela. Assessment of a novel data driven habitat suitability ranking approach for Larus relictus specie using remote sensing and GIS. Ecological Modelling, 432(??):??, September 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002030291X.

Inoue:2022:QEE

[Ino22] Naoto K. Inoue. Quantitative evaluation of the effects of bycatch on native species using mathematical models.

Ecological Modelling, 474(??):??, December 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002200254X.

Jager:2020:RSD

[Jag20] Tjalling Jager. Revisiting simplified DEBtox models for analysing ecotoxicity data. Ecological Modelling, 416(??): ??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019304120.

Jurumai:2023:IPG

[JASW23] La Pande Jurumai, Afri Ahyarky Abidin, Endang Triningsih Soetrisno, and Dwiprayogo Wibowo. Impact of population growth and housing development on the riverine environment: Identifying environmental threat and solution in the Wanggu River, Indonesia. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002703.

Jellouli:2022:IDW

[JB22] Omar Jellouli and Abdes Samed Bernoussi. The impact of dynamic wind flow behavior on forest fire spread using cellular automata: Application to the watershed BOUKHALEF (Morocco). Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000606.

Jimenez:2022:BAC

[JBF22] María Fernanda Jiménez, Gamaliel Blé, and Manuel Falconi. A biocontrol agent as a commensal in a plant-pest interaction. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000758.

Jager:2023:USD

[JGG23] Tjalling Jager, Benoit Goussen, and André Gergs. Using the standard DEB animal model for toxicokinetic-toxicodynamic analysis. Ecological Modelling, 475(??):??, January 2023. CO-

DEN ???? ISSN ???? URL http://www.sciencedirect.
com/science/article/pii/S030438002200285X.

Joks:2023:SMF

[JHKT^{+}23] Madli Jõks, Aveliina Helm, Liis Kasari-Toussaint, Ene Kook, Reimo Lutter, Norbertas Noreika, Ede Oja, Maarja Öpik, Tiina Randlane, Ülle Reier, Kersti Riibak, Andres Saag, Hardi Tullus, Tea Tullus, and Meelis Pärtel. A simulation model of functional habitat connectivity demonstrates the importance of species establishment in older forests. Ecological Modelling, 481(??):??, July 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000893.

Jin:2023:TOM

[JHL23] Ming Jin, Xulong Han, and Mingyu Li. Trade-offs of multiple urban ecosystem services based on land-use scenarios in the Tumen River cross-border area. Ecological Modelling, 482(??): ??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000960.

Jian:2021:NMQ

[Jia21] Fuji Jian. A novel model to quantify ages of organisms and predict development time distribution of their growth stages. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304555.

Jha:2022:COM

[JJN22] Ashish Jha, Praveen J., and P. O. Nameer. Contrasting occupancy models with presence-only models: Does accounting for detection lead to better predictions? Ecological Modelling, 472(??):??, October 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002083.

Jiabu:2023:IDI

[JL23] Duojie Jiabu and Weide Li. Impact of different invasion methods of invasive species on omnivorous food webs. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003155.

Jafarov:2021:LHS

[JLH^{+}21] Elchin E. Jafarov, Louise E. Loudermilk, Kevin J. Hiers, Brett Williams, Rodman Linn, Chas Jones, Samantha C. Hill, and Adam L. Atchley. Linking habitat suitability with a longleaf pine-hardwood model: Building a species-predictive fire-land management framework. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304518.

Jin:2023:SAR

[JLLS23] Hong Jin, Heping Li, Jia Lee, and Weitong Sun. Simulation analysis of rural land use using rate of change driven by population and economic dynamics — a case study of Huangguashan village in Chongqing, China. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003027.

Jha:2021:MOC

[JLV^{+}21] Pramod Jha, Brij Lal Lakaria, AK Vishwakarma, RH Wanjari, M. Mohanty, Nishant K. Sinha, J. Somasundaram, GS Dheri, AK Dwivedi, Raj Paul Sharma, Muneshwar Singh, RC Dalal, AK Biswas, AK Patra, and SK Chaudhari. Modeling the organic carbon dynamics in long-term fertilizer experiments of India using the Rothamsted carbon model. Ecological Modelling, 450(??):??, June 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001320.

Johnson:2022:EDM

[JM22] Bethany Johnson and Stephan B. Munch. An empirical dynamic modeling framework for missing or irregular samples. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000680.

Joetzjer:2022:ETD

[JMC^{+}22] Emilie Joetzjer, Fabienne Maignan, Jérôme Chave, Daniel Goll, Ben Poulter, Jonathan Barichivich, Isabelle Maréchaux, Sebastiaan Luyssaert, Matthieu Guimberteau, Kim Naudts, Damien Bonal, and Philippe Ciais. Effect of tree demography and flexible root water uptake for modeling the carbon and water cycles of Amazonia. Ecological Modelling, 469(??): ??, July 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000862.

Jager:2022:TSE

[JMHH22] Tjalling Jager, Arne M. Malzahn, Andreas Hagemann, and Bjørn Henrik Hansen. Testing a simple energybudget model for yolk-feeding stages of cleaner fish. Ecological Modelling, 469(??):??, July 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001168.

Johnston:2020:ESD

[JMM^{+}20] Alison Johnston, Nick Moran, Andy Musgrove, Daniel Fink, and Stephen R. Baillie. Estimating species distributions from spatially biased citizen science data. Ecological Modelling, 422(??):??, April 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019304351.

Jiang:2024:MSC

[JMPB24] Jiachen Jiang, Almas Mukhametov, Alla Philippova, and Alexander Bakshtanin. Modeling soil carbon accumulation in irrigated agricultural systems. Ecological Modelling, 491(??): ??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000528.

Jessen:2023:LHE

[JOE23] Henrik H. Jessen, Anders F. Opdal, and Katja Enberg. Life-history evolution in response to foraging risk, modelled for Northeast Arctic cod (Gadus morhua). Ecological Modelling, 482(??):??, August 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001096.

Jordan:2022:NPV

[Jor22] Ferenc Jordán. The network perspective: Vertical connections linking organizational levels. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002149.

Ji:2021:ITV

[JPL21] Zhonglin Ji, Yaozhong Pan, and Nan Li. Integrating the temperature vegetation dryness index and meteorology parameters to dynamically predict crop yield with fixed date intervals using an integral regression model. Ecological Modelling, 455(??):??, September 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002118.

James:2023:AIB

[JPMA^{+}23] M. K. James, J. A. Polton, C. G. Mayorga-Adame, K. L. Howell, and A. M. Knights. Assessing the influence of behavioural parameterisation on the dispersal of larvae in marine systems. Ecological Modelling, 476(??):??, February 2023. CODEN EC-

MODT. ISSN ???? URL http://www.sciencedirect.com/
science/article/pii/S0304380022003507.

Junior:2022:EBP

[JRD^{+}22] Emerson Campos Barbosa Júnior, Vitor Passos Rios, Pavel Dodonov, Bruno Vilela, and Hilton F. Japyassú. Effect of behavioural plasticity and environmental properties on the resilience of communities under habitat loss and fragmentation. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001788.

Jimenez:2022:EFN

[JS22] L. Jiménez and J. Soberón. Estimating the fundamental niche: Accounting for the uneven availability of existing climates in the calibration area. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003665.

Joy:2022:FSM

[JSD^{+}22] Ruth Joy, Robert S. Schick, Michael Dowd, Tetyana Margolina, John E. Joseph, and Len Thomas. A fine-scale marine mammal movement model for assessing long-term aggregate noise exposure. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003422.

Joseph:2020:SEM

[JSL^{+}20] Julien Joseph, Fernanda Santibáñez, María Fabiana Laguna, Guillermo Abramson, Marcelo N. Kuperman, and Lucas A. Garibaldi. A spatially extended model to assess the role of landscape structure on the pollination service of Apis mellifera. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302714.

Joly:2022:ADM

[JST^{+}22] Frédéric Joly, Rodolphe Sabatier, Laurent Tatin, Claire Mosnier, Ariell Ahearn, Marc Benoit, Bernard Hubert, and Guillaume Deffuant. Adaptive decision-making on stocking rates improves the resilience of a livestock system exposed to climate shocks. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003446.

Jacobs:2024:LEM

[JTH24] Bas Jacobs, Hilde Tobi, and Geerten M. Hengeveld. Linking error measures to model questions. Ecological Modelling, 487(??):??, January 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002922.

Jeantet:2021:FCN

[JVGC21] Lorène Jeantet, Vincent Vigon, Sébastien Geiger, and Damien Chevallier. Fully convolutional neural network: a solution to infer animal behaviours from multi-sensor data. Ecological Modelling, 450(??):??, June 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001253.

Jacobs:2024:CAS

[JvVvHH24] Bas Jacobs, George van Voorn, Peter van Heijster, and Geerten M. Hengeveld. Consequences of alternative stable states for short-term model-based control of cyanobacterial blooms. Ecological Modelling, 491(??):??, May 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000590.

Jiang:2023:AMP

[JZHT23] Baoying Jiang, Guirong Zhang, Ning He, and Ruiqi Tan. Analytical model for pullout behavior of root system. Ecological Modelling, 479(??):??, May 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000248.

Johnson:2020:UIP

[JZJ^{+}20] Fred A. Johnson, Guthrie S. Zimmerman, Gitte H. Jensen, Kevin K. Clausen, Morten Frederiksen, and Jesper Madsen. Using integrated population models for insights into monitoring programs: an application using pink-footed geese. Ecological Modelling, 415(??):??, January 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303771.

Janiak:2021:AMP

[JZS^{+}21] Kamil Janiak, Bartosz Ziȩba, Ryszard Szetela, Mateusz Muszyński-Huhajło, Stanisław Miodoński, Anna Jurga, and Agnieszka Trusz. Approach to modeling protozoa grazing on the basis of the current state of knowledge. Ecological Modelling, 447(??):??, May 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000740.

Karim:2023:RVM

[KAB23a] Md Aktar Ul Karim, Vikram Aithal, and Amiya Ranjan Bhowmick. Random variation in model parameters: a comprehensive review of stochastic logistic growth equation. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002053.

Khurana:2023:IEM

[KAB^{+}23b] Swamini Khurana, Rose Abramoff, Elisa Bruni, Marta Dondini, Boris Tupek, Bertrand Guenet, Aleksi Lehtonen, and Stefano Manzoni. Interactive effects of microbial functional diversity and carbon availability on decomposition — a theoretical exploration. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002375.

Keevil:2021:MSV

[KABL21] Matthew G. Keevil, Doug P. Armstrong, Ronald J. Brooks, and Jacqueline D. Litzgus. A model of seasonal variation in somatic growth rates applied to two temperate turtle species. Ecological Modelling, 443(??):??, March 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000272.

Kalen:2022:UCD

[KAMS22] Christer Kalén, Henrik Andrén, Johan Månsson, and Håkan Sand. Using citizen data in a population model to estimate population size of moose (Alces alces). Ecological Modelling, 471(??):??, September 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001740.

Kuczynski:2022:KMA

[KAT^{+}22] Anika Kuczynski, Martin T. Auer, William D. Taylor, Steven C. Chapra, and Marcel Dijkstra. Keeping up with the math: Advancing the ecological foundation of the Great Lakes Cladophora model. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002198.

Kuczynski:2020:CEF

[KBAC20] Anika Kuczynski, Ankita Bakshi, Martin T. Auer, and Steven C. Chapra. The canopy effect in filamentous algae: Improved modeling of Cladophora growth via a mechanistic representation of self-shading. Ecological Modelling, 418(??):??, February 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019304144.

Kumar:2023:QEI

[KBSS23] Rohit Kumar, Arvind Bhardwaj, Lakhwinder Pal Singh, and Gurraj Singh. Quantifying ecological impacts: a comparative life cycle assessment of conventional and organic potato cultivation. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002405.

Kilfoil:2021:ISB

[KCHZ21] James P. Kilfoil, Matthew D. Campbell, Michael R. Heithaus, and Yuying Zhang. The influence of shark behavior and environmental conditions on baited remote underwater video survey results. Ecological Modelling, 447(??):??, May 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021000788.

Keesen:2023:NAO

[KCP^{+}23] Fabíola Keesen, Alcides Castro e Silva, Carlos Felipe S. Pinheiro, Everaldo Arashiro, Yan Ligeiro, and Carlos Eduardo de Viveiros Grelle. New applications of an old individual-based model for biological dynamics. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003325.

Kibria:2022:MCA

[KCS22] Abu SMG Kibria, Robert Costanza, and José R. Soto. Modeling the complex associations of human wellbeing dimensions in a coupled human-natural system: In contexts of marginalized communities. Ecological Modelling, 466(??):??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022000114.

Koch:2020:EII

[KD20] Josef Koch and Karel A. C. De Schamphelaere. Estimating inter-individual variability of dynamic energy budget model parameters for the copepod Nitocra spinipes from existing life-history data. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301630.

Kundu:2021:GAK

[KDC^{+}21] Sayani Kundu, Nirjhar Dasgupta, Bratati Chakraborty, Ayan Paul, Santanu Ray, and Sabyasachi Bhattacharya. Growth acceleration is the key for identifying the most favorable food concentration of Artemia sp. Ecological Modelling, 455(??):??, September 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002100199X.

Katin:2021:SAD

[KDH^{+}21] Alexey Katin, Dario Del Giudice, Nathan S. Hall, Hans W. Paerl, and Daniel R. Obenour. Simulating algal dynamics within a Bayesian framework to evaluate controls on estuary productivity. Ecological Modelling, 447(??):??, May 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021000685.

Khelifa:2024:IAH

[KE24] Amira Khelifa and Nadjia El Saadi. The impact of aquatic habitats on the malaria parasite transmission: a view from an agent-based model. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002776.

Kolpas:2020:PMP

[KFJS20] Allison Kolpas, David H. Funk, John K. Jackson, and Bernard W. Sweeney. Phenological modeling of the parthenogenetic mayfly Neocloeon triangulifer (Ephemeroptera: Baetidae) in White Clay Creek. Ecological Modelling, 416(??): ??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019304004.

Kamkar:2023:EHS

[KFM^{+}23] Behnam Kamkar, Mohammad Taghi Feyzbakhsh, Hassan Mokhtarpour, Jelena Barbir, Jasmin Grahić, Sylwester Tabor, and Hossein Azadi. Effect of heat stress during anthesis on the Summer Maize grain formation: Using integrated modelling and multi-criteria GIS-based method. Ecological Modelling, 481(??):??, July 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000467.

Klagkou:2024:DEB

[KGBL24] Evridiki Klagkou, Andre Gergs, Christian U. Baden, and Konstadia Lika. Dynamic Energy Budget approach for modeling growth and reproduction of neotropical stink bugs. Ecological Modelling, 493(??):??, July 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024001285.

Krause:2022:HMS

[KHH^{+}22] Jens Krause, Mike Harfoot, Selwyn Hoeks, Peter Anthoni, Calum Brown, Mark Rounsevell, and Almut Arneth. How more sophisticated leaf biomass simulations can increase the realism of modelled animal populations. Ecological Modelling, 471(??):??, September 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0304380022001697. See corrigendum [KHH^{+}24].

Krause:2024:CHM

[KHH^{+}24] Jens Krause, Mike Harfoot, Selwyn Hoeks, Peter Anthoni, Calum Brown, Mark Rounsevell, and Almut Arneth. Corrigendum to “How more sophisticated leaf biomass simulations can increase the realism of modelled animal populations” [Ecological Modelling 471 (2022) 110061]. Ecological Modelling, 492(??):??, June 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000942. See [KHH^{+}22].

Kane:2023:WAR

[KHR23] Adam Kane, Kevin Healy, and Graeme D. Ruxton. Was Allosaurus really predominantly a scavenger? Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003453. See response [PR23].

Kong:2023:AAW

[KHS^{+}23] Yang Kong, Weijun He, Juqin Shen, Liang Yuan, Xin Gao, Thomas Stephen Ramsey, Qingling Peng, Dagmawi Mulugeta Degefu, and Fuhua Sun. Adaptability analysis of water pollution and advanced industrial structure in Jiangsu Province, China. Ecological Modelling, 481(??):??, July 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000935.

Kakehi:2020:FPS

[KiAM^{+}20] Shigeho Kakehi, Jun ichi Abo, Hiroomi Miyamoto, Taiki Fuji, Kazuyoshi Watanabe, Hideyuki Yamashita, and Satoshi Suyama. Forecasting Pacific saury (Cololabis saira) fishing grounds off Japan using a migration model driven by an ocean circulation model. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302210.

Kooi:2020:CPM

[KK20] B. W. Kooi and S. A. L. M. Kooijman. A cohort projection method to follow DEB-structured populations with periodic, synchronized and iteroparous reproduction. Ecological Modelling, 436(??):??, November 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303689.

Kim:2023:AFS

[KKF^{+}23] Seokmin Kim, Anthony Koop, Glenn Fowler, Kimberly Israel, Yu Takeuchi, and Deah Lieurance. Addition of finer scale data and uncertainty analysis increases precision of geospatial suitability model for non-native plants in the US. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001898.

Kokkonen:2024:ETP

[KKH^{+}24] Eevi Kokkonen, Mikael Kuisma, Pekka Hyvärinen, Anssi Vainikka, Kristiina Vuorio, Tommi Perälä, Laura S. Härkönen, Satu Estlander, and Anna Kuparinen. Effects of top predator re-establishment and fishing on a simulated food web: Allometric Trophic Network model for Lake Oulujärvi. Ecological Modelling, 492(??):??, June 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024001030.

Kherif:2021:AMN

[KKP^{+}21] Omar Kherif, Mohamed Islam Keskes, Marc Pansu, Walid Ouaret, Yacer-Nazih Rebouh, Peter Dokukin, Dmitry Kucher, and Mourad Latati. Agroecological modeling of nitrogen and carbon transfers between decomposer micro-organisms, plant symbionts, soil and atmosphere in an intercropping system. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304543.

Kwon:2020:MCM

[KKPR20] Young Shin Kwon, Hyoun-Woo Kang, Luca Polimene, and Tae Siek Rhee. A marine carbon monoxide (CO) model with a new parameterization of microbial oxidation. Ecological Modelling, 432(??):??, September 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302738.

Kim:2021:UML

[KKW^{+}21] Tae Kyung Kim, Sukyung Kim, Myoungsoo Won, Jong-Hwan Lim, Sukhee Yoon, Keunchang Jang, Kye-Han Lee, Yeong Dae Park, and Hyun Seok Kim. Utilizing machine learning for detecting flowering in mid — range digital repeat photography. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304774.

Kooijman:2020:EBP

[KLA^{+}20] Sebastiaan A. L. M. Kooijman, Konstadia Lika, Starrlight Augustine, Nina Marn, and Bob W. Kooi. The energetic basis of population growth in animal kingdom. Ecological Modelling, 428(??):??, July 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301277.

Kleinhans:2023:VCC

[Kle23] Maarten G. Kleinhans. Viewpoint: Composing complex earth system models. Ecological Modelling, 482(??):??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001345.

Kim:2024:DFL

[KLH^{+}24] Yongeun Kim, Minyoung Lee, Jinsol Hong, Yun-Sik Lee, June Wee, and Kijong Cho. Development of a fuzzy logicembedded system dynamics model to simulate complex socioecological systems. Ecological Modelling, 493(??):??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001261.

Kunegel-Lion:2022:HQD

[KLNMG22] Mélodie Kunegel-Lion, Eric W. Neilson, Nicolas Mansuy, and Devin W. Goodsman. Habitat quality does not predict animal population abundance on frequently disturbed landscapes. Ecological Modelling, 469(??):??, July 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000643.

Kreig:2024:ABM

[KLPJ24] Jasmine A. F. Kreig, Suzanne Lenhart, Eduardo Ponce, and Henriette I. Jager. Agent-based modeling to evaluate the effects of harvesting biomass and hunting on ringnecked pheasant (Phasianus colchicus) populations. Ecological Modelling, 492(??):??, June 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000930.

Kaitaniemi:2020:PLE

[KLS20] Pekka Kaitaniemi, Anna Lintunen, and Risto Sievänen. Power-law estimation of branch growth. Ecological Modelling, 416(??):??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019304089.

Kazanci:2023:RIE

[KMAA23] Caner Kazanci, Qianqian Ma, Aladeen Al Basheer, and Asma Azizi. Resilience, indirect effects and cycling in ecological networks. Ecological Modelling, 481(??):??, July 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000741.

Kalt:2021:EOS

[KMH^{+}21] Gerald Kalt, Andreas Mayer, Helmut Haberl, Lisa Kaufmann, Christian Lauk, Sarah Matej, Elin Röös, Michaela C. Theurl, and Karl-Heinz Erb. Exploring the option space for land system futures at regional to global scales: the diagnostic agro-food, land use and greenhouse gas emission model BioBaM-GHG 2.0. Ecological Modelling, 459(??):??, November 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021002817.

Kiziridis:2023:IPP

[KMP^{+}23] Diogenis A. Kiziridis, Anna Mastrogianni, Magdalini Pleniou, Spyros Tsiftsis, Fotios Xystrakis, and Ioannis Tsiripidis. Improving the predictive performance of CLUE-S by extending demand to land transitions: the trans-CLUE-S model. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000352.

Katz:2020:MMH

[KN20] Keziah Katz and Dhruba Naug. A mechanistic model of how metabolic rate can interact with resource environment to influence foraging success and lifespan. Ecological Modelling, 416(??):??, January 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019304077.

Khangaonkar:2021:PAE

[KNP^{+}21] Tarang Khangaonkar, Adi Nugraha, Lakshitha Premathilake, Julie Keister, and Amy Borde. Projections of algae, eelgrass, and zooplankton ecological interactions in the inner Salish Sea — for future climate, and altered oceanic states. Ecological Modelling, 441(??):??, February 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304786.

Kooijman:2020:CEP

[Koo20a] Sebastiaan A. L. M. Kooijman. The comparative energetics of petrels and penguins. Ecological Modelling, 427(??):??, July 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301241.

Kooijman:2020:SDE

[Koo20b] Sebastiaan A. L. M. Kooijman. The standard dynamic energy budget model has no plausible alternatives. Ecological Modelling, 428(??):??, July 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301782.

Kooijman:2024:WRA

[Koo24a] S. A. L. M. Kooijman. Ways to reduce or avoid juveniledriven cycles in individual-based population models. Ecological Modelling, 490(??):??, April 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000383.

Kooijman:2024:CEB

[Koo24b] Sebastiaan A. L. M. Kooijman. The comparative energetics of branchiopods: Adaptations to volatile environments. Ecological Modelling, 492(??):??, June 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001091.

Kosicki:2022:NSL

[Kos22] Jakub Z. Kosicki. Niche segregation on the landscape scale of two co-existing related congeners in the sympatric zone — modelling approach. Ecological Modelling, 468(??):??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022000783.

Krishna:2024:IEV

[KPLH24] Shubham Krishna, Victoria Peterson, Luisa Listmann, and Jana Hinners. Interactive effects of viral lysis and warming in a coastal ocean identified from an idealized ecosystem model. Ecological Modelling, 487(??):??, January 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002806.

Karppinen:2022:ITU

[KRM^{+}22] Santeri Karppinen, Tuomas Rajala, Samu Mäntyniemi, Ilpo Kojola, and Matti Vihola. Identifying territories using presence-only citizen science data: an application to the Finnish wolf population. Ecological Modelling, 472(??):??, October 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002046.

Kaare-Rasmussen:2023:MFD

[KRMP23] Jakob O. Kaare-Rasmussen, Holly V. Moeller, and Ferdinand Pfab. Modeling food dependent symbiosis in Exaiptasia pallida. Ecological Modelling, 481(??):??, July 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000534.

Kjaer:2022:ELT

[KS22] Lene J. Kjær and Eric M. Schauber. The effect of landscape, transmission mode and social behavior on disease transmission: Simulating the transmission of chronic wasting disease in white-tailed deer (Odocoileus virginianus) populations using a spatially explicit agent-based model. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002162.

Kim:2022:HAB

[KSJ22] Jaeyoung Kim, Dongil Seo, and John R. Jones. Harmful algal bloom dynamics in a tidal river influenced by hydraulic control structures. Ecological Modelling, 467(??):??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022000540.

Karami:2022:IAA

[KT22] Peyman Karami and Sajad Tavakoli. Identification and analysis of areas prone to conflict with wild boar (Sus scrofa) in the vineyards of Malayer County, western Iran. Ecological Modelling, 471(??):??, September 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001491.

Kouadio:2021:PPB

[KTB^{+}21] Louis Kouadio, Philippe Tixier, Vivekananda Byrareddy, Torben Marcussen, Shahbaz Mushtaq, Bruno Rapidel, and Roger Stone. Performance of a process-based model for predicting robusta coffee yield at the regional scale in Vietnam. Ecological Modelling, 443(??):??, March 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000417.

Kerr:2023:MFS

[KTB^{+}23] J. R. Kerr, J. S. Tummers, T. Benson, M. C. Lucas, and P. S. Kemp. Modelling fine scale route choice of upstream migrating fish as they approach an instream structure. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022003088.

Krishna:2021:MBD

[KUK^{+}21] Shubham Krishna, Hugo N. Ulloa, Onur Kerimoglu, Camille Minaudo, Orlane Anneville, and Alfred Wüest. Modelbased data analysis of the effect of winter mixing on primary production in a lake under reoligotrophication. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304634.

Karunarathna:2024:MNR

[KWC24] K. A. N. K. Karunarathna, Konstans Wells, and Nicholas J. Clark. Modelling nonlinear responses of a desert rodent species to environmental change with hierarchical dynamic generalized additive models. Ecological Modelling, 490(??):??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024000371.

Kerkow:2022:LCM

[KWG^{+}22a] Antje Kerkow, Ralf Wieland, Jörn M. Gethmann, Franz Hölker, and Hartmut H. K. Lentz. Linking a compartment model for West Nile virus with a flight simulator for vector mosquitoes. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002100380X.

Kirkow:2022:ICE

[KWG^{+}22b] Velizar Kirkow, Hao Wang, Pablo Venegas Garcia, Shohel Ahmed, and Christopher M. Heggerud. Impacts of a changing environment on a stoichiometric producer-grazer system: a stochastic modelling approach. Ecological Modelling, 469(??): ??, July 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000874.

Koralewski:2023:EPI

[KWG^{+}23] Tomasz E. Koralewski, Hsiao-Hsuan Wang, William E. Grant, Michael J. Brewer, and Norman C. Elliott. Error propagation in an integrated spatially-explicit individual-based model. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003131.

Keretz:2024:HMN

[KWM^{+}24] Shay S. Keretz, Daelyn A. Woolnough, Todd J. Morris, Edward F. Roseman, and David T. Zanatta. Habitat modelling of native freshwater mussels distinguishes river specific differences in the Detroit and St. Clair rivers of the Laurentian Great Lakes. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002673.

Krishna:2023:IGP

[KWP^{+}23] Dyvavani K. Krishna, Taibanganba Watham, Hitendra Padalia, Ritika Srinet, and Subrata Nandy. Improved gross primary productivity estimation using semi empirical (PRE-LES) model for moist Indian sal forest. Ecological Modelling, 475(??):??, January 2023. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002769.

Lusardi:2024:MCT

[LAC^{+}24] Léo Lusardi, Eliot André, Irene Castañeda, Sarah Lemler, Pauline Lafitte, Diane Zarzoso-Lacoste, and Elsa Bonnaud. Methods for comparing theoretical models parameterized with field data using biological criteria and Sobol analysis. Ecological Modelling, 493(??):??, July 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001169.

Lika:2020:UAL

[LAK20] Konstadia Lika, Starrlight Augustine, and Sebastiaan A. L. M. Kooijman. The use of augmented loss functions for estimating dynamic energy budget parameters. Ecological Modelling, 428(??):??, July 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301824.

Landmann:2023:TER

[LAK^{+}23] Tobias Landmann, Komi M. Agboka, Igor Klein, Elfatih M. Abdel-Rahman, Emily Kimathi, Bester T. Mudereri, Benard Malenge, Mahgoub M. Mohamed, and Henri E. Z. Tonnang. Towards early response to desert locust swarming in eastern Africa by estimating timing of hatching. Ecological Modelling, 484(??):??, October 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002065.

Lazebnik:2023:CAE

[Laz23] Teddy Lazebnik. Computational applications of extended SIR models: a review focused on airborne pandemics. Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001539.

Lee:2021:EVG

[LB21] Dong Joo Lee and Mark T. Brown. Estimating the value of global ecosystem structure and productivity: a geographic information system and emergy based approach. Ecological Modelling, 439(??):??, January 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002030377X.

Lett:2020:CAM

[LBB20] Christophe Lett, Nicolas Barrier, and Meissam Bahlali. Converging approaches for modeling the dispersal of propagules in air and sea. Ecological Modelling, 415(??):??, January 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019303667.

Leins:2021:HRP

[LBGD21] Johannes A. Leins, Thomas Banitz, Volker Grimm, and Martin Drechsler. High-resolution PVA along large environmental gradients to model the combined effects of climate change and land use timing: lessons from the large marsh grasshopper. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002030421X.

Li:2020:ERA

[LBI^{+}20] Yan Li, Vicki S. Blazer, Luke R. Iwanowicz, Megan Kepler Schall, Kelly Smalling, Donald E. Tillitt, and Tyler Wagner. Ecological risk assessment of environmental stress and bioactive chemicals to riverine fish populations: an individualbased model of smallmouth bass Micropterus dolomieu. Ecological Modelling, 438(??):??, December 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303914.

Liu:2022:MRI

[LBL22] Xiaoqian Liu, Daniel Bearup, and Jinbao Liao. Metacommunity robustness to invasion in mutualistic and antagonistic networks. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000692.

Lennox:2022:MCF

[LBN^{+}22] Robert J. Lennox, Luke T. Barrett, Cecilie I. Nilsen, Saron Berhe, Bjørn T. Barlaup, and Knut Wiik Vollset. Moving cleaner fish from the wild into fish farms: a zero-sum game? Ecological Modelling, 474(??):??, December 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002502.

Livanou:2020:MNS

[LBPL20] Eleni Livanou, Konstantinos Barsakis, Stella Psarra, and Konstadia Lika. Modelling the nutritional strategies in mixotrophic nanoflagellates. Ecological Modelling, 428(??):??, July 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301253.

Louati:2020:HEI

[LBS20] Dorra Louati, Slimane BenMiled, and Narjés Bellamine Ben Saoud. HermaDEB: an evolutionary IBM for energy allocation in hermaphrodites. Ecological Modelling, 424(??):??, May 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020300806.

Li:2022:HBA

[LBT^{+}22] Jing Li, Melissa Broussard, Nathan Tomer, Mateusz Jochym, Dilini Fonseka, Angela Peace, Linley Jesson, Nilsa A. Bosque-Pérez, David Crowder, Brad G. Howlett, and David Pattemore. Honey bee (Apis mellifera) hive placement is more influential than orchard layout on the fruit set of a dioecious crop. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001776.

Lee:2020:EVG

[LC20] Dong Joo Lee and Moon Bo Choi. Ecological value of global terrestrial plants. Ecological Modelling, 438(??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303987.

Liu:2020:MCB

[LCLG20] Bo Liu, Shuo Chen, Hui Liu, and Yueqiang Guan. Modeling cyanobacteria biomass by surface sediment diatoms in lakes: problems and suggestions. Ecological Modelling, 430(??):??, August 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301289.

Lopolito:2024:CFL

[LCM24] Antonio Lopolito, Rocco Caferra, and Piergiuseppe Morone. Contagion, fast and low: Modeling social influence in socioecological systems. Ecological Modelling, 491(??):??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024000619.

Luquin-Covarrubias:2021:ESG

[LCMB21] Marlene Anaid Luquin-Covarrubias and Enrique Morales-Bojórquez. Effects of stochastic growth on population dynamics and management quantities estimated from an integrated catch-at-length assessment model: Panopea globosa as case study. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304488.

Li:2020:ATS

[LCP^{+}20] Shangge Li, Tae-Soo Chon, Young-Seuk Park, Xiaotao Shi, and Zongming Ren. Application of temporal self-organizing maps to patterning short-time series of fish behavior responding to environmental stress. Ecological Modelling, 433(??):??, October 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303124.

Liu:2021:IVS

[LCW^{+}21] Zhenhai Liu, Bin Chen, Shaoqiang Wang, Qinyi Wang, Jinghua Chen, Weibo Shi, Xiaobo Wang, Yuanyuan Liu, Yongkai Tu, Mei Huang, Junbang Wang, Zhaosheng Wang, Hui Li, and Tongtong Zhu. The impacts of vegetation on the soil surface freezing-thawing processes at permafrost southern edge simulated by an improved process-based ecosystem model. Ecological Modelling, 456(??):??, September 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002222.

Lamonica:2020:FPP

[LDC^{+}20] Dominique Lamonica, Hilaire Drouineau, Hervé Capra, Hervé Pella, and Anthony Maire. A framework for pre-processing individual location telemetry data for freshwater fish in a river section. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020302611.

Lata:2020:TAS

[LDR^{+}20] Trevor D. Lata, Pierre A. Deymier, Keith Runge, François-Michel Le Tourneau, Régis Ferrière, and Falk Huettmann. Topological acoustic sensing of spatial patterns of trees in a model forest landscape. Ecological Modelling, 419(??):??, March 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020300363.

Lopez:2023:CRA

[LDS^{+}23] L. López, C. Dommar, A. San José, L. Meyers, S. Fox, L. Castro, and X. Rodó. Changing risk of arboviral emergence in Catalonia due to higher probability of autochthonous outbreaks. Ecological Modelling, 477(??):??, March 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022003568.

Ling:2023:EVO

[LDX^{+}23] Xiaoxia Ling, Nanyan Deng, Dongliang Xiong, Shen Yuan, Shaobing Peng, Tao Li, and Jianliang Huang. Effect of variation in the observations on the prediction uncertainty in crop model simulation: Use ORYZA (v3) as a case study. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003313.

Lu:2022:MII

[LDXJ22] Yuanming Lu, Donald L. DeAngelis, Junfei Xia, and Jiang Jiang. Modeling the impact of invasive species litter on conditions affecting its spread and potential regime shift. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000801.

Li:2021:EOC

[LDZ21] Dezhi Li, Bingzhen Du, and Jin Zhu. Evaluating old community renewal based on emergy analysis: a case study of nanjing. Ecological Modelling, 449(??):??, June 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001204.

Lang:2023:MFD

[LE23] Dengxiao Lang and Maurits W. Ertsen. Modelling farmland dynamics in response to farmer decisions using an advanced irrigation-related agent-based model. Ecological Modelling, 486(??):??, December 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002300265X.

Li:2023:ELE

[LEC^{+}23] Geng Li, Caroline M. Elliott, Bruce C. Call, Duane C. Chapman, Robert B. Jacobson, and Binbin Wang. Evaluations of Lagrangian egg drift models: From a laboratory flume to large channelized rivers. Ecological Modelling, 475(??):??, January 2023. CODEN ???? ISSN

????  URL http://www.sciencedirect.com/science/
article/pii/S0304380022002988.

Lefebvre:2023:HCD

[Lef23] Jean-Pierre Lefebvre. How can dugongs (Dugong dugon) travel along the water column at low energetic cost? A novel hypothesis. Ecological Modelling, 485(??):??, November 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002351.

Lavaud:2020:DEB

[LFN^{+}20] Romain Lavaud, Ramón Filgueira, André Nadeau, Laura Steeves, and Thomas Guyondet. A dynamic energy budget model for the macroalga Ulva lactuca. Ecological Modelling, 418(??):??, February 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019304302.

Langhammer:2020:MBD

[LG20] Maria Langhammer and Volker Grimm. Mitigating bioenergydriven biodiversity decline: a modelling approach with the European brown hare. Ecological Modelling, 416(??):??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019304223.

Lucey:2020:CRE

[LGA20] Sean M. Lucey, Sarah K. Gaichas, and Kerim Y. Aydin. Conducting reproducible ecosystem modeling using the open source mass balance model Rpath. Ecological Modelling, 427(??):??, July 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301290.

Logan:2021:QTB

[LGA^{+}21] Lauren H. Logan, Rohini S. Gupta, Amy Ando, Cory Suski, and Ashlynn S. Stillwell. Quantifying tradeoffs between electricity generation and fish populations via population habitat duration curves. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020304385.

Lambert:2020:IBM

[LGFT^{+}20] Sébastien Lambert, Emmanuelle Gilot-Fromont, Carole Toïgo, Pascal Marchand, Elodie Petit, Bruno Garin-Bastuji, Dominique Gauthier, Jean-Michel Gaillard, Sophie Rossi, and Anne Thébault. An individual-based model to assess the spatial and individual heterogeneity of Brucella melitensis transmission in Alpine ibex. Ecological Modelling, 425(??):??, June 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020300818.

Ledru:2022:SSN

[LGG^{+}22] Léo Ledru, Jimmy Garnier, Christiane Gallet, Camille Noûs, and Sébastien Ibanez. Spatial structure of natural boxwood and the invasive box tree moth can promote coexistence. Ecological Modelling, 465(??):??, March 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003835.

Logofet:2023:TYM

[LGK^{+}23] Dmitrii O. Logofet, Leonid L. Golubyatnikov, Elena S. Kazantseva, Iya N. Belova, and Nina G. Ulanova. Thirteen years of monitoring an alpine short-lived perennial: Novel methods disprove the former assessment of population viability. Ecological Modelling, 477(??):??, March 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022003064.

Liu:2024:DSM

[LGL^{+}24] Xin Liu, Ruiying Geng, Xin Lv, Xiaoke Hu, Thang Nam Do, and Meng Liu. From dynamics to sustainability: Modeling the shrimp (Penaeus chinensis) fishery in Shandong, China. Ecological Modelling, 489(??):??, March 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003307.

LeGouvello:2020:ESB

[LHDBN20] Diane Z. M. Le Gouvello, Michael G. Hart-Davis, Björn C. Backeberg, and Ronel Nel. Effects of swimming behaviour and oceanography on sea turtle hatchling dispersal at the intersection of two ocean current systems. Ecological Modelling, 431(??):??, September 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020302027.

Gouvello:2024:DCN

[LHH^{+}24] D. Z. M. Le Gouvello, S. Heye, L. R. Harris, J. Temple-Boyer, P. Gaspar, M. G. Hart-Davis, C. Louro, and R. Nel. Dispersal corridors of neonate sea turtles from dominant rookeries in the Western Indian Ocean. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002727.

Liu:2024:ACH

[LHJ^{+}24] Moyang Liu, Serena H. Hamilton, Anthony J. Jakeman, Julien Lerat, Callum Savage, and Barry F. W. Croke. Assessing the contribution of hydrologic and climatic factors on vegetation condition changes in semi-arid wetlands: an analysis for the Narran Lakes. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002983.

Liu:2022:SHT

[LHL22] Sen Liu, Boyang Han, and Weide Li. Self-healing time of population under dynamic disturbance. Ecological Modelling, 470(??):??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380022001260.

Limaki:2021:PES

[LhNA^{+}21] Mahdieh Khalatbari Limaki, Majid Es hagh Nimvari, Seyed Jalil Alavi, Asadollah Mataji, and Farid Kazemnezhad. Potential elevation shift of oriental beech (Fagus orientalis L.) in Hyrcanian mixed forest ecoregion under future global warming. Ecological Modelling, 455(??):??, September 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021001988. See corrigen-

dum [LhNA^{+}22].

Limaki:2022:CPE

[LhNA^{+}22] Mahdieh Khalatbari Limaki, Majid Es hagh Nimvari, Seyed Jalil Alavi, Asadollah Mataji, and Farid Kazemnezhad. Corrigendum to Potential elevation shift of oriental beech (Fagus orientalis L.) in Hyrcanian mixed forest ecoregion under future global warming. Ecological Modelling, 464(??):??, February 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003677. See [LhNA^{+}21].

Liao:2023:ICM

[Lia23] Qingtao Liao. Intelligent classification model of land resource use using deep learning in remote sensing images. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003295.

Luczak:2021:SDT

[ŁJ21] Aleksandra Łuczak and Małgorzata Just. Sustainable development of territorial units: MCDM approach with optimal tail selection. Ecological Modelling, 457(??):??, October 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002325.

Li:2022:MMB

[LJP^{+}22] Shangge Li, Jinfeng Jian, Rama Krishnan Poopal, Xinyu Chen, Yaqi He, Hongbin Xu, Huimin Yu, and Zongming Ren. Mathematical modeling in behavior responses: the tendency-prediction based on a persistence model on real-time data. Ecological Modelling, 464(??):??, February 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003781.

Leandro:2020:MSG

[LJRM^{+}20] Camila Leandro, Pierre Jay-Robert, Bruno Mériguet, Xavier Houard, and Ian W. Renner. Is my SDM good enough? Insights from a citizen science dataset in a point process modeling framework. Ecological Modelling, 438(??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303537.

Langston:2020:MEC

[LK20] Amy K. Langston and David A. Kaplan. Modelling the effects of climate, predation, and dispersal on the poleward range expansion of black mangrove (Avicennia germinans). Ecological Modelling, 434(??):??, October 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002030315X.

Lika:2024:MIB

[LK24a] K. Lika and S. A. L. M. Kooijman. The metabolic interpretation of the von Bertalanffy growth rate. Ecological Modelling, 488(??):??, February 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023003216.

Lika:2024:RBC

[LK24b] Konstadia Lika and Sebastiaan A. L. M. Kooijman. The relationship between confidence intervals and distributions of estimators for parameters of deterministic models. Ecological Modelling, 490(??):??, April 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000346.

Logofet:2020:SBP

[LKO20] Dmitrii O. Logofet, Elena S. Kazantseva, and Vladimir G. Onipchenko. Seed bank as a persistent problem in matrix population models: From uncertainty to certain bounds. Ecological Modelling, 438(??):??, December 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303549.

Lee:2020:ACI

[LKP^{+}20] Dong Hoon Lee, Jin Hwi Kim, Mi-Hyun Park, Michael K. Stenstrom, and Joo-Hyon Kang. Automatic calibration and improvements on an instream chlorophyll a simulation in the HSPF model. Ecological Modelling, 415(??):??, January 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019303436.

Lopucki:2022:HMD

[ŁKPK22] Rafał Łopucki, Adam Kiersztyn, Grzegorz Pitucha, and Ignacy Kitowski. Handling missing data in ecological studies: Ignoring gaps in the dataset can distort the inference. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000825.

Lee:2021:ETT

[LL21] Ying-Chieh Lee and Pei-Ting Liao. The effect of tourism on teleconnected ecosystem services and urban sustainability: an emergy approach. Ecological Modelling, 439(??):??, January 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020304105.

Lapp:2022:NAA

[LL22] Maya Lapp and Colby Long. A new approach to agentbased models of Community Resource Management1 based on the analysis of cheating, monitoring, and sanctioning. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000667.

LefebvreduPrey:2023:AFL

[LLB^{+}23] Marion Lefebvre du Prey, Jérémy Lobry, Anik Brind’Amour, Hervé Le Bris, and Bastien Sadoul. Assessing food limitation for marine juvenile fishes in coastal nurseries using a bioenergetic approach. Ecological Modelling, 482(??):??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001503.

Lavaud:2024:PRA

[LLC^{+}24] Romain Lavaud, Megan K. La Peyre, Brady Couvillion, Jennifer Beseres Pollack, Vincent Brown, Terence A. Palmer, and Barry Keim. Predicting restoration and aquaculture potential of eastern oysters through an eco-physiological mechanistic model. Ecological Modelling, 489(??):??, March 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003332.

Liao:2021:TFM

[LLLZ21] Kaihua Liao, Ligang Lv, Xiaoming Lai, and Qing Zhu. Toward a framework for the multimodel ensemble prediction of soil nitrogen losses. Ecological Modelling, 456(??):??, September 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021002337.

Lu:2021:ERA

[LLW^{+}21] Jingzhao Lu, Hongwei Lu, Weipeng Wang, SanSan Feng, and Kaiwen Lei. Ecological risk assessment of heavy metal contamination of mining area soil based on land type changes: an information network environ analysis. Ecological Modelling, 455(??):??, September 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001952.

Li:2020:NAH

[LLZM20] Lianwei Li, Wendy Li, Quan Zou, and Zhanshan (Sam) Ma. Network analysis of the hot spring microbiome sketches out possible niche differentiations among ecological guilds. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302180.

Lindh:2021:PEA

[LM21] Magnus Lindh and Stefano Manzoni. Plant evolution along the ‘fast–slow’ growth economics spectrum under altered precipitation regimes. Ecological Modelling, 448(??):??, May 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021001010.

Logofet:2023:MCR

[LM23] Dmitrii O. Logofet and Alexander A. Maslov. Markov chain retrospective analysis or how to detect a position of the monitoring period in the course of postfire succession. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002089.

Lawson:2020:ECE

[LMNV20] Laté A. Lawson, Roberto Martino, and Phu Nguyen-Van. Environmental convergence and environmental Kuznets curve: a unified empirical framework. Ecological Modelling, 437(??):??, December 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020303598.

Lin:2023:UCI

[LMW^{+}23a] Zekun Lin, Fanxin Meng, Dongfang Wang, Danqi Liao, Yutong Sun, Jiaqi Hou, Gengyuan Liu, Biagio Fernando Giannetti, Feni Agostinho, and Cecília M. V. B. Almeida. Unfolding carbon inequality across belt and road initiative countries and regions under a global trade network. Ecological Modelling, 482(??):??, August 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001424.

Lucash:2023:BTF

[LMW^{+}23b] Melissa S. Lucash, Adrienne M. Marshall, Shelby A. Weiss, John W. McNabb, Dmitry J. Nicolsky, Gerald N. Flerchinger, Timothy E. Link, Jason G. Vogel, Robert M. Scheller, Rose Z. Abramoff, and Vladimir E. Romanovsky. Burning trees in frozen soil: Simulating fire, vegetation, soil, and hydrology in the boreal forests of Alaska. Ecological Modelling, 481(??): ??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000959.

Lu:2021:HPS

[LNK^{+}21] Na Lu, Jun Niu, Shaozhong Kang, Shailesh Kumar Singh, and Taisheng Du. A hybrid PCA-SEM-ANN model for the prediction of water use efficiency. Ecological Modelling, 460(??):??, November 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003033.

Love:2024:GBA

[Lov24] Joseph W. Love. A GIS-based approach for managing catchand-release fisheries. Ecological Modelling, 491(??):??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024000620.

Lee:2022:EHW

[LP22] Sang-Hee Lee and Cheol-Min Park. The effect of hunter– wild boar interactions and landscape heterogeneity on wild boar population size: a simulation study. Ecological Modelling, 464(??):??, February 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003859.

Lombardi:2021:SAF

[LPA^{+}21] G. V. Lombardi, Silvia Parrini, R. Atzori, G. Stefani, D. Romano, M. Gastaldi, and G. Liu. Sustainable agriculture, food security and diet diversity. The case study of Tuscany, Italy. Ecological Modelling, 458(??):??, October 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002581.

Lorilla:2020:SED

[LPD^{+}20] Roxanne Suzette Lorilla, Konstantinos Poirazidis, Vassilis Detsis, Stamatis Kalogirou, and Christos Chalkias. Socioecological determinants of multiple ecosystem services on the Mediterranean landscapes of the Ionian Islands (Greece). Ecological Modelling, 422(??):??, April 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300661.

Liu:2020:CCA

[LPHO^{+}20] Huan Liu, Diego N. L. Pequeno, Ixchel M. Hernández-Ochoa, Timothy J. Krupnik, Kai Sonder, Wei Xiong, and Yinlong Xu. A consistent calibration across three wheat models to simulate wheat yield and phenology in China. Ecological Modelling, 430(??):??, August 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302040.

Lee:2022:EET

[LPL22a] Sang-Hee Lee, Cheol-Min Park, and Sang-Bin Lee. Exploring the efficiency of termite food transportation in a sinusoidalshaped tunnel. Ecological Modelling, 474(??):??, December 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022002770.

Link:2022:SIC

[LPL22b] Jason S. Link, Fabio Pranovi, and Simone Libralato. Simulations and interpretations of cumulative trophic theory. Ecological Modelling, 463(??):??, January 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003458.

Liu:2021:TMM

[LPS^{+}21] Qiuyu Liu, Changhui Peng, Robert Schneider, Dominic Cyr, Zelin Liu, Xiaolu Zhou, and Daniel Kneeshaw. TRIPLEX-Mortality model for simulating drought-induced tree mortality in boreal forests: Model development and evaluation. Ecological Modelling, 455(??):??, September 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002100212X.

Lecorvaisier:2024:UDM

[LPSF24] Florian Lecorvaisier, Dominique Pontier, Benoît Soubeyrand, and David Fouchet. Using a dynamical model to study the impact of a toxoid vaccine on the evolution of a bacterium: the example of diphtheria. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002995.

Lin:2024:MPB

[LQH24] Jian-Hao Lin, Yu-Juan Quan, and Bo-Ping Han. MetaIBM: a Python-based library for individual-based modelling of eco-evolutionary dynamics in spatial-explicit metacommunities. Ecological Modelling, 492(??):??, June 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001182.

LeGal:2020:MIB

[LRA^{+}20] Antoine Le Gal, Corinne Robert, Francesco Accatino, David Claessen, and Jane Lecomte. Modelling the interactions between landscape structure and spatio-temporal dynamics of pest natural enemies: Implications for conservation biological control. Ecological Modelling, 420(??):??, March 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001930420X.

Leite:2021:MPP

[LSA^{+}21] Maria C. A. Leite, Rebecca Sauchuk, Folashade B. Agusto, Orou G. Gaoue, and Benito Chen-Charpentier. Modeling the persistence of plant populations in fragmented ecosystems. Ecological Modelling, 457(??):??, October 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002398.

Logofet:2021:NCO

[LSG21] Dmitrii O. Logofet and Roberto Salguero-Gómez. Novel challenges and opportunities in the theory and practice of matrix population modelling: an editorial for the special feature: “Theory and Practice in Matrix Population Modelling” of ecological modelling. Ecological Modelling, 443(??):??, March 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021000296.

Liao:2020:RDN

[LSL20] Limei Liao, Yang Shen, and Jinbao Liao. Robustness of dispersal network structure to patch loss. Ecological Modelling, 424(??):??, May 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301083.

Li:2021:MTD

[LSL21] Jie Li, Xuzhu Shen, and YaoTang Li. Modeling the temporal dynamics of gut microbiota from a local community perspective. Ecological Modelling, 460(??):??, November 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002854.

Liu:2021:MWC

[LSS^{+}21] Jinxun Liu, Benjiamin Sleeter, Paul C. Selmants, Jiaojiao Diao, Qiang Zhou, Bruce Worstell, and Monica Moritsch. Modeling watershed carbon dynamics as affected by land cover change and soil erosion. Ecological Modelling, 459(??):??, November 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021002787.

Lai:2024:ABM

[LTR^{+}24] Samantha Lai, Theophilus Zhi En Teo, Arief Rullyanto, Jeffery Low, Karenne Tun, Peter A. Todd, and Siti Maryam Yaakub. An agent-based model approach to assessing the role of vegetative fragments in seagrass connectivity. Ecological Modelling, 487(??):??, January 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002582.

Ly:2024:ILC

[LUT^{+}24] Sophanna Ly, Sovannara Uk, Vouchlay Theng, Vinhteang Kaing, and Chihiro Yoshimura. Integration of life cycle and habitat conditions in modeling fish biomass in the floodplain of the Lower Mekong Basin. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023003356.

Lavaud:2023:PPS

[LUV^{+}23] Romain Lavaud, David S. Ullman, Celeste Venolia, Carol Thornber, Lindsay Green-Gavrielidis, and Austin Humphries. Production potential of seaweed and shellfish integrated aquaculture in Narragansett Bay (Rhode Island, U.S.) using an ecosystem model. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000984.

Lewis:2022:PSD

[LVM^{+}22] Dakota M. Lewis, Tali Vardi, Rebecca L. Maher, Adrienne M. S. Correa, and Geoffrey S. Cook. Predicting shifts in demography of Orbicella franksi following simulated disturbance and restoration. Ecological Modelling, 472(??):??, October 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002071.

Luo:2023:SAN

[LW23] Gongli Luo and Xiaotong Wang. Spatial association network structure of eco-efficiency and its influencing factors: Evidence from the Beijing–Tianjin–Hebei region in China. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003167.

Li:2022:TDL

[LWE^{+}22] Geng Li, Binbin Wang, Caroline M. Elliott, Bruce C. Call, Duane C. Chapman, and Robert B. Jacobson. A threedimensional Lagrangian particle tracking model for predicting transport of eggs of rheophilic-spawning carps in turbulent rivers. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001454.

Liu:2024:MSV

[LWL24] Yingzi Liu, Jing Wu, and Chunbin Li. Model simulation of vegetation canopy precipitation interception in grassland ecosystems on the northeast margin of the Qinghai–Tibet Plateau. Ecological Modelling, 488(??):??, February 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002300306X.

Liu:2023:CBM

[LWZL23] Ming Liu, Jiani Wu, Shuhua Zhang, and Jing Liang. Cyanobacterial blooms management: a modified optimization model for interdisciplinary research. Ecological Modelling, 484(??):??, October 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002107.

Li:2021:DBM

[LYC^{+}21] Yang Li, Lin Yuan, Hao-Bing Cao, Chen-Dong Tang, Xian-Ye Wang, Bo Tian, Shen-Tang Dou, Li-Quan Zhang, and Jian Shen. A dynamic biomass model of emergent aquatic vegetation under different water levels and salinity. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304622.

Lu:2024:SEC

[LYL^{+}24] Yanhua Lu, Lijuan Yan, Jie Li, Yunliang Liang, Chuanjie Yang, Guang Li, Jiangqi Wu, and Hua Xu. Spatiotemporal evolution of county level ecological security based on an emergy ecological footprint model: the case of Dingxi, China. Ecological Modelling, 490(??):??, April 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000498.

Lyu:2023:PCD

[LYM^{+}23] Yanfeng Lyu, Xiangdong Yang, Xiaohan Ma, Hengyu Pan, and Xiaohong Zhang. Promoting coordinated development of the fertilizer production-crop plantation combined system through an integrated approach. Ecological Modelling, 478(??):??, April 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002300008X.

Li:2024:FER

[LYZ^{+}24] Pengcheng Li, Yang Yang, Wenming Zhang, Nan Cong, Ge Yang, and Weiwei Yao. Forebay entrainment risk effects on two fish species in the Williston Reservoir. Ecological Modelling, 487(??):??, January 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023003010.

Li:2024:EFC

[LZD^{+}24] Xinyu Li, Qirui Zhang, Yanfang Diao, Yuzhi Shi, Shuxian Li, Chuanhui Yao, Rui Su, and Shichao Guo. Ecological flow considering hydrological season and habitat suitability for a variety of fish. Ecological Modelling, 489(??):??, March 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024000140.

Liu:2021:MMC

[LZF21] Ningyin Liu, Yan Zhang, and Brian D. Fath. The material metabolism characteristics and growth patterns of the central cities of China’s Beijing–Tianjin–Hebei region. Ecological Modelling, 448(??):??, May 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001022.

Li:2024:DES

[LZF^{+}24] Xuehong Li, Xuebin Zhang, Haoyuan Feng, Yixia Li, Jiale Yu, Yanni Liu, and Hucheng Du. Dynamic evolution and simulation of habitat quality in arid regions: a case study of the Hexi region, China. Ecological Modelling, 493(??):??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001145.

Liu:2022:PCS

[LZL22] Lingchi Liu, Xiaohong Zhang, and Yanfeng Lyu. Performance comparison of sewage treatment plants before and after their upgradation using emergy evaluation combined with economic analysis: a case from Southwest China. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002200182X.

Liang:2023:ESR

[LZL23] Fachao Liang, Runmiao Zhu, and Sheng-Hau Lin. Exploring spatial relationship between landscape configuration and ecosystem services: a case study of Xiamen–Zhangzhou– Quanzhou in China. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002570.

Li:2020:EFL

[LZSQ20] Fei Li, Meijun Zhou, Jiaqi Shao, and Zhangxuan Qin. Evolution forms of land systems based on ascendency and overhead: a case study of Shaanxi Province, China. Ecological Modelling, 419(??):??, March 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300326.

Liu:2020:LUS

[LZW20] Dongya Liu, Xinqi Zheng, and Hongbin Wang. Land-use simulation and decision-support system (LandSDS): Seamlessly integrating system dynamics, agent-based model, and cellular automata. Ecological Modelling, 417(??):??, February 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019304326.

Milne:2023:PMC

[MAB23] Russell Milne, Madhur Anand, and Chris T. Bauch. Preparing for and managing crown-of-thorns starfish outbreaks on reefs under threat from interacting anthropogenic stressors. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001746.

Malard:2020:AFW

[MAD^{+}20] Julien J. Malard, Jan Franklin Adamowski, Marcela Rojas Díaz, Jessica Bou Nassar, Nallusamy Anandaraja, Héctor Tuy, Luís Andrés Arévalo-Rodriguez, and Hugo Ramiro Melgar-Quiñonez. Agroecological food web modelling to evaluate and design organic and conventional agricultural systems. Ecological Modelling, 421(??):??, April 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300338.

Mateo:2022:IBD

[MAGR^{+}22] Rubén G. Mateo, Gabriel Arellano, Virgilio Gómez-Rubio, J. Sebastián Tello, Alfredo F. Fuentes, Leslie Cayola, M. Isabel Loza, Victoria Cala, and Manuel J. Macía. Insights on biodiversity drivers to predict species richness in tropical forests at the local scale. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022002344.

Malard:2020:MPT

[MAN^{+}20] Julien Malard, Jan Adamowski, Jessica Bou Nassar, Nallusamy Anandaraja, Héctor Tuy, and Hugo Melgar-Quiñonez. Modelling predation: Theoretical criteria and empirical evaluation of functional form equations for predator–prey systems. Ecological Modelling, 437(??):??, December 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303343.

Masutomi:2023:AAS

[Mas23] Yuji Masutomi. The appropriate analytical solution for coupled leaf photosynthesis and stomatal conductance models for C3 plants. Ecological Modelling, 481(??):??, July 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000340.

Matyja:2023:SDE

[Mat23] Konrad Matyja. Standard dynamic energy budget model parameter sensitivity. Ecological Modelling, 478(??):??, April 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023000327.

Mendes:2022:GDD

[MB22] Pedro B. Mendes and Walter A. Boeger. Game dynamics as a driver for pathogen spillover pulses. Ecological Modelling, 473(??):??, November 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001995.

Morrice:2020:EBC

[MBB20] Katherine J. Morrice, António M. Baptista, and Brian J. Burke. Environmental and behavioral controls on juvenile Chinook salmon migration pathways in the Columbia River estuary. Ecological Modelling, 427(??):??, July 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300752.

Moullec:2022:USD

[MBD^{+}22] Fabien Moullec, Nicolas Barrier, Sabrine Drira, François Guilhaumon, Tarek Hattab, Myron A. Peck, and Yunne-Jai Shin. Using species distribution models only may underestimate climate change impacts on future marine biodiversity. Ecological Modelling, 464(??):??, February 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003690.

Mattei:2021:GAM

[MBFS21] F. Mattei, E. Buonocore, P. P. Franzese, and M. Scardi. Global assessment of marine phytoplankton primary production: Integrating machine learning and environmental accounting models. Ecological Modelling, 451(??):??, July 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021001435.

Meier:2022:MMM

[MBG^{+}22] Laura Meier, Mario Brauns, Volker Grimm, Markus Weitere, and Karin Frank. MASTIFF: a mechanistic model for cross-scale analyses of the functioning of multiple stressed riverine ecosystems. Ecological Modelling, 470(??):??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022001181.

Mundim:2020:TCQ

[MBMV20] Kleber C. Mundim, Solange Baraldi, Hugo G. Machado, and Fernando M. C. Vieira. Temperature coefficient (Q_{10}) and its applications in biological systems: Beyond the Arrhenius theory. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002030199X.

Molter:2023:DBC

[MBR^{+}23] Alexandre Molter, João I. M. Bezerra, Elvira Rafikova, Dori E. Nava, and Marat Rafikov. Dynamics and biological control of the sugarcane borer with two parasitoids. Ecological Modelling, 481(??):??, July 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000996.

Moscoso:2022:CCB

[MBS22] Jordyn E. Moscoso, Daniele Bianchi, and Andrew L. Stewart. Controls and characteristics of biomass quantization in size-structured planktonic ecosystem models. Ecological Modelling, 468(??):??, June 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000321.

McHuron:2023:MIC

[MCB^{+}23] Elizabeth A. McHuron, Manuel Castellote, Gina K. Himes Boor, Kim E. W. Shelden, Amanda J. Warlick, Tamara L. McGuire, Paul R. Wade, and Kimberly T. Goetz. Modeling the impacts of a changing and disturbed environment on an endangered beluga whale population. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001485.

Marques:2024:UIB

[MCMP^{+}24] Ana Teresa Marques, Tiago Crispim-Mendes, Luís Palma, Ricardo Pita, Francisco Moreira, and Pedro Beja. Using individual-based demographic modelling to estimate the impacts of anthropogenic mortality on territorial predators. Ecological Modelling, 493(??):??, July 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001406.

Mugabi:2023:EDC

[MD23] Francis Mugabi and Kevin J. Duffy. Epidemiological drivers and control strategies for African swine fever transmission cycles at a wildlife-livestock interface. Ecological Modelling, 481(??):??, July 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000728.

Marandure:2020:ASD

[MDB^{+}20] Tawanda Marandure, Kennedy Dzama, James Bennett, Godswill Makombe, and Cletos Mapiye. Application of system dynamics modelling in evaluating sustainability of lowinput ruminant farming systems in Eastern Cape Province, South Africa. Ecological Modelling, 438(??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020303641.

Mamboleo:2021:CMA

[MDP21] Abel Ansporthy Mamboleo, Crile Doscher, and Adrian Paterson. A computational modelling approach to human– elephant interactions in the Bunda District, Tanzania. Ecological Modelling, 443(??):??, March 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000223.

Marchetto:2023:TES

[MDT^{+}23] Elisa Marchetto, Daniele Da Re, Enrico Tordoni, Manuele Bazzichetto, Piero Zannini, Simone Celebrin, Ludovico Chieffallo, Marco Malavasi, and Duccio Rocchini. Testing the effect of sample prevalence and sampling methods on probabilityand favourability-based SDMs. Ecological Modelling, 477(??): ??, March 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022003465.

Meador:2020:IBM

[Mea20] Andrew J. Sánchez Meador. Individual-based Methods in Forest Ecology and Management: Individual-based Methods in Forest Ecology and Management. A. Pommerening and P. Grabarnik. 2019. Springer International Publishing. Switzerland. 411 p. $84.99 (hardcover). ISBN: 978-3-030-24528-3. Ecological Modelling, 428(??):??, July 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301642.

Mattias:2021:MES

[MEGA21] Gaglio Mattias, Tamburini Elena, Castaldelli Giuseppe, and Fano Elisa Anna. Modeling the ecosystem service of agricultural residues provision for bioenergy production: a potential application in the Emilia–Romagna region (Italy). Ecological Modelling, 451(??):??, July 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001368.

Menezes:2022:MVT

[Men22] Jorge Fernando Saraiva Menezes. Marginal value theorem as a special case of the ideal free distribution. Ecological Modelling, 468(??):??, June 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000564.

Moreh:2021:TAP

[MEPJ21] Ágnes Móréh, Anett Endrédi, Sándor Imre Piross, and Ferenc Jordán. Topology of additive pairwise effects in food webs. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304737.

Mendes:2020:EED

[MF20] Pedro B. Mendes and Lucas D. B. Faria. The eco-evolutionary dynamics of a predator–prey system across an r/K continuum. Ecological Modelling, 436(??):??, November 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303392.

Morales:2024:DAB

[MFF^{+}24] Mark M. Morales, Jerome Fiechter, John C. Field, Neosha S Kashef, Elliott L. Hazen, and Mark H. Carr. Development and application of a bioenergetics growth model for multiple early life stages of an ecologically important marine fish. Ecological Modelling, 488(??):??, February 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003058.

Meng:2022:MEA

[MFH^{+}22] Jia-Nan Meng, Hongwei Fang, Lei Huang, Guojian He, Xiaobo Liu, Changyi Xu, Xinghua Wu, and Donald Scavia. Multidimensional ecosystem assessment of Poyang Lake under anthropogenic influences. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002356.

Martinez-Fernandez:2023:SES

[MFPIBGES23] Julia Martínez-Fernández, Irene Pérez-Ibarra, Isabel Banos-González, and Miguel Angel Esteve-Selma. Social-ecological systems modelling to understand the linkages between water, agriculture and rural systems. Ecological Modelling, 482(??): ??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001060.

Ma:2023:CAS

[MFZ23] Wen Ma, Zhuoqiong Fang, and Xiangfeng Zhang. Comparative analysis of structural characteristics of China’s 18 typical urban agglomerations based on flows of various elements. Ecological Modelling, 479(??):??, May 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000364.

Mendiola:2024:EME

[MGC24] Sandra Y. Mendiola, Nicole M. Gerardo, and David J. Civitello. Ecological modeling evaluates the potential impacts of symbionts on plant pathogen vectoring in field populations. Ecological Modelling, 487(??):??, January 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002715.

Montoya:2020:RBC

[MGdM^{+}20] Daniel Montoya, Sabrina Gaba, Claire de Mazancourt, Vincent Bretagnolle, and Michel Loreau. Reconciling biodiversity conservation, food production and farmers’ demand in agricultural landscapes. Ecological Modelling, 416(??):??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019303977.

Martignoni:2020:IIM

[MGHT20] Maria M. Martignoni, Jimmy Garnier, Miranda M. Hart, and Rebecca C. Tyson. Investigating the impact of the mycorrhizal inoculum on the resident fungal community and on plant growth. Ecological Modelling, 438(??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020303896.

Marsh:2023:SPT

[MGK^{+}23] Charles J. Marsh, Yoni Gavish, Mathias Kuemmerlen, Stefan Stoll, Peter Haase, and William E. Kunin. SDM profiling: a tool for assessing the information-content of sampled and unsampled locations for species distribution models. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S030438002200271X.

Monsinjon:2022:TRN

[MGWG22] Jonathan R. Monsinjon, Jean-Michel Guillon, Jeanette Wyneken, and Marc Girondot. Thermal reaction norm for sexualization: the missing link between temperature and sex ratio for temperature-dependent sex determination. Ecological Modelling, 473(??):??, November 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002204.

Maurent:2023:CFM

[MHP^{+}23] Eliott Maurent, Bruno Hérault, Camille Piponiot, Géraldine Derroire, Diego Delgado, Bryan Finegan, Mélaine Aubry Kientz, Bienvenu H. K. Amani, and Marie Ange Ngo Bieng. A common framework to model recovery in disturbed tropical forests. Ecological Modelling, 483(??):??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001497.

Mauritsson:2024:NMM

[MJ24] Karl Mauritsson and Tomas Jonsson. A new mechanistic model for individual growth suggests upregulated maintenance costs when food is scarce in an insect. Ecological Modelling, 491(??):??, May 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000917.

Muniz-Junior:2023:LPD

[MJdORPG23] Gilberto Muniz-Junior, Fábio de Oliveira Roque, Aliny PF. Pires, and Rafael D. Guariento. Are lower pesticide doses better? An evolutionary perspective on integrated pest management. Ecological Modelling, 482(??):??, August 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001394.

Moreh:2024:EJI

[MJS24] Ágnes Móréh, Ferenc Jordán, and István Scheuring. Effects of joint invasion: How co-invaders affect each other’s success in model food webs? Ecological Modelling, 492(??):??, June 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001236.

Maclean:2021:MMM

[MK21] Ilya M. D. Maclean and David H. Klinges. Microclimc: a mechanistic model of above, below and within-canopy microclimate. Ecological Modelling, 451(??):??, July 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001265.

Myllymaki:2024:BAP

[MKB^{+}24] Mari Myllymäki, Mikko Kuronen, Simone Bianchi, Arne Pommerening, and Lauri Mehtätalo. A Bayesian approach to projecting forest dynamics and related uncertainty: an application to continuous cover forests. Ecological Modelling, 491(??): ??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000577.

Meixler:2020:IOL

[MKEH20] Marcia S. Meixler, Christina M. K. Kaunzinger, Jean N. Epiphan, and Steven N. Handel. Identifying opportunities for local assisted expansion of coastal upland vegetation in an urban estuary. Ecological Modelling, 438(??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303793.

Mahmoodi:2023:WNT

[MKN23] Ali Mahmoodi, Mohammad Reza Alizadeh Kiapi, and Hamidreza M. Nick. When nitrate treatment wins the battle against microbial reservoir souring but loses the war. Ecological Modelling, 481(??):??, July 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000571.

Marvuglia:2020:EGR

[MKR20] Antonino Marvuglia, Rembrandt Koppelaar, and Benedetto Rugani. The effect of green roofs on the reduction of mortality due to heatwaves: Results from the application of a spatial microsimulation model to four European cities. Ecological Modelling, 438(??):??, December 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304178.

Malishev:2021:MMM

[MKS21] Matthew Malishev and Stephanie Kramer-Schadt. Movement, models, and metabolism: Individual-based energy budget models as next-generation extensions for predicting animal movement outcomes across scales. Ecological Modelling, 441(??):??, February 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020304725.

Maslov:2020:BVC

[ML20] Alexander A. Maslov and Dmitrii O. Logofet. Bilberry vs. cowberry in a Scots pine boreal forest: III. Another forest, another method, and similar conclusions. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302647.

Mack:2023:WCM

[MLD^{+}23] Sarah K. Mack, Robert R. Lane, Jia Deng, James T. Morris, and Julian J. Bauer. Wetland carbon models: Applications for wetland carbon commercialization. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S030438002200326X.

McDanold:2023:DDU

[MLJ^{+}23] Jenna S. McDanold, Rodman R. Linn, Alex K. Jonko, Adam L. Atchley, Scott L. Goodrick, J. Kevin Hiers, Chad M. Hoffman, E. Louise Loudermilk, Joseph J. O’Brien, Russell A. Parsons, Carolyn H. Sieg, and Julia A. Oliveto. DUET — Distribution of Understory using Elliptical Transport: a mechanistic model of leaf litter and herbaceous spatial distribution based on tree canopy structure. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001564.

Ma:2020:EDC

[MLL20] Wu Ma, Guang Lin, and Jingjing Liang. Estimating dynamics of central hardwood forests using random forests. Ecological Modelling, 419(??):??, March 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300193.

Mahdiyasa:2023:MIM

[MLMI23] Adilan W. Mahdiyasa, David J. Large, Bagus P. Muljadi, and Matteo Icardi. Modelling the influence of mechanicalecohydrological feedback on the nonlinear dynamics of peatlands. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000273.

Mounier:2020:DBPa

[MLP^{+}20] Florence Mounier, Véronique Loizeau, Laure Pecquerie, Hilaire Drouineau, Pierre Labadie, Hélène Budzinski, and Jérémy Lobry. Dietary bioaccumulation of persistent organic pollutants in the common sole Solea solea in the context of global change. Part 2: Sensitivity of juvenile growth and contamination to toxicokinetic parameters uncertainty and environmental conditions variability in estuaries. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302623.

Metzler:2024:ACF

[MLV24] Holger Metzler, Samuli Launiainen, and Giulia Vico. Amount of carbon fixed, transit time and fate of harvested wood products define the climate change mitigation potential of boreal forest management — a model analysis. Ecological Modelling, 491(??):??, May 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000826.

Murray:2021:SMT

[MM21] Alexander G. Murray and Meadhbh Moriarty. A simple modelling tool for assessing interaction with host and local infestation of sea lice from salmonid farms on wild salmonids based on processes operating at multiple scales in space and time. Ecological Modelling, 443(??):??, March 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000314.

Messan:2021:PDV

[MMC^{+}21] Komi Messan, Marisabel Rodriguez Messan, Jun Chen, Gloria DeGrandi-Hoffman, and Yun Kang. Population dynamics of Varroa mite and honeybee: Effects of parasitism with age structure and seasonality. Ecological Modelling, 440(??):??, January 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020304245.

Maciel:2021:DCC

[MMdP^{+}21] Everton A. Maciel, Valeria F. Martins, Mateus D. de Paula, Andreas Huth, Frederico A. G. Guilherme, Rico Fischer, André Giles, Reinaldo I. Barbosa, Osmar Cavassan, and Fernando R. Martins. Defaunation and changes in climate and fire frequency have synergistic effects on aboveground biomass loss in the Brazilian savanna. Ecological Modelling, 454(??):??, August 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021001903.

Mahood:2024:NPP

[MMM^{+}24] Adam L. Mahood, Ranjan Muthukrishnan, Jacob A. Macdonald, David T. Barnett, Eric R. Sokol, and Samuel M. Simkin. neonPlantEcology: an R package for preparing NEON plant data for use in ecological research. Ecological Modelling, 493(??):??, July 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024001388.

Mori:2023:MEZ

[MMMS23] Kensuke Mori, Alessandro Massolo, Danielle Marceau, and Emmanuel Stefanakis. Modelling the epidemiology of zoonotic parasites transmitted through a predator-prey system in urban landscapes: the Calgary Echinococcus multilocularis Coyote Agent-based model (CEmCA). Ecological Modelling, 475(??):??, January 2023. CODEN ???? ISSN

????  URL http://www.sciencedirect.com/science/
article/pii/S0304380022002782.

Melo-Merino:2020:ENM

[MMRBLN20] Sara M. Melo-Merino, Héctor Reyes-Bonilla, and Andrés Lira-Noriega. Ecological niche models and species distribution models in marine environments: a literature review and spatial analysis of evidence. Ecological Modelling, 415(??):??, January 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001930345X.

Manfrin:2024:MCE

[MMS^{+}24] A. Manfrin, G. A. López Moreira Mazacotte, J. W. Spaak, S. E. Osakpolor, C. A. Brühl, V. Lencioni, S. Kolbenschlag, R. B. Schäfer, M. Bundschuh, and R. Schulz. Modeling cumulative effects of acute exposure to toxicants on the life cycle of Chironomidae using Bti as a case study. Ecological Modelling, 494(??):??, August 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002400156X.

Menezes:2021:CIN

[MOS21] Jorge Fernando Saraiva Menezes and Luiz Gustavo Rodrigues Oliveira-Santos. Cautious individuals have noninvadable territories, according to an evolutionary mechanistic model. Ecological Modelling, 449(??):??, June 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001216.

Men:2023:ENI

[MP23] Dan Men and Jinghu Pan. Ecological network identification and connectivity robustness evaluation in the Yellow River Basin under a multi-scenario simulation. Ecological Modelling, 482(??):??, August 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001151.

Movedi:2024:APC

[MPA^{+}24] Ermes Movedi, Livia Paleari, Giovanni Argenti, Fosco M. Vesely, Nicolina Staglianò, Silvia Parrini, and Roberto Confalonieri. The application of a plant community model to evaluate adaptation strategies for alleviating climate change impacts on grassland productivity, biodiversity and forage quality. Ecological Modelling, 488(??):??, February 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003265.

Marzidovsek:2022:BTB

[MPC^{+}22] Martin Marzidovšek, Vid Podpečan, Erminia Conti, Marko Debeljak, and Christian Mulder. BEFANA: a tool for biodiversity-ecosystem functioning assessment by network analysis. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001739.

Moulin:2021:DPB

[MPCG21] Thibault Moulin, Antoine Perasso, Pierluigi Calanca, and François Gillet. DynaGraM: a process-based model to simulate multi-species plant community dynamics in managed grasslands. Ecological Modelling, 439(??):??, January 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304129.

Mounier:2020:DBPb

[MPL^{+}20] Florence Mounier, Laure Pecquerie, Jérémy Lobry, Adriana E. Sardi, Pierre Labadie, Hélène Budzinski, and Véronique Loizeau. Dietary bioaccumulation of persistent organic pollutants in the common sole Solea solea in the context of global change. Part 1: Revisiting parameterisation and calibration of a DEB model to consider interindividual variability in experimental and natural conditions. Ecological Modelling, 433(??):??, October 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302945.

Marull:2023:UTU

[MPMT23] Joan Marull, Joan Pino, Yolanda Melero, and Enric Tello. Using thermodynamics to understand the links between energy, information, structure and biodiversity in a human-transformed landscape. Ecological Modelling, 476 (??):??, February 2023. CODEN ECMODT. ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003556.

McEwan:2023:SMP

[MPPM23] Nicole McEwan, Rich Pawlowicz, Evgeny Pakhomov, and Maria T. Maldonado. Seasonality of modelled planktonic food web structure in the Strait of Georgia, Canada. Ecological Modelling, 482(??):??, August 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001333.

Manshadi:2024:MUG

[MPSS24] Zahra Dehghan Manshadi, Parastoo Parivar, Ahad Sotoudeh, and Ali Morovati Sharifabadi. Modeling urban growth effects on carrying capacity in arid and semi-arid regions using system dynamics. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002958.

Martins:2020:PPA

[MRR^{+}20] Thales G. V. Martins, Geraldo G. Reis, Maria G. F. Reis, Lucas A. A. Telles, Mayara R. Lage, Gleidson G. C. Mendes, Dayane L. Pinto, Nero L. M. Castro, Alexandre S. Lorenzon, Ricardo S. Silva, and Duberlí G. E. Gonzáles. Potential planting areas for native tree species in Minas Gerais State, Brazil, based on environmental variables and wood demand. Ecological Modelling, 432(??):??, September 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020302817.

Mattei:2020:EEK

[MS20] F. Mattei and M. Scardi. Embedding ecological knowledge into artificial neural network training: a marine phytoplankton primary production model case study. Ecological Modelling, 421(??):??, April 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300570.

Mestre:2023:PLR

[MS23] F. Mestre and B. Silva. lconnect R package: a versatile tool for evaluating landscape connectivity and prioritizing habitat patches in conservation research. Ecological Modelling, 484(??):??, October 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002193.

MacPherson:2021:UIB

[MSG21a] Brian MacPherson, Ryan Scott, and Robin Gras. Using individual-based modelling to investigate the possible role that the red tooth effect plays in maintaining sexual reproduction. Ecological Modelling, 459(??):??, November 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002829.

Mishra:2021:CSC

[MSG^{+}21b] Gaurav Mishra, Avishek Sarkar, Krishna Giri, Arun Jyoti Nath, Rattan Lal, and Rosa Francaviglia. Changes in soil carbon stocks under plantation systems and natural forests in Northeast India. Ecological Modelling, 446(??):??, April 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021000715.

MacPherson:2023:UIB

[MSG23] Brian MacPherson, Ryan Scott, and Robin Gras. Using individual-based modelling to investigate a pluralistic explanation for the prevalence of sexual reproduction in animal species. Ecological Modelling, 475(??):??, January 2023. CO-

DEN ???? ISSN ???? URL http://www.sciencedirect.
com/science/article/pii/S0304380022002897.

Madani:2021:MML

[MSL^{+}21] Mohammad Madani, Rajesh Seth, Luis F. Leon, Reza Valipour, and Craig McCrimmon. Microbial modelling of Lake St. Clair: Impact of local tributaries on the shoreline water quality. Ecological Modelling, 458(??):??, October 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002647.

Medeiros-Sousa:2024:MTS

[MSLM^{+}24] Antônio Ralph Medeiros-Sousa, Martin Lange, Luis Filipe Mucci, Mauro Toledo Marrelli, and Volker Grimm. Modelling the transmission and spread of yellow fever in forest landscapes with different spatial configurations. Ecological Modelling, 489(??):??, March 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000176.

Medeiros-Sousa:2022:EDY

[MSLMM22] Antônio Ralph Medeiros-Sousa, Gabriel Zorello Laporta, Luis Filipe Mucci, and Mauro Toledo Marrelli. Epizootic dynamics of yellow fever in forest fragments: an agent-based model to explore the influence of vector and host parameters. Ecological Modelling, 466(??):??, April 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000126.

Mayersohn:2021:ITV

[MSML21] Benjamin Mayersohn, K. Shafer Smith, Inès Mangolte, and Marina Lévy. Intrinsic timescales of variability in a marine plankton model. Ecological Modelling, 443(??):??, March 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002100020X.

Mundetia:2024:GSA

[MSS24] Nitika Mundetia, Devesh Sharma, and Aditya Sharma. Groundwater sustainability assessment under climate change scenarios using integrated modelling approach and multicriteria decision method. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002740.

Shu:2021:DMO

[mSzZK^{+}21] Shu miao Shu, Wan ze Zhu, George Kontsevich, Yang yi Zhao, Wen zhi Wang, Xiao xiang Zhao, and Xiao dan Wang. A discrete model of ontogenetic growth. Ecological Modelling, 460(??):??, November 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002100301X.

Masaka:2023:TMH

[MT23] Kazuhiko Masaka and Takenori Takada. Transition model for the hermaphroditism-dioecy continuum in higher plants. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022002368.

Menard:2022:MMA

[MTBB22] Isabelle Ménard, Evelyne Thiffault, Yan Boulanger, and Jean-François Boucher. Multi-model approach to integrate climate change impact on carbon sequestration potential of afforestation scenarios in Quebec, Canada. Ecological Modelling, 473(??):??, November 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002459.

Mathilde:2023:LVM

[MTL^{+}23] Godefroid Mathilde, Zeimes Tom, Bramanti Lorenzo, Romans Pascal, Bo Marzia, Toma Margherita, Danis Bruno, Dubois Philippe, and Guillaumot Charlène. Low vulnerability of the Mediterranean antipatharian Antipathella subpinnata (Ellis & Solander, 1786) to ocean warming. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003076.

Marull:2020:ELO

[MTP^{+}20] Joan Marull, Parisa Torabi, Roc Padró, Aureli Alabert, Maria José La Rota, and Tarik Serrano. Energy-landscape optimization for land use planning. Application in the Barcelona Metropolitan Area. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302532.

Marques:2020:MDG

[MTS^{+}20] Gonçalo M. Marques, Carlos M. G. L. Teixeira, Tânia Sousa, Tiago G. Morais, Ricardo F. M. Teixeira, and Tiago Domingos. Minimizing direct greenhouse gas emissions in livestock production: the need for a metabolic theory.

Ecological Modelling, 434(??):??, October 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002030329X.

Mosai:2022:CSM

[MTT22] Alseno K. Mosai, Bontle C. Tokwana, and Hlanganani Tutu. Computer simulation modelling of the simultaneous adsorption of Cd, Cu and Cr from aqueous solutions by agricultural clay soil: a PHREEQC geochemical modelling code coupled to parameter estimation (PEST) study. Ecological Modelling, 465(??):??, March 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000011.

Marohn:2022:CBD

[MTW^{+}22] Carsten Marohn, Christian Troost, Benjamin Warth, Christian Bateki, Mink Zijlstra, Faizan Anwar, Benjamin Williams, Katrien Descheemaeker, Thomas Berger, Folkard Asch, Uta Dickhoefer, Regina Birner, and Georg Cadisch. Coupled biophysical and decision-making processes in grassland systems in East African savannahs — a modelling framework. Ecological Modelling, 474(??):??, December 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002150.

Mohammadyari:2023:OLU

[MTZA23] Fatemeh Mohammadyari, Mohsen Tavakoli, Ardavan Zarandian, and Sedighe Abdollahi. Optimization land use based on multi-scenario simulation of ecosystem service for sustainable landscape planning in a mixed urban — forest watershed. Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001710.

Murthy:2024:TBA

[Mur24] Karthik K. Murthy. A trait-based approach to integrate resilience frameworks. Ecological Modelling, 493(??):??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001455.

Mendes:2020:DOS

[MVAD20] Poliana Mendes, Santiago José Elías Velazco, André Felipe Alves de Andrade, and Paulo De Marco. Dealing with overprediction in species distribution models: How adding distance constraints can improve model accuracy. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302519.

Movedi:2022:NAM

[MVC^{+}22] Ermes Movedi, Daniele Valiante, Alessandro Colosio, Luca Corengia, Stefano Cossa, and Roberto Confalonieri. A new approach for modeling crop-weed interaction targeting management support in operational contexts: a case study on the rice weeds barnyardgrass and red rice. Ecological Modelling, 463(??):??, January 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003434.

Martin:2022:DBP

[MYACC22] Gerardo Martín, Carlos Yáñez-Arenas, and Xavier Chiappa-Carrara. Discrepancies between point process models and environmental envelopes identify the niche centroid — geography configuration. Ecological Modelling, 469(??):??, July 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022000898.

Masuda:2020:ICE

[MYH^{+}20] Yoshio Masuda, Yasuhiro Yamanaka, Takafumi Hirata, Hideyuki Nakano, and Takashi S. Kohyama. Inhibition of competitive exclusion due to phytoplankton dispersion: a contribution for solving Hutchinson’s paradox. Ecological Modelling, 430(??):??, August 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301617.

Morales-Zarate:2021:LMT

[MZLRSZ21] M. V. Morales-Zárate, J. A. López-Ramírez, and C. A. Salinas-Zavala. Loggerhead marine turtle (Caretta caretta) ecological facts from a trophic relationship model in a hot spot fishery area: Gulf of Ulloa, Mexico. Ecological Modelling, 439(??):??, January 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020303963.

Nasirzadehdizaji:2022:PPI

[NA22] Rouhollah Nasirzadehdizaji and Dilek Eren Akyuz. Predicting the potential impact of forest fires on runoff and sediment loads using a distributed hydrological modeling approach. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000771.

Nikouei:2022:MLA

[NAM22] Alireza Nikouei, Mohammad Reza Asgharipour, and Zahra Marzban. Modeling land allocation to produce crops under economic and environmental goals in Iran: a multi-objective programming approach. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001703.

Nilsson:2021:ESE

[NBMD21] L. Nilsson, N. Bunnefeld, J. Minderman, and A. B. Duthie. Effects of stakeholder empowerment on crane population and agricultural production. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020304609.

Nakhaei:2021:TDB

[NBML21] Nader Nakhaei, Leon Boegman, Mahyar Mehdizadeh, and Mark Loewen. Three-dimensional biogeochemical modeling of eutrophication in Edmonton stormwater ponds. Ecological Modelling, 456(??):??, September 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002428.

Newman:2023:DEE

[NBU^{+}23] Erica A. Newman, Micah Brush, Kaito Umemura, Meng Xu, and John Harte. Defining an ecological equation of state: Response to Riera et al. 2023. Ecological Modelling, 486(??):??, December 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002624. See [RFHR23, RFHR24].

Chen:2023:IWR

[nCLpW^{+}23] Xiang nan Chen, Fang Li, Feng ping Wu, Xia Xu, and Yue Zhao. Initial water rights allocation of industry in the Yellow River basin driven by high-quality development. Ecological Modelling, 477(??):??, March 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022003702.

Nguyen:2022:SAH

[NDNN22] Anh Duc Nguyen, Dung Ngoc Do, Hung Duc Nguyen, and Thuy Phuong Nguyen. Stability analysis and Hopf bifurcation of a brown planthopper–rice model under the effect of monsoon. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000631.

Nelson:2020:LCS

[NGA^{+}20] Gary A. Nelson, Benjamin I. Gahagan, Michael P. Armstrong, Adrian Jordaan, and Alison Bowden. A life cycle simulation model for exploring causes of population change in alewife (Alosa pseudoharengus). Ecological Modelling, 422(??):??, April 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020300764.

Nakhaii:2024:EES

[NGA^{+}24] Fatemeh Nakhaii, Seyed Ahmad Ghanbari, Mohammad Reza Asgharipour, Esmaeel Seyedabadi, and Enrico Sciubba. Evaluating ecological sustainability of mechanized and traditional systems of damaskrose production using extended exergy analysis. Ecological Modelling, 488(??):??, February 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003253.

Neta:2021:DSP

[NGE^{+}21] Ayana Neta, Roni Gafni, Hilit Elias, Nitsan Bar-Shmuel, Liora Shaltiel-Harpaz, Efrat Morin, and Shai Morin. Decision support for pest management: Using field data for optimizing temperature-dependent population dynamics models. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304658.

Nascimento:2023:MSS

[NHGP23] Marcela C. Nascimento, Berengere Husson, Lilia Guillet, and Torstein Pedersen. Modelling the spatial shifts of functional groups in the Barents Sea using a climate-driven spatial food web model. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023000868.

Nonaka:2023:LES

[NK23] Etsuko Nonaka and Anna Kuparinen. Limited effects of size-selective harvesting and harvesting-induced life-history changes on the temporal variability of biomass dynamics in complex food webs. Ecological Modelling, 476(??): ??, February 2023. CODEN ECMODT. ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002514.

Nakayama:2020:MDI

[NKT^{+}20] K. Nakayama, K. Komai, K. Tada, H. C. Lin, H. Yajima, S. Yano, M. R. Hipsey, and J. W. Tsai. Modeling dissolved inorganic carbon considering submerged aquatic vegetation. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302593.

Nie:2020:HSS

[NL20] Shipeng Nie and Weide Li. How spatial structure of species and disturbance influence the ecological invasion. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302696.

Neta:2023:SFP

[NLMM23] Ayana Neta, Yoav Levi, Efrat Morin, and Shai Morin. Seasonal forecasting of pest population dynamics based on downscaled SEAS5 forecasts. Ecological Modelling, 480(??):??, June 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023000546.

Neil:2020:ABM

[NMC^{+}20] Emily Neil, Jens Koed Madsen, Ernesto Carrella, Nicolas Payette, and Richard Bailey. Agent-based modelling as a tool for elephant poaching mitigation. Ecological Modelling, 427(??):??, July 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301265.

Novoa-Munoz:2021:LVM

[NMGFOB21] Francisco Novoa-Muñoz, Nelly Gómez-Fuentealba, and Florencia Osorio-Baeza. Lotka–Volterra model applied to two sympatric species of Liolaemus in competition. Ecological Modelling, 439(??):??, January 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020304130.

Nabhani:2024:MOO

[NMS24] Abbas Nabhani, Elham Mardaneh, and Hanne K. Sjølie. Multi-objective optimization of forest ecosystem services under uncertainty. Ecological Modelling, 494(??):??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001650.

Najafabadi:2023:NIM

[NMV^{+}23] Mostafa Mardani Najafabadi, Cosimo Magazzino, Donatella Valente, Abbas Mirzaei, and Irene Petrosillo. A new interval meta-goal programming for sustainable planning of agricultural water-land use nexus. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002016.

Newman:2022:ASO

[NND^{+}22] Kevin D. Newman, Jenny L. Nelson, Louise K. Durkin, Jemma K. Cripps, and Michael A. McCarthy. An analytical solution for optimising detections when accounting for site establishment costs. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022002186.

Nakayama:2023:DPB

[NO23]  Tadanobu Nakayama and Masahiro Osako.  Develop-

ment of a process-based eco-hydrology model for evaluating the spatio-temporal dynamics of macro- and microplastics for the whole of Japan. Ecological Modelling, 476 (??):??, February 2023. CODEN ECMODT. ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003416.

Nakayama:2024:PTI

[NO24] Tadanobu Nakayama and Masahiro Osako. Plastic tradeoff: Impact of export and import of waste plastic on plastic dynamics in Asian region. Ecological Modelling, 489(??):??, March 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000139.

Nooteboom:2023:MTA

[NPK^{+}23] Peter D. Nooteboom, Joe Scutt Phillips, Christian Kehl, Simon Nicol, and Erik van Sebille. Modelling of tuna around fish aggregating devices: the importance of ocean flow and prey. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022002861.

Novak:2023:IDF

[NPS^{+}23] Gorazd Novak, Polona Pengal, Ana T. Silva, José M. Domínguez, Angelo Tafuni, Matjaž Četina, and Dušan Žagar. Interdisciplinary design of a fish ramp using migration routes analysis. Ecological Modelling, 475(??):??, January 2023. CO-

DEN ???? ISSN ???? URL http://www.sciencedirect.
com/science/article/pii/S0304380022002873.

Natiello:2020:MPD

[NS20] Mario A. Natiello and Hernán G. Solari. Modelling population dynamics based on experimental trials with genetically modified (RIDL) mosquitoes. Ecological Modelling, 424(??):??, May 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020300582.

Nardella:2023:MPR

[NSS^{+}23] L. Nardella, A. Sebastiani, M. Stafoggia, P. P. Franzese, and F. Manes. Modelling PM_{10} removal in three Italian coastal metropolitan cities along a latitudinal gradient. Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001540.

Nielsen:2023:DPD

[NT23] Julie K. Nielsen and Cindy A. Tribuzio. Development and parameterization of a data likelihood model for geolocation of a bentho-pelagic fish in the North Pacific Ocean. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000108.

Nothaass:2023:MSI

[NTHC23] Dorian Nothaaß, Franziska Taubert, Andreas Huth, and Adam Thomas Clark. Modelling species invasion using a metapopulation model with variable mortality and stochastic birth–death processes. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002300100X.

Nguyen:2022:SMH

[NWJ^{+}22] David Nguyen, Tanay Wakhare, Jing Jiao, Kellen Myers, Oyita Udiani, and Nina H. Fefferman. Seasonality in multihost disease systems. Ecological Modelling, 470(??):??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000886.

Nakayama:2021:EST

[NWO21] Tadanobu Nakayama, Qinxue Wang, and Tomohiro Okadera. Evaluation of spatio-temporal variations in water availability using a process-based eco-hydrology model in arid and semi-arid regions of Mongolia. Ecological Modelling, 440(??): ??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020304671.

Nolte:2020:CSO

[NYH20] Anja Nölte, Rasoul Yousefpour, and Marc Hanewinkel. Changes in sessile oak (Quercus petraea) productivity under climate change by improved leaf phenology in the 3-PG model. Ecological Modelling, 438(??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020303550.

Oliveira:2022:SES

[OBFH22] Bruno M. Oliveira, Roelof Boumans, Brian D. Fath, and Joseph Harari. Socio-ecological systems modelling of coastal urban area under a changing climate — case study for Ubatuba, Brazil. Ecological Modelling, 468(??):??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000734.

Olmos:2023:STI

[OCT^{+}23] Maxime Olmos, Jie Cao, James T. Thorson, André E. Punt, Cole C. Monnahan, Baptiste Alglave, and Cody Szuwalski. A step towards the integration of spatial dynamics in population dynamics models: Eastern Bering Sea snow crab as a case study. Ecological Modelling, 485(??):??, November 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002144.

Ogawa:2024:FRH

[OEC24] Ryo Ogawa, Jan O. Engler, and Anna F. Cord. Functional responses in habitat selection as a management tool to evaluate agri-environment schemes for farmland birds. Ecological Modelling, 494(??):??, August 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024001662.

Ogris:2020:SAC

[OFH^{+}20] Nikica Ogris, Mitja Ferlan, Tine Hauptman, Roman Pavlin, Andreja Kavčič, Maja Jurc, and Maarten de Groot. Sensitivity analysis, calibration and validation of a phenology model for Pityogenes chalcographus (CHAPY). Ecological Modelling, 430(??):??, August 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020302088.

Ojeda:2021:AED

[OHH^{+}21] Jonathan J. Ojeda, Neil Huth, Dean Holzworth, Rubí Raymundo, Robert F. Zyskowski, Sarah M. Sinton, Alexandre J. Michel, and Hamish E. Brown. Assessing errors during simulation configuration in crop models — a global case study using APSIM-Potato. Ecological Modelling, 458(??):??, October 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021002593.

Orning:2023:UST

[OHP^{+}23] Elizabeth K. Orning, Julie A. Heinrichs, David A. Pyke, Peter S. Coates, and Cameron L. Aldridge. Using stateand-transition simulation models to scope post-fire success in restoring greater sage-grouse habitat. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001278.

Omta:2023:TMC

[OHR^{+}23] Anne Willem Omta, Elizabeth A. Heiny, Harshana Rajakaruna, David Talmy, and Michael J. Follows. Trophic model closure influences ecosystem response to enrichment. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022002812.

Osakpolor:2021:MRP

[OKS^{+}21] Stephen E. Osakpolor, Mira Kattwinkel, Jens Schirmel, Alexander Feckler, Alessandro Manfrin, and Ralf B. Schäfer. Mini-review of process-based food web models and their application in aquatic-terrestrial meta-ecosystems. Ecological Modelling, 458(??):??, October 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002659.

Okuyama:2021:DEE

[Oku21] Toshinori Okuyama. Dilution effects enhance variation in parasitism risk among hosts and stabilize host–parasitoid population dynamics. Ecological Modelling, 441(??):??, February 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020304816.

Oliveira:2024:MTM

[OMC^{+}24] Diogo F. Oliveira, Gonçalo M. Marques, Nuno Carolino, José Pais, João M. C. Sousa, and Tiago Domingos. A multi-tier methodology for the estimation of individual-specific parameters of DEB models. Ecological Modelling, 494(??):??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001674.

Ohlmann:2023:QOE

[OMP^{+}23] Marc Ohlmann, Catherine Matias, Giovanni Poggiato, Stéphane Dray, Wilfried Thuiller, and Vincent Miele. Quantifying the overall effect of biotic interactions on species distributions along environmental gradients. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001552.

Olusoji:2021:PCR

[OSH^{+}21] Oluwafemi D. Olusoji, Jurg W. Spaak, Mark Holmes, Thomas Neyens, Marc Aerts, and Frederik De Laender. cyanoFilter: an R package to identify phytoplankton populations from flow cytometry data using cell pigmentation and granularity. Ecological Modelling, 460(??):??, November 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002100291X.

Oubraham:2022:VMP

[OSPZ22] Aichouche Oubraham, Patrick Saint-Pierre, and Georges Zaccour. Viability of a multi-parcel agroecological system. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001314.

OMeara:2021:CCA

[OTR^{+}21] Theresa A. O’Meara, Peter E. Thornton, Daniel M. Ricciuto, Genevieve L. Noyce, Roy L. Rich, and J. Patrick Megonigal. Considering coasts: Adapting terrestrial models to characterize coastal wetland ecosystems. Ecological Modelling, 450(??):??, June 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021001319.

Ortner:2020:ILM

[OW20] Olivia Ortner and Gudrun Wallentin. Integration of landscape metric surfaces derived from vector data improves species distribution models. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302313.

Oliveira:2022:EAM

[OZSU22] M. Oliveira, A. Zucaro, R. Santagata, and S. Ulgiati. Environmental assessment of milk production from local to regional scales. Ecological Modelling, 463(??):??, January 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003410.

Pak:2020:PSO

[PA20] S. Ya. Pak and A. I. Abakumov. Phytoplankton in the Sea of Okhotsk along Western Kamchatka: warm vs cold years. Ecological Modelling, 433(??):??, October 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303148.

Planque:2022:SPD

[PAB^{+}22] Benjamin Planque, Johanna M. Aarflot, Lucie Buttay, JoLynn Carroll, Filippa Fransner, Cecilie Hansen, Bérengère Husson, Øystein Langangen, Ulf Lindstrøm, Torstein Pedersen, Raul Primicerio, Elliot Sivel, Morten D. Skogen, Evelyn Strombom, Leif Christian Stige, Øystein Varpe, and Nigel G. Yoccoz. A standard protocol for describing the evaluation of ecological models. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022001685.

Pinheiro:2024:CVA

[PAdS^{+}24] Antonio Gebson Pinheiro, Cleber Pereira Alves, Carlos André Alves de Souza, George do Nascimento Araújo Júnior, Alexandre Maniçoba da Rosa Ferraz Jardim, José Edson Florentino de Morais, Luciana Sandra Bastos de Souza, Daniela de Carvalho Lopes, Antonio José Steidle Neto, Abelardo Antonio de Assunção Montenegro, João Emanoel Ambrósio Gomes, and Thieres George Freire da Silva. Calibration and validation of the AquaCrop model for production arrangements of forage cactus and grass in a semi-arid environment. Ecological Modelling, 488(??):??, February 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003368.

Pirasteh-Anosheh:2023:BPT

[PAPŁ^{+}23] Hadi Pirasteh-Anosheh, Agnieszka Piernik, Katarzyna Łuczak, Łukasz Mendyk, and Piotr Hulisz. The behavior of potentially toxic elements in the technogenic soil-plant system: a study of Salicornia europaea L. from sites affected by the soda industry. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002478.

Pastor:2021:EME

[Pas21] John Pastor. Ecological Modelling and Ecophysics: Agricultural and Environmental Implications. Hugo Fort. 2020. IOP Publishing Ltd. United Kingdom. 300 p. $142.20. ISBN-13: 978-0-7503-2430-4. Ecological Modelling, 449(??):??, June 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021001125.

Pimenta:2022:OSD

[PAS^{+}22] Mayra Pimenta, André Felipe Alves de Andrade, Fernando Hiago Souza Fernandes, Mayra Pereira de Melo Amboni, Renata Silva Almeida, Ana Hermínia Simões de Bello Soares, Guth Berger Falcon, Daniel Santana Lorenzo Raíces, and Paulo De Marco Júnior. One size does not fit all: Priority areas for real world problems. Ecological Modelling, 470(??): ??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001247.

Papafitsoros:2023:SMB

[PAS23] Kostas Papafitsoros, Lukáš Adam, and Gail Schofield. A social media-based framework for quantifying temporal changes to wildlife viewing intensity. Ecological Modelling, 476 (??):??, February 2023. CODEN ECMODT. ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003210.

Pacini:2023:FMI

[PBF^{+}23] Gaio Cesare Pacini, Piero Bruschi, Lorenzo Ferretti, Margherita Santoni, Francesco Serafini, and Tommaso Gaifami. Fun-Bies, a model for integrated assessment of functional biodiversity of weed communities in agro-ecosystem. Ecological Modelling, 486(??):??, December 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002594.

Panayotova:2022:MEI

[PBH22] Iordanka N. Panayotova and Maila Brucal-Hallare. Modeling ecological issues in fish dynamics: Competing predators, invasiveness, and over-harvesting. Ecological Modelling, 466(??): ??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000138.

Perez:2021:SMS

[PBS21] Alejandra Zubiria Perez, Christopher Bone, and Gordon Stenhouse. Simulating multi-scale movement decision-making and learning in a large carnivore using agent-based modelling. Ecological Modelling, 452(??):??, July 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001289.

Priyadarshi:2022:UPE

[PCK^{+}22] Anupam Priyadarshi, Ram Chandra, Michio J. Kishi, S. Lan Smith, and Hidekatsu Yamazaki. Understanding plankton ecosystem dynamics under realistic micro-scale variability requires modeling at least three trophic levels. Ecological Modelling, 467(??):??, May 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002200059X.

Portalier:2024:LMP

[PCL24] S. M. J. Portalier, J.-N. Candau, and F. Lutscher. Larval mortality from phenological mismatch can affect outbreak frequency and severity of a boreal forest defoliator. Ecological Modelling, 493(??):??, July 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001121.

Peng:2023:CTS

[PCN^{+}23] Qiang Peng, Long Chen, Karl J. Niklas, Weihao Yao, Meng Lian, and Peijian Shi. Comparison of three sigmoidal functions describing the leaf growth of Camptotheca acuminata Decne. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002612.

Perez:2022:CRW

[PDD^{+}22] Géraldine Pérez, Amaël Dupaix, Laurent Dagorn, Jean-Louis Deneubourg, Kim Holland, Sunil Beeharry, and Manuela Capello. Correlated Random Walk of tuna in arrays of Fish Aggregating Devices: a field-based model from passive acoustic tagging. Ecological Modelling, 470(??):??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002200117X.

Pfab:2020:EBT

[PDG^{+}20] Ferdinand Pfab, Graziella V. DiRenzo, Ariel Gershman, Cheryl J. Briggs, and Roger M. Nisbet. Energy budgets for tadpoles approaching metamorphosis. Ecological Modelling, 436(??):??, November 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020303318.

Perennes:2021:HFM

[PDGB21] Marie Perennes, Tim Diekötter, Jens Groß, and Benjamin Burkhard. A hierarchical framework for mapping pollination ecosystem service potential at the local scale. Ecological Modelling, 444(??):??, March 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000557.

Palamara:2022:IEP

[PDH^{+}22] Gian Marco Palamara, Stuart R. Dennis, Corinne Haenggi, Nele Schuwirth, and Peter Reichert. Investigating the effect of pesticides on Daphnia population dynamics by inferring structure and parameters of a stochastic model. Ecological Modelling, 472(??):??, October 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001818.

Pan:2023:IAM

[PDW23] Feng Pan, Zexin Diao, and Lin Wang. The impact analysis of media attention on local environmental governance based on four-party evolutionary game. Ecological Modelling, 478(??): ??, April 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000212.

Petrovskaya:2020:MTU

[PEFW20] Natalia B. Petrovskaya, John R. Ellis, Emily Forbes, and Keith F. A. Walters. Modelling a targeted use of pesticide procedure for pest populations with heterogeneous spatial distributions. Ecological Modelling, 427(??):??, July 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020301319.

Peron:2024:MBC

[Pér24] Guillaume Péron. Movement-based coexistence does not always require a functional trade-off. Ecological Modelling, 487 (??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438002300279X.

Pritchard:2023:TOR

[PF23] Alexander J. Pritchard and Nina H. Fefferman. Tradeoffs in resource access and health by avoidance of selffouling, motivated via disgust. Ecological Modelling, 476 (??):??, February 2023. CODEN ECMODT. ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003234.

Pollesch:2022:DIP

[PFK^{+}22] N. L. Pollesch, K. M. Flynn, S. M. Kadlec, J. A. Swintek, S. Raimondo, and M. A. Etterson. Developing integral projection models for ecotoxicology. Ecological Modelling, 464(??):??, February 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003586.

Panaro:2021:CVS

[PFM^{+}21] D. B. Panaro, L. Frunzo, M. R. Mattei, V. Luongo, and G. Esposito. Calibration, validation and sensitivity analysis of a surface-based ADM1 model. Ecological Modelling, 460(??):??, November 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021002799.

Pal:2023:DCS

[PG23] Saheb Pal and Indrajit Ghosh. Dynamics of a coupled socioenvironmental model: an application to global CO_{2} emissions. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000078.

Pete:2020:BMC

[PGB^{+}20] Romain Pete, Thomas Guyondet, Beatrice Bec, Valérie Derolez, Ludovic Cesmat, Franck Lagarde, Stéphane Pouvreau, Annie Fiandrino, and Marion Richard. A box-model of carrying capacity of the Thau lagoon in the context of ecological status regulations and sustainable shellfish cultures. Ecological Modelling, 426(??):??, June 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301216.

Pollard:2020:PTD

[PGdAM^{+}20] Ciarán P. Pollard, Christine T. Griffin, Rafael de Andrade Moral, Catriona Duffy, Julien Chuche, Michael T. Gaffney, Reamonn M. Fealy, and Rowan Fealy. phenModel: a temperature-dependent phenology/voltinism model for a herbivorous insect incorporating facultative diapause and budburst. Ecological Modelling, 416(??):??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019304181.

Pascual:2022:IES

[PGP^{+}22] Adrián Pascual, Christian P. Giardina, Nicholas A. Povak, Paul F. Hessburg, and Gregory P. Asner. Integrating ecosystem services modeling and efficiencies in decisionsupport models conceptualization for watershed management. Ecological Modelling, 466(??):??, April 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000072.

Puertes:2020:IMU

[PGSL^{+}20] Cristina Puertes, María González-Sanchis, Antonio Lidón, Inmaculada Bautista, Antonio D. del Campo, Cristina Lull, and Félix Francés. Improving the modelling and understanding of carbon-nitrogen-water interactions in a semiarid Mediterranean oak forest. Ecological Modelling, 420(??):??, March 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002030048X.

Panayotova:2023:BAH

[PHK23] Iordanka N. Panayotova, John Herrmann, and Nathan Kolling. Bioeconomic analysis of harvesting within a predator–prey system: a case study in the Chesapeake Bay fisheries. Ecological Modelling, 480(??):??, June 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000583.

Paz:2020:ALU

[PHL20] Diego Bengochea Paz, Kirsten Henderson, and Michel Loreau. Agricultural land use and the sustainability of social-ecological systems. Ecological Modelling, 437(??):??, December 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303823.

Pierre:2023:EFM

[Pie23] Jean-Sébastien Pierre. Ecological forecasting models: Accuracy versus decisional quality. Ecological Modelling, 482(??): ??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001230.

Pakrooh:2023:MPI

[PK23] Parisa Pakrooh and Muhamad Abdul Kamal. Modeling the potential impacts of climate change on wheat yield in Iran: Evidence from national and provincial data analysis. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002430.

Perryman:2023:AEM

[PKB^{+}23] Holly A. Perryman, Isaac C. Kaplan, Julia L. Blanchard, Gavin Fay, Sarah K. Gaichas, Vidette L. McGregor, Hem Nalini Morzaria-Luna, Javier Porobic, Howard Townsend, and Elizabeth A. Fulton. Atlantis Ecosystem Model Summit 2022: Report from a workshop. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001734.

Petter:2021:MLT

[PKO^{+}21] Gunnar Petter, Holger Kreft, Yongzhi Ong, Gerhard Zotz, and Juliano Sarmento Cabral. Modelling the long-term dynamics of tropical forests: From leaf traits to whole-tree growth patterns. Ecological Modelling, 460(??):??, November 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021002866.

Prakash:2022:MCB

[PM22] Shivendra Prakash and Corey D. Markfort. A Monte-Carlo based 3-D ballistics model for guiding bat carcass surveys using environmental and turbine operational data. Ecological Modelling, 470(??):??, August 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001399.

Pilowsky:2022:SHM

[PMB^{+}22] Julia A. Pilowsky, Andrea Manica, Stuart Brown, Carsten Rahbek, and Damien A. Fordham. Simulations of human migration into North America are more sensitive to demography than choice of palaeoclimate model. Ecological Modelling, 473(??):??, November 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002174.

Pointurier:2021:IBM

[PMP^{+}21] Olivia Pointurier, Delphine Moreau, Loïc Pagès, Jacques Caneill, and Nathalie Colbach. Individual-based 3D modelling of root systems in heterogeneous plant canopies at the multiannual scale. Case study with a weed dynamics model. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304415.

Paleari:2021:SAU

[PMZ^{+}21] Livia Paleari, Ermes Movedi, Michele Zoli, Andrea Burato, Irene Cecconi, Jabir Errahouly, Eleonora Pecollo, Carla Sorvillo, and Roberto Confalonieri. Sensitivity analysis using Morris: Just screening or an effective ranking method? Ecological Modelling, 455(??):??, September 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002088.

Plekhanova:2021:HDL

[PNGESS21] Elena Plekhanova, Pascal A. Niklaus, Jean-Philippe Gastellu-Etchegorry, and Gabriela Schaepman-Strub. How does leaf functional diversity affect the light environment in forest canopies? An in-silico biodiversity experiment. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304580.

Porto:2021:TPE

[Por21] Lito Elio Porto. “Teaching ‘process ecology’ at the undergraduate level: a curricular response to a latent yearning for an ecologically-based understanding of transdisciplinary realities”. Ecological Modelling, 462(??):??, December 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003264.

Perin:2021:STL

[PPH^{+}21] Jérôme Perin, Mikhail Pitchugin, Jacques Hébert, Yves Brostaux, Philippe Lejeune, and Gauthier Ligot. SIMREG, a tree-level distance-independent model to simulate forest dynamics and management from national forest inventory (NFI) data. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304464.

Podani:2022:CAB

[PPSS22] János Podani, Katalin Patonai, Péter Szabó, and András Szilágyi. Coefficients of association between nominal and fully ranked ordinal variables with applications to ecological network analysis. Ecological Modelling, 466(??):??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022000023.

Pahl:2021:CAS

[PR21] Cameron C. Pahl and Luis A. Ruedas. Carnosaurs as apex scavengers: Agent-based simulations reveal possible vulture analogues in late Jurassic dinosaurs. Ecological Modelling, 458(??):??, October 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021002611.

Poppe:2022:AFI

[PR22] Katrina L. Poppe and John M. Rybczyk. Assessing the future of an intertidal seagrass meadow in response to sea level rise with a hybrid ecogeomorphic model of elevation change. Ecological Modelling, 469(??):??, July 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000904.

Pahl:2023:AWP

[PR23] Cameron C. Pahl and Luis A. Ruedas. Allosaurus was predominantly a scavenger. Ecological Modelling, 477(??):??, March 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022003593. See [KHR23].

Paulin:2020:ILK

[PRdN^{+}20] M. J. Paulin, M. Rutgers, T. de Nijs, A. J. Hendriks, K. R. Koopman, T. Van Buul, M. Frambach, G. Sardano, and A. M. Breure. Integration of local knowledge and data for spatially quantifying ecosystem services in the Hoeksche Waard, the Netherlands. Ecological Modelling, 438(??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020303999.

Pretzsch:2022:FCR

[Pre22] Hans Pretzsch. Facilitation and competition reduction in tree species mixtures in Central Europe: Consequences for growth modeling and forest management. Ecological Modelling, 464(??):??, February 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003574.

Pauer:2024:MNT

[PRM^{+}24] James J. Pauer, Mark D. Rowe, Wilson Melendez, Thomas P. Hollenhorst, Lisa L. Lowe, Peter J. Alsip, Dale M. Robertson, and Steven A. Pothoven. Modeling nearshore total phosphorus in lake Michigan using linked hydrodynamic and water quality models. Ecological Modelling, 493(??):??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001066.

Peirce:2023:IPM

[PSBE23] J. Peirce, G. Sandland, B. Bennie, and R. Erickson. An integral projection model for gizzard shad (Dorosoma cepedianum) utilizing density-dependent age-0 survival. Ecological Modelling, 477(??):??, March 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003581.

Pastor:2021:SFG

[PSG21] Juan Manuel Pastor, Luciano Stucchi, and Javier Galeano. Study of a factored general logistic model of population dynamics with inter- and intraspecific interactions. Ecological Modelling, 444(??):??, March 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000466.

Pandi:2022:RPD

[PSG^{+}22] S. Senthil Pandi, A. Senthilselvi, J. Gitanjali, K. ArivuSelvan, Jagadeesh Gopal, and J. Vellingiri. Rice plant disease classification using dilated convolutional neural network with global average pooling. Ecological Modelling, 474(??):??, December 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002678.

Poggi:2021:DRG

[PSM^{+}21] Sylvain Poggi, Mike Sergent, Youcef Mammeri, Manuel Plantegenest, Ronan Le Cointe, and Yoann Bourhis. Dynamic role of grasslands as sources of soil-dwelling insect pests: New insights from in silico experiments for pest management strategies. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304439.

Pasquali:2022:MED

[PSM22] S. Pasquali, C. Soresina, and E. Marchesini. Mortality estimate driven by population abundance field data in a stage-structured demographic model. The case of Lobesia botrana. Ecological Modelling, 464(??):??, February 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003823.

Pommerening:2022:STI

[PSW22] Arne Pommerening, Hubert Sterba, and Philip West. Sampling theory inspires quantitative forest ecology: the story of the relascope kernel function. Ecological Modelling, 467(??): ??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000473.

Pommerening:2020:NNN

[PSZ20] Arne Pommerening, Janusz Szmyt, and Gongqiao Zhang. A new nearest-neighbour index for monitoring spatial size diversity: the hyperbolic tangent index. Ecological Modelling, 435(??):??, November 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020303021.

Pinto:2020:DPE

[PTBFP^{+}20] Carlos F. Pinto, Daniel Torrico-Bazoberry, Luis Flores-Prado, Ramiro O. Bustamante, and Hermann M. Niemeyer. Demographic and performance effects of alternative host use in a neotropical treehopper (Hemiptera: Membracidae). Ecological Modelling, 416(??):??, January 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019304132.

Petsas:2023:CCS

[PTD^{+}23] Panagiotis Petsas, Marianna Tzivanopoulou, Aggeliki Doxa, Sévrine F. Sailley, and Antonios D. Mazaris. Climate change on sea currents is not expected to alter contemporary migration routes of loggerhead sea turtles. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003180.

Paff:2023:CWL

[PTF23] K. Paff, D. Timlin, and D. H. Fleisher. A comparison of wheat leaf-appearance rate submodules for DSSAT CROPSIM-CERES (CSCER). Ecological Modelling, 482(??): ??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001370.

Perryman:2020:RDM

[PTG^{+}20] Holly A. Perryman, Joseph H. Tarnecki, Arnaud Grüss, Elizabeth A. Babcock, Skyler R. Sagarese, Cameron H. Ainsworth, and Alisha M. Gray DiLeone. A revised diet matrix to improve the parameterization of a West Florida Shelf Ecopath model for understanding harmful algal bloom impacts. Ecological Modelling, 416(??):??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303989.

Puts:2020:IIH

[PTNR^{+}20] Miriam Püts, Marc Taylor, Ismael Núñez-Riboni, Jeroen Steenbeek, Moritz Stäbler, Christian Möllmann, and Alexander Kempf. Insights on integrating habitat preferences in process-oriented ecological models — a case study of the southern North Sea. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002030260X.

Pashanejad:2023:ASM

[PTRP23] Ehsan Pashanejad, Hugo Thierry, Brian E. Robinson, and Lael Parrott. The application of semantic modelling to map pollination service provisioning at large landscape scales. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001837.

Pappalardo:2021:TIM

[PVS^{+}21] Sonia Pappalardo, María Villa, Sónia A. P. Santos, Jacinto Benhadi-Marín, José Alberto Pereira, and Ezio Venturino. A tritrophic interaction model for an olive tree pest, the olive moth — Prays oleae (Bernard). Ecological Modelling, 462(??):??, December 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003240.

Pan:2024:SLE

[PWN^{+}24] Guoliang Pan, Ho Yi Wan, David R. Nash, Kun Shi, and Samuel Cushman. Snow leopards exhibit non-stationarity in scale-dependent habitat selection between two national protected areas in China. Ecological Modelling, 494(??):??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001479.

Pata:2021:CIE

[PYD^{+}21a] Patrick R. Pata, Aletta T. Yñiguez, Josephine Dianne L. Deauna, Asuncion B. De Guzman, Cesaria R. Jimenez, Roselle T. Borja-Del Rosario, and Cesar L. Villanoy. Corrigendum to Insights into the environmental conditions contributing to variability in the larval recruitment of the tropical sardine Sardinella lemuru. Ecological Modelling, 456(??):??, September 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021002350.

Pata:2021:IEC

[PYD^{+}21b] Patrick R. Pata, Aletta T. Yñiguez, Josephine Dianne L. Deauna, Asuncion B. De Guzman, Cesaria R. Jimenez, Roselle T. Borja-Del Rosario, and Cesar L. Villanoy. Insights into the environmental conditions contributing to variability in the larval recruitment of the tropical sardine Sardinella lemuru. Ecological Modelling, 451(??):??, July 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021001290.

Ping:2021:MTD

[PZH^{+}21] Jiaye Ping, Jian Zhou, Kun Huang, Xiaoying Sun, Huanfa Sun, and Jianyang Xia. Modeling the typhoon disturbance effect on ecosystem carbon storage dynamics in a subtropical forest of China’s coastal region. Ecological Modelling, 455(??):??, September 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021001976.

Qin:2023:MIU

[QS23] Zilong Qin and Zongyao Sha. Modeling the impact of urbanization and climate changes on terrestrial vegetation productivity in China by a neighborhood substitution analysis. Ecological Modelling, 482(??):??, August 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001369.

Qi:2021:DRI

[QZSV21] Man Qi, Feng Zhao, Tao Sun, and Alexey Voinov. Disentangling the relative influence of regeneration processes on marsh plant assembly with a stage-structured plant assembly model. Ecological Modelling, 455(??):??, September 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021002064.

Rufino:2021:WSI

[RAB21] Marta M. Rufino, Camille Albouy, and Anik Brind’Amour. Which spatial interpolators I should use? A case study applying to marine species. Ecological Modelling, 449(??):??, June 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021000727.

Rouby:2024:ATT

[RAC^{+}24] Etienne Rouby, Matthieu Authier, Emmanuelle Cam, Ursula Siebert, and Floriane Plard. Addressing temporal trends in survivorship from cross-sectional sampling designs: a modelling framework with applications for megafauna conservation. Ecological Modelling, 490(??):??, April 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002400036X.

Rodriguez:2021:ASE

[RBEC21] Maykol Rodriguez, Antonio Bodini, Francisco J. Escobedo, and Nicola Clerici. Analyzing socio-ecological interactions through qualitative modeling: Forest conservation and implications for sustainability in the peri-urban Bogota (Colombia). Ecological Modelling, 439(??):??, January 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304117.

Rohan:2021:MEM

[RBEH21] Sean K. Rohan, David A. Beauchamp, Timothy E. Essington, and Adam G. Hansen. Merging empirical and mechanistic approaches to modeling aquatic visual foraging using a generalizable visual reaction distance model. Ecological Modelling, 457(??):??, October 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002465.

Rana:2023:MES

[RBGB23] Sourav Rana, Ayanendranath Basu, Sinchan Ghosh, and Sabyasachi Bhattacharya. Moths exhibit strong memory among cooperative species of other taxonomic groups: an empirical study. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT. ISSN ???? URL http://www.

sciencedirect.com/science/article/pii/S0304380022003337.

Rai:2023:MGR

[RBLS23] Anu Rai, Tawqir Bashir, Elio Guarionex Lagunes–Díaz, and Bibek Shrestha. Modeling Ganges river dolphin distribution and prioritizing areas for efficient conservation planning — a range-wide assessment. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002300090X.

Rawlings:2020:FIC

[RBMRR20] E. Soraya Rawlings, J. C. Barrera-Martinez, and Vicente Rico-Ramirez. Fisher information calculation in a complex ecological model: an optimal control-based approach. Ecological Modelling, 416(??):??, January 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303539.

Ruiz-Benito:2020:AMD

[RBVL^{+}20] Paloma Ruiz-Benito, Giorgio Vacchiano, Emily R. Lines, Christopher P. O. Reyer, Sophia Ratcliffe, Xavier Morin, Florian Hartig, Annikki Mäkelä, Rasoul Yousefpour, Jimena E. Chaves, Alicia Palacios-Orueta, Marta Benito-Garzón, Cesar Morales-Molino, J. Julio Camarero, Alistair S. Jump, Jens Kattge, Aleksi Lehtonen, Andreas Ibrom, Harry J. F. Owen, and Miguel A. Zavala. Available and missing data to model impact of climate change on European forests. Ecological Modelling, 416(??):??, January 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303783.

Ramirez-Carrasco:2022:MMI

[RCCLMGV22] C. Ramirez-Carrasco, F. Córdova-Lepe, F. N. Moreno-Gómez, and N. A. Velásquez. A mathematical model for the impact of noise on population dynamics of a single species experiencing Lombard effect. Ecological Modelling, 470(??):??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022001326.

Rossini:2020:RDD

[RCSS20] Luca Rossini, Mario Contarini, Maurizio Severini, and Stefano Speranza. Reformulation of the distributed delay model to describe insect pest populations using count variables. Ecological Modelling, 436(??):??, November 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303562.

Rincon:2024:PIP

[REGI^{+}24] Diego F. Rincon, Evan D. Esch, Javier Gutierrez-Illan, Melissa Tesche, and David W. Crowder. Predicting insect population dynamics by linking phenology models and monitoring data. Ecological Modelling, 493(??):??, July 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001510.

Ratte-Fortin:2022:RMP

[RFCEL22] Claudie Ratté-Fortin, Karem Chokmani, Anas El Alem, and Isabelle Laurion. A regional model to predict the occurrence of natural events: Application to phytoplankton blooms in continental waterbodies. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022002381.

Rodigheri:2024:TBW

[RFdL^{+}24] Grazieli Rodigheri, Denise Cybis Fontana, Luana Becker da Luz, Genei Antonio Dalmago, Lucimara Wolfarth Schirmbeck, Juliano Schirmbeck, Jorge Alberto de Gouvêa, and Gilberto Rocca da Cunha. TVDI-based water stress coefficient to estimate net primary productivity in soybean areas. Ecological Modelling, 490(??):??, April 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000255.

Rius:2023:HFD

[RFF^{+}23] Bianca Fazio Rius, João Paulo Darela Filho, Katrin Fleischer, Florian Hofhansl, Carolina Casagrande Blanco, Anja Rammig, Tomas Ferreira Domingues, and David Montenegro Lapola. Higher functional diversity improves modeling of Amazon forest carbon storage. Ecological Modelling, 481(??): ??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000510.

Riera:2023:CRP

[RFHR23] Rodrigo Riera, Brian D. Fath, Ada M. Herrera, and Ricardo A. Rodríguez. Concerns regarding the proposal for an ecological equation of state: an assessment starting from the organic biophysics of ecosystems (OBEC). Ecological Modelling, 484(??):??, October 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0304380023001916. See response [NBU^{+}23] and reply [RFHR24].

Riera:2024:SRI

[RFHR24] Rodrigo Riera, Brian D. Fath, Ada M. Herrera, and Ricardo A. Rodríguez. A strategic roadmap for interdisciplinary modeling in ecology: the result of reading Defining an ecological equation of state: Response to Riera et al. 2023 (Newman et al., 2023). Ecological Modelling, 490(??):??, April 2024.

CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000449. See [RFHR23, NBU^{+}23].

Rau:2022:TIT

[RFJ^{+}22] E-Ping Rau, Fabian Fischer, Émilie Joetzjer, Isabelle Maréchaux, I Fang Sun, and Jérôme Chave. Transferability of an individual- and trait-based forest dynamics model: a test case across the tropics. Ecological Modelling, 463(??):??, January 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002100346X.

Ratte-Fortin:2023:PVN

[RFPRC23] Claudie Ratté-Fortin, Jean-François Plante, Alain N. Rousseau, and Karem Chokmani. Parametric versus nonparametric machine learning modelling for conditional density estimation of natural events: Application to harmful algal blooms. Ecological Modelling, 482(??):??, August 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001461.

Roy:2022:NGC

[RGB22] Trina Roy, Sinchan Ghosh, and Sabyasachi Bhattacharya. A new growth curve model portraying the stress response regulation of fish: Illustration through particle motion and real data. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001120.

Rouger:2023:SAC

[RGB^{+}23] Baptiste Rouger, Isabelle Goldringer, Pierre Barbillon, Anne Miramon, Abdel Kader Naino Jika, and Mathieu Thomas. Sensitivity analysis of a crop metapopulation model. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022002757.

Rodriguez:2020:MSS

[RGD20] Andres F. Rodriguez, Stefan Gerber, and Samira H. Daroub. Modeling soil subsidence in a subtropical drained peatland.

The case of the Everglades Agricultural Area. Ecological Modelling, 415(??):??, January 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019303679.

Richter:2022:CSB

[RGD^{+}22] Ian A. Richter, Henrique Corrêa Giacomini, Derrick Tupper De Kerckhove, Donald A. Jackson, and Nicholas Edward Jones. Correcting for size bias in electrofishing removal samples. Ecological Modelling, 467(??):??, May 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000527.

Rodriguez-Gonzalez:2020:EFL

[RGRMRR20] Pablo T. Rodriguez-Gonzalez, Ramiro Rico-Martinez, and Vicente Rico-Ramirez. Effect of feedback loops on the sustainability and resilience of human-ecosystems. Ecological Modelling, 426(??):??, June 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300909.

Ren:2020:DEB

[RJY^{+}20] Jeffrey S. Ren, Xianshi Jin, Tao Yang, Sebastiaan A. L. M. Kooijman, and Xiujuan Shan. A dynamic energy budget model for small yellow croaker Larimichthys polyactis: Parameterisation and application in its main geographic distribution waters. Ecological Modelling, 427(??):??, July 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002030123X.

Rastetter:2020:AMR

[RK20] Edward B. Rastetter and Bonnie L. Kwiatkowski. An approach to modeling resource optimization for substitutable and interdependent resources. Ecological Modelling, 425(??): ??, June 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301058.

Rhomad:2021:MIN

[RKN^{+}21] Hanane Rhomad, Karima Khalil, Ramiro Neves, Blaid Bougadir, and Khalid Elkalay. Modeling investigation of the nutrients and phytoplankton dynamics in the Moroccan Atlantic coast: a case study of Agadir coast. Ecological Modelling, 447(??):??, May 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021000818.

Rotter:2020:WDF

[RLdM20] Pavel Rotter, Michel Loreau, and Claire de Mazancourt. Why do forests respond differently to nitrogen deposition? A modelling approach. Ecological Modelling, 425(??):??, June 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002030106X.

Richard:2024:TSR

[RLH24] Romain Richard, Jude Antoine Lugsanay, and Shu-Ping Huang. The temperature-size rule in the context of Dynamic Energy Budget theory. Ecological Modelling, 493(??):??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001492.

Ricouard:2023:MYS

[RLM23] Antoine Ricouard, Sigrid Lehuta, and Stéphanie Mahévas. Are maximum yields sustainable? Effect of intra-annual time-scales on MSY, stability and resilience. Ecological Modelling, 479(??):??, May 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000492.

Rizzuto:2024:AVN

[RLS^{+}24] Matteo Rizzuto, Shawn J. Leroux, Oswald J. Schmitz, Eric Vander Wal, Yolanda F. Wiersma, and Travis R. Heckford. Animal-vectored nutrient flows across resource gradients influence the nature of local and meta-ecosystem functioning. Ecological Modelling, 488(??):??, February 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003009.

Romanov:2020:LBP

[RM20] Michael S. Romanov and Vladimir B. Masterov. Low breeding performance of the Steller’s sea eagle (Haliaeetus pelagicus) causes the populations to decline. Ecological Modelling, 420(??):??, March 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019303850.

Rondon:2024:BMS

[RMA^{+}24] Diego Rondon, Samu Mäntyniemi, Jouni Aspi, Laura Kvist, and Mikko J. Sillanpää. A Bayesian multi-state model with data augmentation for estimating population size and effect of inbreeding on survival. Ecological Modelling, 490(??):??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000504.

Revilla:2021:PCU

[RMK21] Tomás A. Revilla, Thomas Marcou, and Vlastimil Křivan. Plant competition under simultaneous adaptation by herbivores and pollinators. Ecological Modelling, 455(??):??, September 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380021001939.

Ramakhanna:2022:CSP

[RMO22] Selebalo Joseph Ramakhanna, Botle Esther Mapeshoane, and Christian Thine Omuto. Carbon sequestration potential in croplands in Lesotho. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001624.

Raapoto:2024:ARC

[RMV^{+}24] H. Raapoto, C. J. Monaco, S. Van Wynsberge, R. Le Gendre, and J. Le Luyer. Assessing regional connectivity patterns of bivalvia in fragmented archipelagos: Insights from biophysical modeling in French Polynesia. Ecological Modelling, 489(??): ??, March 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000152.

Ricotta:2022:NPM

[RP22] Carlo Ricotta and Sandrine Pavoine. A new parametric measure of functional dissimilarity: Bridging the gap between the Bray–Curtis dissimilarity and the Euclidean distance. Ecological Modelling, 466(??):??, April 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000084.

Ruiz-Perez:2020:WFC

[RPARM20] María Rocío Ruíz-Pérez, M. Desirée Alba-Rodríguez, and Madelyn Marrero. The water footprint of city naturalisation. Evaluation of the water balance of city gardens. Ecological Modelling, 424(??):??, May 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301034.

Ramirez:2023:MPD

[RPB^{+}23] Jorge A. Ramirez, Nadav Peleg, Andy J. Baird, Dylan M. Young, Paul J. Morris, Marie Larocque, and Michelle Garneau. Modelling peatland development in high-boreal Quebec, Canada, with DigiBog Boreal. Ecological Modelling, 478(??): ??, April 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000261.

Rakel:2020:IBD

[RPG20] Kim J. Rakel, Thomas G. Preuss, and André Gergs. Individual-based dynamic energy budget modelling of earthworm life-histories in the context of competition. Ecological Modelling, 432(??):??, September 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020302921.

Rotllan-Puig:2021:DMB

[RPT21] Xavier Rotllan-Puig and Anna Traveset. Determining the minimal background area for species distribution models: MinBAR package. Ecological Modelling, 439(??):??, January 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304191.

Ren:2020:OAD

[RRCZ20] Jeffrey S. Ren, Norman L. C. Ragg, Vonda J. Cummings, and Jihong Zhang. Ocean acidification and dynamic energy budget models: Parameterisation and simulations for the greenlipped mussel. Ecological Modelling, 426(??):??, June 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020301411.

Ruiz-Rodriguez:2022:RWB

[RRFLV^{+}22] Carmen Ruiz-Rodríguez, Javier Fernández-López, Joaquín Vicente, José A. Blanco-Aguiar, and Pelayo Acevedo. Revisiting wild boar spatial models based on hunting yields to assess their predictive performance on interpolation and extrapolation areas. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002200151X.

Reichert:2023:MCC

[RRP^{+}23] Tatiana Reichert, Anja Rammig, Phillip Papastefanou, Laynara F. Lugli, João P. Darela Filho, Konstantin Gregor, Lucia Fuchslueger, Carlos A. Quesada, and Katrin Fleischer. Modeling the carbon costs of plant phosphorus acquisition in Amazonian forests. Ecological Modelling, 485(??):??, November 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002211.

Rossini:2021:GOB

[RRSG21] Luca Rossini, Nicolás Bono Rosselló, Stefano Speranza, and Emanuele Garone. A general ODE-based model to describe the physiological age structure of ectotherms: Description and application to Drosophila suzukii. Ecological Modelling, 456(??):??, September 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002313.

Raj:2023:SPM

[RS23] Alok Raj and Laxmi Kant Sharma. Spatial E-PSR modelling for ecological sensitivity assessment for arid rangeland resilience and management. Ecological Modelling, 478(??):??, April 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002300011X.

Rafiei-Sardooi:2022:LTA

[RSASP22] Elham Rafiei-Sardooi, Ali Azareh, Sharif Joorabian Shooshtari, and Eric J. R. Parteli. Long-term assessment of land-use and climate change on water scarcity in an arid basin in Iran. Ecological Modelling, 467(??):??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000576.

Raju:2023:VHR

[RSB^{+}23] Anjumol Raju, S. Sijikumar, Pramit Kumar Deb Burman, Vinu Valsala, Yogesh K. Tiwari, Sandipan Mukherjee, Priyanka Lohani, and Kireet Kumar. Very highresolution Net Ecosystem Exchange over India using Vegetation Photosynthesis and Respiration Model (VPRM) simulations. Ecological Modelling, 481(??):??, July 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000686.

Ricci:2021:MTR

[RSC^{+}21] P. Ricci, L. Sion, F. Capezzuto, G. Cipriano, G. D’Onghia, S. Libralato, P. Maiorano, A. Tursi, and R. Carlucci. Modelling the trophic roles of the demersal Chondrichthyes in the Northern Ionian Sea (Central Mediterranean Sea). Ecological Modelling, 444(??):??, March 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000405.

Ricci:2023:IFC

[RSC^{+}23] P. Ricci, N. Serpetti, D. Cascione, G. Cipriano, G. D’Onghia, D. De Padova, C. Fanizza, M. Ingrosso, and R. Carlucci. Investigating fishery and climate change effects on the conservation status of odontocetes in the Northern Ionian Sea (Central Mediterranean Sea). Ecological Modelling, 485(??):??, November 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002302.

Rana:2024:DBS

[RSC^{+}24] Divyashree Rana, Caroline Charão Sartor, Luca Chiaverini, Samuel Alan Cushman, Żaneta Kaszta, Uma Ramakrishnan, and David W. Macdonald. Differentially biased sampling strategies reveal the non-stationarity of species distribution models for Indian small felids. Ecological Modelling, 493(??): ??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001376.

Ramampiandra:2023:CML

[RSWS23] Emma Chollet Ramampiandra, Andreas Scheidegger, Jonas Wydler, and Nele Schuwirth. A comparison of machine learning and statistical species distribution models: Quantifying overfitting supports model interpretation. Ecological Modelling, 481(??):??, July 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000819.

Rodrigues:2021:MST

[RTH^{+}21] Vinícius Barros Rodrigues, Elio Tuci, Horst Holstein, Miriam S. Bowen, Diogo Andrade Costa, Lucy T. Takahashi, Paulo Fellipe Cristaldo, and Og DeSouza. Macro-symbiosis triggered by a defence-reproduction tradeoff in a social insect. Ecological Modelling, 462(??):??, December 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003276.

Renu:2023:AIV

[RUTY23] Renu, Ranjit Kumar Upadhyay, S. P. Tiwari, and R. P. Yadav. Analysis of interval-valued model for interaction between plankton-fish population in marine ecosystem. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001795.

Rao:2022:MMI

[RWF22] B. Veena Shankara Narayana Rao, J. R. Walton, and M. Fujiwara. A mathematical model to investigate the effects of fishing zone configurations and mass dependent rates on biomass yield: Application to brown shrimp in Gulf of Mexico. Ecological Modelling, 463(??):??, January 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003288.

Russell:2021:SMH

[RWS^{+}21] Robin E. Russell, Daniel P. Walsh, Michael D. Samuel, Martin D. Grunnill, and Tonie E. Rocke. Space matters: host spatial structure and the dynamics of plague transmission. Ecological Modelling, 443(??):??, March 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000235.

Sierra-Altamiranda:2020:SCP

[SACE^{+}20] Alvaro Sierra-Altamiranda, Hadi Charkhgard, Mitchell Eaton, Julien Martin, Simeon Yurek, and Bradley J. Udell. Spatial conservation planning under uncertainty using modern portfolio theory and Nash bargaining solution. Ecological Modelling, 423(??):??, May 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300880.

Sillero:2021:WMS

[SACEU^{+}21] Neftalí Sillero, Salvador Arenas-Castro, Urtzi Enriquez-Urzelai, Cândida Gomes Vale, Diana Sousa-Guedes, Fernando Martínez-Freiría, Raimundo Real, and A. Márcia Barbosa. Want to model a species niche? A step-by-step guideline on correlative ecological niche modelling. Ecological Modelling, 456(??):??, September 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002301.

Scarpelin:2022:VLD

[SAdA^{+}22] Juliano Scarpelin, Feni Dalano Roosevelt Agostinho, Cecília Maria Villas Bôas de Almeida, Biagio Fernando Giannetti, and Lívia Cristina Pinto Dias. Valuation of losses and damages resulting from the Fundão’s dam failure: an emergy perspective. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001612.

Sakiyama:2023:SIM

[Sak23] Tomoko Sakiyama. Spatial inconsistency of memorized positions produces different types of movements. Ecological Modelling, 481(??):??, July 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002300087X.

Salimova:2022:MME

[SAL^{+}22] Guzel Salimova, Alisa Ableeva, Tatiana Lubova, Aidar Sharafutdinov, and Irek Araslanbaev. Multidimensional modeling of the economy of forest management and reforestation. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002010.

Sebastiani:2021:MAQ

[SBF^{+}21] A. Sebastiani, E. Buonocore, P. P. Franzese, A. Riccio, E. Chianese, L. Nardella, and F. Manes. Modeling air quality regulation by green infrastructure in a Mediterranean coastal urban area: the removal of PM_{10} in the Metropolitan City of Naples (Italy). Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304476.

Somodi:2024:CCD

[SBFBDM24] Imelda Somodi, Ákos Bede-Fazekas, Zoltán Botta-Dukát, and Zsolt Molnár. Confidence and consistency in discrimination: a new family of evaluation metrics for potential distribution models. Ecological Modelling, 491(??):??, May 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000553.

Seixas:2022:EER

[SBM^{+}22] Hugo Tameirão Seixas, Nathaniel A. Brunsell, Elisabete Caria Moraes, Gabriel de Oliveira, and Guilherme Mataveli. Exploring the ecosystem resilience concept with land surface model scenarios. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003616.

Serodio:2023:LNC

[SBMF23] João Serôdio, Alexandra Bastos, Jérôme Morelle, and Silja Frankenbach. Light niche construction: Motility of sedimentinhabiting diatoms determines the experienced light environment. Ecological Modelling, 481(??):??, July 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001102.

Steen:2024:HSS

[SBMG24] Bart Steen, Olivier Broennimann, Luigi Maiorano, and Antoine Guisan. How sensitive are species distribution models to different background point selection strategies? A test with species at various equilibrium levels. Ecological Modelling, 493(??):??, July 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002400142X.

Suppo:2020:TPD

[SBR20] Christelle Suppo, Audrey Bras, and Christelle Robinet. A temperature- and photoperiod-driven model reveals complex temporal population dynamics of the invasive box tree moth in Europe. Ecological Modelling, 432(??):??, September 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020302994.

Sadchatheeswaran:2020:MCT

[SBS^{+}20] Saachi Sadchatheeswaran, George M. Branch, Lynne J. Shannon, Coleen L. Moloney, Marta Coll, and Tamara B. Robinson. Modelling changes in trophic and structural impacts of alien ecosystem engineers on a rocky-shore island. Ecological Modelling, 433(??):??, October 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302970.

Sadchatheeswaran:2021:NAE

[SBS^{+}21] Saachi Sadchatheeswaran, George M. Branch, Lynne J. Shannon, Marta Coll, and Jeroen Steenbeek. A novel approach to explicitly model the spatiotemporal impacts of structural complexity created by alien ecosystem engineers in a marine benthic environment. Ecological Modelling, 459(??):??, November 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021002830.

Stanley:2024:MUE

[SBWB24] Rebecca E. Stanley, Matthew V. Bilskie, C. Brock Woodson, and James E. Byers. A model for understanding the effects of flow conditions on oyster reef development and impacts to wave attenuation. Ecological Modelling, 489(??):??, March 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000164.

Soares:2021:SIA

[SC21] L. M. V. Soares and M. C. Calijuri. Sensitivity and identifiability analyses of parameters for water quality modeling of subtropical reservoirs. Ecological Modelling, 458(??):??, October 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002100274X.

Srivastava:2023:DDM

[SC23] Vivek Srivastava and Allan L. Carroll. Dynamic distribution modelling using a native invasive species, the mountain pine beetle. Ecological Modelling, 482(??):??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001400.

Schartel:2024:BSC

[SC24a] Tyler E. Schartel and Yong Cao. Background selection complexity influences Maxent predictive performance in freshwater systems. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023003228.

Sikk:2024:FSS

[SC24b] Kaarel Sikk and Geoffrey Caruso. Framing settlement systems as spatial adaptive systems. Ecological Modelling, 490(??):??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000413.

Sillero:2023:CLP

[SCACB23] Neftalí Sillero, João Carlos Campos, Salvador Arenas-Castro, and A. Márcia Barbosa. A curated list of R packages for ecological niche modelling. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT. ISSN

????  URL http://www.sciencedirect.com/science/
article/pii/S0304380022003404.

Santostasi:2020:ADH

[SCB^{+}20] Nina Luisa Santostasi, Paolo Ciucci, Giovanni Bearzi, Silvia Bonizzoni, and Olivier Gimenez. Assessing the dynamics of hybridization through a matrix modelling approach. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301927.

Sato:2023:DSS

[SCG^{+}23] Hiromitsu Sato, Emeline Chaste, Martin P. Girardin, Jed O. Kaplan, Christelle Hély, Jean-Noël Candau, and Stephen J. Mayor. Dynamically simulating spruce budworm in Eastern Canada and its interactions with wildfire. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001436.

Sinha:2021:MLS

[SCH^{+}21] Paramita Sinha, Robert C. Coville, Satoshi Hirabayashi, Brian Lim, Theodore A. Endreny, and David J. Nowak. Modeling lives saved from extreme heat by urban tree cover. Ecological Modelling, 449(??):??, June 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002100123X.

Senapati:2024:DPA

[SD24] Ujjal Senapati and Tapan Kumar Das. Delineation of potential alternative agriculture region using RS and AHP-based GIS techniques in the drought prone upper Dwarakeswer river basin, West Bengal, India. Ecological Modelling, 490(??):??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000395.

Serrouya:2020:PER

[SDD^{+}20] R. Serrouya, M. Dickie, C. DeMars, M. J. Wittmann, and S. Boutin. Predicting the effects of restoring linear features on woodland caribou populations. Ecological Modelling, 416(??):??, January 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019303990.

Speich:2021:SMC

[SDH21] Matthias Speich, Carsten F. Dormann, and Florian Hartig. Sequential Monte-Carlo algorithms for Bayesian model calibration — a review and method comparison. Ecological Modelling, 455(??):??, September 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001708.

Sanz:2020:SMM

[SdlP20] Luis Sanz and Rafael Bravo de la Parra. Stochastic matrix metapopulation models with fast migration: Re-scaling survival to the fast scale. Ecological Modelling, 418(??):??, February 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019303370.

Sun:2021:MPP

[SDP21] Detong Sun, Peter H. Doering, and Edward J. Phlips. Modeling primary production: Developing the BZI model from the production–irradiance (P–I) curves. Ecological Modelling, 458(??):??, October 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021002623.

Steen:2023:EMO

[SDP23] Valerie A. Steen, Adam Duarte, and James T. Peterson. An evaluation of multistate occupancy models for estimating relative abundance and population trends. Ecological Modelling, 478(??):??, April 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000315.

Sturbois:2021:ECT

[SDSP^{+}21] A. Sturbois, M. De Cáceres, M. Sánchez-Pinillos, G. Schaal, O. Gauthier, P. Le Mao, A. Ponsero, and N. Desroy. Extending community trajectory analysis: New metrics and representation. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304646.

Salas-Eljatib:2020:CHG

[SE20a] Christian Salas-Eljatib. Corrigendum to “Height growth– rate at a given height: a mathematical perspective for forest productivity”. Ecological Modelling 431 (2020) 109198. Ecological Modelling, 438(??):??, December 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303811.

Salas-Eljatib:2020:HGR

[SE20b] Christian Salas-Eljatib. Height growth–rate at a given height: a mathematical perspective for forest productivity.

Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302684.

Singh:2021:GSC

[SE21] Abhyudai Singh and Brooks Emerick. Generalized stability conditions for host–parasitoid population dynamics: Implications for biological control. Ecological Modelling, 456(??):??, September 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021002167.

Singh:2022:CCG

[SE22] Abhyudai Singh and Brooks Emerick. Coexistence conditions in generalized discrete-time models of insect population dynamics. Ecological Modelling, 474(??):??, December 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022002496.

Schmid:2023:PSM

[SFGB23] Ueli Schmid, Monika Frehner, Jonas Glatthorn, and Harald Bugmann. ProForM: a simulation model for the management of mountain protection forests. Ecological Modelling, 478(??): ??, April 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002300025X.

Scott:2020:SSS

[SG20] Ryan Scott and Robin Gras. A simulation study shows impacts of genetic diversity on establishment success of digital invaders in heterogeneous environments. Ecological Modelling, 431(??):??, September 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020302441.

Smith:2022:QPC

[SGA22] Phoebe Smith, Chris Guiver, and Ben Adams. Quantifying the per-capita contribution of all components of a migratory cycle: a modelling framework. Ecological Modelling, 471(??):??, September 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002200165X.

Segurado:2022:SGC

[SGCFB22] Pedro Segurado, Cayetano Gutiérrez-Cánovas, Teresa Ferreira, and Paulo Branco. Stressor gradient coverage affects interaction identification. Ecological Modelling, 472(??):??, October 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001946.

Shen:2023:FGP

[SGH^{+}23] Anglu Shen, Shufei Gao, Christopher M. Heggerud, Hao Wang, Zengling Ma, and Sanling Yuan. Fluctuation of growth and photosynthetic characteristics in Prorocentrum shikokuense under phosphorus limitation: Evidence from field and laboratory. Ecological Modelling, 479(??):??, May 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000388.

Sadoul:2020:MWH

[SGLK20] Bastien Sadoul, Benjamin Geffroy, Stephane Lallement, and Michael Kearney. Multiple working hypotheses for hyperallometric reproduction in fishes under metabolic theory. Ecological Modelling, 433(??):??, October 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302982.

Savoca:2020:MPP

[SGP^{+}20] S. Savoca, G. Grifó, G. Panarello, M. Albano, S. Giacobbe, G. Capillo, N. Spanó, and G. Consolo. Modelling prey-predator interactions in Messina beachrock pools. Ecological Modelling, 434(??):??, October 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302763.

Schrenk:2022:QRP

[SGPSzC22] Hannah Schrenk, Carlos Garcia-Perez, Nico Schreiber, and Wolfgang zu Castell. QtAC: an R-package for analyzing complex systems development in the framework of the adaptive cycle metaphor. Ecological Modelling, 466(??):??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021003987.

Stankova:2024:PEM

[SGRDA^{+}24] Tatiana V. Stankova, Miguel Ángel González-Rodríguez, Ulises Diéguez-Aranda, Angel Ferezliev, Proletka Dimitrova, Kristiyan Kolev, and Penka Stefanova. Productivityenvironment models for Scots pine plantations in Bulgaria: an interaction of anthropogenic origin peculiarities and climate change. Ecological Modelling, 490(??):??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000437.

Schneider:2021:MMA

[SGTS21] Lisa K. Schneider, Nathalie Gypens, Tineke A. Troost, and Willem Stolte. Modeling mixoplankton along the biogeochemical gradient of the Southern North Sea. Ecological Modelling, 459(??):??, November 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0304380021002477. See corrigendum [SGTS23].

Schneider:2023:CLM

[SGTS23] Lisa K. Schneider, Natalie Gypens, Tineke A. Troost, and Willem Stolte. Corrigendum to ‘Modeling mixoplankton along the biogeochemical gradient of the Southern North Sea’ Ecological Modelling 458 (2021) 0304-3800/109690. Ecological Modelling, 481(??):??, July 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001059. See [SGTS21].

Sorel:2024:UFP

[SGV^{+}24] Maeva Sorel, Pierre-Emmanuel Gay, Camille Vernier, Sory Cissé, and Cyril Piou. Upwind flight partially explains the migratory routes of locust swarms. Ecological Modelling, 489(??): ??, March 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000115.

Skartsaeterhagen:2024:EEE

[SHF24] Maria Skartsæterhagen, Cecilie Hansen, and Elizabeth A. Fulton. Exploring ecosystem effects of underwater noise in the Nordic Seas, using the NoBa–Atlantis E2E model. Ecological Modelling, 492(??):??, June 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000929.

Schmid:2021:ITP

[SHT21] Julia S. Schmid, Andreas Huth, and Franziska Taubert. Influences of traits and processes on productivity and functional composition in grasslands: a modeling study. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304592.

Sohma:2022:MLC

[SINS22] Akio Sohma, Riku Imada, Tetsuya Nishikawa, and Hisashi Shibuki. Modeling the life cycle of four types of phytoplankton and their bloom mechanisms in a benthic-pelagic coupled ecosystem. Ecological Modelling, 467(??):??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022000102.

Shinohara:2024:ESD

[SIOM24] Kengo Shinohara, Yuji Ito, Suguru Okunishi, and Hiroto Maeda. Elucidating the seasonal dominance replacement mechanism of harmful raphidophytes Chattonella marina and Heterosigma akashiwo using the Lotka–Volterra model. Ecological Modelling, 491(??):??, May 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000486.

Santibanez:2022:DCP

[SJA^{+}22] Fernanda Santibañez, Julien Joseph, Guillermo Abramson, Marcelo N. Kuperman, María Fabiana Laguna, and Lucas A. Garibaldi. Designing crop pollination services: a spatially explicit agent-based model for real agricultural landscapes. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001983.

Sridharan:2023:SMD

[SJH^{+}23] Vamsi Krishna Sridharan, Doug Jackson, Andrew M. Hein, Russell W. Perry, Adam C. Pope, Noble Hendrix, Eric M. Danner, and Steven T. Lindley. Simulating the migration dynamics of juvenile salmonids through rivers and estuaries using a hydrodynamically driven enhanced particle tracking model. Ecological Modelling, 482(??):??, August 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001242.

Streib:2020:HDH

[SKH^{+}20] Lucas Streib, Mira Kattwinkel, Henriette Heer, Stefan Ruzika, and Ralf B. Schäfer. How does habitat connectivity influence the colonization success of a hemimetabolous aquatic insect? — a modeling approach. Ecological Modelling, 416(??):??, January 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438001930417X.

Scheiter:2024:MDM

[SKPL24] Simon Scheiter, Dushyant Kumar, Mirjam Pfeiffer, and Liam Langan. Modeling drought mortality and resilience of savannas and forests in tropical Asia. Ecological Modelling, 494(??): ??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001716.

Simon:2023:IHL

[SKW23] Alois Simon, Klaus Katzensteiner, and Gudrun Wallentin. The integration of hierarchical levels of scale in tree species distribution models of silver fir (Abies alba Mill.) and European beech (Fagus sylvatica L.) in mountain forests. Ecological Modelling, 485(??):??, November 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002296.

Schickele:2020:MES

[SLB^{+}20a] Alexandre Schickele, Boris Leroy, Gregory Beaugrand, Eric Goberville, Tarek Hattab, Patrice Francour, and Virginie Raybaud. Modelling European small pelagic fish distribution: Methodological insights. Ecological Modelling, 416(??):??, January 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380019304107.

Shen:2020:MII

[SLB20b] Chunqi Shen, Qian Liao, and Harvey A. Bootsma. Modelling the influence of invasive mussels on phosphorus cycling in Lake Michigan. Ecological Modelling, 416(??):??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019304284.

Spring:2023:RDM

[SLB^{+}23] Daniel Spring, Thao P. Le, Samuel Adam Bloom, Jonathan M. Keith, and Tom Kompas. Reconstructing the dynamics of managed populations to estimate the impact of citizen surveillance. Ecological Modelling, 475(??):??, January 2023. CO-

DEN ???? ISSN ???? URL http://www.sciencedirect.
com/science/article/pii/S0304380022003039.

Sun:2020:RPF

[SLC^{+}20] Jingyao Sun, Xinrong Li, Ning Chen, Yanli Wang, and Guang Song. Regular pattern formation regulates population dynamics: Logistic growth in cellular automata. Ecological Modelling, 418(??):??, February 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303862.

Sauterey:2023:PAR

[SLC^{+}23] Boris Sauterey, Guillaume Le Gland, Pedro Cermeño, Olivier Aumont, Marina Lévy, and Sergio M. Vallina. Phytoplankton adaptive resilience to climate change collapses in case of extreme events — a modeling study. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001680.

Sanchez-Lopez:2023:SEM

[SLHB^{+}23] Nuria Sánchez-López, Andrew T. Hudak, Luigi Boschetti, Carlos A. Silva, Kevin Robertson, E. Louise Loudermilk, Benjamin C. Bright, Mac A. Callaham, and Melanie K. Taylor. A spatially explicit model of tree leaf litter accumulation in fire maintained longleaf pine forests of the southeastern US. Ecological Modelling, 481(??):??, July 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000972.

Simons-Legaard:2021:PCI

[SLLW21] Erin Simons-Legaard, Kasey Legaard, and Aaron Weiskittel. Projecting complex interactions between forest harvest and succession in the northern Acadian Forest Region. Ecological Modelling, 456(??):??, September 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002179.

Shoot:2024:DCD

[SLM24] Tanya T. Shoot, Tristan A. F. Long, and Noam Miller. Debugging collective digging in Drosophila melanogaster using an agent-based model. Ecological Modelling, 494(??):??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001698.

Shah:2021:SAA

[SLY^{+}21] Syed Mahboob Shah, Gengyuan Liu, Qing Yang, Marco Casazza, Feni Agostinho, and Biagio F. Giannetti. Sustainability assessment of agriculture production systems in Pakistan: a provincial-scale energy-based evaluation. Ecological Modelling, 455(??):??, September 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002143.

Sangare:2020:IEV

[SLYL^{+}20] Nathanaël Sangare, Alain Lo-Yat, Gilles Le Moullac, Laure Pecquerie, Yoann Thomas, Sébastien Lefebvre, Romain Le Gendre, Benoît Beliaeff, and Serge Andréfouët. Impact of environmental variability on Pinctada margaritifera life-history traits: a full life cycle DEB modeling approach. Ecological Modelling, 423(??):??, May 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300788.

Sells:2020:ETS

[SM20] Sarah N. Sells and Michael S. Mitchell. The economics of territory selection. Ecological Modelling, 438(??):??, December 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303975.

Sharif:2022:CIE

[SM22a] Umi Syahirah Binti Umar Sharif and Mohd Hafiz Mohd. Combined influences of environmental enrichment and harvesting mediate rich dynamics in an intraguild predation fishery system. Ecological Modelling, 474(??):??, December 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022002411.

Sutton:2022:TMM

[SM22b] G. F. Sutton and G. D. Martin. Testing MaxEnt model performance in a novel geographic region using an intentionally introduced insect. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002200240X.

Simioni:2020:NFD

[SMD^{+}20] Guillaume Simioni, Guillaume Marie, Hendrik Davi, Nicolas Martin-St Paul, and Roland Huc. Natural forest dynamics have more influence than climate change on the net ecosystem production of a mixed Mediterranean forest. Ecological Modelling, 416(??):??, January 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019304296.

Sun:2020:MEE

[SML^{+}20] Yuanyuan Sun, Xianqiang Mao, Gengyuan Liu, Xinan Yin, and Yanwei Zhao. Modelling the effects of energy taxes on ecological footprint transfers in China’s foreign trade. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302702.

Shaukat:2022:PME

[SMM^{+}22] Muhammad Shaukat, Sher Muhammad, Ellen D. V. L. Maas, Tasneem Khaliq, and Ashfaq Ahmad. Predicting methane emissions from paddy rice soils under biochar and nitrogen addition using DNDC model. Ecological Modelling, 466(??): ??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000230.

Smith:2020:BCA

[SMP20] Nicola J. Smith, Garry W. McDonald, and Murray G. Patterson. Biogeochemical cycling in the Anthropocene: Quantifying global environment–economy exchanges. Ecological Modelling, 418(??):??, February 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303242.

Szewczyk:2024:MSK

[SMS^{+}24] Tim M. Szewczyk, Pippa J. Moore, Dan A. Smale, Thomas Adams, and Michael T. Burrows. Mechanistic simulations of kelp populations in a dynamic landscape of light, temperature, and winter storms. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023003204.

Stubbs:2020:SGR

[SMV^{+}20] Jessica L. Stubbs, Nina Marn, Mathew A. Vanderklift, Sabrina Fossette, and Nicola J. Mitchell. Simulated growth and reproduction of green turtles (Chelonia mydas) under climate change and marine heatwave scenarios. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302568.

Silva:2023:IPC

[SNM^{+}23] Gabriela Cristina Costa Silva, Júlio César Lima Neves, Gustavo Eduardo Marcatti, Carlos Pedro Boechat Soares, Natalino Calegario, Carlos Alberto Araújo Júnior, Duberlí Geomar Elera Gonzáles, José Marinaldo Gleriani, Daniel Henrique Breda Binoti, Haroldo Nogueira de Paiva, and Helio Garcia Leite. Improving 3-PG calibration and parameterization using artificial neural networks. Ecological Modelling, 479(??):??, May 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000297.

Srinet:2023:SFC

[SNP^{+}23] Ritika Srinet, Subrata Nandy, N. R. Patel, Hitendra Padalia, Taibanganba Watham, Sanjeev K. Singh, and Prakash Chauhan. Simulation of forest carbon fluxes by integrating remote sensing data into biome-BGC model. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022002836.

Schippers:2022:FAF

[SOP22] Peter Schippers, Wim A. Ozinga, and Rogier Pouwels. Factors affecting functional diversity of grassland vegetations. Ecological Modelling, 472(??):??, October 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001831.

Shen:2021:SEM

[SOV21] Yunyi Shen, Erik R. Olson, and Timothy R. Van Deelen. Spatially explicit modeling of community occupancy using Markov random field models with imperfect observation: Mesocarnivores in Apostle Islands National Lakeshore. Ecological Modelling, 459(??):??, November 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002660.

Scott:2024:MTF

[SPB^{+}24] Rebecca L. Scott, Nathan F. Putman, R. Taylor Beyea, Hallie C. Repeta, and Cameron H. Ainsworth. Modeling transport and feeding of juvenile Kemp’s ridley sea turtles on the West Florida shelf. Ecological Modelling, 490(??):??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024000474.

Sinha:2021:MSI

[SPPC21] Sanjiv K. Sinha, Hitendra Padalia, N. R. Patel, and Prakash Chauhan. Modelling sun-induced fluorescence for improved evaluation of forest carbon flux (GPP): Case study of tropical deciduous forest, India. Ecological Modelling, 449(??):??, June 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021001228.

Stock:2020:ICP

[SPV^{+}20] Michiel Stock, Niels Piot, Sarah Vanbesien, Bernard Vaissière, Clémentine Coiffait-Gombault, Guy Smagghe, and Bernard De Baets. Information content in pollination network reveals missing interactions. Ecological Modelling, 431(??): ??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380020302325.

Stock:2021:PLP

[SPV^{+}21] Michiel Stock, Niels Piot, Sarah Vanbesien, Joris Meys, Guy Smagghe, and Bernard De Baets. Pairwise learning for predicting pollination interactions based on traits and phylogeny. Ecological Modelling, 451(??):??, July 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002100079X.

Shang:2021:BAG

[SQDM21] Zuchong Shang, Yuanhua Qiao, Lijuan Duan, and Jun Miao. Bifurcation analysis and global dynamics in a predator–prey system of Leslie type with an increasing functional response. Ecological Modelling, 455(??):??, September 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002192.

Salgado-Rojas:2020:MIP

[SRÁMH^{+}20] José Salgado-Rojas, Eduardo Álvarez-Miranda, Virgilio Hermoso, Jordi Garcia-Gonzalo, and Andrés Weintraub. A mixed integer programming approach for multi-action planning for threat management. Ecological Modelling, 418(??):??, February 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019304090.

Sun:2020:DGM

[SRB^{+}20] Ke Sun, Jeffrey S. Ren, Tao Bai, Jihong Zhang, Qing Liu, Wenguang Wu, Yunxia Zhao, and Yi Liu. A dynamic growth model of Ulva prolifera: Application in quantifying the biomass of green tides in the Yellow Sea, China. Ecological Modelling, 428(??):??, July 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301447.

Skov:2020:FWR

[SRK^{+}20] Henrik Skov, Erik Kock Rasmussen, Jonne Kotta, Anne Lise Middelboe, Thomas Uhrenholdt, and Ramunas Žydelis. Food web responses to eutrophication control in a coastal area of the Baltic Sea. Ecological Modelling, 435(??):??, November 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020303197.

Schmidt:2020:RSU

[SRTS20] Heiko Schmidt, Johannes Radinger, Daniel Teschlade, and Stefan Stoll. The role of spatial units in modelling freshwater fish distributions: Comparing a subcatchment and river network approach using MaxEnt. Ecological Modelling, 418(??):??, February 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300089.

Srisunont:2022:DMS

[SSB22] Chayarat Srisunont, Treeranut Srisunont, and Sandhya Babel. Development of models for sustainable green mussel cultivation under climate change events. Ecological Modelling, 473(??):??, November 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002423.

Small:2023:SFC

[SSJC23] Gaston E. Small, Marisa Smedsrud, Ivan Jimenez, and Eric Chapman. Simulating the fate of compost-derived nutrients in an urban garden. Ecological Modelling, 483(??):??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023001722.

Sasaki:2024:NMD

[SSM24] Fumiya Sasaki, Takuya Shiba, and Keiichiro Matsukura. Novel method of determining parameters for the effective accumulated temperature model by using seasonal pest occurrence data. Ecological Modelling, 490(??):??, April 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000401.

Soares:2024:USM

[ST24] Ana Jacinta Soares and Giorgio Tosato. An urban system mathematical approach with human factor: the case of pedestrianization in a consolidated area. Ecological Modelling, 489(??):??, March 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000085.

Steingruber:2020:IEM

[Ste20] Sandra Martina Steingruber. Improved empirical models for predicting nitrogen retention in lakes and reservoirs. Ecological Modelling, 416(??):??, January 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303618.

Sandhu:2022:SAB

[STQ^{+}22] Rimple Sandhu, Charles Tripp, Eliot Quon, Regis Thedin, Michael Lawson, David Brandes, Christopher J. Farmer, Tricia A. Miller, Caroline Draxl, Paula Doubrawa, Lindy Williams, Adam E. Duerr, Melissa A. Braham, and Todd Katzner. Stochastic agent-based model for predicting turbinescale raptor movements during updraft-subsidized directional flights. Ecological Modelling, 466(??):??, April 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000047.

Srivastava:2024:DOD

[STRC24] Vivek Srivastava, Amberly R. Tai, Jeanne A. Robert, and Allan L. Carroll. A dynamic outbreak distribution model (DODM) for an irruptive folivore: the western spruce budworm. Ecological Modelling, 492(??):??, June 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002400125X.

Sinnott:2023:ESS

[STW^{+}23] Emily A. Sinnott, Frank R. Thompson, Mitch D. Weegman, Thomas R. Thompson, Alisha R. Mosloff, R. Kyle Hedges, and Frank L. Loncarich. Evaluation of seasonal site-level demography and management for northern bobwhite using integrated population models. Ecological Modelling, 475(??):??, January 2023. CODEN ???? ISSN

????  URL http://www.sciencedirect.com/science/
article/pii/S030438002200309X.

Schouten:2020:IDP

[SVK20] Rafael Schouten, Peter A. Vesk, and Michael R. Kearney. Integrating dynamic plant growth models and microclimates for species distribution modelling. Ecological Modelling, 435(??):??, November 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002030332X.

Searle:2021:AMM

[SvV21] K. D. Searle and J. H. van Vuuren. An abstract mathematical model for sustainable harvesting of a biological species on the boundary of a protected habitat. Ecological Modelling, 452(??):??, July 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438002100154X.

Silva:2021:HMA

[SWF21] Tammy L. Silva, David N. Wiley, and Gavin Fay. A hierarchical modelling approach to estimating humpback whale abundance from sand lance abundance. Ecological Modelling, 456(??):??, September 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021002210.

Shi:2022:YEL

[SWHZ22] Yinfang Shi, Zhaoyang Wang, Cheng Hou, and Puhan Zhang. Yield estimation of Lycium barbarum L. based on the WOFOST model. Ecological Modelling, 473(??):??, November 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022002472.

Slade:2023:IEP

[SWL^{+}23] Andrew Slade, Andy White, Peter W. W. Lurz, Craig Shuttleworth, David G. Tosh, and Joshua P. Twining. Indirect effects of pine marten recovery result in benefits to native prey through suppression of an invasive species and a shared pathogen. Ecological Modelling, 476(??): ??, February 2023. CODEN ECMODT. ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0304380022003143.

Song:2023:MLU

[SWP^{+}23] Youcheng Song, Haijun Wang, Xiaotao Peng, Duan Sun, and Rui Chen. Modeling land use change prediction using multimodel fusion techniques: a case study in the Pearl River Delta, China. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002752.

Song:2023:IAM

[SWW^{+}23] Lili Song, Yingying Wu, Moyu Wu, Jie Ma, and Wei Cao. An integrated approach to model connectivity and identify modules for habitat networks. Ecological Modelling, 483(??):??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001771.

Sikazwe:2024:MWD

[SYL^{+}24] Geofrey Sikazwe, Rosita E. E. Yocgo, Pietro Landi, David M. Richardson, and Cang Hui. Managing whitefly development to control cassava brown streak virus coinfections. Ecological Modelling, 493(??):??, July 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024001418.

Song:2020:STP

[SZ20] Chenyu Song and Haiping Zhang. Study on turbidity prediction method of reservoirs based on long short term memory neural network. Ecological Modelling, 432(??):??, September 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302805.

Sun:2020:EII

[SZL^{+}20] Ke Sun, Jihong Zhang, Fan Lin, Jeffrey S. Ren, Yunxia Zhao, Wenguang Wu, and Yi Liu. Evaluating the influences of integrated culture on pelagic ecosystem by a numerical approach: a case study of Sungo Bay, China.

Ecological Modelling, 415(??):??, January 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303680.

Su:2021:ESS

[SZL^{+}21] Xuemeng Su, Zhengchao Zhou, Jun’e Liu, Liguo Cao, Junyang Liu, and Peipei Wang. Estimating slope stability by the root reinforcement mechanism of Artemisia sacrorum on the Loess Plateau of China. Ecological Modelling, 444(??):??, March 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021000454.

Shen:2023:ILE

[SZPW23] Jiake Shen, Wenjia Zhu, Zhenwei Peng, and Yuncai Wang. Improving landscape ecological network connectivity in urbanizing areas from dual dimensions of structure and function. Ecological Modelling, 482(??):??, August 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001114.

Toorians:2020:CCT

[TA20] Marjolein E. M. Toorians and Thomas P. Adams. Critical connectivity thresholds and the role of temperature in parasite metapopulations. Ecological Modelling, 435(??):??, November 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303288.

Taghikhah:2022:DDA

[TBCV22] Firouzeh Taghikhah, Justin Borevitz, Robert Costanza, and Alexey Voinov. DAESim: a dynamic agro-ecosystem simulation model for natural capital assessment. Ecological Modelling, 468(??):??, June 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000539.

Thompson:2024:IBM

[TBM^{+}24] Noelle E. Thompson, David J. Butts, Michael S. Murillo, Daniel J. O’Brien, Sonja A. Christensen, William F. Porter, and Gary J. Roloff. An individual-based model for direct and indirect transmission of chronic wasting disease in freeranging white-tailed deer. Ecological Modelling, 491(??):??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000851.

Takahashi:2023:CRM

[TBP^{+}23] Amane Takahashi, Syuhei Ban, Rey Donne S. Papa, Dino T. Tordesillas, and Gaël Dur. Cumulative reproduction model to quantify the production of the invasive species Arctodiaptomus dorsalis (Calanoida, Copepoda). Ecological Modelling, 482(??):??, August 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001291.

Tang:2022:RPW

[TBZ^{+}22] Jianzhao Tang, Huizi Bai, Xinjun Zhang, Rende Wang, Fenghua Guo, Dengpan Xiao, and Haitao Zhou. Reducing potato water footprint by adjusting planting date in the agro-pastoral ecotone in North China. Ecological Modelling, 474(??):??, December 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002563.

Thurner:2021:MOE

[TCB21] Stephanie D. Thurner, Sarah J. Converse, and Trevor A. Branch. Modeling opportunistic exploitation: increased extinction risk when targeting more than one species. Ecological Modelling, 454(??):??, August 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021001733.

Tellez:2020:CST

[TCG^{+}20] M. M. Téllez, T. Cabello, M. Gámez, F. J. Burguillo, and E. Rodríguez. Comparative study of two predatory mites Amblyseius swirskii Athias-Henriot and Transeius montdorensis (Schichaa) by predator–prey models for improving biological control of greenhouse cucumber. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302672.

Triana:2021:AIA

[TCS21] Juan S. Acero Triana, Maria L. Chu, and Jeffrey A. Stein. Assessing the impacts of agricultural conservation practices on freshwater biodiversity under changing climate. Ecological Modelling, 453(??):??, August 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001666.

Toda:2020:BIF

[TDI^{+}20] Motomu Toda, Kazuki Doi, Masae I. Ishihara, Wakana A. Azuma, and Masayuki Yokozawa. A Bayesian inversion framework to evaluate parameter and predictive inference of a simple soil respiration model in a cool-temperate forest in western Japan. Ecological Modelling, 418(??):??, February 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019304260.

Tadiello:2023:NMS

[TGB^{+}23] Tommaso Tadiello, Mara Gabbrielli, Marco Botta, Marco Acutis, Luca Bechini, Giorgio Ragaglini, Andrea Fiorini, Vincenzo Tabaglio, and Alessia Perego. A new module to simulate surface crop residue decomposition: Description and sensitivity analysis. Ecological Modelling, 480(??):??, June 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023000558.

Turkeltaub:2022:RMS

[TGL^{+}22] Tuvia Turkeltaub, Kate Gongadze, Yihe Lü, Mingbin Huang, Xiaoxu Jia, Huiyi Yang, Ming’an Shao, Andrew Binley, Paul Harris, and Lianhai Wu. A review of models for simulating the soil-plant interface for different climatic conditions and land uses in the Loess Plateau, China. Ecological Modelling, 474(??):??, December 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002745.

Tikuye:2023:MIC

[TGRM23] Birhan Getachew Tikuye, Laurence Gill, Milos Rusnak, and Busnur R. Manjunatha. Modelling the impacts of changing land use and climate on sediment and nutrient retention in Lake Tana Basin, Upper Blue Nile River Basin, Ethiopia. Ecological Modelling, 482(??):??, August 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002300114X.

Takada:2020:AEV

[TK20a] Takenori Takada and Yuka Kawai. An analysis of elasticity vector distribution specific to semelparous species using randomly generated population projection matrices and the COMPADRE plant matrix database. Ecological Modelling, 431(??):??, September 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301976.

Turner:2020:SSD

[TK20b] Benjamin L. Turner and Srinadh Kodali. Soil system dynamics for learning about complex, feedback-driven agricultural resource problems: model development, evaluation, and sensitivity analysis of biophysical feedbacks. Ecological Modelling, 428(??):??, July 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301228.

Tsujimoto:2022:QGR

[TKO^{+}22] K. Tsujimoto, N. Kuriya, T. Ohta, K. Homma, and M. So Im. Quantifying the GCM-related uncertainty for climate change impact assessment of rainfed rice production in Cambodia by a combined hydrologic–rice growth model. Ecological Modelling, 464(??):??, February 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003604.

Tan:2022:MFR

[TLC^{+}22] Kangming Tan, Yuliang Li, Yun Chen, Fangdan Liu, Jingmin Ou, Yuhan Zhang, and Xiaolong Wang. Modified framework to reflect contribution of soil storage in emergy synthesis under different agricultural practices at farm level. Ecological Modelling, 465(??):??, March 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003938.

Tian:2020:PLE

[TLH^{+}20] Fenghao Tian, Mingyu Li, Xulong Han, Hui Liu, and Boxian Mo. A production–living–ecological space model for landuse optimisation: a case study of the core Tumen River region in China. Ecological Modelling, 437(??):??, December 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002030380X.

Tardy:2023:RTA

[TLL^{+}23] Olivia Tardy, Christophe Lenglos, Sandra Lai, Dominique Berteaux, and Patrick A. Leighton. Rabies transmission in the Arctic: an agent-based model reveals the effects of broad-scale movement strategies on contact risk between Arctic foxes. Ecological Modelling, 476(??):??, February 2023. CODEN EC-

MODT. ISSN ???? URL http://www.sciencedirect.com/
science/article/pii/S0304380022003052.

Tan:2022:ABG

[TMFV22] Tjui Yeuw Tan, Marcel Câmara Miraldo, Roberto Fioravanti Carelli Fontes, and Fabio Stucchi Vannucchi. Assessing bivalve growth using bio-energetic models. Ecological Modelling, 473(??):??, November 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001764.

Tomkins:2020:PAC

[TMNA20] Melissa Tomkins, Adrian P. Martin, A. J. George Nurser, and Thomas R. Anderson. Phytoplankton acclimation to changing light intensity in a turbulent mixed layer: a Lagrangian modelling study. Ecological Modelling, 417(??):??, February 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019304259.

Theng:2024:IRM

[TPD^{+}24] Meryl Theng, Thomas A. A. Prowse, Steven Delean, Phillip Cassey, and Chloe Bracis. Integrating resource memory and cue-based territoriality to simulate movement dynamics: a process-explicit and pattern-oriented approach. Ecological Modelling, 487(??):??, January 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002909.

Tang:2024:QGG

[TQM^{+}24] Yujie Tang, Yunfa Qiao, Yinzheng Ma, Weiliang Huang, Khan Komal, and Shujie Miao. Quantifying greenhouse gas emissions in agricultural systems: a comparative analysis of process models. Ecological Modelling, 490(??):??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024000358.

Thomas:2020:MSI

[TRMB20] Yoann Thomas, Ntsoa Rakoto Razafimahefa, Alain Ménesguen, and Cédric Bacher. Multi-scale interaction processes modulate the population response of a benthic species to global warming. Ecological Modelling, 436(??):??, November 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020303653.

Tunca:2023:DMS

[TSBE23] Mehmet Can Tunca, Ali Kerem Saysel, Masoud Babaei, and Günay Erpul. A dynamic model for salinity and sodicity management on agricultural lands: Interactive simulation approach. Ecological Modelling, 482(??):??, August 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002300131X.

Tourinho:2021:LDD

[TSdOCV21] Luara Tourinho, Barry Sinervo, Gabriel Henrique de Oliveira Caetano, and Mariana M. Vale. A less data demanding ecophysiological niche modeling approach for mammals with comparison to conventional correlative niche modeling. Ecological Modelling, 457(??):??, October 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021002453.

Theng:2023:PPT

[TSUY23] Vouchlay Theng, Ratino Sith, Sovannara Uk, and Chihiro Yoshimura. Phytoplankton productivity in a tropical lakefloodplain system revealed by a process-based primary production model. Ecological Modelling, 479(??):??, May 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000455.

Tsydenov:2022:MBP

[Tsy22] Bair O. Tsydenov. Modeling biogeochemical processes in a freshwater lake during the spring thermal bar. Ecological Modelling, 465(??):??, March 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000059.

Tremblay:2021:PPH

[TTPA21] Raymond L. Tremblay, Andrew J. Tyre, Maria-Eglée Pérez, and James D. Ackerman. Population projections from holey matrices: Using prior information to estimate rare transition events. Ecological Modelling, 447(??):??, May 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000971.

Turner:2020:MLQ

[Tur20] Benjamin L. Turner. Model laboratories: a quick-start guide for design of simulation experiments for dynamic systems models. Ecological Modelling, 434(??):??, October 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303161.

Tang:2022:ECC

[TWL^{+}22] Yuzhi Tang, Mengdi Wang, Qian Liu, Zhongwen Hu, Jie Zhang, Tiezhu Shi, Guofeng Wu, and Fenzhen Su. Ecological carrying capacity and sustainability assessment for coastal zones: a novel framework based on spatial scene and three-dimensional ecological footprint model. Ecological Modelling, 466(??):??, April 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022000096.

Uchmanski:2023:CDH

[Uch23] Janusz Uchmański. Catastrophic dynamics: How to avoid population extinction? Ecological Modelling, 483(??):??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001692.

Ulrich:2024:RPE

[UGSG24] Werner Ulrich, Nicholas J. Gotelli, Giovanni Strona, and William Godsoe. Reconsidering the Price equation: Benchmarking the analytical power of additive partitioning in ecology. Ecological Modelling, 491(??):??, May 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000838.

Ustaoglu:2021:DAS

[USA21] E. Ustaoglu, S. Sisman, and A. C. Aydınoglu. Determining agricultural suitable land in peri-urban geography using GIS and Multi Criteria Decision Analysis (MCDA) techniques. Ecological Modelling, 455(??):??, September 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021001721.

Vasechkina:2020:NRB

[Vas20] E. F. Vasechkina. Nonlinear relationships between phytoplankton nutrient utilization traits and environmental factors. Ecological Modelling, 433(??):??, October 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303033.

Vasechkina:2023:MEI

[Vas23] E. F. Vasechkina. Modeling the environmental impact of mussels’ cultivation in the coastal zone of Crimea. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S030438002200343X.

Vortkamp:2020:EAE

[VBH20] Irina Vortkamp, Frédéric Barraquand, and Frank M. Hilker. Ecological Allee effects modulate optimal strategies for conservation in agricultural landscapes. Ecological Modelling, 435(??):??, November 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020302787.

vanDaalen:2020:VLH

[vDC20] Silke van Daalen and Hal Caswell. Variance as a life history outcome: Sensitivity analysis of the contributions of stochasticity and heterogeneity. Ecological Modelling, 417 (??):??, February 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019303643.

vanderHeijden:2020:QFW

[vdHNH^{+}20] L. H. van der Heijden, N. Niquil, M. Haraldsson, R. M. Asmus, S. R. Pacella, M. Graeve, J. Rzeznik-Orignac, H. Asmus, B. Saint-Béat, and B. Lebreton. Quantitative food web modeling unravels the importance of the microphytobenthos-meiofauna pathway for a high trophic transfer by meiofauna in soft-bottom intertidal food webs. Ecological Modelling, 430(??):??, August 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302015.

vanderMeer:2020:PPN

[vdMvDSL20] Jaap van der Meer, Susanne van Donk, Alejandro Sotillo, and Luc Lens. Predicting post-natal energy intake of lesser black-backed gull chicks by dynamic energy budget modeling. Ecological Modelling, 423(??):??, May 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300776.

Velander:2023:DTM

[VJK^{+}23] Taylor B. Velander, Michael J. Joyce, Angela M. Kujawa, Robert L. Sanders, Paul W. Keenlance, and Ron A. Moen. A dynamic thermal model for predicting internal temperature of tree cavities and nest boxes. Ecological Modelling, 478(??): ??, April 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000303.

Vegh:2024:MSD

[VK24] Lea Végh and Tomomichi Kato. Modified SEIB-DGVM enables simulation of masting in a temperate forest. Ecological Modelling, 488(??):??, February 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023003071.

Venolia:2020:MGS

[VLGG^{+}20] Celeste T. Venolia, Romain Lavaud, Lindsay A. Green-Gavrielidis, Carol Thornber, and Austin T. Humphries. Modeling the growth of sugar kelp (Saccharina latissima) in aquaculture systems using dynamic energy budget theory. Ecological Modelling, 430(??):??, August 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302222.

VanEupen:2021:IDQ

[VMH^{+}21] Camille Van Eupen, Dirk Maes, Marc Herremans, Kristijn R. R. Swinnen, Ben Somers, and Stijn Luca. The impact of data quality filtering of opportunistic citizen science data on species distribution model performance. Ecological Modelling, 444(??):??, March 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000260.

Virgillito:2021:MAA

[VMM^{+}21] Chiara Virgillito, Mattia Manica, Giovanni Marini, Beniamino Caputo, Alessandra della Torre, and Roberto Rosà. Modelling arthropod active dispersal using partial differential equations: the case of the mosquito Aedes albopictus.

Ecological Modelling, 456(??):??, September 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002180.

Vasechkina:2023:VPP

[VN23] E. F. Vasechkina and I. P. Naumenko. Variability of photosynthetic parameters of macroalgae and seagrasses based on laboratory experiments. Ecological Modelling, 486(??):??, December 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002429.

vanOort:2023:CSF

[vOVvdW23] P. A. J. van Oort, A. Verhagen, and A. K. van der Werf. Can seaweeds feed the world? Modelling world offshore seaweed production potential. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002168.

Valero:2022:EAT

[VPAV22] Antonio Valero, Bárbara Palacino, Sonia Ascaso, and Alicia Valero. Exergy assessment of topsoil fertility. Ecological Modelling, 464(??):??, February 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003471.

Veiskarami:2021:HSS

[VPV21] Zahra Veiskarami, Babak Pilehvar, and Gholam Hassan Veiskarami. How sampling scheme and species-area scale contribute to select the best statistical functions for upscaling plant species richness. Ecological Modelling, 457(??):??, October 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021002507.

Valle:2024:PAD

[VRRI^{+}24] Mireia Valle, Eduardo Ramírez-Romero, Leire Ibaibarriaga, Leire Citores, Jose A. Fernandes-Salvador, and Guillem Chust. Pan-Atlantic 3D distribution model incorporating water column for commercial fish. Ecological Modelling, 490(??):

??, April 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024000218.

Vermeiren:2020:IUP

[VRS20] Peter Vermeiren, Peter Reichert, and Nele Schuwirth. Integrating uncertain prior knowledge regarding ecological preferences into multi-species distribution models: Effects of model complexity on predictive performance. Ecological Modelling, 420(??):??, March 15, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020300284.

VanBuskirk:2021:MHA

[VRW^{+}21] Amanda N. Van Buskirk, Christopher S. Rosenberry, Bret D. Wallingford, Emily Just Domoto, Marc E. McDill, Patrick J. Drohan, and Duane R. Diefenbach. Modeling how to achieve localized areas of reduced white-tailed deer density. Ecological Modelling, 442(??):??, February 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304579.

Voinson:2022:HDH

[VSB22] Marina Voinson, Charline Smadi, and Sylvain Billiard. How does the host community structure affect the epidemiological dynamics of emerging infectious diseases? Ecological Modelling, 472(??):??, October 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001958.

VivesiBatlle:2020:MEI

[VSK^{+}20] J. Vives i Batlle, T. Sazykina, A. Kryshev, M. D. Wood, K. Smith, D. Copplestone, and G. Biermans. Modelling the effects of ionising radiation on a vole population from the Chernobyl Red Forest in an ecological context. Ecological Modelling, 438(??):??, December 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303768.

Viswanathan:2022:BML

[VSS^{+}22] Michelle Viswanathan, Andreas Scheidegger, Thilo Streck, Sebastian Gayler, and Tobias K. D. Weber. Bayesian multi-level calibration of a process-based maize phenology model. Ecological Modelling, 474(??):??, December 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002551.

Vaugeois:2020:PCM

[VVH^{+}20] Maxime Vaugeois, Paul A. Venturelli, Stephanie L. Hummel, Chiara Accolla, and Valery E. Forbes. Population context matters: Predicting the effects of metabolic stress mediated by food availability and predation with an agent- and energy budget-based model. Ecological Modelling, 416(??):??, January 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019304119.

vanWeerden:2020:MNA

[vWVH20] J. Fransje van Weerden, Rineke Verbrugge, and Charlotte K. Hemelrijk. Modelling non-attentional visual information transmission in groups under predation. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301459.

Whitehouse:2020:AST

[WA20] George A. Whitehouse and Kerim Y. Aydin. Assessing the sensitivity of three Alaska marine food webs to perturbations: an example of Ecosim simulations using Rpath. Ecological Modelling, 429(??):??, August 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301460.

Wang:2024:ASI

[WAHA^{+}24] Wei Wang, Arwa AL-Huqail, Elimam Ali, Mohamed Abbas, and Hamid Assilzadeh. Analysis of the sustainability index for ecologically low-input integrated farming: a comprehensive assessment of environmental, economic, and social impact. Ecological Modelling, 493(??):??, July 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000899.

Watson:2022:IEV

[WBD^{+}22a] Joseph W. Watson, Robin Boyd, Ritabrata Dutta, Georgios Vasdekis, Nicola D. Walker, Shovonlal Roy, Richard Everitt, Kieran Hyder, and Richard M. Sibly. Incorporating environmental variability in a spatially-explicit individual-based model of European sea bass. Ecological Modelling, 466(??): ??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000060.

Wongnak:2022:HBA

[WBD^{+}22b] Phrutsamon Wongnak, Séverine Bord, Sophie Donnet, Thierry Hoch, Frederic Beugnet, and Karine Chalvet-Monfray. A hierarchical Bayesian approach for incorporating expert opinions into parametric survival models: a case study of female Ixodes ricinus ticks exposed to various temperature and relative humidity conditions. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021003653.

Walker:2020:SEI

[WBW^{+}20] Nicola D. Walker, Robin Boyd, Joseph Watson, Max Kotz, Zachary Radford, Lisa Readdy, Richard Sibly, Shovonlal Roy, and Kieran Hyder. A spatially explicit individualbased model to support management of commercial and recreational fisheries for European sea bass Dicentrarchus labrax. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302507.

Wu:2022:TAE

[WCF^{+}22] Zijian Wu, Mingli Chen, Xinxi Fu, Linnan Ouyang, and Xiaofu Wu. Thermodynamic analysis of an ecologically restored plant community: Ecological niche. Ecological Modelling, 464(??):??, February 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003811.

Wu:2024:SCN

[WCG^{+}24] Dongshao Wu, Min Cao, Wei Gao, Zhongzhao Duan, and Yuan Zhang. Simulating critical nutrient loadings of regime shift in the shallow plateau Lake Dianchi. Ecological Modelling, 491(??):??, May 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000772.

Wenxiang:2022:ODL

[WCSX22] Ding Wenxiang, Zhang Caiyun, Shang Shaoping, and Li Xueding. Optimization of deep learning model for coastal chlorophyll a dynamic forecast. Ecological Modelling, 467(??):??, May 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000370.

Wohner:2022:IMO

[WDW^{+}22] Patti J. Wohner, Adam Duarte, John Wikert, Brad Cavallo, Steven C. Zeug, and James T. Peterson. Integrating monitoring and optimization modeling to inform flow decisions for Chinook salmon smolts. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001673.

Wang:2024:ACE

[WF24] Geng Wang and Yan Feng. Analysis of carbon emission drivers and peak carbon forecasts for island economies. Ecological Modelling, 489(??):??, March 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023003411.

Wszola:2023:IHB

[WFD23] Lyndsie S. Wszola, Zachary S. Feiner, and John P. DeLong. Integrating human behavior into structured population models. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002685.

Wei:2024:VES

[WFW^{+}24] Ruhao Wei, Yanmin Fan, Hongqi Wu, Kai Zheng, Jie Fan, Zhuo Liu, Junwei Xuan, and Jien Zhou. The value of ecosystem services in arid and semi-arid regions: a multiscenario analysis of land use simulation in the Kashgar region of Xinjiang. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023003095.

Wang:2021:RTS

[WG21] Hsiao-Hsuan Wang and William E. Grant. Reflections of two systems ecologists on modelling coupled human and natural (socio-ecological, socio-environmental) systems. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002030466X.

Wang:2022:MCE

[WGAH22] Hao Wang, Pablo Venegas Garcia, Shohel Ahmed, and Christopher M. Heggerud. Mathematical comparison and empirical review of the Monod and Droop forms for resourcebased population dynamics. Ecological Modelling, 466(??):??, April 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002200014X.

Wang:2021:SMW

[WGK^{+}21] Hsiao-Hsuan Wang, William E. Grant, Tomasz E. Koralewski, Michael J. Brewer, and Norman C. Elliott. Simulating migration of wind-borne pests: “Deconstructing” representation of the emigration process. Ecological Modelling, 460(??):??, November 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021002921.

Wies:2023:HBT

[WGMR23] Germán Wies, Jeroen C. J. Groot, and Miguel Martinez-Ramos. In highly-biodiverse tropical landscapes, multipleobjective optimization reveals opportunities for increasing both conservation and agricultural production. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001667.

Wadoux:2021:SCV

[WHdBB21] Alexandre M. J.-C. Wadoux, Gerard B. M. Heuvelink, Sytze de Bruin, and Dick J. Brus. Spatial cross-validation is not the right way to evaluate map accuracy. Ecological Modelling, 457(??):??, October 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021002489.

Wadkin:2024:CER

[WHE^{+}24a] Laura E. Wadkin, John Holden, Rammile Ettelaie, Melvin J. Holmes, James Smith, Andrew Golightly, Nick G. Parker, and Andrew W. Baggaley. Corrigendum to “Estimating the reproduction number, R_{0}, from individual-based models of tree disease spread” [Ecological Modelling 489 (2024) 110630]. Ecological Modelling, 490(??):??, April 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000231. See [WHE^{+}24b].

Wadkin:2024:ERN

[WHE^{+}24b] Laura E. Wadkin, John Holden, Rammile Ettelaie, Melvin J. Holmes, James Smith, Andrew Golightly, Nick G. Parker, and Andrew W. Baggaley. Estimating the reproduction number, R_{0}, from individual-based models of tree disease spread. Ecological Modelling, 489(??):??, March 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S030438002400019X. See corrigendum [WHE^{+}24a].

Wang:2023:CST

[WHL^{+}23] Qinying Wang, Hong S. He, Kai Liu, Shengwei Zong, and Haibo Du. Comparing simulated tree biomass from daily, monthly, and seasonal climate input of terrestrial ecosystem model. Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001515.

Qinyi:2023:EIO

[WHW^{+}23] Qinyi Wang, Mei Huang, Shaoqiang Wang, Bin Chen, Zhenhai Liu, Zhaosheng Wang, Shiliang Chen, Hui Li, Tongtong Zhu, Donghui Li, Yuelin Li, Hu Lin, and Leigang Sun. Evaluation of the impacts of ozone on the vegetation productivity of woodland and grassland ecosystems in China. Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001576.

Wieland:2021:CCR

[WKL^{+}21] Ralf Wieland, Katrin Kuhls, Hartmut H. K. Lentz, Franz Conraths, Helge Kampen, and Doreen Werner. Combined climate and regional mosquito habitat model based on machine learning. Ecological Modelling, 452(??):??, July 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021001563.

Wang:2020:BPC

[WLLW20] Shuran Cindy Wang, Xueqin Liu, Yong Liu, and Hongzhu Wang. Benthic-pelagic coupling in lake energetic food webs. Ecological Modelling, 417(??):??, February 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019304363.

Wang:2023:MFE

[WLQ^{+}23] Chenchen Wang, Jing Li, Chunsheng Qiu, Lei Wang, Xiao Su, Panpan Huang, Na He, Liping Sun, Yaohui Bai, Chaocan Li, and Yufei Wang. Multimedia fates and ecological risk control strategies of phthalic acid esters in a lake recharged by reclaimed water using the QWASI fugacity model. Ecological Modelling, 475(??):??, January 2023. CODEN ????

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003209.

Wu:2023:SFC

[WLT^{+}23] Mia M. Wu, Yu Liang, Franziska Taubert, Andreas Huth, Min Zhang, and Xugao Wang. Sensitivity of forest composition and productivity to climate change in mixed broadleaved-Korean pine forest of Northeastern China. Ecological Modelling, 483(??):??, September 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023001655.

Wang:2023:PCS

[WLWL23] Lin Wang, Jia Li, Ting Wang, and Jinbao Liao. A positive complexity-stability relationship emerges in pollinatorplant-consumer tripartite networks disturbed by plant invasion. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001928.

Wang:2023:MCM

[WLY^{+}23] Xin Wang, Shuo Liu, Yifan Yu, Shengzhi Yue, Ying Liu, Fumin Zhang, and Yuanshan Lin. Modeling collective motion for fish schooling via multi-agent reinforcement learning. Ecological Modelling, 477(??):??, March 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002200357X.

Wang:2023:ECD

[WLZ^{+}23a] Qianqian Wang, Pengcheng Li, Wenming Zhang, Nan Cong, Yuqian Xi, Lirong Xiao, Yihang Wang, and Weiwei Yao. Evaluating the cascade dam construction effects on endemic fish habitat and population status in spawning sites of Lancang River (in Tibet), China. Ecological Modelling, 483(??):??, September 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001813.

Wu:2023:ICG

[WLZ^{+}23b] Genan Wu, Xinchen Lu, Wei Zhao, Ruochen Cao, Wenqi Xie, Liyun Wang, Qiuhong Wang, Jiexuan Song, Shaobo Gao, Shenggong Li, and Zhongmin Hu. The increasing contribution of greening to the terrestrial evapotranspiration in China. Ecological Modelling, 477(??):??, March 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000017.

Warth:2021:ISP

[WMA21] Benjamin Warth, Carsten Marohn, and Folkard Asch. Improved simulation of plant-animal interactions in African savannas with the extended land use change model LU-CIA. Ecological Modelling, 446(??):??, April 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000673.

Wang:2023:MMP

[WPL^{+}23] Yanping Wang, Zhaoliang Peng, Gang Liu, Hui Zhang, Xiangqian Zhou, and Weiping Hu. A mathematical model for phosphorus interactions and transport at the sedimentwater interface in a large shallow lake. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003520.

Walker:2021:IDD

[WRC21] Melody Walker, Michael A. Robert, and Lauren M. Childs. The importance of density dependence in juvenile mosquito development and survival: a model-based investigation. Ecological Modelling, 440(??):??, January 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304221.

Wongchai:2022:FMD

[WrJP^{+}22] Anupong Wongchai, Durga rao Jenjeti, A. Indira Priyadarsini, Nabamita Deb, Arpit Bhardwaj, and Pradeep Tomar. Farm monitoring and disease prediction by classification based on deep learning architectures in sustainable agriculture.

Ecological Modelling, 474(??):??, December 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002200268X.

Woodstock:2021:EWS

[WSFZ21] Matthew S. Woodstock, Tracey T. Sutton, Tamara Frank, and Yuying Zhang. An early warning sign: trophic structure changes in the oceanic Gulf of Mexico from 2011– 2018. Ecological Modelling, 445(??):??, April 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000806.

Wang:2024:SAS

[WSHZ24] Zhaoyang Wang, Yinfang Shi, Cheng Hou, and Puhan Zhang. Sensitivity analysis of simulated Lycium barbarum L. yield in the WOFOST model under different climate conditions. Ecological Modelling, 488(??):??, February 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003320.

Williamson:2022:CDH

[WSL^{+}22] Laura D. Williamson, Beth E. Scott, Megan Laxton, Janine B. Illian, Victoria L. G. Todd, Peter I. Miller, and Kate L. Brookes. Comparing distribution of harbour porpoise using generalized additive models and hierarchical Bayesian models with integrated nested Laplace approximation. Ecological Modelling, 470(??):??, August 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001223.

Whitford:2024:IND

[WSM24] Anna M. Whitford, Benjamin R. Shipley, and Jenny L. McGuire. The influence of the number and distribution of background points in presence-background species distribution models. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023003344.

Wood:2021:PAH

[WSN^{+}21] Kevin A. Wood, Richard A. Stillman, Julia L. Newth, Rascha J. M. Nuijten, Geoff M. Hilton, Bart A. Nolet, and Eileen C. Rees. Predicting avian herbivore responses to changing food availability and competition. Ecological Modelling, 441(??):??, February 1, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020304798.

Weerasena:2023:SAR

[WST^{+}23] Lakmali Weerasena, Douglas Shier, David Tonkyn, Mark McFeaters, and Christopher Collins. A sequential approach to reserve design with compactness and contiguity considerations. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000091.

Wang:2023:MWD

[WSW23] Binbin Wang, Lauren L. Sullivan, and Jeffrey D. Wood. Modeling wind-driven seed dispersal using a coupled Lagrangian particle tracking and 1-D k–ϵ turbulence model. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002338.

Wang:2024:MES

[WSZ^{+}24] Binbin Wang, Brandon J. Sansom, Wenyu Zhu, James Kunz, M. Christopher Barnhart, Henry Brown, Stephen Mc-Murray, Andrew D. Roberts, Christopher Shulse, Caleb J. Knerr, Kathleen Trauth, Jeffery A. Steevens, and Baolin Deng. A model for evaluation of sediment exposure and burial for freshwater mussels from heavy particle sedimentation. Ecological Modelling, 493(??):??, July 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002400139X.

Wiltshire:2020:CME

[WT20] Kathryn H. Wiltshire and Jason E. Tanner. Comparing maximum entropy modelling methods to inform aquaculture site selection for novel seaweed species. Ecological Modelling, 429(??):??, August 1, 2020. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380020301435.

Wirth:2021:DDM

[WTT^{+}21] Stephen Björn Wirth, Franziska Taubert, Britta Tietjen, Christoph Müller, and Susanne Rolinski. Do details matter? Disentangling the processes related to plant species interactions in two grassland models of different complexity. Ecological Modelling, 460(??):??, November 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002878.

Wu:2024:EPI

[Wu24] Haoran Wu. ecode: an R package to investigate community dynamics in ordinary differential equation systems. Ecological Modelling, 491(??):??, May 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024000644.

West:2022:MEM

[WWA^{+}22] Benjamin M. West, Mark L. Wildhaber, Kevin J. Aagaard, Wayne E. Thogmartin, Adrian P. Moore, and Michael J. Hooper. Migration and energetics model predicts delayed migration and likely starvation in oiled waterbirds. Ecological Modelling, 474(??):??, December 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002393.

Wang:2022:DSP

[WWC^{+}22] Quan Wang, Haijun Wang, Ruihan Chang, Haoran Zeng, and Xuepiao Bai. Dynamic simulation patterns and spatiotemporal analysis of land-use/land-cover changes in the Wuhan metropolitan area, China. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021003884.

Wu:2024:AEB

[WWCT24] Chen-Fa Wu, Hsiao-Hsuan Wang, Szu-Hung Chen, and Luu Van Thong Trac. Assessing the efficiency of bird habitat conservation strategies in farmland ecosystems. Ecological Modelling, 492(??):??, June 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380024001200.

Weller:2023:ABM

[WWF^{+}23] Florian G. Weller, Elisabeth B. Webb, Sean Fogenburg, William S. Beatty, Dylan Kesler, Robert H. Blenk, Kevin M. Ringelman, Matt L. Miller, Céline Arzel, and John M. Eadie. An agent-based model to quantify energetics, movement and habitat selection of mid-continent mallards in the Mississippi Alluvial Valley. Ecological Modelling, 485(??):??, November 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023002181.

Whitlock:2020:MPA

[WWP20] Steven L. Whitlock, Jamie N. Womble, and James T. Peterson. Modelling pinniped abundance and distribution by combining counts at terrestrial sites and in-water sightings. Ecological Modelling, 420(??):??, March 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300375.

Wang:2022:MER

[WXH^{+}22] Wenhui Wang, Feng Xu, Yunxin Huang, Hongqiang Feng, and Peng Wan. Modeling the evolution of resistance in cotton bollworm to concurrently planted Bt cotton and Bt maize in China. Ecological Modelling, 467(??):??, May 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000369.

Wu:2020:SVG

[WXT^{+}20] Dandan Wu, Xianhong Xie, Juxiu Tong, Shanshan Meng, and Yibing Wang. Sensitivity of vegetation growth to precipitation in a typical afforestation area in the Loess Plateau: Plantwater coupled modelling. Ecological Modelling, 430(??):??, August 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302003.

Wang:2023:IEI

[WXX^{+}23] Jing Wang, Lianqing Xue, Chenguang Xiang, Xinghan Li, and Lei Xie. Improving ecological indicators of arid zone deserts through simulation. Ecological Modelling, 485(??):??, November 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002363.

Wanlong:2023:ITB

[WYBX23] Sun Wanlong, Son Yowhan, He Baishuo, and Liu Xuehua. An individual tree-based model for estimating regional and temporal carbon storage of Abies chensiensis forest ecosystem in the Qinling Mountains, China. Ecological Modelling, 479(??): ??, May 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000339.

Wang:2022:RAE

[WYJ^{+}22] Zhiqiang Wang, Li Ye, Jingyi Jiang, Yida Fan, and Xiaoran Zhang. Review of application of EPIC crop growth model. Ecological Modelling, 467(??):??, May 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000722.

Wang:2021:SPC

[WZHP21] Hongxiang Wang, Xiaohong Zhang, Yanbo Hu, and Arne Pommerening. Spatial patterns of correlation between conspecific species and size diversity in forest ecosystems. Ecological Modelling, 457(??):??, October 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002362.

Wang:2024:EEC

[WZLC24] Lin Wang, Junsan Zhao, Yilin Lin, and Guoping Chen. Exploring ecological carbon sequestration advantage and economic responses in an ecological security pattern: a naturebased solutions perspective. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023003277.

Wang:2020:MSA

[WZWF20] Xin Wang, Lan Zhao, Rui-Wu Wang, and Sheng-Guo Fang. Mate success affects sex ratio strategies in structured population. Ecological Modelling, 429(??):??, August 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301769.

Wu:2023:RPU

[WZZF23] Dongdong Wu, Yan Zhang, Xiaolin Zhang, and Brain D. Fath. Research progress of urban nitrogen cycle and metabolism. Ecological Modelling, 486(??):??, December 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002387.

Xiao:2021:CMS

[XCH^{+}21] Hui Xiao, Iadine Chadès, Narelle Hill, Nicholas Murray, Richard A. Fuller, and Eve McDonald-Madden. Conserving migratory species while safeguarding ecosystem services. Ecological Modelling, 442(??):??, February 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002100017X.

Xing:2020:EII

[XCZ^{+}20] Lei Xing, Yong Chen, Chongliang Zhang, Bai Li, Kisei R. Tanaka, Robert Boenish, and Yiping Ren. Evaluating impacts of imprecise parameters on the performance of an ecosystem model OSMOSE-JZB. Ecological Modelling, 419(??):??, March 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019304314.

Xie:2021:PDG

[XFK21] Congbo Xie, Meng Fan, and Yun Kang. Population dynamics of the giant jellyfish Nemopilema nomurai with age structure. Ecological Modelling, 441(??):??, February 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020304713.

Xu:2024:LML

[XGT^{+}24] Linhao Xu, Ashley B. C. Goode, Philip W. Tipping, Melissa C. Smith, Lyn A. Gettys, Brittany K. Knowles, Eileen Pokorny, Luz Salinas, and Donald L. DeAngelis. Less is more: Less herbicide does more when biological control is present in Pontederia crassipes. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002300296X.

Xu:2024:COE

[XHS^{+}24] Aokang Xu, Mengjun Hu, Jing Shi, Qingzhu Bai, and Xuehong Li. Construction and optimization of ecological network in inland river basin based on circuit theory, complex network and ecological sensitivity: a case study of Gansu section of Heihe River Basin. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023003083.

Xu:2024:DDD

[XHZH24] Peng Xu, Lintao Huang, Yuyang Zhang, and Hui Huang. Different degrees of decline in China’s fishing efforts are beneficial to the sustainable development of coral reef fisheries resources on Yongxing Island China. Ecological Modelling, 494(??):??, August 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380024001546.

Xu:2021:IMS

[XJW21] Meng Xu, Mengke Jiang, and Hua-Feng Wang. Integrating metabolic scaling variation into the maximum entropy theory of ecology explains Taylor’s law for individual metabolic rate in tropical forests. Ecological Modelling, 455(??):??, September 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021002155.

Xu:2023:SAE

[XLM^{+}23] Duo Xu, Gengyuan Liu, Fanxin Meng, Ningyu Yan, Hui Li, Feni Agostinho, Cecilia MVB Almeida, and Biagio F. Giannetti. Sector aggregation effect on embodied carbon emission based on city-centric global multi-region input-output (CCG-MRIO) model. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002300217X.

Xie:2020:ICC

[XLS20] Yalin Xie, Xiangdong Lei, and Jingning Shi. Impacts of climate change on biological rotation of Larix olgensis plantations for timber production and carbon storage in northeast China using the 3-PG_{mix} model. Ecological Modelling, 435(??):??, November 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303379.

Xue:2023:IKL

[XMWL23] Shaobo Xue, Bo Ma, Chenguang Wang, and Zhanbin Li. Identifying key landscape pattern indices influencing the NPP: a case study of the upper and middle reaches of the Yellow River. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001886.

Xie:2024:EZD

[XWX^{+}24] Yuqi Xie, Shu Wang, Shujiang Xiang, Zifang Wang, Yanyun Li, Zhongshu Wang, Mi Zhou, Ying Wang, and Ming Gao. Ecological zoning and dynamic assessment of effectiveness in the Three Gorges Reservoir Area, China. Ecological Modelling, 487(??):??, January 2024. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023002934.

Xie:2022:ABE

[XWZ^{+}22] Tian Xie, Meie Wang, Yuan Zhang, Changfeng Liu, Fei Lu, Shoukang Ding, Weiping Chen, and Suriyanarayanan Sarvajayakesavalu. Ambient background estimation of PAHs in urban soils: a case study in Macau, China. Ecological Modelling, 472(??):??, October 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022002113.

Xia:2021:MSS

[XYZR21] Shujuan Xia, Takashi Yamakawa, Chongliang Zhang, and Yiping Ren. A multispecies size-structured matrix model incorporating seasonal dynamics. Ecological Modelling, 453(??):??, August 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380021001745.

Xia:2022:ABP

[XZZ^{+}22] Rui Xia, Lei Zou, Yuan Zhang, Yongyong Zhang, Yan Chen, Chengjian Liu, Zhongwen Yang, and Shuqin Ma. Algal bloom prediction influenced by the Water Transfer Project in the Middle-lower Hanjiang River. Ecological Modelling, 463(??):??, January 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003598.

Young:2024:ROW

[YA24] Jesse D. Young and Alan A. Ager. Resource objective wildfire leveraged to restore old growth forest structure while stabilizing carbon stocks in the southwestern United States. Ecological Modelling, 488(??):??, February 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003034.

Yao:2021:ETA

[Yao21] Weiwei Yao. Ecohydraulic tools for aquatic fauna habitat and population status assessment, analysis and monitoring aimed at promoting integrated river management. Ecological Modelling, 456(??):??, September 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021002404.

Yang:2022:SDC

[YB22] Yinghui Yang and Liping Bao. Scale-dependent changes in species richness caused by invader competition. Ecological Modelling, 469(??):??, July 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001107.

Yu:2020:ISD

[YCI20] Hao Yu, Arthur R. Cooper, and Dana M. Infante. Improving species distribution model predictive accuracy using species abundance: Application with boosted regression trees. Ecological Modelling, 432(??):??, September 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302726.

Yang:2020:RPW

[YCK20] Xuebin Yang, Kelley A. Crews, and Peter Kedron. Response of potential woody cover of Texas savanna to climate change in the 21st century. Ecological Modelling, 431(??):??, September 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302489.

Yu:2020:CBM

[YD20] Zhe Yu and Qianbin Di. The coordination between maritime economies and marine carrying capacity and their spatiotemporal evolution in the cities of the Bohai Rim in China. Ecological Modelling, 438(??):??, December 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020302635.

Yurek:2024:VWB

[YDLT24] Simeon Yurek, Donald L. DeAngelis, Hyo Won Lee, and Stephen Tennenbaum. Visualizing wading bird optimal foraging decisions with aggregation behaviors using individualbased modeling. Ecological Modelling, 493(??):??, July 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000905.

Yan:2023:SEM

[YDQ^{+}23] Zhigang Yan, Buddhi Dayananda, Wei Qin, Jie Gao, Xueli Wang, and Igor Popovic. System evolution model of human disturbances in panda habitats. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380023002132.

Yurek:2021:MSM

[YEL^{+}21] Simeon Yurek, Mitchell J. Eaton, Romain Lavaud, R. Wilson Laney, Donald L. DeAngelis, William E. Pine, Megan La Peyre, Julien Martin, Peter Frederick, Hongqing Wang, Michael R. Lowe, Fred Johnson, Edward V. Camp, and Rua Mordecai. Modeling structural mechanics of oyster reef selforganization including environmental constraints and community interactions. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020304531.

Yin:2021:MDO

[YFW^{+}21] Liang Yin, Lijiang Fu, Hao Wu, Qian Xia, Yongnian Jiang, Jinglu Tan, and Ya Guo. Modeling dissolved oxygen in a crab pond. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002030449X.

Yan:2021:KFA

[YG21] Renhua Yan and Junfeng Gao. Key factors affecting discharge, soil erosion, nitrogen and phosphorus exports from agricultural polder. Ecological Modelling, 452(??):??, July 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021001514.

Yang:2021:AHS

[YHC^{+}21] Lu Yang, Jingming Hou, Long Cheng, Pan Wang, Zhanpeng Pan, Tian Wang, Yongyong Ma, Gao Xujun, Sun Jixin, and Na Liu. Application of habitat suitability model coupling with high-precision hydrodynamic processes. Ecological Modelling, 462(??):??, December 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003392.

Yuan:2024:SWC

[YJC^{+}24] Xin Yuan, Liang Jiao, Xichen Che, Jingjing Wu, Xuli Zhu, and Qian Li. Study on the water–carbon coupling coordination function on the eastern edge of the Qinghai–Tibet plateau. Ecological Modelling, 487(??):??, January 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023003022.

Yonow:2023:PMP

[YKZ^{+}23] Tania Yonow, Darren J. Kriticos, Myron P. Zalucki, Rory J. Mc Donnell, and Valerie Caron. Population modelling for pest management: a case study using a pest land snail and its fly parasitoid in Australia. Ecological Modelling, 482(??):??, August 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023001448.

Yacine:2022:SCP

[YL22]  Youssef Yacine and Nicolas Loeuille.  Stable coexis-

tence in plant-pollinator–herbivore communities requires balanced mutualistic vs antagonistic interactions. Ecological Modelling, 465(??):??, March 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021003951.

Yao:2023:MIC

[YLLL23] Yao Yao, Guang Li, Yanhua Lu, and Shuainan Liu. Modelling the impact of climate change and tillage practices on soil CO_{2} emissions from dry farmland in the Loess Plateau of China. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000042.

Ying:2021:SRD

[YLS^{+}21] Zhixia Ying, Yongjie Liu, Chen Shi, Zhichun Lan, Huan Zhang, Yizhen Liu, and Gang Ge. Simulating root distribution of plant individual with a three-dimensional model. Ecological Modelling, 455(??):??, September 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002100209X.

Yee:2021:PEL

[YPS^{+}21] Susan H. Yee, E. Paulukonis, C. Simmons, M. Russell, R. Fulford, L. Harwell, and L. M. Smith. Projecting effects of land use change on human well-being through changes in ecosystem services. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020304233.

Yang:2020:DEB

[YRK^{+}20] Tao Yang, Jeffrey S. Ren, Sebastiaan A. L. M. Kooijman, Xiujuan Shan, and Harry Gorfine. A dynamic energy budget model of Fenneropenaeus chinensis with applications for aquaculture and stock enhancement. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002030257X.

Yu:2023:ETS

[YSFX23] Shiyang Yu, Da Song, Meng Fan, and Congbo Xie. Effects of temperature and salinity on growth of Aurelia aurita. Ecological Modelling, 476(??):??, February 2023. CODEN ECMODT.

ISSN ???? URL http://www.sciencedirect.com/science/
article/pii/S0304380022003271.

Yeste:2024:SLT

[YSIB24] Antonio Yeste, Brad Seely, J. Bosco Imbert, and Juan A. Blanco. Sensitivity of long-term productivity estimations in mixed forests to uncertain parameters related to fine roots. Ecological Modelling, 490(??):??, April 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000589.

Yu:2020:IEN

[Yu20] Rui Yu. An improved estimation of net primary productivity of grassland in the Qinghai–Tibet region using light use efficiency with vegetation photosynthesis model. Ecological Modelling, 431(??):??, September 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020301939.

Yang:2020:SMM

[YXH20] Xinbin Yang, Xinming Xu, and Dawei Hu. Succession mechanism of microbial community with high species diversity in nutrient-deficient environments with low-dose ionizing radiation. Ecological Modelling, 435(??):??, November 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303409.

Yi:2020:SHM

[YXY^{+}20] Yujun Yi, Hongyi Xie, Yufeng Yang, Yang Zhou, and Zhifeng Yang. Suitable habitat mathematical model of common reed (Phragmites australis) in shallow lakes with coupling cellular automaton and modified logistic function. Ecological Modelling, 419(??):??, March 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300090.

Yang:2022:ESI

[YYS^{+}22] Songxi Yang, Jian Yang, Shuo Shi, Shalei Song, Yangyang Zhang, Yi Luo, and Lin Du. An exploration of solarinduced chlorophyll fluorescence (SIF) factors simulated by SCOPE for capturing GPP across vegetation types. Ecological Modelling, 472(??):??, October 2022. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380022001843.

Ye:2020:EIV

[YZF20] Sufen Ye, Luoping Zhang, and Huan Feng. Ecosystem intrinsic value and its evaluation. Ecological Modelling, 430(??):??, August 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020302039.

Yang:2022:CSC

[YZZF22] Meng Yang, Xiaolin Zhang, Yan Zhang, and Brian D. Fath. Consistence of structural changes in food nitrogen consumption between rural and urban residents in the context of rapid urbanization. Ecological Modelling, 471(??):??, September 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380022001661.

Zajac:2020:ICW

[ZA20] P. Zaja̧c and A. Avdiushchenko. The impact of converting waste into resources on the regional economy, evidence from Poland. Ecological Modelling, 437(??):??, December 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020303690.

Zhang:2022:SEG

[ZBH^{+}22] Hongyuan Zhang, William D. Batchelor, Kelin Hu, Hao Liang, Hui Han, and Ji Li. Simulation of N_{2}O emissions from greenhouse vegetable production under different management systems in North China. Ecological Modelling, 470(??):??, August 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022001302.

Zhang:2024:EAL

[ZBW^{+}24] Yufei Zhang, Zhenxing Bian, Shuai Wang, Xiaoyu Guo, and Wei Zhou. Effect of agricultural landscape pattern on the qualitative food web of epigaeic arthropods in low hilly areas of northern China. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023003046.

Zhang:2020:IIM

[ZCJ^{+}20] Minxia Zhang, Shulin Chen, Hong Jiang, Changhui Peng, Jinmeng Zhang, and Guomo Zhou. The impact of intensive management on net ecosystem productivity and net primary productivity of a Lei bamboo forest. Ecological Modelling, 435(??):??, November 1, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303185.

Zhu:2023:MDT

[ZFC^{+}23] Shihua Zhu, Xia Fang, Liangzhong Cao, Xin Hang, Xiaoping Xie, Liangxiao Sun, and Yachun Li. Multivariate drives and their interactive effects on the ratio of transpiration to evapotranspiration over Central Asia ecosystems. Ecological Modelling, 478(??):??, April 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000224.

Zhu:2021:CWI

[ZFP21] Gengping Zhu, Jingyu Fan, and A. Townsend Peterson. Cautions in weighting individual ecological niche models in ensemble forecasting. Ecological Modelling, 448(??):??, May 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021000739.

Zhang:2020:PCC

[ZFT20] Jien Zhang, Benjamin S. Felzer, and Tara J. Troy. Projected changes of carbon balance in mesic grassland ecosystems in response to warming and elevated CO_{2} using CMIP5 GCM results in the Central Great Plains, USA. Ecological Modelling, 434(??):??, October 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020303173.

Zheng:2022:SGZ

[ZGG22] Yanlin Zheng, Xiang Gong, and Huiwang Gao. Selective grazing of zooplankton on phytoplankton defines rapid algal succession and blooms in oceans. Ecological Modelling, 468(??): ??, June 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000679.

Zuluaga-Guerra:2023:SEM

[ZGMFESD23] Paula Andrea Zuluaga-Guerra, Julia Martinez-Fernandez, Miguel Angel Esteve-Selma, and Jampel Dell’Angelo. A socio-ecological model of the Segura River basin, Spain. Ecological Modelling, 478(??):??, April 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000121.

Zhu:2023:TCC

[ZHC23] Benhui Zhu, Shizuka Hashimoto, and Samuel A. Cushman. A two concentric circles model incorporating availability of ecosystem services and affordability of humans to clarify the ecological security concept. Ecological Modelling, 481(??):??, July 2023. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023000716.

Zhang:2023:CES

[ZHJQ23] Ze Zhang, Baoqing Hu, Weiguo Jiang, and Haihong Qiu. Construction of ecological security pattern based on ecological carrying capacity assessment 1990–2040: a case study of the Southwest Guangxi Karst — Beibu Gulf. Ecological Modelling, 479(??):??, May 2023. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380023000509.

Zhou:2023:NST

[ZHWZ23] Ye Zhou, Chen Huang, Tao Wu, and Mingyue Zhang. A novel spatio-temporal cellular automata model coupling partitioning with CNN-LSTM to urban land change simulation. Ecological Modelling, 482(??):??, August 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023001254.

Zare:2024:BPS

[ZJEG24] Fateme Zare, Anthony J. Jakeman, Sondoss Elsawah, and Joseph H. A. Guillaume. Bridging practice and science in socio-environmental systems research and modelling: a design science approach. Ecological Modelling, 492(??):??, June 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380024001078.

Zhang:2024:ISP

[ZL24] Xiaojun Zhang and Baohuan Li. Information sharing promotes bacterial diversity in oligotrophic environment with low-dose X-ray radiation based on modeling and simulation of agent-based model. Ecological Modelling, 488(??):??, February 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380023003319.

Zhang:2023:CMF

[ZLL^{+}23] Pengyi Zhang, Yu Liang, Bo Liu, Tianxiao Ma, and Mia M. Wu. A coupled modelling framework for predicting tree species’ altitudinal migration velocity in montane forest. Ecological Modelling, 484(??):??, October 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023002119.

Zhang:2024:DSP

[ZLLZ24] Zhengrong Zhang, Xuemei Li, Xinyu Liu, and Kaixin Zhao. Dynamic simulation and projection of land use change using system dynamics model in the Chinese Tianshan mountainous region, central Asia. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380023002946.

Zheng:2020:EFE

[ZLML20] Hongmei Zheng, Aimin Li, Fanxin Meng, and Gengyuan Liu. Energy flows embodied in China’s interregional trade: Case study of Hebei Province. Ecological Modelling, 428(??):??, July 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020301332.

Zhu:2024:DSS

[ZLZ^{+}24] Zhizhuo Zhu, Long Li, Shangxuan Zhang, Peng Zhang, Yanan Ren, and Yu Zhang. Dynamic simulation study of soil erosion intensity on slopes with different vegetation patterns in pisha sandstone area. Ecological Modelling, 491(??):??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438002400053X.

Zhang:2020:INA

[ZMF^{+}20] Yujin Zhang, Minna Ma, Huajun Fang, Dahe Qin, Shulan Cheng, and Wenping Yuan. Impacts of nitrogen addition on nitrous oxide emission: Comparison of five nitrous oxide modules or algorithms. Ecological Modelling, 421(??):??, April 1, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020300351.

Zaatour:2021:ABM

[ZMG^{+}21] Wajdi Zaatour, Nicolas Marilleau, Patrick Giraudoux, Nadège Martiny, Abdesslem Ben Haj Amara, and Slimane Ben Miled. An agent-based model of a cutaneous leishmaniasis reservoir host, Meriones shawi. Ecological Modelling, 443(??):??, March 1, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380021000259.

Zadehdabagh:2022:SAI

[ZMK^{+}22] Nasim Zadehdabagh, Seyed Masoud Monavari, Nargess Kargari, Lobat Taghavi, and Saeid Pirasteh. Sustainability of agroecosystems by indices: a comparative study between indicators of ecological footprint sustainability and emergy analysis; a case study in Dez catchment, Iran. Ecological Modelling, 474(??):??, December 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022002666.

Zakharova:2020:CTI

[ZMS20] L. Zakharova, K. M. Meyer, and M. Seifan. Combining traitand individual-based modelling to understand desert plant community dynamics. Ecological Modelling, 434(??):??, October 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020303306.

Zamponi:2023:UAD

[ZOJ^{+}23] Virginia Zamponi, Kevin O’Brien, Erik Jensen, Brandon Feldhaus, Russell Moore, Christopher J. Lynch, and Ross Gore. Understanding and assessing demographic (in)equity resulting from extreme heat and direct sunlight exposure due to lack of tree canopies in Norfolk, VA using agent-based modeling. Ecological Modelling, 483(??):??, September 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002300176X.

Zhang:2022:ENN

[ZPZ^{+}22] Kerou Zhang, Changhui Peng, Qiuan Zhu, Mingxu Li, Zhongqing Yan, Meng Li, Liang Yan, Xiaodong Zhang, Jinzhi Wang, Yong Li, Enze Kang, Hanxiong Song, and Xiaoming Kang. Estimating natural nitrous oxide emissions from the Qinghai–Tibetan Plateau using a processbased model: Historical spatiotemporal patterns and future trends. Ecological Modelling, 466(??):??, April 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438002200028X.

Zhao:2020:MSP

[ZRMW^{+}20] Xiaochen Zhao, Victor H. Rivera-Monroy, Hongqing Wang, Z. George Xue, Cheng-Feng Tsai, Clinton S. Willson, Edward Castañeda-Moya, and Robert R. Twilley. Modeling soil porewater salinity in mangrove forests (Everglades, Florida, USA) impacted by hydrological restoration and a warming climate. Ecological Modelling, 436(??):??, November 15, 2020. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380020303628.

Zylstra:2020:BSS

[ZS20] Erin R. Zylstra and Robert J. Steidl. A Bayesian statespace model for seasonal growth of terrestrial vertebrates. Ecological Modelling, 420(??):??, March 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380020300478.

Zhang:2020:MGD

[ZSS20] Gengxi Zhang, Xiaoling Su, and Vijay P. Singh. Modelling groundwater-dependent vegetation index using entropy theory. Ecological Modelling, 416(??):??, January 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019304247.

Zhang:2021:SDC

[ZSS21] Qian Zhang, Juqin Shen, and Fuhua Sun. Spatiotemporal differentiation of coupling coordination degree between economic development and water environment and its influencing factors using GWR in China’s province. Ecological Modelling, 462(??):??, December 15, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003409.

Zhang:2021:BNF

[ZWL^{+}21] Quanzhong Zhang, Haiyan Wei, Jing Liu, Zefang Zhao, Qiao Ran, and Wei Gu. A Bayesian network with fuzzy mathematics for species habitat suitability analysis: a case with limited Angelica sinensis (Oliv.) Diels data. Ecological Modelling, 450(??):??, June 15, 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380021001307.

Zou:2020:SAB

[ZWZ^{+}20] Rui Zou, Zhen Wu, Lei Zhao, James J. Elser, Yanhong Yu, Yihui Chen, and Yong Liu. Seasonal algal blooms support sediment release of phosphorus via positive feedback in a eutrophic lake: Insights from a nutrient flux tracking modeling. Ecological Modelling, 416(??):??, January 15, 2020. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303898.

Zheng:2024:NIC

[ZXK^{+}24] Zhoumin Zheng, Nuo Xu, Mohsin Khan, Michael Pedersen, Tarteel Abdalgader, and Lai Zhang. Nonlinear impacts of climate change on dengue transmission in mainland China: Underlying mechanisms and future projection. Ecological Modelling, 492(??):??, June 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001224.

Zhou:2022:DUG

[ZYK22] Ting Zhou, Xi Yang, and Xinli Ke. Delimitation of urban growth boundaries by integratedly incorporating ecosystem conservation, cropland protection and urban compactness. Ecological Modelling, 468(??):??, June 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022000813.

Zhao:2021:RBA

[ZYM^{+}21] Chunli Zhao, Yan Yan, Wenyong Ma, Xu Shang, Jianguo Chen, Yuejing Rong, Tian Xie, and Yuan Quan. RESTRENDbased assessment of factors affecting vegetation dynamics on the Mongolian Plateau. Ecological Modelling, 440(??):??, January 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380020304749.

Zeng:2022:EIP

[ZYZ22] Yong Zeng, Wei Yang, and Yanwei Zhao. Ecological impact of polycyclic aromatic hydrocarbons on Baiyangdian Lake based on an ecosystem model. Ecological Modelling, 472(??):??, October 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438002200206X.

Zhu:2024:MCE

[ZZG^{+}24] Ke-Hua Zhu, Jian Zeng, Zhen-Ming Ge, Yin Zuo, Shi-Hua Li, Lei-Hua Zhao, Yu Han, Hai-Feng Cheng, and Pei Xin. A model coupling ecological and hydrodynamic processes for simulating the biogeomorphology of a coastal salt marsh. Ecological Modelling, 493(??):??, July 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024001467.

Zhou:2024:SGP

[ZZLF24] Junzhi Zhou, Ting Zhang, Jianzhu Li, and Ping Feng. Simulation of gross primary productivity and impact of drought in Liulin watershed of Taihang mountains over 2000– 2020. Ecological Modelling, 489(??):??, March 2024. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380024000188.

Zhang:2023:OSL

[ZZQ23] Yi Zhang, Minrui Zheng, and Bo Qin. Optimization of spatial layout based on ESV-FLUS model from the perspective of “Production-Living-Ecological”: a case study of Wuhan City. Ecological Modelling, 481(??):??, July 2023. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380023000844.

Zhang:2022:EMI

[ZZŠ^{+}22] Chen Zhang, Zixuan Zhu, Maria Špoljar, Natalia Kuczyńska-Kippen, Tvrtko Dražina, Matija Cvetnić, and Mirosław Mleczek. Ecosystem models indicate zooplankton biomass response to nutrient input and climate warming is related to lake size. Ecological Modelling, 464(??):??, February 2022. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021003793.

Zhang:2022:RDP

[ZZW^{+}22] Yajuan Zhang, Lijin Zhang, Huan Wang, Yueyao Wang, Jiaqi Ding, Jiashu Shen, Zheng Wang, Yinglu Liu, Chenyu Liang, and Shuangcheng Li. Reconstructing deforestation patterns in China from 2000 to 2019. Ecological Modelling, 465(??):??, March 2022. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380022000035.

Zhang:2021:RPH

[ZZWF21] Xiaolin Zhang, Yan Zhang, Yifan Wang, and Brian D. Fath. Research progress and hotspot analysis for reactive nitrogen flows in macroscopic systems based on a CiteSpace analysis. Ecological Modelling, 443(??):??, March 1, 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380021000284.

Zhang:2022:DIT

[ZZZD22] Xinyi Zhang, Qiande Zhu, Aifeng Zhai, and Xiaowen Ding. Development of interval transient pollution distribution model and its application in the fenghuangshan drinking water source. Ecological Modelling, 471(??):??, September 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380022001478.